Programming in Emacs Lisp An Introduction

Programming in Emacs Lisp

An Introduction

Edition 1.05, 21 October 1997

by Robert J. Chassell

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1990, ’91, ’92, ’93, ’94, ’95, ’97 Free Software Foundation, Inc.

Published by the Free Software Foundation, Inc. 59 Temple Place, Suite 330 Boston, MA 02111-1307 USA

ISBN-1882114-41-8

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies.

Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one.

Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by the Free Software Foundation.

Preface

Most of the GNU Emacs text editor is written in the programming language called Emacs Lisp. The code written in this programming language is the software—the sets of instructions— that tell the computer what to do when you give it commands. Emacs is designed so that you can write new code in Emacs Lisp and easily install it as an extension to the editor. This is why Emacs is called the “extensible editor”.

(Indeed, since Emacs does so much more than provide editing capabilities, it should perhaps be called an “extensible computing environment”, but that phrase is quite a mouthful. Also, everything you do in Emacs—find the Mayan date and phases of the moon, simplify polynomials, debug code, manage files, read letters, write books—all these activities are kinds of editing in the most general sense of the word.)

Although Emacs Lisp is usually thought of in association with the text editor, it is a full computer programming language. You can use it as you would any other programming language.

Perhaps you want to understand programming; perhaps you want to extend Emacs; or perhaps you want to become a programmer. This introduction to Emacs Lisp is designed to get you started: to guide you in learning the fundamentals of programming, and more importantly, to show you how you can teach yourself to go further.

All through this document, you will see little sample programs you can run inside of Emacs. If you read this document in Info inside of GNU Emacs, you can run the programs as they appear. (This is easy to do and is explained when the examples are presented.) Alternatively, you can read this introduction as a printed book while sitting beside a computer running Emacs. (This is what I like to do; I like printed books.) If you don’t have a running Emacs beside you, you can still read this book, but in this case, it is best to treat it as a novel or as a travel guide to a country not yet visited: interesting, but not the same as being there.

Much of this introduction is dedicated to walk-throughs or guided tours of code used in GNU Emacs. These tours are designed for two purposes: first, to give you familiarity with real, working code (code you use every day); and, second, to give you familiarity with the way Emacs works. It is interesting to see how an editor is implemented. Also, I hope that you will pick up the habit of browsing through source code. You can learn from it and mine it for ideas. Having GNU Emacs is like having a dragon’s cave of treasures.

In addition to learning about Emacs as an editor and Emacs Lisp as a programming language, the examples and guided tours will give you an opportunity to get acquainted with Emacs as a Lisp programming environment. GNU Emacs supports programming and provides tools that you will want to become comfortable using, such as M-. (the key which invokes the find-tag command). You will also learn about buffers and other objects that are part of the editing environment. Learning about these features of Emacs is like learning new routes around your home town.

Finally, I hope to convey some of the skills for using Emacs to learn aspects of programming that you don’t know. You can often use Emacs to help you understand what puzzles you or to find out how to do something new. This self-reliance is not only a pleasure, but an advantage.

For Whom This is Written

This text is written as an elementary introduction for people who are not programmers. If you are a programmer, you may not be satisfied with this primer. The reason is that you may have become expert at reading reference manuals and be put off by the way this text is organized.

An expert programmer who reviewed this text said to me:

I prefer to learn from reference manuals. I “dive into” each paragraph, and “come up for air” between paragraphs.

When I get to the end of a paragraph, I assume that that subject is done, finished, that I know everything I need (with the possible exception of the case when the next paragraph starts talking about it in more detail). I expect that a well written reference manual will not have a lot of redundancy, and that it will have excellent pointers to the (one) place where the information I want is.

This introduction is not written for this person!

Firstly, I try to say everything at least three times: first, to introduce it; second, to show it in context; and third, to show it in a different context, or to review it.

Secondly, I hardly ever put all the information about a subject in one place, much less in one paragraph. To my way of thinking, that imposes too heavy a burden on the reader. Instead I try to explain only what you need to know at the time. (Sometimes I include a little extra information so you won’t be surprised later when the additional information is formally introduced.)

When you read this text, you are not expected to learn everything the first time. Frequently, you need only make, as it were, a ‘nodding acquaintance’ with some of the items mentioned. My hope is that I have structured the text and given you enough hints that you will be alert to what is important, and concentrate on it.

You will need to “dive into” some paragraphs; there is no other way to read them. But I have tried to keep down the number of such paragraphs. This book is intended as an approachable hill, rather than as a daunting mountain.

This introduction to Programming in Emacs Lisp has a companion document, The GNU Emacs Lisp Reference Manual. The reference manual has more detail than this introduction. In the reference manual, all the information about one topic is concentrated in one place. You should turn to it if you are like the programmer quoted above. And, of course, after you have read this Introduction, you will find the Reference Manual useful when you are writing your own programs.

Lisp History

Lisp was first developed in the late 1950s at the Massachusetts Institute of Technology for research in artificial intelligence. The great power of the Lisp language makes it superior for other purposes as well, such as writing editor commands.

GNU Emacs Lisp is largely inspired by Maclisp, which was written at MIT in the 1960’s. It is somewhat inspired by Common Lisp, which became a standard in the 1980s. However, Emacs Lisp is much simpler than Common Lisp. (The standard Emacs distribution contains an optional extensions file, ‘cl.el’, that adds many Common Lisp features to Emacs Lisp.)

A Note for Novices

If you don’t know GNU Emacs, you can still read this document profitably. However, I recommend you learn Emacs, if only to learn to move around your computer screen. You can teach yourself how to use Emacs with the on-line tutorial. To use it, type C-h t. (This means you press and release the ^{⟨}\underline{\overline{CTRL}}^{⟩} key and the h at the same time, and then press and release t.)

Also, I often refer to one of Emacs’s standard commands by listing the keys which you press to invoke the command and then giving the name of the command in parentheses, like this: M-C-\ (indent-region). What this means is that the indent-region command is customarily invoked by typing M-C-\. (You can, if you wish, change the keys that are typed to invoke the command; this is called rebinding. See Section 16.11, “Keymaps”, page 157.) The abbreviation M-C-\ means that you type your ^{⟨}\underline{\overline{META}}^{⟩} key, ^{⟨}\underline{\overline{CTRL}}^{⟩} key and ^{⟨}\underline{\overline{\}}^{⟩} key all at the same time. Sometimes a combination like this is called a keychord, since it is similar to the way you play a chord on a piano. If your keyboard does not have a ^{⟨}\underline{\overline{META}}^{⟩} key, the ^{⟨}\underline{\overline{ESC}}^{⟩} key prefix is used in place of it. In this case, M-C-\ means that you press and release your ^{⟨}\underline{\overline{ESC}}^{⟩} key and then type the ^{⟨}\underline{\overline{CTRL}}^{⟩} key and the ^{⟨}\underline{\overline{\}}^{⟩} key at the same time.

If you are reading this in Info using GNU Emacs, you can read through this whole document just by pressing the space bar, ^{⟨}\underline{\overline{SPC}}^{⟩}. (To learn about Info, type C-h i and then select Info.)

A note on terminology: when I use the word Lisp alone, I am usually referring to the various dialects of Lisp in general, but when I speak of Emacs Lisp, I am referring to GNU Emacs Lisp in particular.

Thank You

My thanks to all who helped me with this book. My especial thanks to Jim Blandy, Noah Friedman, Jim Kingdon, Roland McGrath, Frank Ritter, Randy Smith, Richard M. Stallman, and Melissa Weisshaus. My thanks also go to both Philip Johnson and David Stampe for their patient encouragement. My mistakes are my own.

1 List Processing

To the untutored eye, Lisp is a strange programming language. In Lisp code there are parentheses everywhere. Some people even claim that the name stands for ‘Lots of Isolated Silly Parentheses’. But the claim is unwarranted. Lisp stands for LISt Processing and the programming language handles lists (and lists of lists) by putting them between parentheses. The parentheses mark the boundaries of the list. Sometimes a list is preceded by a single apostrophe or quotation mark, ‘'’. Lists are the basis of Lisp.

1.1 Lisp Lists

In Lisp, a list looks like this: '(rose violet daisy buttercup). This list is preceded by a single apostrophe. It could just as well be written as follows, which looks more like the kind of list you are likely to be familiar with:

     '(rose violet daisy buttercup)

The elements of this list are the names of the four different flowers, separated from each other by whitespace and surrounded by parentheses, like flowers in a field with a stone wall around them.

Lists can also have numbers in them, as in this list: (+ 2 2). This list has a plus-sign, ‘+’, followed by two ‘2’s, each separated by whitespace.

In Lisp, both data and programs are represented the same way; that is, they are both lists of words, numbers, or other lists, separated by whitespace and surrounded by parentheses. (Since a program looks like data, one program may easily serve as data for another; this is a very powerful feature of Lisp.) (Incidentally, these two parenthetical remarks are not Lisp lists, because they contain ‘;’ and ‘.’ as punctuation marks.)

Here is another list, this time with a list inside of it:

     '(this list has (a list inside of it))

The components of this list are the words ‘this’, ‘list’, ‘has’, and the list ‘(a list inside of it)’. The interior list is made up of the words ‘a’, ‘list’, ‘inside’, ‘of’, ‘it’.

1.1.1 Lisp Atoms

In Lisp, what we have been calling words are called atoms. This term comes from the historical meaning of the word atom, which means ‘indivisible’. As far as Lisp is concerned, the words we have been using in the lists cannot be divided into any smaller parts and still mean the same thing as part of a program; likewise with numbers and single character symbols like ‘+’. On the other hand, unlike an atom, a list can be split into parts. (See Chapter 7, “car cdr & cons Fundamental Functions”, page 59.)

In a list, atoms are separated from each other by whitespace. They can be right next to a parenthesis.

Technically speaking, a list in Lisp consists of parentheses surrounding atoms separated by whitespace or surrounding other lists or surrounding both atoms and other lists. A list can have just one atom in it or have nothing in it at all. A list with nothing in it looks like this: (), and is called the empty list. Unlike anything else, an empty list is considered both an atom and a list at the same time.

The printed representation of both atoms and lists are called symbolic expressions or, more concisely, s-expressions. The word expression by itself can refer to either the printed representation, or to the atom or list as it is held internally in the computer. Often, people use the term expression indiscriminately. (Also, in many texts, the word form is used as a synonym for expression.)

Incidentally, the atoms that make up our universe were named such when they were thought to be indivisible; but it has been found that physical atoms are not indivisible. Parts can split off an atom or it can fission into two parts of roughly equal size. Physical atoms were named prematurely, before their truer nature was found. In Lisp, certain kinds of atom, such as an array, can be separated into parts; but the mechanism for doing this is different from the mechanism for splitting a list. As far as list operations are concerned, the atoms of a list are unsplittable.

As in English, the meanings of the component letters of a Lisp atom are different from the meaning the letters make as a word. For example, the word for the South American sloth, the ‘ai’, is completely different from the two words, ‘a’, and ‘i’.

There are many kinds of atom in nature but only a few in Lisp: for example, numbers, such as 37, 511, or 1729, and symbols, such as ‘+’, ‘foo’, or ‘forward-line’. The words we have listed in the examples above are all symbols. In everyday Lisp conversation, the word “atom” is not often used, because programmers usually try to be more specific about what kind of atom they are dealing with. Lisp programming is mostly about symbols (and sometimes numbers) within lists. (Incidentally, the preceding three word parenthetical remark is a proper list in Lisp, since it consists of atoms, which in this case are symbols, separated by whitespace and enclosed by parentheses, without any non-Lisp punctuation.)

In addition, text between double quotation marks—even sentences or paragraphs—is an atom. Here is an example:

     '(this list includes "text between quotation marks.")

In Lisp, all of the quoted text including the punctuation mark and the blank spaces is a single atom. This kind of atom is called a string (for ‘string of characters’) and is the sort of thing that is used for messages that a computer can print for a human to read. Strings are a different kind of atom than numbers or symbols and are used differently.

1.1.2 Whitespace in Lists

The amount of whitespace in a list does not matter. From the point of view of the Lisp language,

     '(this list
        looks like this)

is exactly the same as this:

     '(this list looks like this)

Both examples show what to Lisp is the same list, the list made up of the symbols ‘this’, ‘list’, ‘looks’, ‘like’, and ‘this’ in that order.

Extra whitespace and newlines are designed to make a list more readable by humans. When Lisp reads the expression, it gets rid of all the extra whitespace (but it needs to have at least one space between atoms in order to tell them apart.)

Odd as it seems, the examples we have seen cover almost all of what Lisp lists look like! Every other list in Lisp looks more or less like one of these examples, except that the list may be longer and more complex. In brief, a list is between parentheses, a string is between quotation marks, a symbol looks like a word, and a number looks like a number. (For certain situations, square brackets, dots and a few other special characters may be used; however, we will go quite far without them.)

1.1.3 GNU Emacs Helps You Type Lists

If you type a Lisp expression in GNU Emacs using either Lisp Interaction mode or Emacs Lisp mode, you will have available to you several commands to format the Lisp expression so it is easy to read. For example, pressing the ^{⟨}\underline{\overline{TAB}}^{⟩} key automatically indents the line the cursor is on by the right amount. A command to properly indent the code in a region is customarily bound to M-C-\. Indentation is designed so that you can see which elements of a list belongs to which list—elements of a sub-list are indented more than the elements of the enclosing list.

In addition, when you type a closing parenthesis, Emacs momentarily jumps the cursor back to the matching opening parenthesis, so you can see which one it is. This is very useful, since every list you type in Lisp must have its closing parenthesis match its opening parenthesis. (See section “Major Modes” in The GNU Emacs Manual, for more information about Emacs’ modes.)

1.2 Run a Program

A list in Lisp—any list—is a program ready to run. If you run it (for which the Lisp jargon is evaluate), the computer will do one of three things: do nothing except return to you the list itself; send you an error message; or, treat the first symbol in the list as a command to do something. (Usually, of course, it is the last of these three things that you really want!)

The single apostrophe, ', that I put in front of some of the example lists in preceding sections is called a quote; when it precedes a list, it tells Lisp to do nothing with the list, other than take it as it is written. But if there is no quote preceding a list, the first item of the list is special: it is a command for the computer to obey. (In Lisp, these commands are called functions.) The list (+ 2 2) shown above did not have a quote in front of it, so Lisp understands that the + is an instruction to do something with the rest of the list; in this case, to add the numbers that follow.

If you are reading this inside of GNU Emacs in Info, here is how you can evaluate such a list: place your cursor immediately after the right hand parenthesis of the following list and then

type C-x C-e:
     (+ 2 2)

You will see the number 4 appear in the echo area. (In the jargon, what you have just done is “evaluate the list.” The echo area is the line at the bottom of the screen that displays or “echoes” text.) Now try the same thing with a quoted list: place the cursor right after the following list and type C-x C-e:

     '(this is a quoted list)

In this case, you will see (this is a quoted list) appear in the echo area.

In both cases, what you are doing is giving a command to the program inside of GNU Emacs called the Lisp interpreter—giving the interpreter a command to evaluate the expression. The name of the Lisp interpreter comes from the word for the task done by a human who comes up with the meaning of an expression—who “interprets” it.

You can also evaluate an atom that is not part of a list—one that is not surrounded by parentheses; again, the Lisp interpreter translates from the humanly readable expression to the language of the computer. But before discussing this (see Section 1.7, “Variables”, page 7), we will discuss what the Lisp interpreter does when you make an error.

1.3 Generate an Error Message

Partly so you won’t worry if you do it accidentally, we will now give a command to the Lisp interpreter that generates an error message. This is a harmless activity; and indeed, we will often try to generate error messages intentionally. Once you understand the jargon, error messages can be informative. Instead of being called “error” messages, they should be called “help” messages. They are like signposts to a traveller in a strange country; decyphering them can be hard, but once understood, they can point the way.

What we will do is evaluate a list that is not quoted and does not have a meaningful command as its first element. Here is a list almost exactly the same as the one we just used, but without the single-quote in front of it. Position the cursor right after it and type C-x C-e:

     (this is an unquoted list)

This time, you will see the following appear in the echo area:

     Symbol's function definition is void: this

(Also, your terminal may beep at you—some do, some don’t; and others blink. This is just a device to get your attention.) The message goes away as soon as you type another key, even just to move the cursor.

Based on what we already know, we can almost read this error message. We know the meaning of the word ‘Symbol’. In this case, it refers to the first atom of the list, the word ‘this’. The word ‘function’ was mentioned once before. It is a very important word. For our purposes, we can define it by saying that a function is a set of instructions to the computer that tell the computer to do something. (Technically, the symbol tells the computer where to find the instructions, but this is a complication we can ignore for the moment.)

Now we can begin to understand the error message: ‘Symbol's function definition is void: this’. The symbol (that is, the word ‘this’) does not have a definition of any set of instructions for the computer to carry out.

The slightly odd wording of the message, ‘function definition is void’, is designed to cover the way Emacs Lisp is implemented, which is that when the symbol does not have a function definition attached to it, the place that should contain the instructions is ‘void’.

On the other hand, since we were able to add 2 plus 2 successfully, by evaluating (+ 2 2), we can infer that the symbol + must have a set of instructions for the computer to obey and those instructions must be to add the numbers that follow the +.

1.4 Symbol Names and Function Definitions

We can articulate another characteristic of Lisp based on what we have discussed so far—an important characteristic: a symbol, like +, is not itself the set of instructions for the computer to carry out. Instead, the symbol is used, perhaps temporarily, as a way of locating the definition or set of instructions. What we see is the name through which the instructions can be found. Names of people work the same way. I can be referred to as ‘Bob’; however, I am not the letters ‘B’, ‘o’, ‘b’ but am the consciousness consistently associated with a particular life-form. The name is not me, but it can be used to refer to me.

In Lisp, one set of instructions can be attached to several names. For example, the computer instructions for adding numbers can be linked to the symbol plus as well as to the symbol + (and are in some dialects of Lisp). Among humans, I can be referred to as ‘Robert’ as well as ‘Bob’ and by other words as well.

On the other hand, a symbol can have only one function definition attached to it at a time. Otherwise, the computer would be confused as to which definition to use. If this were the case among people, only one person in the world could be named ‘Bob’. However, the function definition to which the name refers can be changed readily. (See Section 3.2, “Install a Function Definition”, page 23.)

Since Emacs Lisp is large, it is customary to name symbols in a way that identifies the part of Emacs to which the function belongs. Thus, all the names for functions that deal with Texinfo start with ‘texinfo-’ and those for functions that deal with reading mail start with ‘rmail-’.

1.5 The Lisp Interpreter

Based on what we have seen, we can now start to figure out what the Lisp interpreter does when we command it to evaluate a list. First, it looks to see whether there is a quote before the list; if there is, the interpreter just gives us the list. On the other hand, if there is no quote, the interpreter looks at the first element in the list and sees whether it has a function definition. If it does, the interpreter carries out the instructions in the function definition. Otherwise, the interpreter prints an error message.

This is how Lisp works. Simple. There are added complications which we will get to in a minute, but these are the fundamentals. Of course, to write Lisp programs, you need to know how to write function definitions and attach them to names, and how to do this without confusing either yourself or the computer.

Now, for the first complication. In addition to lists, the Lisp interpreter can evaluate a symbol that is not quoted and does not have parentheses around it. In this case, the Lisp interpreter will attempt to determine the symbol’s value as a variable. This situation is described in the section on variables. (See Section 1.7, “Variables”, page 7.)

The second complication occurs because some functions are unusual and do not work in the usual manner. Those that don’t are called special forms. They are used for special jobs, like defining a function, and there are not many of them. In the next few chapters, you will be introduced to several of the more important special forms.

The third and final complication is this: if the function that the Lisp interpreter is looking at is not a special form, and if it is part of a list, the Lisp interpreter looks to see whether the list has a list inside of it. If there is an inner list, the Lisp interpreter first figures out what it should do with the inside list, and then it works on the outside list. If there is yet another list embedded inside the inner list, it works on that one first, and so on. It always works on the innermost list first. The interpreter works on the innermost list first in order to find out the result of doing that. The result may be used by the enclosing expression.

Otherwise, the interpreter works left to right, from one expression to the next.

1.5.1 Byte Compiling

One other aspect of interpreting: the Lisp interpreter is able to interpret two kinds of entity: humanly readable code, on which we will focus exclusively, and specially processed code, called byte compiled code, which is not humanly readable. Byte compiled code runs faster than humanly readable code.

You can transform humanly readable code into byte compiled code by running one of the compile commands such as byte-compile-file. Byte compiled code is usually stored in a file that ends with a ‘.elc’ extension rather than a ‘.el’ extension. You will see both kinds of file in the ‘emacs/lisp’ directory; the files to read are those with ‘.el’ extensions.

As a practical matter, for most things you might do to customize or extend Emacs, you do not need to byte compile; and I will not discuss the topic here. See section “Byte Compilation” in The GNU Emacs Lisp Reference Manual, for a full description of byte compilation.

1.6 Evaluation

When the Lisp interpreter works on an expression, the term for the activity is called evaluation. We say that the interpreter ‘evaluates the expression’. I’ve used this term several times before. The word comes from its use in everyday language, ‘to ascertain the value or amount of; to appraise’, according to Webster’s New Collegiate Dictionary.

After evaluating an expression, the Lisp interpreter will most likely return the value that the computer produces by carrying out the instructions it found in the function definition, or perhaps it will give up on that function and produce an error message. (The interpreter may also find itself tossed, so to speak, to a different function or it may attempt to repeat continually what it is doing for ever and ever in what is called an ‘infinite loop’. These actions are less common; and we can ignore them.) Most frequently, the interpreter returns a value.

At the same time the interpreter returns a value, it may do something else as well, such as move a cursor or copy a file; this other kind of action is called a side effect. Actions that we humans think are important, such as printing results, are often “side effects” to the Lisp interpreter. The jargon can sound peculiar, but it turns out that it is fairly easy to learn to use side effects.

In summary, evaluating a symbolic expression most commonly causes the Lisp interpreter to return a value and perhaps carry out a side effect; or else produce an error.

1.6.1 Evaluating Inner Lists

If evaluation applies to a list that is inside another list, the outer list may use the value returned by the first evaluation as information when the outer list is evaluated. This explains why inner expressions are evaluated first: the values they return are used by the outer expressions.

We can investigate this process by evaluating another addition example. Place your cursor after the following expression and type C-x C-e:

     (+ 2 (+ 3 3))

The number 8 will appear in the echo area.

What happens is that the Lisp interpreter first evaluates the inner expression, (+ 3 3), for which the value 6 is returned; then it evaluates the outer expression as if it were written (+ 2 6), which returns the value 8. Since there are no more enclosing expressions to evaluate, the interpreter prints that value in the echo area.

Now it is easy to understand the name of the command invoked by the keystrokes C-x C-e: the name is eval-last-sexp. The letters sexp are an abbreviation for ‘symbolic expression’, and eval is an abbreviation for ‘evaluate’. The command means ‘evaluate last symbolic expression’.

As an experiment, you can try evaluating the expression by putting the cursor at the beginning of the next line immediately following the expression, or inside the expression.

Here is another copy of the expression:

     (+ 2 (+ 3 3))

If you place the cursor at the beginning of the blank line that immediately follows the expression and type C-x C-e, you will still get the value 8 printed in the echo area. Now try putting the cursor inside the expression. If you put it right after the next to last parenthesis (so it appears to sit on top of the last parenthesis), you will get a 6 printed in the echo area! This is because the command evaluates the expression (+ 3 3).

Now put the cursor immediately after a number. Type C-x C-e and you will get the number itself. In Lisp, if you evaluate a number, you get the number itself—this is how numbers differ from symbols. If you evaluate a list starting with a symbol like +, you will get a value returned that is the result of the computer carrying out the instructions in the function definition attached to that name. If a symbol by itself is evaluated, something different happens, as we will see in the next section.

1.7 Variables

In Lisp, a symbol can have a value attached to it just as it can have a function definition attached to it. The two are different. The function definition is a set of instructions that a computer will obey. A value, on the other hand, is something, such as number or a name, that can vary (which is why such a symbol is called a variable). The value of a symbol can be any expression in Lisp, such as a symbol, number, list, or string. A symbol that has a value is often called a variable.

A symbol can have both a function definition and a value attached to it at the same time. The two are separate. This is somewhat similar to the way the name Cambridge can refer to the city in Massachusetts and have some information attached to the name as well, such as “great programming center”.

Another way of thinking of this is to imagine a symbol as being a chest of drawers. The function definition is put in one drawer, the value in another, and so on. What is put in the drawer holding the value can be changed without affecting the contents of the drawer holding the function definition, and vice-versa.

The variable fill-column illustrates a symbol with a value attached to it: in every GNU Emacs buffer, this symbol is set to some value, usually 72 or 70, but sometimes to some other value. To find the value of this symbol, evaluate it by itself. If you are reading this in Info inside of GNU Emacs, you can do this by putting the cursor after the symbol and typing C-x C-e:

     fill-column

After I typed C-x C-e, Emacs printed the number 72 in my echo area. This is the value for which fill-column is set for me as I write this. It may be different for you in your Info buffer. Notice that the value returned as a variable is printed in exactly the same way as the value returned by a function carrying out its instructions. From the point of view of the Lisp interpreter, a value returned is a value returned. What kind of expression it came from ceases to matter once the value is known.

A symbol can have any value attached to it or, to use the jargon, we can bind the variable to a value: to a number, such as 72; to a string, "such as this"; to a list, such as (spruce pine oak); we can even bind a variable to a function definition.

A symbol can be bound to a value in several ways. See Section 1.9, “Setting the Value of a Variable”, page 12, for information about one way to do this.

Notice that there were no parentheses around the word fill-column when we evaluated it to find its value. This is because we did not intend to use it as a function name. If fill-column were the first or only element of a list, the Lisp interpreter would attempt to find the function definition attached to it. But fill-column has no function definition. Try evaluating this:

     (fill-column)

You will produce an error message that says:

     Symbol's function definition is void: fill-column

1.7.1 Error Message for a Symbol Without a Value

If you attempt to evaluate a symbol that does not have a value bound to it, you will receive an error message. You can see this by experimenting with our 2 plus 2 addition. In the following expression, put your cursor right after the +, before the first number 2, type C-x C-e:

     (+ 2 2)

You will get an error message that says:

     Symbol's value as variable is void: +

This is different from the first error message we saw, which said, ‘Symbol's function definition is void: this’. In this case, the symbol does not have a value as a variable; in the other case, the symbol (which was the word ‘this’) did not have a function definition.

In this experiment with the +, what we did was cause the Lisp interpreter to evaluate the + and look for the value of the variable instead of the function definition. We did this by placing the cursor right after the symbol rather than after the parenthesis of the enclosing list as we did before. As a consequence, the Lisp interpreter evaluated the preceding s-expression, which in this case was the + by itself.

Since + does not have a value bound to it, just the function definition, the error message reported that the symbol’s value as a variable was void.

1.8 Arguments

To see how information is passed to functions, let’s look again at our old standby, the addition of two plus two. In Lisp, this is written as follows:

     (+ 2 2)

If you evaluate this expression, the number 4 will appear in your echo area. What the Lisp interpreter does is add the numbers that follow the +.

The numbers added by + are called the arguments of the function +. These numbers are the information that is given to or passed to the function.

The word ‘argument’ comes from the way it is used in mathematics and does not refer to a disputation between two people; instead it refers to the information presented to the function, in this case, to the +. In Lisp, the arguments to a function are the atoms or lists that follow the function. The values returned by the evaluation of these atoms or lists are passed to the function. Different functions require different numbers of arguments; some functions require none at all.^{1}

1 It is curious to track the path by which the word ‘argument’ came to have two different meanings, one in mathematics and the other in everyday English. According to the Oxford English Dictionary, the word derives from the Latin for ‘to make clear, prove’; thus it came to mean, by one thread of derivation, ‘the evidence offered as proof’, which is to say, ‘the information offered’, which led to its meaning in Lisp. But in the other thread of derivation, it came to mean ‘to assert in a manner against which others may make counter assertions’, which led to the meaning of the word as a disputation. (Note here that the English word

1.8.1 Arguments’ Data Types

The type of data that should be passed to a function depends on what kind of information it uses. The arguments to a function such as + must have values that are numbers, since + adds numbers. Other functions use different kinds of data for their arguments.

For example, the concat function links together or unites two or more strings of text to produce a string. The arguments are strings. Concatenating the two character strings abc, def produces the single string abcdef. This can be seen by evaluating the following:

     (concat "abc" "def")

The value produced by evaluating this expression is "abcdef".

A function such as substring uses both a string and numbers as arguments. The function returns a part of the string, a substring of the first argument. This function takes three arguments. Its first argument is the string of characters, the second and third arguments are numbers that indicate the beginning and end of the substring. The numbers are a count of the number of characters (including spaces and punctuations) from the beginning of the string.

For example, if you evaluate the following:

     (substring "The quick brown fox jumped." 16 19)

you will see "fox" appear in the echo area. The arguments are the string and the two numbers.

Note that the string passed to substring is a single atom even though it is made up of several words separated by spaces. Lisp counts everything between the two quotation marks as part of the string, including the spaces. You can think of the substring function as a kind of ‘atom smasher’ since it takes an otherwise indivisible atom and extracts a part. However, substring is only able to extract a substring from an argument that is a string, not from another type of atom such as a number or symbol.

1.8.2 An Argument as the Value of a Variable or List

An argument can be a symbol that returns a value when it is evaluated. For example, when the symbol fill-column by itself is evaluated, it returns a number. This number can be used in an addition. Position the cursor after the following expression and type C-x C-e:

     (+ 2 fill-column)

The value will be a number two more than what you get by evaluating fill-column alone. For me, this is 74, because the value of fill-column is 72.

As we have just seen, an argument can be a symbol that returns a value when evaluated. In addition, an argument can be a list that returns a value when it is evaluated. For example, in the following expression, the arguments to the function concat are the strings "The " and

" red foxes." and the list (+ 2 fill-column).
     (concat "The " (+ 2 fill-column) " red foxes.")

If you evaluate this expression, "The 74 red foxes." will appear in the echo area. (Note that you must put spaces after the word ‘The’ and before the word ‘red’ so they will appear in the final string.)

has two different definitions attached to it at the same time. By contrast, in Emacs Lisp, a symbol cannot have two different function definitions at the same time.)

1.8.3 Variable Number of Arguments

Some functions, such as concat, + or *, take any number of arguments. (The * is the symbol for multiplication.) This can be seen by evaluating each of the following expressions in the usual way. What you will see in the echo area is printed in this text after ‘⇒’, which you may read as ‘evaluates to’.

In the first set, the functions have no arguments:

     (+)  ⇒ 0 (*)  ⇒ 1

In this set, the functions have one argument each:

     (+ 3)  ⇒ 3 (* 3)  ⇒ 3

In this set, the functions have three arguments each:

     (+ 3 4 5) ⇒ 12 (* 3 4 5) ⇒ 60

1.8.4 Using the Wrong Type Object as an Argument

When a function is passed an argument of the wrong type, the Lisp interpreter produces an error message. For example, the + function expects the values of its arguments to be numbers. As an experiment we can pass it the quoted symbol hello instead of a number. Position the cursor after the following expression and type C-x C-e:

     (+ 2 'hello)

When you do this you will generate an error message. What has happened is that + has tried to add the 2 to the value returned by 'hello, but the value returned by 'hello is the symbol hello, not a number. Only numbers can be added. So + could not carry out its addition.

As usual, the error message tries to be helpful and makes sense after you learn how to read it. What it says is this:

     Wrong type argument: integer-or-marker-p, hello

The first part of the error message is straightforward; it says ‘Wrong type argument’. Next comes the mysterious jargon word ‘integer-or-marker-p’. This word is trying to tell you what kind of argument the + expected.

The symbol integer-or-marker-p says that the Lisp interpreter is trying to determine whether the information presented it (the value of the argument) is an integer (that is, a whole number) or a marker (a special object representing a buffer position). What it does is test to see whether the + is being given whole numbers to add. It also tests to see whether the argument is something called a marker, which is a specific feature of Emacs Lisp. (In Emacs, locations in a buffer are recorded as markers. When the mark is set with the C-@ or C-^{⟨}\underline{\overline{SPC}}^{⟩} command, its position is kept as a marker. The mark can be considered a number—the number of characters the location is from the beginning of the buffer.) In Emacs Lisp, + can be used to add the numeric value of marker positions as numbers.

The ‘p’ of integer-or-marker-p is the embodiment of a practice started in the early days of Lisp programming. The ‘p’ stands for ‘predicate’. In the jargon used by the early Lisp researchers, a predicate refers to a function to determine whether some property is true or false.

So the ‘p’ tells us that integer-or-marker-p is the name of a function that determines whether it is true or false that the argument supplied is an integer or a marker. Other Lisp symbols that end in ‘p’ include zerop, a function that tests whether its argument has the value of zero, and listp, a function that tests whether its argument is a list.

Finally, the last part of the error message is the symbol hello. This is the value of the argument that was passed to +. If the addition had been passed the correct type of object, the value passed would have been a number, such as 37, rather than a symbol like hello. But then you would not have got the error message.

1.8.5 The message Function

Like +, the message function takes a variable number of arguments. It is used to send messages to the user and is so useful that we will describe it here.

A message is printed in the echo area. For example, you can print a message in your echo area by evaluating the following list:

     (message "This message appears in the echo area!")

The whole string between double quotation marks is a single argument and is printed in toto. (Note that in this example, the message itself will appear in the echo area within double quotes; that is because you see the value returned by the message function. In most uses of message in programs that you write, the text will be printed in the echo area as a side-effect, without the quotes. See Section 3.3.1, “multiply-by-seven in detail”, page 25, for an example of this.)

However, if there is a ‘%s’ in the quoted string of characters, the message function does not print the ‘%s’ as such, but looks to the argument that follows the string. It evaluates the second argument and prints the value in the location in the string where the ‘%s’ is.

You can see this by positioning the cursor after the following expression and typing C-x C-e:

     (message "The name of this buffer is: %s." (buffer-name))

In Info, "The name of this buffer is: *info*." will appear in the echo area. The function buffer-name returns the name of the buffer as a string, which the message function inserts in place of %s.

To print a value as a decimal number, use ‘%d’ in the same way as ‘%s’. For example, to print a message in the echo area that states the value of the fill-column, evaluate the following:

     (message "The value of fill-column is %d." fill-column)

On my system, when I evaluate this list, "The value of fill-column is 72." appears in my echo area.

If there is more than one ‘%s’ in the quoted string, the value of the first argument following the quoted string is printed at the location of the first ‘%s’ and the value of the second argument is printed at the location of the second ‘%s’, and so on. For example, if you evaluate the following,

     (message "There are %d %s in the office!"
              (- fill-column 14) "pink elephants")

a rather whimsical message will appear in your echo area. On my system it says, "There are

58 pink elephants in the office!".

The expression (- fill-column 14) is evaluated and the resulting number is inserted in place of the ‘%d’; and the string in double quotes, "pink elephants", is treated as a single argument and inserted in place of the ‘%s’. (That is to say, a string between double quotes evaluates to itself, like a number.)

Finally, here is a somewhat complex example that not only illustrates the computation of a number, but also shows how you can use an expression within an expression to generate the text that is substituted for ‘%s’:

     (message "He saw %d %s"
              (- fill-column 34)
              (concat "red "
                      (substring
                       "The quick brown foxes jumped." 16 21)
                      " leaping."))

In this example, message has three arguments: the string, "He saw %d %s", the expression, (- fill-column 32), and the expression beginning with the function concat. The value resulting from the evaluation of (- fill-column 32) is inserted in place of the ‘%d’; and the value returned by the expression beginning with concat is inserted in place of the ‘%s’.

When I evaluate the expression, the message, "He saw 38 red foxes leaping.", appears in my echo area.

1.9 Setting the Value of a Variable

There are several ways by which a variable can be given a value. One of the ways is to use either the function set or the function setq. Another way is to use let (see Section 3.6, “let”, page 27). (The jargon for this process is to bind a variable to a value.)

The following sections not only describe how set and setq work but also illustrate how arguments are passed.

1.9.1 Using set

To set the value of the symbol flowers to the list '(rose violet daisy buttercup), evaluate the following expression by positioning the cursor after the expression and typing C-x

C-e.
     (set 'flowers '(rose violet daisy buttercup))

The list (rose violet daisy buttercup) will appear in the echo area. This is what is returned by the set function. As a side effect, the symbol flowers is bound to the list ; that is, the symbol flowers, which can be viewed as a variable, is given the list as its value. (This process, by the way, illustrates how a side effect to the Lisp interpreter, setting the value, can be the primary effect that we humans are interested in. This is because every Lisp function must return a value if it does not get an error, but it will only have a side effect if it is designed to have one.)

After evaluating the set expression, you can evaluate the symbol flowers and it will return the value you just set. Here is the symbol. Place your cursor after it and type C-x C-e.

     flowers

When you evaluate flowers, the list (rose violet daisy buttercup) appears in the echo area.

Incidentally, if you evaluate 'flowers, the variable with a quote in front of it, what you will see in the echo area is the symbol itself, flowers. Here is the quoted symbol, so you can try this:

     'flowers

Note also, that when you use set, you need to quote both arguments to set, unless you want them evaluated. In this case, we do not want either argument evaluated, neither the variable flowers nor the list (rose violet daisy buttercup), so both are quoted. (When you use set without quoting its first argument, the first argument is evaluated before anything else is done. If you did this and flowers did not have a value already, you would get an error message that the ‘Symbol's value as variable is void’; on the other hand, if flowers did return a value after it was evaluated, the set would attempt to set the value that was returned. There are situations where this is the right thing for the function to do; but such situations are rare.)

1.9.2 Using setq

As a practical matter, you almost always quote the first argument to set. The combination of set and a quoted first argument is so common that it has its own name: the special form setq. This special form is just like set except that the first argument is quoted automatically, so you don’t need to type the quote mark yourself. Also, as an added convenience, setq permits you to set several different variables to different values, all in one expression.

To set the value of the variable carnivores to the list '(lion tiger leopard) using setq, the following expression is used:

     (setq carnivores '(lion tiger leopard))

This is exactly the same as using set except the first argument is automatically quoted by setq. (The ‘q’ in setq means quote.) With set, the expression would look like this:

     (set 'carnivores '(lion tiger leopard))

Also, setq can be used to assign different values to different variables. The first argument is bound to the value of the second argument, the third argument is bound to the value of the fourth argument, and so on. For example, you could use the following to assign a list of trees to the symbol trees and a list of herbivores to the symbol herbivores:

     (setq trees '(pine fir oak maple)
           herbivores '(gazelle antelope zebra))

(The expression could just as well have been on one line, but it might not have fit on a page; and humans find it easier to read nicely formatted lists.)

Although I have been using the term ‘assign’, there is another way of thinking about the workings of set and setq; and that is to say that set and setq make the symbol point to the list. This latter way of thinking is very common and in forthcoming chapters we shall come upon at least one symbol that has ‘pointer’ as part of its name. The name is chosen because the symbol has a value, specifically a list, attached to it; or, expressed in this other way, the symbol is set to “point” to the list.

1.9.3 Counting

Here is an example that shows how to use setq in a counter. You might use this to count how many times a part of your program repeats itself. First set a variable to zero; then add one to the number each time the program repeats itself. To do this, you need a variable that serves as a counter, and two expressions: an initial setq expression that sets the counter variable to zero; and a second setq expression that increments the counter each time it is evaluated.

(setq counter 0) ; Let’s call this the initializer.

     (setq counter (+ counter 1))    ; This is the incrementer.

counter ; This is the counter.

(The text following the ‘;’ are comments. See Section 3.2.1, “Change a Function Definition”, page 23.)

If you evaluate the first of these expressions, the initializer, (setq counter 0), and then evaluate the third expression, counter, the number 0 will appear in the echo area. If you then evaluate the second expression, the incrementer, (setq counter (+ counter 1)), the counter will get the value 1. So if you again evaluate counter, the number 1 will appear in the echo area. Each time you evaluate the second expression, the value of the counter will be incremented.

When you evaluate the incrementer, (setq counter (+ counter 1)), the Lisp interpreter first evaluates the innermost list; this is the addition. In order to evaluate this list, it must evaluate the variable counter and the number 1. When it evaluates the variable counter, it receives its current value. It passes this value and the number 1 to the + which adds them together. The sum is then returned as the value of the inner list and passed to the setq which sets the variable counter to this new value. Thus, the value of the variable, counter, is changed.

1.10 Summary

Learning Lisp is like climbing a hill in which the first part is the steepest. You have now climbed the most difficult part; what remains becomes easier as you progress onwards.

In summary,

• Lisp programs are made up of expressions, which are lists or single atoms.

• Lists are made up of zero or more atoms or inner lists, separated by whitespace and surrounded by parentheses. A list can be empty.

• Atoms are multi-character symbols, like forward-paragraph, single character symbols like +, strings of characters between double quotation marks, or numbers.

• A number evaluates to itself.

• A string between double quotes also evaluates to itself.

• When you evaluate a symbol by itself, its value is returned.

• When you evaluate a list, the Lisp interpreter looks at the first symbol in the list and then at the function definition bound to that symbol. Then the instructions in the function definition are carried out.

• A single-quote, ', tells the Lisp interpreter that it should return the following expression as written, and not evaluate it as it would if the quote were not there.

• Arguments are the information passed to a function. The arguments to a function are computed by evaluating the rest of the elements of the list of which the function is the first element.

• A function always returns a value when it is evaluated (unless it gets an error); in addition, it may also carry out some action called a “side effect”. In many cases, a function’s primary purpose is to create a side effect.

1.11 Exercises

A few simple exercises:

• Generate an error message by evaluating an appropriate symbol that is not within parentheses.

• Generate an error message by evaluating an appropriate symbol that is between parentheses.

• Create a counter that increments by two rather than one.

• Write an expression that prints a message in the echo area when evaluated.

2 Practicing Evaluation

Before learning how to write a function definition in Emacs Lisp, it is useful to spend a little time evaluating various expressions that have already been written. These expressions will be lists with the functions as their first (and often only) element. Since some of the functions associated with buffers are both simple and interesting, we will start with those. In this section, we will evaluate a few of these. In another section, we will study the code of several other buffer-related functions, to see how they were written.

Whenever you give an editing command to Emacs Lisp, such as the command to move the cursor or to scroll the screen, you are evaluating an expression, the first element of which is the function. This is how Emacs works.

When you type keys, you cause the Lisp interpreter to evaluate an expression and that is how you get your results. Even typing plain text involves evaluating an Emacs Lisp function, in this case, one that uses self-insert-command, which simply inserts the character you typed. The functions you evaluate by typing keystrokes are called interactive functions, or commands; how you make a function interactive will be illustrated in the chapter on how to write function definitions. See Section 3.3, “Making a Function Interactive”, page 24.

In addition to typing keyboard commands, we have seen a second way to evaluate an expression: by positioning the cursor after a list and typing C-x C-e. This is what we will do in the rest of this section. There are other ways to evaluate an expression as well; these ways will be described in other sections as we come to them.

Besides being used for practicing evaluation, the functions shown in the next few sections are important in their own right. A study of these functions makes clear the distinction between buffers and files, how to switch to a buffer, and how to determine a location within it.

2.1 Buffer Names

The two functions, buffer-name and buffer-file-name, show the difference between a file and a buffer. When you evaluate the following expression, (buffer-name), the name of the buffer appears in the echo area. When you evaluate (buffer-file-name), the name of the file to which the buffer refers appears in the echo area. Usually, the name returned by (buffer-name) is the same as the name of the file to which it refers, and the name returned by (buffer-file-name) is the full path-name of the file.

A file and a buffer are two different entities. A file is information recorded permanently in the computer (unless you delete it). A buffer, on the other hand, is information inside of Emacs that will vanish at the end of the editing session (or when you kill the buffer). Usually, a buffer contains information that you have copied from a file; we say the buffer is visiting that file. This copy is what you work on and modify. Changes to the buffer do not change the file, until you save the buffer. When you save the buffer, the buffer is copied to the file and is thus saved permanently.

If you are reading this in Info inside of GNU Emacs, you can evaluate each of the following expressions by positioning the cursor after it and typing C-x C-e.

     (buffer-name)
     (buffer-file-name)

When I do this, ‘"introduction.texinfo"’ is the value returned by evaluating (buffer-

name), and ‘"/gnu/work/intro/introduction.texinfo"’ is the value returned by evaluating

(buffer-file-name). The former is the name of the buffer and the latter is the name of the file.

(In the expressions, the parentheses tell the Lisp interpreter to treat buffer-name and bufferfile-name as functions; without the parentheses, the interpreter would attempt to evaluate the symbols as variables. See Section 1.7, “Variables”, page 7.)

In spite of the distinction between files and buffers, you will often find that people refer to a file when they mean a buffer and vice-versa. Indeed, most people say, “I am editing a file,” rather than saying, “I am editing a buffer which I will soon save to a file.” It is almost always clear from context what people mean. When dealing with computer programs, however, it is important to keep the distinction in mind, since the computer is not as smart as a person.

The word ‘buffer’, by the way, comes from the meaning of the word as a cushion that deadens the force of a collision. In early computers, a buffer cushioned the interaction between files and the computer’s central processing unit. The drums or tapes that held a file and the central processing unit were pieces of equipment that were very different from each other, working at their own speeds, in spurts. The buffer made it possible for them to work together effectively. Eventually, the buffer grew from being an intermediary, a temporary holding place, to being the place where work is done. This transformation is rather like that of a small seaport that grew into a great city: once it was merely the place where cargo was warehoused temporarily before being loaded onto ships; then it became a business and cultural center in its own right.

Not all buffers are associated with files. For example, when you start an Emacs session by typing the command emacs alone, without naming any files, Emacs will start with the ‘*scratch*’ buffer on the screen. This buffer is not visiting any file. Similarly, a ‘*Help*’ buffer is not associated with any file.

If you switch to the ‘*scratch*’ buffer, type (buffer-name), position the cursor after it, and type C-x C-e to evaluate the expression, the name "*scratch*" is returned and will appear in the echo area. "*scratch*" is the name of the buffer. However, if you type (buffer-file-name) in the ‘*scratch*’ buffer and evaluate that, nil will appear in the echo area. nil is from the Latin word for ‘nothing’; in this case, it means that the ‘*scratch*’ buffer is not associated with any file. (In Lisp, nil is also used to mean ‘false’ and is a synonym for the empty list, ().)

Incidentally, if you are in the ‘*scratch*’ buffer and want the value returned by an expression to appear in the ‘*scratch*’ buffer itself rather than in the echo area, type C-u C-x C-e instead of C-x C-e. This causes the value returned to appear after the expression. The buffer will look like this:

     (buffer-name)"*scratch*"

You cannot do this in Info since Info is read-only and it will not allow you to change the contents of the buffer. But you can do this in any buffer you can edit; and when you write code or documentation (such as this manual), this feature is very useful.

2.2 Getting Buffers

The buffer-name function returns the name of the buffer; to get the buffer itself, a different function is needed: the current-buffer function. If you use this function in code, what you get is the buffer itself.

A name and the object or entity to which the name refers are different from each other. You are not your name. You are a person to whom others refer by name. If you ask to speak to George and someone hands you a card with the letters ‘G’, ‘e’, ‘o’, ‘r’, ‘g’, and ‘e’ written on it, you might be amused, but you would not be satisfied. You do not want to speak to the name, but to the person to whom the name refers. A buffer is similar: the name of the scratch buffer is ‘*scratch*’, but the name is not the buffer. To get a buffer itself, you need to use a function

such as current-buffer.

However, there is a slight complication: if you evaluate current-buffer in an expression on its own, as we will do here, what you see is a printed representation of the name of the buffer without the contents of the buffer. Emacs works this way for two reasons: the buffer may be thousands of lines long—too long to be conveniently displayed; and, another buffer may have the same contents but a different name, and it is important to distinguish between them.

Here is an expression containing the function:

     (current-buffer)

If you evaluate the expression in the usual way, ‘#<buffer *info*>’ appears in the echo area. The special format indicates that the buffer itself is being returned, rather than just its name.

Incidentally, while you can type a number or symbol into a program, you cannot do that with the printed representation of a buffer: the only way to get a buffer itself is with a function

such as current-buffer.

A related function is other-buffer. This returns the most recently selected buffer other than the one you are in currently. If you have recently switched back and forth from the ‘*scratch*’ buffer, other-buffer will return that buffer.

You can see this by evaluating the expression:

     (other-buffer)

You should see ‘#<buffer *scratch*>’ appear in the echo area, or the name of whatever other buffer you switched back from most recently.

2.3 Switching Buffers

The other-buffer function actually provides a buffer when it is used as an argument to a function that requires one. We can see this by using other-buffer and switch-to-buffer to switch to a different buffer.

But first, a brief introduction to the switch-to-buffer function. When you switched back and forth from Info to the ‘*scratch*’ buffer to evaluate (buffer-name), you most likely typed C-x b and then typed ‘*scratch*’ when prompted in the minibuffer for the name of the buffer to which you wanted to switch. The keystrokes, C-x b, cause the Lisp interpreter to evaluate the interactive Emacs Lisp function switch-to-buffer. As we said before, this is how Emacs works: different keystrokes call or run different functions. For example, C-f calls forward-char,

M-e calls forward-sentence, and so on.

By writing switch-to-buffer in an expression, and giving it a buffer to switch to, we can switch buffers just the way C-x b does.

Here is the Lisp expression:

     (switch-to-buffer (other-buffer))

The symbol switch-to-buffer is the first element of the list, so the Lisp interpreter will treat it as a function and carry out the instructions that are attached to it. But before doing that, the interpreter will note that other-buffer is inside parentheses and work on that symbol first. other-buffer is the first (and in this case, the only) element of this list, so the Lisp interpreter calls or runs the function. It returns another buffer. Next, the interpreter runs switch-to-buffer, passing to it, as an argument, the other buffer, which is what Emacs will switch to. If you are reading this in Info, try this now. Evaluate the expression. (To get back, type C-x b ⟨\underline{\overline{RET}}^{⟩}.)

In the programming examples in later sections of this document, you will see the function setbuffer more often than switch-to-buffer. This is because of a difference between computer programs and humans: humans have eyes and expect to see the buffer on which they are working on their computer terminals. This is so obvious, it almost goes without saying. However, programs do not have eyes. When a computer program works on a buffer, that buffer does not need to be visible on the screen.

switch-to-buffer is designed for humans and does two different things: it switches the buffer to which Emacs attention is directed; and it switches the buffer displayed in the window to the new buffer. set-buffer, on the other hand, does only one thing: it switches the attention of the computer program to a different buffer. The buffer on the screen remains unchanged (of course, normally nothing happens there until the command finishes running).

Also, we have just introduced another jargon term, the word call. When you evaluate a list in which the first symbol is a function, you are calling that function. The use of the term comes from the notion of the function as an entity that can do something for you if you ‘call’ it—just as a plumber is an entity who can fix a leak if you call him or her.

2.4 Buffer Size and the Location of Point

Finally, let’s look at several rather simple functions, buffer-size, point, point-min, and point-max. These give information about the size of a buffer and the location of point within it.

The function buffer-size tells you the size of the current buffer; that is, the function returns a count of the number of characters in the buffer.

     (buffer-size)

You can evaluate this in the usual way, by positioning the cursor after the expression and typing

C-x C-e.

In Emacs, the current position of the cursor is called point. The expression (point) returns a number that tells you where the cursor is located as a count of the number of characters from the beginning of the buffer up to point.

You can see the character count for point in this buffer by evaluating the following expression in the usual way:

     (point)

As I write this, the value of point is 65724. The point function is frequently used in some of the examples later in this manual.

The value of point depends, of course, on its location within the buffer. If you evaluate point in this spot, the number will be larger:

     (point)

For me, the value of point in this location is 66043, which means that there are 319 characters (including spaces) between the two expressions.

The function point-min is somewhat similar to point, but it returns the value of the minimum permissible value of point in the current buffer. This is the number 1 unless narrowing is in effect. (Narrowing is a mechanism whereby you can restrict yourself, or a program, to operations on just a part of a buffer. See Chapter 6, “Narrowing and Widening”, page 55.) Likewise, the function point-max returns the value of the maximum permissible value of point in the current buffer.

Exercise 19

2.5 Exercise

Find a file with which you are working and move towards its middle. Find its buffer name, file name, length, and your position in the file.

20 Chapter 2: Practicing Evaluation

3 How To Write Function Definitions

When the Lisp interpreter evaluates a list, it looks to see whether the first symbol on the list has a function definition attached to it; or, put another way, whether the symbol points to a function definition. If it does, the computer carries out the instructions in the definition. A symbol that has a function definition is called, simply, a function (although, properly speaking, the definition is the function and the symbol refers to it.)

All functions are defined in terms of other functions, except for a few primitive functions that are written in the C programming language. When you write functions’ definitions, you will write them in Emacs Lisp and use other functions as your building blocks. Some of the functions you will use will themselves be written in Emacs Lisp (perhaps by you) and some will be primitives written in C. The primitive functions are used exactly like those written in Emacs Lisp and behave like them. They are written in C so we can easily run GNU Emacs on any computer that has sufficient power and can run C.

Let me re-emphasize this: when you write code in Emacs Lisp, you do not distinguish between the use of functions written in C and the use of functions written in Emacs Lisp. The difference is irrelevant. I mention the distinction only because it is interesting to know. Indeed, unless you investigate, you won’t know whether an already-written function is written in Emacs Lisp or C.

3.1 The defun Special Form

In Lisp, a symbol such as mark-whole-buffer has code attached to it that tells the computer what to do when the function is called. This code is called the function definition and is created by evaluating a Lisp expression that starts with the symbol defun (which is an abbreviation for define function). Because defun does not evaluate its arguments in the usual way, it is called a special form.

In subsequent sections, we will look at function definitions from the Emacs source code, such as mark-whole-buffer. In this section, we will describe a simple function definition so you can see how it looks. This function definition uses arithmetic because it makes for a simple example. Some people dislike examples using arithmetic; however, if you are such a person, do not despair. Hardly any of the code we will study in the remainder of this introduction involves arithmetic or mathematics. The examples mostly involve text in one way or another.

A function definition has up to five parts following the word defun:

1. The name of the symbol to which the function definition should be attached.

2. A list of the arguments that will be passed to the function. If no arguments will be passed to the function, this is an empty list, ().

3. Documentation describing the function. (Technically optional, but strongly recommended.)

4. Optionally, an expression to make the function interactive so you can use it by typing M-x and then the name of the function; or by typing an appropriate key or keychord.

5. The code that instructs the computer what to do: the body of the function definition.

It is helpful to think of the five parts of a function definition as being organized in a template, with slots for each part:

(defun function-name (arguments...) "optional-documentation..." (interactive argument-passing-info) ; optional

       body...)

As an example, here is the code for a function that multiplies its argument by 7. (This example is not interactive. See Section 3.3, “Making a Function Interactive”, page 24, for that information.)

     (defun multiply-by-seven (number)
       "Multiply NUMBER by seven."
       (* 7 number))

This definition begins with a parenthesis and the symbol defun, followed by the name of the function.

The name of the function is followed by a list that contains the arguments that will be passed to the function. This list is called the argument list. In this case, the list has only one element, the symbol, number. When the function is used, the symbol will be bound to the value that is used as the argument to the function.

Instead of choosing the word number for the name of the argument, I could have picked any other name. For example, I could have chosen the word multiplicand. I picked the word ‘number’ because it tells what kind of value is intended for this slot; but I could just as well have chosen the word ‘multiplicand’ to indicate the role that the value placed in this slot will play in the workings of the function. I could have called it foogle, but that would have been a bad choice because it would not tell humans what it means. The choice of name is up to the programmer and should be chosen to make the meaning of the function clear.

Indeed, you can choose any name you wish for a symbol in an argument list, even the name of a symbol used in some other function: the name you use in an argument list is private to that particular definition. In that definition, the name refers to a different entity than any use of the same name outside the function definition. Suppose you have a nick-name ‘Shorty’ in your family; when your family members refer to ‘Shorty’, they mean you. But outside your family, in a movie, for example, the name ‘Shorty’ refers to someone else. Because a name in an argument list is private to the function definition, you can change the value of such a symbol inside the body of a function without changing its value outside the function. The effect is similar to that produced by a let expression. (See Section 3.6, “let”, page 27.)

The argument list is followed by the documentation string that describes the function. This is what you see when you type C-h f and the name of a function. Incidentally, when you write a documentation string like this, you should make the first line a complete sentence since some commands, such as apropos, print only the first line of a multi-line documentation string. Also, you should not indent the second line of a documentation string, if you have one, because that looks odd when you use C-h f (describe-function). The documentation string is optional, but it is so useful, it should be included in almost every function you write.

The third line of the example consists of the body of the function definition. (Most functions’ definitions, of course, are longer than this.) In this case, the body is the list, (* 7 number), which says to multiply the value of number by 7. (In Emacs Lisp, * is the function for multiplication, just as + is the function for addition.)

When you use the multiply-by-seven function, the argument number evaluates to the actual number you want used. Here is an example that shows how multiply-by-seven is used; but don’t try to evaluate this yet!

     (multiply-by-seven 3)

The symbol number, specified in the function definition in the next section, is given or “bound to” the value 3 in the actual use of the function. Note that although number was inside parentheses in the function definition, the argument passed to the multiply-by-seven function is not in parentheses. The parentheses are written in the function definition so the computer can figure out where the argument list ends and the rest of the function definition begins.

If you evaluate this example, you are likely to get an error message. (Go ahead, try it!) This is because we have written the function definition, but not yet told the computer about the definition—we have not yet installed (or ‘loaded’) the function definition in Emacs. Installing a function is the process that tells the Lisp interpreter the definition of the function. Installation is described in the next section.

3.2 Install a Function Definition

If you are reading this inside of Info in Emacs, you can try out the multiply-by-seven function by first evaluating the function definition and then evaluating (multiply-by-seven 3). A copy of the function definition follows. Place the cursor after the last parenthesis of the function definition and type C-x C-e. When you do this, multiply-by-seven will appear in the echo area. (What this means is that when a function definition is evaluated, the value it returns is the name of the defined function.) At the same time, this action installs the function definition.

     (defun multiply-by-seven (number)
       "Multiply NUMBER by seven."
       (* 7 number))

By evaluating this defun, you have just installed multiply-by-seven in Emacs. The function is now just as much a part of Emacs as forward-word or any other editing function you use. (multiply-by-seven will stay installed until you quit Emacs. To reload code automatically whenever you start Emacs, see Section 3.5, “Installing Code Permanently”, page 26.)

You can see the effect of installing multiply-by-seven by evaluating the following sample. Place the cursor after the following expression and type C-x C-e. The number 21 will appear in the echo area.

     (multiply-by-seven 3)

If you wish, you can read the documentation for the function by typing C-h f (describefunction) and then the name of the function, multiply-by-seven. When you do this, a ‘*Help*’ window will appear on your screen that says:

     multiply-by-seven:
     Multiply NUMBER by seven.

(To return to a single window on your screen, type C-x 1.)

3.2.1 Change a Function Definition

If you want to change the code in multiply-by-seven, just rewrite it. To install the new version in place of the old one, evaluate the function definition again. This is how you modify code in Emacs. It is very simple.

As an example, you can change the multiply-by-seven function to add the number to itself seven times instead of multiplying the number by seven. The produces the same answer, but by a different path. At the same time, we will add a comment to the code; a comment is text that the Lisp interpreter ignores, but that a human reader may find useful or enlightening. In this case the comment is that this is the “second version”.

     (defun multiply-by-seven (number)       ; Second version.
       "Multiply NUMBER by seven."
       (+ number number number number number number number))

The comment follows a semicolon, ‘;’. In Lisp, everything on a line that follows a semicolon is a comment. The end of the line is the end of the comment. To stretch a comment over two or more lines, begin each line with a semicolon.

See Section 16.3, “Beginning a ‘.emacs’ File”, page 150, and section “Comments” in The GNU Emacs Lisp Reference Manual, for more about comments.

You can install this version of the multiply-by-seven function by evaluating it in the same way you evaluated the first function: place the cursor after the last parenthesis and type C-x

C-e.

In summary, this is how you write code in Emacs Lisp: you write a function; install it; test it; and then make fixes or enhancements and install it again.

3.3 Make a Function Interactive

You make a function interactive by placing a list that begins with the special form interactive immediately after the documentation. A user can invoke an interactive function by typing M-x and then the name of the function; or by typing the keys to which it is bound, for example, by typing C-n for next-line or C-x h for mark-whole-buffer.

Interestingly, when you call an interactive function interactively, the value returned is not automatically displayed in the echo area. This is because you often call an interactive function for its side effects, such as moving forward by a word or line, and not for the value returned. If the returned value were displayed in the echo area each time you typed a key, it would be very distracting.

Both the use of the special form interactive and one way to display a value in the echo area can be illustrated by creating an interactive version of multiply-by-seven.

Here is the code:

     (defun multiply-by-seven (number)       ; Interactive version.
       "Multiply NUMBER by seven."
       (interactive "p")
       (message "The result is %d" (* 7 number)))

You can install this code by placing your cursor after it and typing C-x C-e. The name of the function will appear in your echo area. Then, you can use this code by typing C-u and a number and then typing M-x multiply-by-seven and pressing ^{⟨}\underline{\overline{RET}}^{⟩}. The phrase ‘The result is ...’ followed by the product will appear in the echo area.

Speaking more generally, you invoke a function like this in either of two ways:

1. By typing a prefix argument that contains the number to be passed, and then typing M-x and the name of the function, as with C-u 3 M-x forward-sentence; or,

2. By typing whatever key or keychord the function is bound to, as with C-u 3 M-e.

Both the examples just mentioned work identically to move point forward three sentences. (Since multiply-by-seven is not bound to a key, it could not be used as an example of key binding.)

(See Section 16.7, “Some Keybindings”, page 153, to learn how to bind a command to a key.)

A prefix argument is passed to an interactive function by typing the ^{⟨}\underline{\overline{META}}^{⟩} key followed by a number, for example, M-3 M-e, or by typing C-u and then a number, for example, C-u 3 M-e (if you type C-u without a number, it defaults to 4).

3.3.1 An Interactive multiply-by-seven.

Let’s look at the use of the special form interactive and then at the function message in the interactive version of multiply-by-seven. You will recall that the function definition looks like this:

     (defun multiply-by-seven (number)       ; Interactive version.
       "Multiply NUMBER by seven."
       (interactive "p")
       (message "The result is %d" (* 7 number)))

In this function, the expression, (interactive "p"), is a list of two elements. The "p" tells Emacs to pass the prefix argument to the function and use its value for the argument of the function.

The argument will be a number. This means that the symbol number will be bound to a number in the line:

     (message "The result is %d" (* 7 number))

For example, if your prefix argument is 5, the Lisp interpreter will evaluate the line as if it were:

     (message "The result is %d" (* 7 5))

(If you are reading this in GNU Emacs, you can evaluate this expression yourself.) First, the interpreter will evaluate the inner list, which is (* 7 5). This returns a value of 35. Next, it will evaluate the outer list, passing the values of the second and subsequent elements of the list to the function message.

As we have seen, message is an Emacs Lisp function especially designed for sending a one line message to a user. (See Section 1.8.5, “The message function”, page 11.) In summary, the message function prints its first argument in the echo area as is, except for occurrences of ‘%d’, ‘%s’, or ‘%c’. When it sees one of these control sequences, the function looks to the second and subsequent arguments and prints the value of the argument in the location in the string where the control sequence is located.

In the interactive multiply-by-seven function, the control string is ‘%d’, which requires a number, and the value returned by evaluating (* 7 5) is the number 35. Consequently, the number 35 is printed in place of the ‘%d’ and the message is ‘The result is 35’.

(Note that when you call the function multiply-by-seven, the message is printed without quotes, but when you call message, the text is printed in double quotes. This is because the value returned by message is what appears in the echo area when you evaluate an expression whose first element is message; but when embedded in a function, message prints the text as a side effect without quotes.)

3.4 Different Options for interactive

In the example, multiply-by-seven used "p" as the argument to interactive. This argument told Emacs to interpret your typing either C-u followed by a number or ^{⟨}\underline{\overline{META}}^{⟩} followed by a number as a command to pass that number to the function as its argument. Emacs has more than twenty characters predefined for use with interactive. In almost every case, one or other of these options will enable you to pass the right information interactively to a function. (See section “Code Characters for interactive” in The GNU Emacs Lisp Reference Manual.)

For example, the character ‘r’ causes Emacs to pass the beginning and end of the region (the current values of point and mark) to the function as two separate arguments. It is used as follows:

     (interactive "r")

On the other hand, a ‘B’ tells Emacs to ask for the name of a buffer that will be passed to the function. In this case, Emacs will ask for the name by prompting the user in the minibuffer, using a string that follows the ‘B’, as in "BAppend to buffer: ". Not only will Emacs prompt for the name, but Emacs will complete the name if you type enough of it and press ^{⟨}\underline{\overline{TAB}}^{⟩}.

A function with two or more arguments can have information passed to each argument by adding parts to the string that follows interactive. When you do this, the information is passed to each argument in the same order it is specified in the interactive list. In the string, each part is separated from the next part by a ‘\n’, which is a newline. For example, you could follow "BAppend to buffer: " with a ‘\n’) and an ‘r’. This would cause Emacs to pass the values of point and mark to the function as well as prompt you for the buffer—three arguments in all.

In this case, the function definition would look like the following, where buffer, start, and end are the symbols to which interactive binds the buffer and the current values of the beginning and ending of the region:

     (defun name-of-function (buffer start end)

"documentation..."

       (interactive "BAppend to buffer: \nr")

body-of-function...)

(The space after the colon in the prompt makes it look better when you are prompted. The append-to-buffer function looks exactly like this. See Section 4.4, “The Definition of

append-to-buffer”, page 40.)

If a function does not have arguments, then interactive does not require any. Such a function contains the simple expression (interactive). The mark-whole-buffer function is like this.

Alternatively, if the special letter-codes are not right for your application, you can pass your own arguments to interactive as a list. See section “Using Interactive” in The GNU Emacs Lisp Reference Manual, for more information about this advanced technique.

3.5 Install Code Permanently

When you install a function definition by evaluating it, it will stay installed until you quit Emacs. The next time you start a new session of Emacs, the function will not be installed unless you evaluate the function definition again.

At some point, you may want to have code installed automatically whenever you start a new session of Emacs. There are several ways of doing this:

• If you have code that is just for yourself, you can put the code for the function definition in your ‘.emacs’ initialization file. When you start Emacs, your ‘.emacs’ file is automatically evaluated and all the function definitions within it are installed. See Chapter 16, “Your ‘.emacs’ File”, page 149.

• Alternatively, you can put the function definitions that you want installed in one or more files of their own and use the load function to cause Emacs to evaluate and thereby install each of the functions in the files. See Section 16.8, “Loading Files”, page 154.

• On the other hand, if you have code that your whole site will use, it is usual to put it in a file called ‘site-init.el’ that is loaded when Emacs is built. This makes the code available to everyone who uses your machine. (See the ‘INSTALL’ file that is part of the Emacs distribution.)

Finally, if you have code that everyone who uses Emacs may want, you can post it on a computer network or send a copy to the Free Software Foundation. (When you do this, please put a copyleft notice on the code before posting it.) If you send a copy of your code to the Free Software Foundation, it may be included in the next release of Emacs. In large part, this is how Emacs has grown over the past years, by donations.

3.6 let

The let expression is a special form in Lisp that you will need to use in most function definitions. Because it is so common, let will be described in this section.

let is used to attach or bind a symbol to a value in such a way that the Lisp interpreter will not confuse the variable with a variable of the same name that is not part of the function. To understand why this special form is necessary, consider the situation in which you own a home that you generally refer to as ‘the house’, as in the sentence, “The house needs painting.” If you are visiting a friend and your host refers to ‘the house’, he is likely to be referring to his house, not yours, that is, to a different house. If he is referring to his house and you think he is referring to your house, you may be in for some confusion. The same thing could happen in Lisp if a variable that is used inside of one function has the same name as a variable that is used inside of another function, and the two are not intended to refer to the same value.

The let special form prevents this kind of confusion. let creates a name for a local variable that overshadows any use of the same name outside the let expression. This is like understanding that whenever your host refers to ‘the house’, he means his house, not yours. (Symbols used in argument lists work the same way. See Section 3.1, “The defun Special Form”, page 21.)

Local variables created by a let expression retain their value only within the let expression itself (and within expressions called within the let expression); the local variables have no effect outside the let expression.

let can create more than one variable at once. Also, let gives each variable it creates an initial value, either a value specified by you, or nil. (In the jargon, this is called ‘binding the variable to the value’.) After let has created and bound the variables, it executes the code in the body of the let, and returns the value of the last expression in the body, as the value of the whole let expression. (‘Execute’ is a jargon term that means to evaluate a list; it comes from the use of the word meaning ‘to give practical effect to’ (Oxford English Dictionary). Since you evaluate an expression to perform an action, ‘execute’ has evolved as a synonym to ‘evaluate’.)

3.6.1 The Parts of a let Expression

A let expression is a list of three parts. The first part is the symbol let. The second part is a list, called a varlist, each element of which is either a symbol by itself or a two-element list, the first element of which is a symbol. The third part of the let expression is the body of the let. The body usually consists of one or more lists.

A template for a let expression looks like this:

(let varlist body...)

The symbols in the varlist are the variables that are given initial values by the let special form. Symbols by themselves are given the initial value of nil; and each symbol that is the first element of a two-element list is bound to the value that is returned when the Lisp interpreter evaluates the second element.

Thus, a varlist might look like this: (thread (needles 3)). In this case, in a let expression, Emacs binds the symbol thread to an initial value of nil, and binds the symbol needles to an initial value of 3.

When you write a let expression, what you do is put the appropriate expressions in the slots of the let expression template.

If the varlist is composed of two-element lists, as is often the case, the template for the let expression looks like this:

(let ((variable value)

(variable value)

           ...)
           body...)

3.6.2 Sample let Expression

The following expression creates and gives initial values to the two variables zebra and tiger. The body of the let expression is a list which calls the message function.

     (let ((zebra 'stripes)
           (tiger 'fierce))
       (message "One kind of animal has %s and another is %s."
                zebra tiger))
   Here, the varlist is ((zebra 'stripes) (tiger 'fierce)).

The two variables are zebra and tiger. Each variable is the first element of a two-element list and each value is the second element of its two-element list. In the varlist, Emacs binds the variable zebra to the value stripes, and binds the variable tiger to the value fierce. In this case, both values are symbols preceded by a quote. The values could just as well have been another list or a string. The body of the let follows after the list holding the variables. In this case, the body is a list that uses the message function to print a string in the echo area.

You may evaluate the example in the usual fashion, by placing the cursor after the last parenthesis and typing C-x C-e. When you do this, the following will appear in the echo area:

     "One kind of animal has stripes and another is fierce."

As we have seen before, the message function prints its first argument, except for ‘%s’. In this case, the value of the variable zebra is printed at the location of the first ‘%s’ and the value of the variable tiger is printed at the location of the second ‘%s’.

3.6.3 Uninitialized Variables in a let Statement

If you do not bind the variables in a let statement to specific initial values, they will automatically be bound to an initial value of nil, as in the following expression:

     (let ((birch 3)
           pine
           fir
           (oak 'some))
       (message
        "Here are %d variables with %s, %s, and %s value."
        birch pine fir oak))
Here, the varlist is ((birch 3) pine fir (oak 'some)).

If you evaluate this expression in the usual way, the following will appear in your echo area:

     "Here are 3 variables with nil, nil, and some value."

In this case, Emacs binds the symbol birch to the number 3, binds the symbols pine and fir to nil, and binds the symbol oak to the value some.

Note that in the first part of the let, the variables pine and fir stand alone as atoms that are not surrounded by parentheses; this is because they are being bound to nil, the empty list. But oak is bound to some and so is a part of the list (oak 'some). Similarly, birch is bound to the number 3 and so is in a list with that number. (Since a number evaluates to itself, the number does not need to be quoted. Also, the number is printed in the message using a ‘%d’ rather than a ‘%s’.) The four variables as a group are put into a list to delimit them from the body of the let.

3.7 The if Special Form

A third special form, in addition to defun and let, is the conditional if. This form is used to instruct the computer to make decisions. You can write function definitions without using if, but it is used often enough, and is important enough, to be included here. It is used, for example, in the code for the function beginning-of-buffer.

The basic idea behind an if, is that “if a test is true, then an expression is evaluated.” If the test is not true, the expression is not evaluated. For example, you might make a decision such as, “if it is warm and sunny, then go to the beach!”

An if expression written in Lisp does not use the word ‘then’; the test and the action are the second and third elements of the list whose first element is if. Nonetheless, the test part of an if expression is often called the if-part and the second argument is often called the then-part.

Also, when an if expression is written, the true-or-false-test is usually written on the same line as the symbol if, but the action to carry out if the test is true, the “then-part”, is written on the second and subsequent lines. This makes the if expression easier to read.

(if true-or-false-test action-to-carry-out-if-test-is-true)

The true-or-false-test will be an expression that is evaluated by the Lisp interpreter.

Here is an example that you can evaluate in the usual manner. The test is whether the number 5 is greater than the number 4. Since it is, the message ‘5 is greater than 4!’ will be printed.

     (if (> 5 4)                             ; if-part
         (message "5 is greater than 4!"))   ; then-part

(The function > tests whether its first argument is greater than its second argument and returns true if it is.)

Of course, in actual use, the test in an if expression will not be fixed for all time as it is by the expression (> 5 4). Instead, at least one of the variables used in the test will be bound to a value that is not known ahead of time. (If the value were known ahead of time, we would not need to run the test!)

For example, the value may be bound to an argument of a function definition. In the following function definition, the character of the animal is a value that is passed to the function. If the value bound to characteristic is fierce, then the message, ‘It's a tiger!’ will be printed; otherwise, nil will be returned.

(defun type-of-animal (characteristic)
  "Print message in echo area depending on CHARACTERISTIC.
If the CHARACTERISTIC is the symbol `fierce',
then warn of a tiger."
  (if (equal characteristic 'fierce)
      (message "It's a tiger!")))

If you are reading this inside of GNU Emacs, you can evaluate the function definition in the usual way to install it in Emacs, and then you can evaluate the following two expressions to see the results:

(type-of-animal 'fierce)
(type-of-animal 'zebra)

When you evaluate (type-of-animal 'fierce), you will see the following message printed in the echo area: "It's a tiger!"; and when you evaluate (type-of-animal 'zebra) you will see nil printed in the echo area.

3.7.1 The type-of-animal Function in Detail

Let’s look at the type-of-animal function in detail.

The function definition for type-of-animal was written by filling the slots of two templates, one for a function definition as a whole, and a second for an if expression.

The template for every function that is not interactive is:

(defun name-of-function (argument-list)

"documentation..."

  body...)

The parts of the function that match this template look like this:

(defun type-of-animal (characteristic)
  "Print message in echo area depending on CHARACTERISTIC.
If the CHARACTERISTIC is the symbol `fierce',
then warn of a tiger."

body: the if expression)

In this case, the name of function is type-of-animal; it is passed the value of one argument. The argument list is followed by a multi-line documentation string. The documentation string is included in the example because it is a good habit to write documentation string for every function definition. The body of the function definition consists of the if expression.

The template for an if expression looks like this:

(if true-or-false-test action-to-carry-out-if-the-test-returns-true)

In the type-of-animal function, the actual code for the if looks like this:

(if (equal characteristic 'fierce)
    (message "It's a tiger!")))

Here, the true-or-false-test is the expression:

(equal characteristic 'fierce)

In Lisp, equal is a function that determines whether its first argument is equal to its second argument. The second argument is the quoted symbol 'fierce and the first argument is the value of the symbol characteristic—in other words, the argument passed to this function.

In the first exercise of type-of-animal, the argument fierce is passed to type-of-animal.

Since fierce is equal to fierce, the expression, (equal characteristic 'fierce), returns a

value of true. When this happens, the if evaluates the second argument or then-part of the if:

(message "It's tiger!").

On the other hand, in the second exercise of type-of-animal, the argument zebra is passed to type-of-animal. zebra is not equal to fierce, so the then-part is not evaluated and nil is returned by the if expression.

3.8 If–then–else Expressions

An if expression may have an optional third argument, called the else-part, for the case when the true-or-false-test returns false. When this happens, the second argument or then-part of the overall if expression is not evaluated, but the third or else-part is evaluated. You might think of this as the cloudy day alternative for the decision ‘if it is warm and sunny, then go to the beach, else read a book!”.

The word “else” is not written in the Lisp code; the else-part of an if expression comes after the then-part. In the written Lisp, the else-part is usually written to start on a line of its own and is indented less than the then-part:

(if true-or-false-test

action-to-carry-out-if-the-test-returns-true) action-to-carry-out-if-the-test-returns-false)

For example, the following if expression prints the message ‘4 is not greater than 5!’ when you evaluate it in the usual way:

     (if (> 4 5)                             ; if-part
         (message "5 is greater than 4!")    ; then-part
       (message "4 is not greater than 5!")) ; else-part

Note that the different levels of indentation make it easy to distinguish the then-part from the else-part. (GNU Emacs has several commands that automatically indent if expressions correctly. See Section 1.1.3, “GNU Emacs Helps You Type Lists”, page 3.)

We can extend the type-of-animal function to include an else-part by simply incorporating an additional part to the if expression.

You can see the consequences of doing this if you evaluate the following version of the typeof-animal function definition to install it and then evaluate the two subsequent expressions to pass different arguments to the function.

     (defun type-of-animal (characteristic)  ; Second version.
       "Print message in echo area depending on CHARACTERISTIC.
     If the CHARACTERISTIC is the symbol `fierce',
     then warn of a tiger;
     else say it's not fierce."
       (if (equal characteristic 'fierce)
           (message "It's a tiger!")
         (message "It's not fierce!")))
     (type-of-animal 'fierce)

     (type-of-animal 'zebra)

When you evaluate (type-of-animal 'fierce), you will see the following message printed in the echo area: "It's a tiger!"; but when you evaluate (type-of-animal 'zebra), you will

see "It's not fierce!".

(Of course, if the characteristic were ferocious, the message "It's not fierce!" would be printed; and it would be misleading! When you write code, you need to take into account the possibility that some such argument will be tested by the if and write your program accordingly.)

3.9 Truth and Falsehood in Lisp

There is an important aspect to the truth test in an if expression. So far, we have spoken of ‘true’ and ‘false’ as values of predicates as if they were new kinds of Lisp objects. In fact, ‘false’ is just our old friend nil. Anything else—anything at all—is ‘true’.

The expression that tests for truth is interpreted as true if the result of evaluating it is a value that is not nil. In other words, the result of the test is considered true if the value returned is a number such as 47, a string such as "hello", or a symbol (other than nil) such as flowers, or a list, or even a buffer!

Before illustrating this, we need an explanation of nil.

In Lisp, the symbol nil has two meanings. First, it means the empty list. Second, it means false and is the value returned when a true-or-false-test tests false. nil can be written as an empty list, (), or as nil. As far as the Lisp interpreter is concerned, () and nil are the same. Humans, however, tend to use nil for false and () for the empty list.

In Lisp, any value that is not nil—is not the empty list—is considered true. This means that if an evaluation returns something that is not an empty list, an if expression will test true. For example, if a number is put in the slot for the test, it will be evaluated and will return itself, since that is what numbers do when evaluated. In this case, the if expression will test true. The expression tests false only when nil, an empty list, is returned by evaluating the expression.

You can see this by evaluating the two expressions in the following examples.

In the first example, the number 4 is evaluated as the test in the if expression and returns itself; consequently, the then-part of the expression is evaluated and returned: ‘true’ appears in the echo area. In the second example, the nil indicates false; consequently, the else-part of the expression is evaluated and returned: ‘false’ appears in the echo area.

     (if 4 'true 'false)

     (if nil 'true 'false)

Incidentally, if some other useful value is not available for a test that returns true, then the Lisp interpreter will return the symbol t for true. For example, the expression (> 5 4) returns t when evaluated, as you can see by evaluating it in the usual way:

     (> 5 4)

On the other hand, this function returns nil if the test is false.

     (> 4 5)

3.10 save-excursion

The save-excursion function is the fourth and final special form that we will discuss in this chapter.

In Emacs Lisp programs used for editing, the save-excursion function is very common. It saves the location of point and mark, executes the body of the function, and then restores point and mark to their previous positions if their locations were changed. Its primary purpose is to keep the user from being surprised and disturbed by unexpected movement of point or mark.

Before discussing save-excursion, however, it may be useful first to review what point and mark are in GNU Emacs. Point is the current location of the cursor. Wherever the cursor is, that is point. More precisely, on terminals where the cursor appears to be on top of a character, point is immediately before the character. In Emacs Lisp, point is an integer. The first character in a buffer is number one, the second is number two, and so on. The function point returns the current position of the cursor as a number. Each buffer has its own value for point.

The mark is another position in the buffer; its value can be set with a command such as C-^{⟨}\underline{\overline{SPC}}^{⟩} (set-mark-command). If a mark has been set, you can use the command C-x C-x (exchange-point-and-mark) to cause the cursor to jump to the mark and set the mark to be the previous position of point. In addition, if you set another mark, the position of the previous mark is saved in the mark ring. Many mark positions can be saved this way. You can jump the cursor to a saved mark by typing C-u C-^{⟨}\underline{\overline{SPC}}^{⟩} one or more times.

The part of the buffer between point and mark is called the region. Numerous commands

work on the region, including center-region, count-lines-region, kill-region, and print-
region.

The save-excursion special form saves the locations of point and mark and restores those positions after the code within the body of the special form is evaluated by the Lisp interpreter. Thus, if point were in the beginning of a piece of text and some code moved point to the end of the buffer, the save-excursion would put point back to where it was before, after the expressions in the body of the function were evaluated.

In Emacs, a function frequently moves point as part of its internal workings even though a user would not expect this. For example, count-lines-region moves point. To prevent the user from being bothered by jumps that are both unexpected and (from the user’s point of view) unnecessary, save-excursion is often used to keep point and mark in the location expected by the user. The use of save-excursion is good housekeeping.

To make sure the house stays clean, save-excursion restores the values of point and mark even if something goes wrong in the code inside of it (or, to be more precise and to use the proper jargon, “in case of abnormal exit”). This feature is very helpful.

In addition to recording the values of point and mark, save-excursion keeps track of the current buffer, and restores it, too. This means you can write code that will change the buffer and have save-excursion switch you back to the original buffer. This is how save-excursion is used in append-to-buffer. (See Section 4.4, “The Definition of append-to-buffer”, page 40.)

3.10.1 Template for a save-excursion Expression

The template for code using save-excursion is simple:

     (save-excursion
       body...)

The body of the function is one or more expressions that will be evaluated in sequence by the Lisp interpreter. If there is more than one expression in the body, the value of the last one will be returned as the value of the save-excursion function. The other expressions in the body are evaluated only for their side effects; and save-excursion itself is used only for its side effect (which is restoring the positions of point and mark).

In more detail, the template for a save-excursion expression looks like this:

(save-excursion

first-expression-in-body second-expression-in-body third-expression-in-body

...

last-expression-in-body)

An expression, of course, may be a symbol on its own or a list.

In Emacs Lisp code, a save-excursion expression often occurs within the body of a let expression. It looks like this:

(let varlist

(save-excursion
body...))

3.11 Review

In the last few chapters we have introduced a fair number of functions and special forms. Here they are described in brief, along with a few similar functions that have not been mentioned yet.

eval-last-sexp

Evaluate the last symbolic expression before the current location of point. The value is printed in the echo area unless the function is invoked with an argument; in that case, the output is printed in the current buffer. This command is normally bound

to C-x C-e.

defun Define function. This special form has up to five parts: the name, a template for the arguments that will be passed to the function, documentation, an optional interactive declaration, and the body of the definition. For example:

     (defun back-to-indentation ()
       "Point to first visible character on line."
       (interactive)
       (beginning-of-line 1)
       (skip-chars-forward " \t"))
interactive

Declare to the interpreter that the function can be used interactively. This special form may be followed by a string with one or more parts that pass the information to the arguments of the function, in sequence. These parts may also tell the interpreter to prompt for information. Parts of the string are separated by newlines, ‘\n’. Common code characters are:

b The name of an existing buffer.

f The name of an existing file.

p The numeric prefix argument. (Note that this ‘p’ is lower case.) r Point and the mark, as two numeric arguments, smallest first. This is

the only code letter that specifies two successive arguments rather than one.

See section “Code Characters for ‘interactive’” in The GNU Emacs Lisp Reference Manual, for a complete list of code characters.

let Declare that a list of variables is for use within the body of the let and give them an initial value, either nil or a specified value; then evaluate the rest of the expressions in the body of the let and return the value of the last one. Inside the body of the let, the Lisp interpreter does not see the values of the variables of the same names that are bound outside of the let. For example,

     (let ((foo (buffer-name))
           (bar (buffer-size)))
       (message
        "This buffer is %s and has %d characters."
        foo bar))
save-excursion

Record the values of point and mark and the current buffer before evaluating the body of this special form. Restore the values of point and mark and buffer afterward. For example,

     (message "We are %d characters into this buffer."
              (- (point)
                 (save-excursion
                   (goto-char (point-min)) (point))))

if Evaluate the first argument to the function; if it is true, evaluate the second argument; else evaluate the third argument, if there is one. The if special form is called a conditional. There are other conditionals in Emacs Lisp, but if is perhaps the most commonly used. For example,

     (if (string= (int-to-string 19)
                  (substring (emacs-version) 10 12))
         (message "This is version 19 Emacs")
       (message "This is not version 19 Emacs"))
equal

eq Test whether two objects are the same. equal returns true if the two objects have a similar structure and contents. Another function, eq, returns true if both arguments are actually the same object.

<
>
<=

>= The < function tests whether the first argument is smaller than the second argument. A corresponding function, >, tests whether the first argument is greater than the second. Likewise, <= tests whether the first argument is less than or equal to the second and >= tests whether the first argument is greater than or equal to the second. In all cases, both arguments must be numbers.

message Print a message in the echo area. The message can be only one line long. The first argument is a string that can contain ‘%s’, ‘%d’, or ‘%c’ to print the value of arguments that follow the string. The argument used by ‘%s’ must be a string or a symbol; the argument used by ‘%d’ must be a number. The argument used by ‘%c’ must be a number; it will be printed as the character with that ascii code.

setq

set The setq function sets the value of its first argument to the value of the second argument. The first argument is automatically quoted by setq. It does the same for succeeding pairs of arguments. Another function, set, takes only two arguments and evaluates both of them before setting the value returned by its first argument to the value returned by its second argument.

buffer-name

Without an argument, return the name of the buffer, as a string.

buffer-file-name

Without an argument, return the name of the file the buffer is visiting.

current-buffer

Return the buffer in which Emacs is active; it may not be the buffer that is visible on the screen.

other-buffer

Return the most recently selected buffer (other than the buffer passed to otherbuffer as an argument and other than the current buffer).

switch-to-buffer

Select a buffer for Emacs to be active in and display it in the current window so users can look at it. Usually bound to C-x b.

set-buffer

Switch Emacs’s attention to a buffer on which programs will run. Don’t alter what the window is showing.

buffer-size

Return the number of characters in the current buffer.

point Return the value of the current position of the cursor, as an integer counting the number of characters from the beginning of the buffer.

point-min

Return the minimum permissible value of point in the current buffer. This is 1, unless narrowing is in effect.

point-max

Return the value of the maximum permissible value of point in the current buffer. This is the end of the buffer, unless narrowing is in effect.

3.12 Exercises

• Write a non-interactive function that doubles the value of its argument, a number. Make that function interactive.

• Write a function that tests whether the current value of fill-column is greater than the argument passed to the function, and if so, prints an appropriate message.

4 A Few Buffer–Related Functions

In this chapter we study in detail several of the functions used in GNU Emacs. This is called a “walk-through”. These functions are used as examples of Lisp code, but are not imaginary examples; with the exception of the first, simplified function definition, these functions show the actual code used in GNU Emacs. You can learn a great deal from these definitions. The functions described here are all related to buffers. Later, we will study other functions.

4.1 Finding More Information

In this walk-through, I will describe each new function as we come to it, sometimes in detail and sometimes briefly. If you are interested, you can get the full documentation of any Emacs Lisp function at any time by typing C-h f and then the name of the function (and then ^{⟨}\underline{\overline{RET}}^{⟩}). Similarly, you can get the full documentation for a variable by typing C-h v and then the name of the variable (and then ^{⟨}\underline{\overline{RET}}^{⟩}).

Also, if you want to see a function in its original source file, you can use the find-tags function to jump to it. Type M-. (i.e., type the ^{⟨}\underline{\overline{META}}^{⟩} and the period key at the same time, or else type the ^{⟨}\underline{\overline{ESC}}^{⟩} key and then type the period key), and then, at the prompt, type in the name of the function whose source code you want to see, such as mark-whole-buffer, and then type ⟨\underline{\overline{RET}}^{⟩}. Emacs will switch buffers and display the source code for the function on your screen. To switch back to this buffer, type C-x b ^{⟨}\underline{\overline{RET}}^{⟩}.

Depending on how the initial default values of your copy of Emacs are set, you may also need to specify a ‘tags table’, which is a file called ‘TAGS’. The one you will most likely want to specify is in the ‘emacs/src’ directory; thus you would use the M-x visit-tags-table command and specify a pathname such as ‘/usr/local/lib/emacs/19.23/src/TAGS’. See section “Tag Tables” in The GNU Emacs Manual. Also, see Section 12.5, “Create Your Own ‘TAGS’ File”, page 112, for how to create your own.

After you become more familiar with Emacs Lisp, you will find that you will frequently use find-tags to navigate your way around source code; and you will create your own ‘TAGS’ tables.

Incidentally, the files that contain Lisp code are conventionally called libraries. The metaphor is derived from that of a specialized library, such as a law library or an engineering library, rather than a general library. Each library, or file, contains functions that relate to a particular topic or activity, such as ‘abbrev.el’ for handling abbreviations and other typing shortcuts, and ‘help.el’ for on-line help. (Sometimes several libraries provide code for a single activity, as the various ‘rmail...’ files provide code for reading electronic mail.) In The GNU Emacs Manual, you will see sentences such as “The C-h p command lets you search the standard Emacs Lisp libraries by topic keywords.”

4.2 A Simplified beginning-of-buffer Definition

The beginning-of-buffer command is a good function to start with since you are likely to be familiar with it and it is easy to understand. Used as an interactive command, beginningof-buffer moves the cursor to the beginning of the buffer, leaving the mark at the previous position. It is generally bound to M-<.

In this section, we will discuss a shortened version of the function that shows how it is most frequently used. This shortened function works as written, but it does not contain the code for a complex option. In another section, we will describe the entire function. (See Section 5.3, “Complete Definition of beginning-of-buffer”, page 50.)

Before looking at the code, let’s consider what the function definition has to contain: it must include an expression that makes the function interactive so it can be called by typing M-x beginning-of-buffer or by typing a keychord such as C-<; it must include code to leave a mark at the original position in the buffer; and it must include code to move the cursor to the beginning of the buffer.

Here is the complete text of the shortened version of the function:

     (defun simplified-beginning-of-buffer ()
       "Move point to the beginning of the buffer;
     leave mark at previous position."
       (interactive)
       (push-mark)
       (goto-char (point-min)))

Like all function definitions, this definition has five parts following the special form defun:

 1. The name: in this case, simplified-beginning-of-buffer.

2. A list of the arguments: in this case, an empty list, (),

3. The documentation string.

4. The interactive expression.

5. The body.

In this function definition, the argument list is empty; this means that this function does not require any arguments. (When we look at the definition for the complete function, we will see that it may be passed an optional argument.)

The interactive expression tells Emacs that the function is intended to be used interactively. In this case, interactive does not have an argument because simplified-beginning-of-buffer does not require one.

The body of the function consists of the two lines:

     (push-mark)
     (goto-char (point-min))

The first of these lines is the expression, (push-mark). When this expression is evaluated by the Lisp interpreter, it sets a mark at the current position of the cursor, wherever that may be. The position of this mark is saved in the mark ring.

The next line is (goto-char (point-min)). This expression jumps the cursor to the minimum point in the buffer, that is, to the beginning of the buffer (or to the beginning of the accessible portion of the buffer if it is narrowed. See Chapter 6, “Narrowing and Widening”, page 55.)

The push-mark command sets a mark at the place where the cursor was located before it was moved to the beginning of the buffer by the (goto-char (point-min)) expression. Consequently, you can, if you wish, go back to where you were originally by typing C-x C-x.

That is all there is to the function definition!

When you are reading code such as this and come upon an unfamiliar function, such as gotochar, you can find out what it does by using the describe-function command. To use this command, type C-h f and then type in the name of the function and press^{⟨}\underline{\overline{RET}}^{⟩}. The describefunction command will print the function’s documentation string in a ‘*Help*’ window. For example, the documentation for goto-char is:

Body of mark-whole-buffer  39

     One arg, a number.  Set point to that number.
     Beginning of buffer is position (point-min),
     end is (point-max).

(The prompt for describe-function will offer you the symbol preceding the cursor, so you can save typing by positioning the cursor right after the function and then typing C-h f ^{⟨}\underline{\overline{RET}}^{⟩}.)

The end-of-buffer function definition is written in the same way as the beginning-of-buffer definition except that the body of the function contains the expression (goto-char

(point-max)) in place of (goto-char (point-min)).

4.3 The Definition of mark-whole-buffer

The mark-whole-buffer function is no harder to understand than the simplifiedbeginning-of-buffer function. In this case, however, we will look at the complete function, not a shortened version.

   The mark-whole-buffer function is not as commonly used as the beginning-of-buffer

function, but is useful nonetheless: it marks a whole buffer as a region by putting point at the beginning and a mark at the end of the buffer. It is generally bound to C-x h.

The code for the complete function looks like this:

     (defun mark-whole-buffer ()
       "Put point at beginning and mark at end of buffer."
       (interactive)
       (push-mark (point))
       (push-mark (point-max))
       (goto-char (point-min)))

Like all other functions, the mark-whole-buffer function fits into the template for a function definition. The template looks like this:

(defun name-of-function (argument-list)

"documentation..."

(interactive-expression...)

       body...)

Here is how the function works: the name of the function is mark-whole-buffer; it is followed by an empty argument list, ‘()’, which means that the function does not require arguments. The documentation comes next.

The next line is an (interactive) expression that tells Emacs that the function will be used interactively. These details are similar to the simplified-beginning-of-buffer function described in the previous section.

4.3.1 Body of mark-whole-buffer

The body of the mark-whole-buffer function consists of three lines of code:

     (push-mark (point))
     (push-mark (point-max))
     (goto-char (point-min))

The first of these lines is the expression, (push-mark (point)).

This line does exactly the same job as the first line of the body of the simplified-beginning-of-buffer function, which is written (push-mark). In both cases, the Lisp interpreter sets a mark at the current position of the cursor.

I don’t know why the expression in mark-whole-buffer is written (push-mark (point)) and the expression in beginning-of-buffer is written (push-mark). Perhaps whoever wrote the code did not know that the argument for push-mark is optional and that if push-mark is not passed an argument, the function automatically sets mark at the location of point by default. Or perhaps the expression was written so as to parallel the structure of the next line. In any case, the line causes Emacs to determine the position of point and set a mark there.

   The next line of mark-whole-buffer is (push-mark (point-max)). This expression sets a

mark at the point in the buffer that has the highest number. This will be the end of the buffer (or, if the buffer is narrowed, the end of the accessible portion of the buffer. See Chapter 6, “Narrowing and Widening”, page 55, for more about narrowing.) After this mark has been set, the previous mark, the one set at point, is no longer set, but Emacs remembers its position, just as all other recent marks are always remembered. This means that you can, if you wish, go back to that position by typing C-u C-^{⟨}\underline{\overline{SPC}}^{⟩} twice.

Finally, the last line of the function is (goto-char (point-min))). This is written exactly the same way as it is written in beginning-of-buffer. The expression moves the cursor to the minimum point in the buffer, that is, to the beginning of the buffer (or to the beginning of the accessible portion of the buffer). As a result of this, point is placed at the beginning of the buffer and mark is set at the end of the buffer. The whole buffer is, therefore, the region.

4.4 The Definition of append-to-buffer

The append-to-buffer command is very nearly as simple as the mark-whole-buffer command. What it does is copy the region (that is, the part of the buffer between point and mark) from the current buffer to a specified buffer.

   The append-to-buffer command uses the insert-buffer-substring function to copy the

region. insert-buffer-substring is described by its name: it takes a string of characters from part of a buffer, a “substring”, and inserts them into another buffer. Most of append-to-buffer is concerned with setting up the conditions for insert-buffer-substring to work: the code must specify both the buffer to which the text will go and the region that will be copied. Here is the complete text of the function:

     (defun append-to-buffer (buffer start end)
       "Append to specified buffer the text of the region.
     It is inserted into that buffer before its point.
     When calling from a program, give three arguments:
     a buffer or the name of one, and two character numbers
     specifying the portion of the current buffer to be copied."
       (interactive "BAppend to buffer: \nr")
       (let ((oldbuf (current-buffer)))
         (save-excursion
           (set-buffer (get-buffer-create buffer))
           (insert-buffer-substring oldbuf start end))))

The function can be understood by looking at it as a series of filled-in templates.

The outermost template is for the function definition. In this case, it looks like this (with several slots filled in):

The Body of append-to-buffer  41

     (defun append-to-buffer (buffer start end)

"documentation..."

       (interactive "BAppend to buffer: \nr")
       body...)

The first line of the function includes its name and three arguments. The arguments are the buffer to which the text will be copied, and the start and end of the region in the current buffer that will be copied.

The next part of the function is the documentation, which is clear and complete.

4.4.1 The append-to-buffer Interactive Expression

Since the append-to-buffer function will be used interactively, the function must have an interactive expression. (For a review of interactive, see Section 3.3, “Making a Function Interactive”, page 24.) The expression reads as follows:

     (interactive "BAppend to buffer: \nr")

This expression has an argument inside of quotation marks and that argument has two parts, separated by ‘\n’.

The first part is ‘BAppend to buffer: ’. Here, the ‘B’ tells Emacs to ask for the name of the buffer that will be passed to the function. Emacs will ask for the name by prompting the user in the minibuffer, using the string following the ‘B’, which is the string ‘Append to buffer: ’. Emacs then binds the variable buffer in the function’s argument list to the specified buffer.

The newline, ‘\n’, separates the first part of the argument from the second part. It is followed by an ‘r’ that tells Emacs to bind the two arguments that follow the symbol buffer in the function’s argument list (that is, start and end) to the values of point and mark.

4.4.2 The Body of append-to-buffer

The body of the append-to-buffer function begins with let.

As we have seen before (see Section 3.6, “let”, page 27), the purpose of a let expression is to create and give initial values to one or more variables that will only be used within the body of the let. This means that such a variable will not be confused with any variable of the same name outside the let expression.

We can see how the let expression fits into the function as a whole by showing a template for append-to-buffer with the let expression in outline:

     (defun append-to-buffer (buffer start end)

"documentation..."

       (interactive "BAppend to buffer: \nr")

(let ((variable value))

             body...)

The let expression has three elements:

1. The symbol let;

2. A varlist containing, in this case, a single two-element list, (variable value);

3. The body of the let expression.

In the append-to-buffer function, the varlist looks like this:

     (oldbuf (current-buffer))

In this part of the let expression, the one variable, oldbuf, is bound to the value returned by the (current-buffer) expression. The variable, oldbuf, is used to keep track of the buffer in which you are working.

The element or elements of a varlist are surrounded by a set of parentheses so the Lisp interpreter can distinguish the varlist from the body of the let. As a consequence, the twoelement list within the varlist is surrounded by a circumscribing set of parentheses. The line looks like this:

     (let ((oldbuf (current-buffer)))
       ... )

The two parentheses before oldbuf might surprise you if you did not realize that the first parenthesis before oldbuf marks the boundary of the varlist and the second parenthesis marks the beginning of the two-element list, (oldbuf (current-buffer)).

4.4.3 save-excursion in append-to-buffer

The body of the let expression in append-to-buffer consists of a save-excursion expression.

The save-excursion function saves the locations of point and mark, and restores them to those positions after the expressions in the body of the save-excursion complete execution. In addition, save-excursion keeps track of the original buffer, and restores it. This is how

save-excursion is used in append-to-buffer.

Incidentally, it is worth noting here that a Lisp function is normally formatted so that everything that is enclosed in a multi-line spread is indented more to the right than the first symbol. In this function definition, the let is indented more than the defun, and the saveexcursion is indented more than the let, like this:

     (defun ...
       ...
       ...
       (let...
         (save-excursion
           ...

This formatting convention makes it easy to see that the two lines in the body of the saveexcursion are enclosed by the parentheses associated with save-excursion, just as the saveexcursion itself is enclosed by the parentheses associated with the let:

     (let ((oldbuf (current-buffer)))
       (save-excursion
         (set-buffer (get-buffer-create buffer))
         (insert-buffer-substring oldbuf start end))))

The use of the save-excursion function can be viewed as a process of filling in the slots of a template:

     (save-excursion

first-expression-in-body second-expression-in-body

        ...

last-expression-in-body)

In this function, the body of the save-excursion contains only two expressions. The body looks like this:

     (set-buffer (get-buffer-create buffer))
     (insert-buffer-substring oldbuf start end)

When the append-to-buffer function is evaluated, the two expressions in the body of the save-excursion are evaluated in sequence. The value of the last expression is returned as the value of the save-excursion function; the other expression is evaluated only for its side effects.

The first line in the body of the save-excursion uses the set-buffer function to change the current buffer to the one specified in the first argument to append-to-buffer. (Changing the buffer is the side effect; as we have said before, in Lisp, a side effect is often the primary thing we want.) The second line does the primary work of the function.

The set-buffer function changes Emacs’ attention to the buffer to which the text will be copied and from which save-excursion will return. The line looks like this:

     (set-buffer (get-buffer-create buffer))

The innermost expression of this list is (get-buffer-create buffer). This expression uses the get-buffer-create function, which either gets the named buffer, or if it does not exist, creates one with the given name. This means you can use append-to-buffer to put text into a buffer that did not previously exist.

get-buffer-create also keeps set-buffer from getting an unnecessary error: set-buffer needs a buffer to go to; if you were to specify a buffer that does not exist, Emacs would baulk. Since get-buffer-create will create a buffer if none exists, set-buffer is always provided with a buffer.

The last line of append-to-buffer does the work of appending the text:

     (insert-buffer-substring oldbuf start end)

The insert-buffer-substring function copies a string from the buffer specified as its first argument and inserts the string into the present buffer. In this case, the argument to insertbuffer-substring is the value of the variable created and bound by the let, namely the value of oldbuf, which was the current buffer when you gave the append-to-buffer command.

After insert-buffer-substring has done its work, save-excursion will restore the action to the original buffer and append-to-buffer will have done its job.

Written in skeletal form, the workings of the body look like this:

     (let (bind-oldbuf-to-value-of-current-buffer)

(save-excursion ; Keep track of buffer. change-buffer insert-substring-from-oldbuf-into-buffer)

change-back-to-original-buffer-when-finished let-the-local-meaning-of-oldbuf-disappear-when-finished

In summary, append-to-buffer works as follows: it saves the value of the current buffer in the variable called oldbuf. It gets the new buffer, creating one if need be, and switches Emacs to it. Using the value of oldbuf, it inserts the region of text from the old buffer into the new buffer; and then using save-excursion, it brings you back to your original buffer.

In looking at append-to-buffer, you have explored a fairly complex function. It shows how to use let and save-excursion, and how to change to and come back from another buffer. Many function definitions use let, save-excursion, and set-buffer this way.

4.5 Review

Here is a brief summary of the various functions discussed in this chapter.

describe-function
describe-variable

Print the documentation for a function or variable. Conventionally bound to C-h f

and C-h v.

find-tag Find the file containing the source for a function or variable and switch buffers to it, positioning point at the beginning of the item. Conventionally bound to M-. (that’s a period following the ^{⟨}\underline{\overline{META}}^{⟩} key).

save-excursion

Save the location of point and mark and restore their values after the arguments to save-excursion have been evaluated. Also, remember the current buffer and return to it.

push-mark

Set mark at a location and record the value of the previous mark on the mark ring. The mark is a location in the buffer that will keep its relative position even if text is added to or removed from the buffer.

goto-char

Set point to the location specified by the value of the argument, which can be a number, a marker, or an expression that returns the number of a position, such as

(point-min).
insert-buffer-substring

Copy a region of text from a buffer that is passed to the function as an argument and insert the region into the current buffer.

mark-whole-buffer

Mark the whole buffer as a region. Normally bound to C-x h.

set-buffer

Switch the attention of Emacs to another buffer, but do not change the window being displayed. Used when the program rather than a human is to work on a different buffer.

get-buffer-create
get-buffer

Find a named buffer or create one if a buffer of that name does not exist. The get-buffer function returns nil if the named buffer does not exist.

4.6 Exercises

• Write your own simplified-end-of-buffer function definition; then test it to see whether it works.

• Use if and get-buffer to write a function that prints a message telling you whether a buffer exists.

• Using find-tag, find the source for the copy-to-buffer function.

5 A Few More Complex Functions

In this chapter, we build on what we have learned in previous chapters by looking at more complex functions. The copy-to-buffer function illustrates use of two save-excursion expressions in one definition, while the insert-buffer function illustrates use of^{⟨}\underline{\overline{*}}^{⟩} in an interactive expression, use of or, and the important distinction between a name and the object to which the name refers.

5.1 The Definition of copy-to-buffer

After understanding how append-to-buffer works, it is easy to understand copy-to-buffer. This function copies text into a buffer, but instead of adding to the second buffer, it replaces the previous text in the second buffer. The code for the copy-to-buffer function is almost the same as the code for append-to-buffer, except that erase-buffer and a second save-excursion are used. (See Section 4.4, “The Definition of append-to-buffer”, page 40, for the description of append-to-buffer.)

The body of copy-to-buffer looks like this

     ...
     (interactive "BCopy to buffer: \nr")
       (let ((oldbuf (current-buffer)))
         (save-excursion
           (set-buffer (get-buffer-create buffer))
           (erase-buffer)
           (save-excursion
             (insert-buffer-substring oldbuf start end)))))

This code is similar to the code in append-to-buffer: it is only after changing to the buffer to which the text will be copied that the definition for this function diverges from the definition for append-to-buffer: the copy-to-buffer function erases the buffer’s former contents. (This is what is meant by ‘replacement’; to replace text, Emacs erases the previous text and then inserts new text.) After erasing the previous contents of the buffer, save-excursion is used for a second time and the new text is inserted.

Why is save-excursion used twice? Consider again what the function does.

In outline, the body of copy-to-buffer looks like this:

     (let (bind-oldbuf-to-value-of-current-buffer)
       (save-excursion         ; First use of save-excursion.

change-buffer

           (erase-buffer)
           (save-excursion     ; Second use of save-excursion.

insert-substring-from-oldbuf-into-buffer)))

The first use of save-excursion returns Emacs to the buffer from which the text is being copied. That is clear, and is just like its use in append-to-buffer. Why the second use? The reason is that insert-buffer-substring always leaves point at the end of the region being inserted. The second save-excursion causes Emacs to leave point at the beginning of the text being inserted. In most circumstances, users prefer to find point at the beginning of inserted text. (Of course, the copy-to-buffer function returns the user to the original buffer when done—but if the user then switches to the copied-to buffer, point will go to the beginning of the text. Thus, this use of a second save-excursion is a little nicety.)

5.2 The Definition of insert-buffer

insert-buffer is yet another buffer-related function. This command copies another buffer into the current buffer. It is the reverse of append-to-buffer or copy-to-buffer, since they copy a region of text from the current buffer to another buffer.

In addition, this code illustrates the use of interactive with a buffer that might be read-only and the important distinction between the name of an object and the object actually referred to. Here is the code:

     (defun insert-buffer (buffer)
       "Insert after point the contents of BUFFER.
     Puts mark after the inserted text.
     BUFFER may be a buffer or a buffer name."
       (interactive "*bInsert buffer: ")
       (or (bufferp buffer)
           (setq buffer (get-buffer buffer)))
       (let (start end newmark)
         (save-excursion
           (save-excursion
             (set-buffer buffer)
             (setq start (point-min) end (point-max)))
           (insert-buffer-substring buffer start end)
           (setq newmark (point)))
         (push-mark newmark)))

As with other function definitions, you can use a template to see an outline of the function:

     (defun insert-buffer (buffer)

"documentation..."

       (interactive "*bInsert buffer: ")
       body...)

5.2.1 The Interactive Expression in insert-buffer

In insert-buffer, the argument to the interactive declaration has two parts, an asterisk,

‘*’, and ‘bInsert buffer: ’.

A Read-only Buffer

The asterisk is for the situation when the buffer is a read-only buffer—a buffer that cannot be modified. If insert-buffer is called on a buffer that is read-only, a message to this effect is printed in the echo area and the terminal may beep or blink at you; you will not be permitted to insert anything into current buffer. The asterisk does not need to be followed by a newline to separate it from the next argument.

‘b’ in an Interactive Expression

The next argument in the interactive expression starts with a lower case ‘b’. (This is different from the code for append-to-buffer, which uses an upper-case ‘B’. See Section 4.4, “The Definition of append-to-buffer”, page 40.) The lower-case ‘b’ tells the Lisp interpreter that the argument for insert-buffer should be an existing buffer or else its name. (The upper-case ‘B’ option provides for the possibility that the buffer does not exist.) Emacs will prompt you for the name of the buffer, offering you a default buffer, with name completion enabled. If the buffer does not exist, you receive a message that says “No match”; your terminal may beep at you as well.

5.2.2 The Body of the insert-buffer Function

The body of the insert-buffer function has two major parts: an or expression and a let expression. The purpose of the or expression is to ensure that the argument buffer is bound to a buffer and not just the name of a buffer. The body of the let expression contains the code which copies the other buffer into the current buffer.

In outline, the two expressions fit into the insert-buffer function like this:

     (defun insert-buffer (buffer)

"documentation..."

       (interactive "*bInsert buffer: ")
       (or ...
           ...

(let (varlist)

body-of-let... )

To understand how the or expression ensures that the argument buffer is bound to a buffer and not to the name of a buffer, it is first necessary to understand the or function.

Before doing this, let me rewrite this part of the function using if so that you can see what is done in a manner that will be familiar.

5.2.3 insert-buffer With an if Instead of an or

The job to be done is to make sure the value of buffer is a buffer itself and not the name of a buffer. If the value is the name, then the buffer itself must be got.

You can imagine yourself at a conference where an usher is wandering around holding a list with your name on it and looking for you: the usher is “bound” to your name, not to you; but when the usher finds you and takes your arm, the usher becomes “bound” to you.

In Lisp, you might describe this situation like this:

     (if (not (holding-on-to-guest))
         (find-and-take-arm-of-guest))

We want to do the same thing with a buffer—if we do not have the buffer itself, we want to get it.

Using a predicate called bufferp that tells us whether we have a buffer (rather than its name), we can write the code like this:

     (if (not (bufferp buffer))              ; if-part
         (setq buffer (get-buffer buffer)))  ; then-part

Here, the true-or-false-test of the if expression is (not (bufferp buffer)); and the then-part

is the expression (setq buffer (get-buffer buffer)).

In the test, the function bufferp returns true if its argument is a buffer—but false if its argument is the name of the buffer. (The last character of the function name bufferp is the character ‘p’; as we saw earlier, such use of ‘p’ is a convention that indicates that the function is a predicate, which is a term that means that the function will determine whether some property is true or false. See Section 1.8.4, “Using the Wrong Type Object as an Argument”, page 10.)

The function not precedes the expression (bufferp buffer), so the true-or-false-test looks like this:

     (not (bufferp buffer))

not is a function the returns true if its argument is false and false if its argument is true. So if (bufferp buffer) returns true, the not expression returns false and vice-versa: what is “not true” is false and what is “not false” is true.

Using this test, the if expression works as follows: when the value of the variable buffer is actually a buffer rather then its name, the true-or-false-test returns false and the if expression does not evaluate the then-part. This is fine, since we do not need to do anything to the variable buffer if it really is a buffer.

On the other hand, when the value of buffer is not a buffer itself, but the name of a buffer, the true-or-false-test returns true and the then-part of the expression is evaluated. In this case,

the then-part is (setq buffer (get-buffer buffer)). This expression uses the get-buffer

function to return an actual buffer itself, given its name. The setq then sets the variable buffer to the value of the buffer itself, replacing its previous value (which was the name of the buffer).

5.2.4 The or in the Body

The purpose of the or expression in the insert-buffer function is to ensure that the argument buffer is bound to a buffer and not just the name of a buffer. The previous section shows how the job could have been done using an if expression. However, the insert-buffer function actually uses or. To understand this, it is necessary to understand how or works.

An or function can have any number of arguments. It evaluates each argument in turn and returns the value of the first of its arguments that is not nil. Also, and this is a crucial feature of or, it does not evaluate any subsequent arguments after returning the first non-nil value.

The or expression looks like this:

     (or (bufferp buffer)
         (setq buffer (get-buffer buffer)))

The first argument to or is the expression (bufferp buffer). This expression returns true (a non-nil value) if the buffer is actually a buffer, and not just the name of a buffer. In the or expression, if this is the case, the or expression returns this true value and does not evaluate the next expression—and this is fine with us, since we do not want to do anything to the value of buffer if it really is a buffer.

On the other hand, if the value of (bufferp buffer) is nil, which it will be if the value of buffer is the name of a buffer, the Lisp interpreter evaluates the next element of the or expression. This is the expression (setq buffer (get-buffer buffer)). This expression returns a non-nil value, which is the value to which it sets the variable buffer—and this value is a buffer itself, not the name of a buffer.

The result of all this is that the symbol buffer is always bound to a buffer itself rather than the name of a buffer. All this is necessary because the set-buffer function in a following line only works with a buffer itself, not with the name to a buffer.

Incidentally, using or, the situation with the usher would be written like this:

     (or (holding-on-to-guest) (find-and-take-arm-of-guest))

5.2.5 The let Expression in insert-buffer

After ensuring that the variable buffer refers to a buffer itself and not just to the name of a buffer, the insert-buffer function continues with a let expression. This specifies three local variables, start, end, and newmark and binds them to the initial value nil. These variables are used inside the remainder of the let and temporarily hide any other occurrence of variables of the same name in Emacs until the end of the let.

The body of the let contains two save-excursion expressions. First, we will look at the inner save-excursion expression in detail. The expression looks like this:

     (save-excursion
       (set-buffer buffer)
       (setq start (point-min) end (point-max)))

The expression (set-buffer buffer) changes Emacs’s attention from the current buffer to the one from which the text will copied. In that buffer, the variables start and end are set to the beginning and end of the buffer, using the commands point-min and point-max. Note that we have here an illustration of how setq is able to set two variables in the same expression. setq’s first argument is set to the value of its second and its third argument is set to the value of its fourth.

After the body of the inner save-excursion is evaluated, the save-excursion restores the original buffer, but start and end remain set to the values of the beginning and end of the buffer from which the text will be copied.

The outer save-excursion expression looks like this:

     (save-excursion

(inner-save-excursion-expression

(go-to-new-buffer-and-set-start-and-end)

       (insert-buffer-substring buffer start end)
       (setq newmark (point)))

The insert-buffer-substring function copies the text into the current buffer from the region indicated by start and end in buffer. Since the whole of the second buffer lies between start and end, the whole of the second buffer is copied into the buffer you are editing. Next, the value of point, which will be at the end of the inserted text, is recorded in the variable newmark.

After the body of the outer save-excursion is evaluated, point and mark are relocated to their original places.

However, it is convenient to locate a mark at the end of the newly inserted text and locate point at its beginning. The newmark variable records the end of the inserted text. In the last line of the let expression, the (push-mark newmark) expression function sets a mark to this location. (The previous location of the mark is still accessible; it is recorded on the mark ring and you can go back to it with C-u C-^{⟨}\underline{\overline{SPC}}^{⟩}.) Meanwhile, point is located at the beginning of the inserted text, which is where it was before you called the insert function.

The whole let expression looks like this:

     (let (start end newmark)
       (save-excursion
         (save-excursion
           (set-buffer buffer)
           (setq start (point-min) end (point-max)))
         (insert-buffer-substring buffer start end)
         (setq newmark (point)))
       (push-mark newmark))

   Like the append-to-buffer function, the insert-buffer function uses let, save-

excursion, and set-buffer. In addition, the function illustrates one way to use or. All these functions are building blocks that we will find and use again and again.

5.3 Complete Definition of beginning-of-buffer

The basic structure of the beginning-of-buffer function has already been discussed. (See Section 4.2, “A Simplified beginning-of-buffer Definition”, page 37.) This section describes the complex part of the definition.

As previously described, when invoked without an argument, beginning-of-buffer moves the cursor to the beginning of the buffer, leaving the mark at the previous position. However, when the command is invoked with a number between one and ten, the function considers that number to be a fraction of the length of the buffer, measured in tenths, and Emacs moves the cursor that fraction of the way from the beginning of the buffer. Thus, you can either call this function with the key command M-<, which will move the cursor to the beginning of the buffer, or with a key command such as C-u 7 M-< which will move the cursor to a point 70% of the way through the buffer. If a number bigger than ten is used for the argument, it moves to the end of the buffer.

The beginning-of-buffer function can be called with or without an argument. The use of the argument is optional.

5.3.1 Optional Arguments

Unless told otherwise, Lisp expects that a function with an argument in its function definition will be called with a value for that argument. If that does not happen, you get an error and a

message that says ‘Wrong number of arguments’.

However, optional arguments are a feature of Lisp: a keyword may be used to tell the Lisp interpreter that an argument is optional. The keyword is &optional. (The ‘&’ in front of ‘optional’ is part of the keyword.) In a function definition, if an argument follows the keyword &optional, a value does not need to be passed to that argument when the function is called.

The first line of the function definition of beginning-of-buffer therefore looks like this:

     (defun beginning-of-buffer (&optional arg)

In outline, the whole function looks like this:

     (defun beginning-of-buffer (&optional arg)

"documentation..."

       (interactive "P")
       (push-mark)
       (goto-char

(if-there-is-an-argument

figure-out-where-to-go else-go-to

           (point-min))))
   The function is similar to simplified-beginning-of-buffer except that the interactive

expression has "P" as an argument and the goto-char function is followed by an if-then-else expression that figures out where to put the cursor if there is an argument.

The "P" in the interactive expression tells Emacs to pass a prefix argument, if there is one, to the function. A prefix argument is made by typing the ^{⟨}\underline{\overline{META}}^{⟩} key followed by a number, or by typing C-u and then a number (if you don’t type a number, C-u defaults to 4).

The true-or-false-test of the if expression is simple: it is simply the argument arg. If arg has a value that is not nil, which will be the case if beginning-of-buffer is called with an argument, then this true-or-false-test will return true and the then-part of the if expression will be evaluated. On the other hand, if beginning-of-buffer is not called with an argument, the value of arg will be nil and the else-part of the if expression will be evaluated. The else-part is simply point-min, and when this is the outcome, the whole goto-char expression is (gotochar (point-min)), which is how we saw the beginning-of-buffer function in its simplified form.

5.3.2 beginning-of-buffer with an Argument

When beginning-of-buffer is called with an argument, an expression is evaluated which calculates what value to pass to goto-char. This expression is rather complicated at first sight. It includes an inner if expression and much arithmetic. It looks like this:

     (if (> (buffer-size) 10000)

;; Avoid overflow for large buffer sizes!

         (* (prefix-numeric-value arg) (/ (buffer-size) 10))
       (/
        (+ 10
           (*
            (buffer-size) (prefix-numeric-value arg))) 10))

Like other complex-looking expressions, this one can be distangled by looking at it as parts of a template, in this case, the template for an if-then-else expression. When in skeletal form, the expression looks like this:

(if (buffer-is-large

divide-buffer-size-by-10-and-multiply-by-arg else-use-alternate-calculation

The true-or-false-test of this inner if expression checks the size of the buffer. The reason for this is that version 18 Emacs Lisp uses numbers that are no bigger than eight million or so (bigger numbers are not needed) and in the computation that follows, Emacs might try to use over-large numbers if the buffer were large. The term ‘overflow’, mentioned in the comment, means numbers that are over large.

There are two cases: if the buffer is large and if it is not.

What happens in a large buffer

In beginning-of-buffer, the inner if expression tests whether the size of the buffer is greater than 10,000 characters. To do this, it uses the > function and the buffer-size function. The line looks like this:

     (if (> (buffer-size) 10000)

When the buffer is large, the then-part of the if expression is evaluated. It reads like this (after formatting for easy reading):

     (*
       (prefix-numeric-value arg)
       (/ (buffer-size) 10))

This expression is a multiplication, with two arguments to the function *.

The first argument is (prefix-numeric-value arg). When "P" is used as the argument for interactive, the value passed to the function as its argument is passed a “raw prefix argument”, and not a number. (It is a number in a list.) To perform the arithmetic, a conversion is necessary,

and prefix-numeric-value does the job.

The second argument is (/ (buffer-size) 10). This expression divides the numeric value of the buffer by ten. This produces a number that tells how many characters make up one tenth of the buffer size. (In Lisp, / is used for division, just as * is used for multiplication.)

In the multiplication expression as a whole, this amount is multiplied by the value of the prefix argument—the multiplication looks like this:

(* numeric-value-of-prefix-arg number-of-characters-in-one-tenth-of-the-buffer)

If, for example, the prefix argument is ‘7’, the one-tenth value will be multiplied by 7 to give a position 70% of the way through the buffer.

The result of all this is that if the buffer is large, the goto-char expression reads like this:

     (goto-char (* (prefix-numeric-value arg)
                   (/ (buffer-size) 10)))

This puts the cursor where we want it.

What happens in a small buffer

If the buffer contains fewer than 10,000 characters, a slightly different computation is performed. You might think this is not necessary, since the first computation could do the job. However, in a small buffer, the first method may not put the cursor on exactly the desired line; the second method does a better job.

The code looks like this:

     (/ (+ 10 (* (buffer-size) (prefix-numeric-value arg))) 10))

This is code in which you figure out what happens by discovering how the functions are embedded in parentheses. It is easier to read if you reformat it with each expression indented more deeply than its enclosing expression:

       (/
        (+ 10
           (*
            (buffer-size)
            (prefix-numeric-value arg)))
        10))

Looking at parentheses, we see that the innermost operation is (prefix-numeric-value arg), which converts the raw argument to a number. This number is multiplied by the buffer size in the following expression:

     (* (buffer-size) (prefix-numeric-value arg)

This multiplication creates a number that may be larger than the size of the buffer—seven times larger if the argument is 7, for example. Ten is then added to this number and finally the large number is divided by ten to provide a value that is one character larger than the percentage position in the buffer.

The number that results from all this is passed to goto-char and the cursor is moved to that point.

5.3.3 The Complete beginning-of-buffer

Here is the complete text of the beginning-of-buffer function:

     (defun beginning-of-buffer (&optional arg)
       "Move point to the beginning of the buffer;
     leave mark at previous position.
     With arg N, put point N/10 of the way
     from the true beginning.
     Don't use this in Lisp programs!
     \(goto-char (point-min)) is faster
     and does not set the mark."
       (interactive "P")
       (push-mark)
       (goto-char
        (if arg
            (if (> (buffer-size) 10000)

;; Avoid overflow for large buffer sizes!

                (* (prefix-numeric-value arg)
                   (/ (buffer-size) 10))
              (/ (+ 10 (* (buffer-size)
                          (prefix-numeric-value arg)))
                 10))
          (point-min)))
       (if arg (forward-line 1)))

Except for two small points, the previous discussion shows how this function works. The first point deals with a detail in the documentation string, and the second point concerns the last line of the function.

In the documentation string, there is reference to an expression:

     \(goto-char (point-min))

A ‘\’ is used before the first parenthesis of this expression. This ‘\’ tells the Lisp interpreter that the expression should be printed as shown in the documentation rather than evaluated as a symbolic expression, which is what it looks like.

Finally, the last line of the beginning-of-buffer command says to move point to the beginning of the next line if the command is invoked with an argument:

     (if arg (forward-line 1)))

This puts the cursor at the beginning of the first line after the appropriate tenths position in the buffer. This is a flourish that means that the cursor is always located at least the requested tenths of the way through the buffer, which is a nicety that is, perhaps, not necessary, but which, if it did not occur, would be sure to draw complaints.

5.4 Review

Here is a brief summary of some of the topics covered in this chapter.

or Evaluate each argument in sequence, and return the value of the first argument that is not nil; if none return a value that is not nil, return nil. In brief, return the first true value of the arguments; return a true value if one or any of the other are true. and Evaluate each argument in sequence, and if any are nil, return nil; if none are

nil, return the value of the last argument. In brief, return a true value only if all the arguments are true; return a true value if one and each of the others is true.

&optional

A keyword used to indicate that an argument to a function definition is optional; this means that the function can be evaluated without the argument, if desired.

prefix-numeric-value

Convert the ‘raw prefix argument’ produced by (interactive "P") to a numeric value.

forward-line

Move point forward to the beginning of the next line, or if the argument is greater than one, forward that many lines. If it can’t move as far forward as it is supposed to, forward-line goes forward as far as it can and then returns a count of the number of additional lines it was supposed to move but couldn’t.

erase-buffer

Delete the entire contents of the current buffer.

bufferp Return t if its argument is a buffer; otherwise return nil.

5.5 &optional Argument Exercise

Write an interactive function with an optional argument that tests whether its argument, a number, is greater or less than the value of fill-column, and tells you which, in a message. However, if you do not pass an argument to the function, use 56 as a default value.

what-line  55

6 Narrowing and Widening

Narrowing is a feature of Emacs that makes it possible for you to focus on a specific part of a buffer, and work without accidentally changing other parts. Narrowing is normally disabled since it can confuse novices.

With narrowing, the rest of a buffer is made invisible, as if it weren’t there. This is an advantage if, for example, you want to replace a word in one part of a buffer but not in another: you narrow to the part you want and the replacement is carried out only in that section, not in the rest of the buffer. Searches will only work within a narrowed region, not outside of one, so if you are fixing a part of a document, you can keep yourself from accidentally finding parts you do not need to fix by narrowing just to the region you want.

However, narrowing does make the rest of the buffer invisible, which can scare people who inadvertently invoke narrowing and think they have deleted a part of their file. Moreover, the undo command (which is usually bound to C-x u) does not turn off narrowing (nor should it), so people can become quite desperate if they do not know that they can return the rest of a buffer to visibility with the widen command. (In Emacs version 18, the key binding for widen is C-x w; in version 19, it is C-x n w.)

Narrowing is just as useful to the Lisp interpreter as to a human. Often, an Emacs Lisp function is designed to work on just part of a buffer; or conversely, an Emacs Lisp function needs to work on all of a buffer that has been narrowed. The what-line function, for example, removes the narrowing from a buffer, if it has any narrowing and when it has finished its job, restores the narrowing to what it was. On the other hand, the count-lines function, which is called by what-line, uses narrowing to restrict itself to just that portion of the buffer in which it is interested and then restores the previous situation.

6.1 The save-restriction Special Form

In Emacs Lisp, you can use the save-restriction special form to keep track of whatever narrowing is in effect, if any. When the Lisp interpreter meets with save-restriction, it executes the code in the body of the save-restriction expression, and then undoes any changes to narrowing that the code caused. If, for example, the buffer is narrowed and the code that follows save-restriction gets rid of the narrowing, save-restriction returns the buffer to its narrowed region afterwards. In the what-line command, any narrowing the buffer may have is undone by the widen command that immediately follows the save-restriction command. Any original narrowing is restored just before the completion of the function.

The template for a save-restriction expression is simple:

     (save-restriction
       body... )

The body of the save-restriction is one or more expressions that will be evaluated in sequence by the Lisp interpreter.

Finally, a point to note: when you use both save-excursion and save-restriction, one right after the other, you should use save-excursion outermost. If you write them in reverse order, you may fail to record narrowing in the buffer to which Emacs switches after calling save-

excursion. Thus, when written together, save-excursion and save-restriction should be

written like this:

     (save-excursion
       (save-restriction
         body...))

6.2 what-line

The what-line command tells you the number of the line in which the cursor is located. The function illustrates the use of the save-restriction and save-excursion commands. Here is the text of the function in full:

     (defun what-line ()
       "Print the current line number (in the buffer) of point."
       (interactive)
       (save-restriction
         (widen)
         (save-excursion
           (beginning-of-line)
           (message "Line %d"
                    (1+ (count-lines 1 (point)))))))

The function has a documentation line and is interactive, as you would expect. The next two lines use the functions save-restriction and widen.

The save-restriction special form notes whatever narrowing is in effect, if any, in the current buffer and restores that narrowing after the code in the body of the save-restriction has been evaluated.

The save-restriction special form is followed by widen. This function undoes any narrowing the current buffer may have had when what-line was called. (The narrowing that was there is the narrowing that save-restriction remembers.) This widening makes it possible for the line counting commands to count from the beginning of the buffer. Otherwise, they would have been limited to counting within the accessible region. Any original narrowing is restored just before the completion of the function by the save-restriction special form.

The call to widen is followed by save-excursion, which saves the location of the cursor (i.e., of point) and of the mark, and restores them after the code in the body of the save-excursion uses the beginning-of-line function to move point.

(Note that the (widen) expression comes between save-restriction and save-excursion. When you write the two save- ... expressions in sequence, write save-excursion outermost.)

The last two lines of the what-line function are functions to count the number of lines in the buffer and then print the number in the echo area.

     (message "Line %d"
              (1+ (count-lines 1 (point)))))))

The message function prints a one-line message at the bottom of the Emacs screen. The first argument is inside of quotation marks and is printed as a string of characters. However, it may contain ‘%d’, ‘%s’, or ‘%c’ to print arguments that follow the string. ‘%d’ prints the argument as a decimal, so the message will say something such as ‘Line 243’.

The number that is printed in place of the ‘%d’ is computed by the last line of the function:

     (1+ (count-lines 1 (point)))

What this does is count the lines from the first position of the buffer, indicated by the 1, up to (point), and then add one to that number. (The 1+ function adds one to its argument.) We add one to it because line 2 has only one line before it, and count-lines counts only the lines before the current line.

After count-lines has done it job, and the message has been printed in the echo area, the save-excursion restores point and mark to their original positions; and save-restriction restores the original narrowing, if any.

6.3 Exercise with Narrowing

Write a function that will display the first 60 characters of the current buffer, even if you have narrowed the buffer to its latter half so that the first line is inaccessible. Restore point, mark, and narrowing. For this exercise, you need to use save-restriction, widen, goto-char, point-min, buffer-substring, message, and other functions, a whole potpourri.

58 Chapter 6: Narrowing and Widening

car and cdr  59

7 car, cdr, cons: Fundamental Functions

In Lisp, car, cdr, and cons are fundamental functions. The cons function is used to construct lists, and the car and cdr functions are used to take them apart.

In the walk through of the copy-region-as-kill function, we will see cons as well as two variants on cdr, namely, setcdr and nthcdr. (See Section 8.5, “copy-region-as-kill”, page 73.)

The name of the cons function is not unreasonable: it is an abbreviation of the word ‘construct’. The origins of the names for car and cdr, on the other hand, are esoteric: car is an acronym from the phrase ‘Contents of the Address part of the Register’; and cdr (pronounced ‘could-er’) is an acronym from the phrase ‘Contents of the Decrement part of the Register’. These phrases refer to specific pieces of hardware on the very early computer on which the original Lisp was developed. Besides being obsolete, the phrases have been completely irrelevant for more than 25 years to anyone thinking about Lisp. Nonetheless, although a few brave scholars have begun to use more reasonable names for these functions, the old terms are still in use. In particular, since the terms are used in the Emacs Lisp source code, we will use them in this introduction.

7.1 car and cdr

The car of a list is, quite simply, the first item in the list. Thus the car of the list (rose

violet daisy buttercup) is rose.

If you are reading this in Info in GNU Emacs, you can see this by evaluating the following:

     (car '(rose violet daisy buttercup))

After evaluating the expression, rose will appear in the echo area.

Clearly, a more reasonable name for the car function would be first and this is often suggested.

car does not remove the first item from the list; it only reports what it is. After car has been applied to a list, the list is still the same as it was. In the jargon, car is ‘non-destructive’. This feature turns out to be important.

The cdr of a list is the rest of the list, that is, the cdr function returns the part of the list that follows the first item. Thus, while the car of the list '(rose violet daisy buttercup) is rose, the rest of the list, the value returned by cdr, is (violet daisy buttercup).

You can see this by evaluating the following in the usual way:

     (cdr '(rose violet daisy buttercup))

When you evaluate this, (violet daisy buttercup) will appear in the echo area.

Like car, cdr does not remove any elements from the list—it just returns a report of what the second and subsequent elements are.

Incidentally, in the example, the list of flowers is quoted. If it were not, the Lisp interpreter would try to evaluate the list by calling rose as a function. In this example, we do not want to do that.

Clearly, a more reasonable name for cdr would be rest.

(There is a lesson here: when you name new functions, consider very carefully about what you are doing, since you may be stuck with the names for far longer than you expect. The reason this document perpetuates these names is that the Emacs Lisp source code uses them, and if I did not use them, you would have a hard time reading the code; but do please try to avoid using these terms yourself. The people who come after you will be grateful to you.)

When car and cdr are applied to a list made up of symbols, such as the list (pine fir oak maple), the element of the list returned by the function car is the symbol pine without any parentheses around it. pine is the first element in the list. However, the cdr of the list is a list itself, (fir oak maple), as you can see by evaluating the following expressions in the usual way:

     (car '(pine fir oak maple))
     (cdr '(pine fir oak maple))

On the other hand, in a list of lists, the first element is itself a list. car returns this first element as a list. For example, the following list contains three sub-lists, a list of carnivores, a list of herbivores and a list of sea mammals:

     (car '((lion tiger cheetah)
            (gazelle antelope zebra)
            (whale dolphin seal)))

In this case, the first element or car of the list is the list of carnivores, (lion tiger cheetah),

and the rest of the list is ((gazelle antelope zebra) (whale dolphin seal)).
     (cdr '((lion tiger cheetah)
            (gazelle antelope zebra)
            (whale dolphin seal)))

It is worth saying again that car and cdr are non-destructive—that is, they do not modify or change lists to which they are applied. This is very important for how they are used.

Also, in the first chapter, in the discussion about atoms, I said that in Lisp, “certain kinds of atom, such as an array, can be separated into parts; but the mechanism for doing this is different from the mechanism for splitting a list. As far as Lisp is concerned, the atoms of a list are unsplittable.” (See Section 1.1.1, “Lisp Atoms”, page 1.) The car and cdr functions are used for splitting lists and are considered fundamental to Lisp. Since they cannot split or gain access to the parts of an array, an array is considered an atom. Conversely, the other fundamental function, cons, can put together or construct a list, but not an array. (Arrays are handled by array-specific functions. See section “Arrays” in The GNU Emacs Lisp Reference Manual.)

7.2 cons

The cons function constructs lists; it is the inverse of car and cdr. For example, cons can be used to make a four element list from the three element list, (fir oak maple):

     (cons 'pine '(fir oak maple))

After evaluating this list, you will see

     (pine fir oak maple)

appear in the echo area. cons puts a new element at the beginning of a list; it attaches or pushes elements onto the list.

cons must have a list to attach to.^{1} You cannot start from absolutely nothing. If you are building a list, you need to provide at least an empty list at the beginning. Here is a series of cons’s that build up a list of flowers. If you are reading this in Info in GNU Emacs, you can evaluate each of the expressions in the usual way; the value is printed in this text after ‘⇒’, which you may read as ‘evaluates to’.

     (cons 'buttercup ())
          ⇒ (buttercup)
     (cons 'daisy '(buttercup))
          ⇒ (daisy buttercup)
     (cons 'violet '(daisy buttercup))
          ⇒ (violet daisy buttercup)
     (cons 'rose '(violet daisy buttercup))
          ⇒ (rose violet daisy buttercup)

In the first example, the empty list is shown as () and a list made up of buttercup followed by the empty list is constructed. As you can see, the empty list is not shown in the list that was constructed. All that you see is (buttercup). The empty list is not counted as an element of a list because there is nothing in an empty list. Generally speaking, an empty list is invisible.

The second example, (cons 'daisy '(buttercup)) constructs a new, two element list by putting daisy in front of buttercup; and the third example constructs a three element list by

putting violet in front of daisy and buttercup.

7.2.1 Find the Length of a List: length

You can find out how many elements there are in a list by using the Lisp function length, as in the following examples:

     (length '(buttercup))
          ⇒ 1
     (length '(daisy buttercup))
          ⇒ 2
     (length (cons 'violet '(daisy buttercup)))
          ⇒ 3

In the third example, the cons function is used to construct a three element list which is then passed to the length function as its argument.

We can also use length to count the number of elements in an empty list:

     (length ()) ⇒ 0

As you would expect, the number of elements in an empty list is zero.

An interesting experiment is to find out what happens if you try to find the length of no list at all; that is, if you try to call length without giving it an argument, not even an empty list:

     (length )

What you see, if you evaluate this, is the error message

1 Actually, you can cons an element to an atom to produce a dotted pair. Dotted pairs are not discussed here; see section “Dotted Pair Notation” in The GNU Emacs Lisp Reference Manual.

     Wrong number of arguments: #<subr length>, 0

This means that the function receives the wrong number of arguments, zero, when it expects some other number of arguments. In this case, one argument is expected, the argument being a list whose length the function is measuring. (Note that one list is one argument, even if the list has many elements inside it.)

The part of the error message that says ‘#<subr length>’ is the name of the function. This is written with a special notation, ‘#<subr’, that indicates that the function length is one of the primitive functions written in C rather than in Emacs Lisp. (‘subr’ is an abbreviation for ‘subroutine’.) See section “What Is a Function?” in The GNU Emacs Lisp Reference Manual, for more about subroutines.

7.3 nthcdr

The nthcdr function is associated with the cdr function. What it does is take the cdr of a list repeatedly.

If you take the cdr of the list (pine fir oak maple), you will be returned the list (fir oak maple). If you repeat this on what was returned, you will be returned the list (oak maple). (Of course, repeated cdring on the original list will just give you the original cdr since the function does not change the list. You need to evaluate the cdr of the cdr and so on.) If you continue this, eventually you will be returned an empty list, which in this case, instead of being shown as () is shown as nil.

For review, here is a series of repeated cdrs, the text following the ‘⇒’ shows what is returned.

     (cdr '(pine fir oak maple))
          ⇒(fir oak maple)
     (cdr '(fir oak maple))
          ⇒ (oak maple)
     (cdr '(oak maple))
          ⇒(maple)
     (cdr '(maple))
          ⇒ nil
     (cdr 'nil) ⇒ nil
     (cdr ()) ⇒ nil

You can also do several cdrs without printing the values in between, like this:

     (cdr (cdr '(pine fir oak maple)))
          ⇒ (oak maple)

In this case, the Lisp interpreter evaluates the innermost list first. The innermost list is quoted, so it just passes the list as it is to the innermost cdr. This cdr passes a list made up of the second and subsequent elements of the list to the outermost cdr, which produces a list composed of the third and subsequent elements of the original list. In this example, the cdr function is repeated and returns a list that consists of the original list without its first two elements.

The nthcdr function does the same as repeating the call to cdr. In the following example, the argument 2 is passed to the function nthcdr, along with the list, and the value returned is the list without its first two items, which is exactly the same as repeating cdr twice on the list:

setcar  63

     (nthcdr 2 '(pine fir oak maple))
          ⇒ (oak maple)

Using the original four element list, we can see what happens when various numeric arguments are passed to nthcdr, including 0, 1, and 5:

;; Leave the list as it was.

     (nthcdr 0 '(pine fir oak maple))
          ⇒ (pine fir oak maple)

;; Return a copy without the first element.

     (nthcdr 1 '(pine fir oak maple))
          ⇒ (fir oak maple)

;; Return a copy of the list without three elements.

     (nthcdr 3 '(pine fir oak maple))
          ⇒ (maple)

;; Return a copy lacking all four elements.

     (nthcdr 4 '(pine fir oak maple))
          ⇒ nil

;; Return a copy lacking all elements.

     (nthcdr 5 '(pine fir oak maple))
          ⇒ nil

It is worth mentioning that nthcdr, like cdr, does not change the original list—the function is non-destructive. This is in sharp contrast to the setcar and setcdr functions.

7.4 setcar

As you might guess from their names, the setcar and setcdr functions set the car or the cdr of a list to a new value. They actually change the original list, unlike car and cdr which leave the original list as it was. One way to find out how this works is to experiment. We will start with the setcar function.

First, we can make a list and then set the value of a variable to the list, using the setq function. Here is a list of animals:

     (setq animals '(giraffe antelope tiger lion))

If you are reading this in Info inside of GNU Emacs, you can evaluate this expression in the usual fashion, by positioning the cursor after the expression and typing C-x C-e. (I’m doing this right here as I write this. This is one of the advantages of having the interpreter built into the computing environment.)

When we evaluate the variable animals, we see that it is bound to the list (giraffe antelope

tiger lion):
     animals
          ⇒ (giraffe antelope tiger lion)

Put another way, the variable animals points to the list (giraffe antelope tiger lion).

Next, evaluate the function setcar while passing it two arguments, the variable animals and the quoted symbol hippopotamus; this is done by writing the three element list (setcar animals 'hippopotamus) and then evaluating it in the usual fashion:

     (setcar animals 'hippopotamus)

After evaluating this expression, evaluate the variable animals again. You will see that the list of animals has changed:

     animals
          ⇒ (hippopotamus antelope tiger lion)

The first element on the list, giraffe is replaced by hippopotamus.

So we can see that setcar did not add a new element to the list as cons would have; it replaced giraffe with hippopotamus; it changed the list.

7.5 setcdr

The setcdr function is similar to the setcar function, except that the function replaces the second and subsequent elements of a list rather than the first element.

To see how this works, set the value of the variable to a list of domesticated animals by evaluating the following expression:

     (setq domesticated-animals '(horse cow sheep goat))

If you now evaluate the list, you will be returned the list (horse cow sheep goat):

     domesticated-animals
          ⇒ (horse cow sheep goat)

Next, evaluate setcdr with two arguments, the name of the variable which has a list as its value, and the list to which the cdr of the first list will be set;

     (setcdr domesticated-animals '(cat dog))

If you evaluate this expression, the list (cat dog) will appear in the echo area. This is the value returned by the function. The result we are interested in is the “side effect”, which we can see by evaluating the variable domesticated-animals:

     domesticated-animals
          ⇒ (horse cat dog)

Indeed, the list is changed from (horse cow sheep goat) to (horse cat dog). The cdr of the

list is changed from (cow sheep goat) to (cat dog).

7.6 Exercise

Construct a list of four birds by evaluating several expressions with cons. Find out what happens when you cons a list onto itself. Replace the first element of the list of four birds with a fish. Replace the rest of that list with a list of other fish.

zap-to-char  65

8 Cutting and Storing Text

Whenever you cut or clip text out of a buffer with a ‘kill’ command in GNU Emacs, it is stored in a list and you can bring it back with a ‘yank’ command.

(The use of the word ‘kill’ in Emacs for processes which specifically do not destroy the values of the entities is an unfortunate historical accident. A much more appropriate word would be ‘clip’ since that is what the kill commands do; they clip text out of a buffer and put it into storage from which it can be brought back. I have often been tempted to replace globally all occurrences of ‘kill’ in the Emacs sources with ‘clip’ and all occurrences of ‘killed’ with ‘clipped’.)

When text is cut out of a buffer, it is stored on a list. Successive pieces of text are stored on the list successively, so the list might look like this:

     ("a piece of text" "last piece")

The function cons can be used to add a piece of text to the list, like this:

     (cons "another piece"
           '("a piece of text" "last piece"))

If you evaluate this expression, a list of three elements will appear in the echo area:

     ("another piece" "a piece of text" "last piece")

With the car and nthcdr functions, you can retrieve whichever piece of text you want. For example, in the following code, nthcdr 1 ... returns the list with the first item removed; and the car returns the first element of that remainder—the second element of the original list:

     (car (nthcdr 1 '("another piece"
                      "a piece of text"
                      "last piece")))
          ⇒ "a piece of text"

The actual functions in Emacs are more complex than this, of course. The code for cutting and retrieving text has to be written so that Emacs can figure out which element in the list you want—the first, second, third, or whatever. In addition, when you get to the end of the list, Emacs should give you the first element of the list, rather than nothing at all.

The list that holds the pieces of text is called the kill ring. This chapter leads up to a description of the kill ring and how it is used by first tracing how the zap-to-char function works. This function uses (or ‘calls’) a function that invokes a function that manipulates the kill ring. Thus, before reaching the mountains, we climb the foothills.

A subsequent chapter describes how text that is cut from the buffer is retrieved. See Chapter 10, “Yanking Text Back”, page 85.

8.1 zap-to-char

The zap-to-char function is written differently in GNU Emacs version 18 and version 19. The version 19 implementation is simpler, and works slightly differently. We will first show the function as it is written for version 19 and then for version 18.

The Emacs version 19 implementation of the interactive zap-to-char function removes the text in the region between the location of the cursor (i.e., of point) up to and including the next occurrence of a specified character. The text that zap-to-char removes is put in the kill ring; and it can be retrieved from the kill ring by typing C-y (yank). If the command is given an argument, it removes text through that number of occurrences. Thus, if the cursor were at the beginning of this sentence and the character were ‘s’, ‘Thus’ would be removed. If the argument were two, ‘Thus, if the curs’ would be removed, up to and including the ‘s’ in ‘cursor’.

The Emacs version 18 implementation removes the text from point up to but not including the specified character. Thus, in the example shown in the previous paragraph, the ‘s’ would not be removed.

In addition, the version 18 implementation will go to the end of the buffer if the specified character is not found; but the version 19 implementation will simply generate an error (and not remove any text).

In order to determine how much text to remove, both versions of zap-to-char use a search function. Searches are used extensively in code that manipulates text, and it is worth focusing attention on the search function as well as on the deletion command.

Here is the complete text of the version 19 implementation of the function:

     (defun zap-to-char (arg char)  ; version 19 implementation
       "Kill up to and including ARG'th occurrence of CHAR.
     Goes backward if ARG is negative; error if CHAR not found."
       (interactive "*p\ncZap to char: ")
       (kill-region (point)
                    (progn
                      (search-forward
                       (char-to-string char) nil nil arg)
                      (point))))

8.1.1 The interactive Expression

The interactive expression in the zap-to-char command looks like this:

     (interactive "*p\ncZap to char: ")

The part within quotation marks, "*p\ncZap to char: ", specifies three different things. First, and most simply, the asterisk, ‘*’, causes an error to be signalled if the buffer is read-only. This means that if you try zap-to-char in a read-only buffer you will not be able to remove text, and you will receive a message that says “Buffer is read-only”; your terminal may beep at you as well.

The second part of "*p\ncZap to char: " is the ‘p’. This part is ended by a newline, ‘\n’. The ‘p’ means that the first argument to the function will be passed the value of a ‘processed prefix’. The prefix argument is passed by typing C-u and a number, or M- and a number. If the function is called interactively without a prefix, 1 is passed to this argument.

The third part of "*p\ncZap to char: " is ‘cZap to char: ’. In this part, the lower case ‘c’ indicates that interactive expects a prompt and that the argument will be a character. The prompt follows the ‘c’ and is the string ‘Zap to char: ’ (with a space after the colon to make it look good).

What all this does is prepare the arguments to zap-to-char so they are of the right type, and give the user a prompt.

8.1.2 The Body of zap-to-char

The body of the zap-to-char function contains the code that kills (that is, removes) the text in the region from the current position of the cursor up to and including the specified character. The first part of the code looks like this:

 (kill-region (point) ...

(point) is the current position of the cursor.

The next part of the code is an expression using progn. The body of the progn consists of

calls to search-forward and point.

It is easier to understand how progn works after learning about search-forward, so we will

look at search-forward and then at progn.

8.1.3 The search-forward Function

The search-forward function is used to locate the zapped-for-character in zap-to-char. If the search is successful, search-forward leaves point immediately after the last character in the target string. (In this case the target string is just one character long.) If the search is backwards, search-forward leaves point just before the first character in the target. Also, search-forward returns t for true. (Moving point is therefore a ‘side effect’.)

In zap-to-char, the search-forward function looks like this:

 (search-forward (char-to-string char) nil nil arg)

The search-forward function takes four arguments:

1. The first argument is the target, what is searched for. This must be a string, such as ‘"z"’.

As it happens, the argument passed to zap-to-char is a single character. Because of the way computers are built, the Lisp interpreter treats a single character as being different from a string of characters. Inside the computer, a single character has a different electronic format than a string of one character. (A single character can often be recorded in the computer using exactly one byte; but a string may be longer or shorter, and the computer needs to be ready for this.) Since the search-forward function searches for a string, the character that the zap-to-char function receives as its argument must be converted inside the computer from one format to the other; otherwise the search-forward function will fail. The char-to-string function is used to make this conversion.

2. The second argument bounds the search; it is specified as a position in the buffer. In this case, the search can go to the end of the buffer, so no bound is set and the second argument

is nil.

3. The third argument tells the function what it should do if the search fails—it can signal an error (and print a message) or it can return nil. A nil as the third argument causes the function to signal an error when the search fails.

4. The fourth argument to search-forward is the repeat count—how many occurrences of the string to look for. This argument is optional and if the function is called without a repeat count, this argument is passed the value 1. If this argument is negative, the search goes backwards.

In template form, a search-forward expression looks like this:

 (search-forward "target-string"

limit-of-search

what-to-do-if-search-fails repeat-count)

We will look at progn next.

8.1.4 The progn Function

progn is a function that causes each of its arguments to be evaluated in sequence and then returns the value of the last one. The preceding expressions are evaluated only for the side effects they perform. The values produced by them are discarded.

The template for a progn expression is very simple:

     (progn body...)

In zap-to-char, the progn expression has to do two things: put point in exactly the right position; and return the location of point so that kill-region will know how far to kill to.

The first argument to the progn is search-forward. When search-forward finds the string, the function leaves point immediately after the last character in the target string. (In this case the target string is just one character long.) If the search is backwards, search-forward leaves point just before the first character in the target. The movement of point is a side effect.

The second and last argument to progn is the expression (point). This expression returns the value of point, which in this case will be the location to which it has been moved by search-forward. This value is returned by the progn expression and is passed to kill-region as kill-region’s second argument.

8.1.5 Summing up zap-to-char

Now that we have seen how search-forward and progn work, we can see how the zap-to-char function works as a whole.

The first argument to kill-region is the position of the cursor when the zap-to-char command is given—the value of point at that time. Within the progn, the search function then moves point to just after the zapped-to-character and point returns the value of this location. The kill-region function puts together these two values of point, the first one as the beginning of the region and the second one as the end of the region, and removes the region.

The progn function is necessary because the kill-region command takes two arguments; and it would fail if search-forward and point expressions were written in sequence as two additional arguments. The progn expression is a single argument to kill-region and returns the one value that kill-region needs for its second argument.

8.1.6 The Version 18 Implementation

The version 18 implementation of zap-to-char is slightly different from the version 19 implementation: it zaps the text up to but not including the zapped-to-character; and zaps to the end of the buffer if the specified character is not found.

The difference is in the second argument to the kill-region command. Where the version 19 implementation looks like this:

     (progn
       (search-forward (char-to-string char) nil nil arg)
       (point))

The version 18 implementation looks like this:

     (if (search-forward (char-to-string char) nil t arg)
         (progn (goto-char
                 (if (> arg 0) (1- (point)) (1+ (point))))
                (point))
       (if (> arg 0)
           (point-max)
         (point-min)))

This looks considerably more complicated, but the code can be readily understood if it is looked at part by part.

The first part is:

     (if (search-forward (char-to-string char) nil t arg)

This fits into an if expression that does the following job, as we shall see:

(if able-to-locate-zapped-for-character-and-move-point-to-it then-move-point-to-the-exact-spot-and-return-this-location else-move-to-end-of-buffer-and-return-that-location)

Evaluation of the if expression specifies the second argument to kill-region. Since the first argument is point, this process makes it possible for kill-region to remove the text between point and the zapped-to location.

We have already described how search-forward moves point as a side effect. The value that search-forward returns is t if the search is successful and either nil or an error message depending on the value of the third argument to search-forward. In this case, t is the third argument and it causes the function to return nil when the search fails. As we will see, it is easy to write the code for handling the case when the search returns nil.

In the version 18 implementation of zap-to-char, the search takes place because the if causes the search expression to be evaluated as its true-or-false-test. If the search is successful, Emacs evaluates the then-part of the if expression. On the other hand, if the search fails, Emacs evaluates the else-part of the if expression.

In the if expression, when the search succeeds, a progn expression is executed—which is to say, it is run as a program.

As we said earlier, progn is a function that causes each of its arguments to be evaluated in sequence and then returns the value of the last one. The preceding expressions are evaluated only for the side effects they perform. The values produced by them are discarded.

In this version of zap-to-char, the progn expression is executed when the search-forward function finds the character for which it is searching. The progn expression has to do two things: put point in exactly the right position; and return the location of point so that kill-region will know how far to kill to.

The reason for all the code in the progn is that when search-forward finds the string it is looking for, it leaves point immediately after the last character in the target string. (In this case the target string is just one character long.) If the search is backwards, search-forward leaves point just before the first character in the target.

However, this version of the zap-to-char function is designed so that it does not remove the character being zapped to. For example, if zap-to-char is to remove all the text up to a ‘z’, this version will not remove the ‘z’ as well. So point has to be moved just enough that the zapped-to character is not removed.

The body of the progn expression

The body of the progn consists of two expressions. Spread out to delineate the different parts more clearly, and with comments added, the progn expression looks like this:

     (progn

(goto-char ; First expression in progn.

             (if (> arg 0)       ; If arg is positive,

(1- (point)) ; move back one character; (1+ (point)))) ; else move forward one character.

(point)) ; Second expression in progn: ; return position of point.

The progn expression does this: when the search is forward (arg is positive), Emacs leaves point just after the searched-for character. By moving point back one position, the character is uncovered. In this case, the expression in the progn reads as follows: (goto-char (1- (point))). This moves point one character back. (The 1- function subtracts one from its argument, just as 1+ adds ones to its argument.) On the other hand, if the argument to zap-to-character is negative, the search will be backwards. The if detects this and the expression reads: (goto-char (1+ (point))). (The 1+ function adds one to its argument.)

The second and last argument to progn is the expression (point). This expression returns the value of the position to which point is moved by the first argument to progn. This value is then returned by the if expression of which it is a part and is passed to kill-region as kill-region’s second argument.

In brief, the function works like this: the first argument to kill-region is the position of the cursor when the zap-to-char command is given—the value of point at that time. The search function then moves point if the search is successful. The progn expression moves point just enough so the zapped to character is not removed, and returns the value of point after all this is done. The kill-region function then removes the region.

Finally, the else-part of the if expression takes care of the situation in which the zappedtowards character is not found. If the argument to the zap-to-char function is positive (or if none is given) and the zapped-to character is not found, all the text is removed from the current position of point to the end of the accessible region of the buffer (the end of the buffer if there is no narrowing). If the arg is negative and the zapped-to character is not found, the operation goes to the beginning of the accessible region. The code for this is a simple if clause:

     (if (> arg 0) (point-max) (point-min))

This says that if arg is a positive number, return the value of point-max, otherwise, return the value of point-min.

For review, here is the code involving kill-region, with comments:

     (kill-region

(point) ; beginning-of-region

      (if (search-forward
           (char-to-string char) ; target

nil ; limit-of-search: none t ; Return nil if fail. arg) ; repeat-count.

(progn ; then-part

            (goto-char
             (if (> arg 0)
                 (1- (point))
               (1+ (point))))
            (point))
        (if (> arg 0)            ; else-part
            (point-max)
          (point-min))))

As you can see, the version 19 implementation does slightly less than the version 18 implementation, but is much simpler.

8.2 kill-region

The zap-to-char function uses the kill-region function. This function is very simple; leaving out part of its documentation string, it looks like this:

     (defun kill-region (beg end)
       "Kill between point and mark.
     The text is deleted but saved in the kill ring."
       (interactive "*r")
       (copy-region-as-kill beg end)
       (delete-region beg end))

The main point to note is that it uses the delete-region and copy-region-as-kill functions which are described in following sections.

8.3 delete-region: A Digression into C

The zap-to-char command uses the kill-region function, which in turn uses two other

functions, copy-region-as-kill and delete-region. The copy-region-as-kill function

will be described in a following section; it puts a copy of the region in the kill ring so it can be yanked back. (See Section 8.5, “copy-region-as-kill”, page 73.)

The delete-region function removes the contents of a region and you cannot get it back.

Unlike the other code discussed here, delete-region is not written in Emacs Lisp; it is written in C and is one of the primitives of the GNU Emacs system. Since it is very simple, I will digress briefly from Lisp and describe it here.

Like many of the other Emacs primitives, delete-region is written as an instance of a C macro, a macro being a template for code. The first section of the macro looks like this:

     DEFUN ("delete-region", Fdelete_region, Sdelete_region, 2, 2, "r",
       "Delete the text between point and mark.\n\
     When called from a program, expects two arguments,\n\
     character numbers specifying the stretch to be deleted.")

Without getting into the details of the macro writing process, let me point out that this macro starts with the word DEFUN. The word DEFUN was chosen since the code serves the same purpose as defun does in Lisp. The word DEFUN is followed by seven parts inside of parentheses:

• The first part is the name given to the function in Lisp, in this case, delete-region.

• The second part is the name of the function in C, Fdelete_region. By convention, it starts with ‘F’. Since C does not use hyphens in names, an underscore is used instead.

• The third part is the name for the C constant structure that records information on this function for internal use. It is the name of the function in C but begins with an ‘S’ instead of an ‘F’.

• The fourth and fifth parts specify the minimum and maximum number of arguments the function can have. In this case, the function demands exactly 2 arguments.

• The sixth part is just like the argument that follows the interactive declaration in a function written in Lisp: a letter followed, perhaps, by a prompt. In this case, the letter is "r" which indicates that the two arguments to the function will be the position of the beginning and end of a region in the buffer. In this code, there isn’t any prompt.

• The seventh part is a documentation string, just like the one for a function written in Emacs Lisp, except that every newline must be written explicitly as ‘\n’ followed by a backslash and carriage return.

The formal parameters come next, with a statement of what kind of object they are, and then what might be called the ‘body’ of the macro. For delete-region the ‘body’ consists of the following three lines:

     validate_region (&b, &e);
     del_range (XINT (b), XINT (e));
     return Qnil;

The first function, validate_region checks whether the values passed as the beginning and end of the region are the proper type and are within range. The second function, del_range, actually deletes the text. If the function completes its work without error, the third line returns Qnil to indicate this.

del_range is a complex function we will not look into. It updates the buffer and does other things. However, it is worth looking at the two arguments passed to del_range. These are XINT (b) and XINT (e). As far as the C language is concerned, b and e are two thirty-two bit integers that mark the beginning and end of the region to be deleted. But like other numbers in Emacs Lisp, only twenty-four bits of the thirty-two bits are used for the number; the remaining eight bits are used for keeping track of the type of information and other purposes. (On certain machines, only six bits are so used.) In this case, the eight bits are used to indicate that these numbers are for buffer positions. When bits of a number are used this way, they are called a tag. The use of the eight bit tag on each thirty-two bit integer made it possible to write Emacs to run much faster than it would otherwise. On the other hand, with numbers limited to twenty-four bits, Emacs buffers are limited to approximately eight megabytes. (You can sharply increase the maximum buffer size by adding defines for VALBITS and GCTYPEBITS in the ‘emacs/src/config.h’ file before compiling. See the note in the ‘emacs/etc/FAQ’ file that is part of the Emacs distribution.)

‘XINT’ is a C macro that extracts the 24 bit number from the thirty-two bit Lisp object; the eight bits used for other purposes are discarded. So del_range (XINT (b), XINT (e)) deletes the region between the beginning position, b, and the ending position, e.

From the point of view of the person writing Lisp, Emacs is all very simple; but hidden underneath is a great deal of complexity to make it all work.

8.4 Initializing a Variable with defvar

Unlike the delete-region function, the copy-region-as-kill function is written in Emacs Lisp. It copies a region in a buffer and saves it in a variable called the kill-ring. This section describes how this variable is created and initialized.

copy-region-as-kill  73

(Again we note that the term kill-ring is a misnomer. The text that is clipped out of the buffer can be brought back; it is not a ring of corpses, but a ring of resurrectable text.)

In Emacs Lisp, a variable such as the kill-ring is created and given an initial value by using the defvar special form. The name comes from “define variable”.

The defvar special form is similar to setq in that it sets the value of a variable. It is unlike setq in two ways: first, it only sets the value of the variable if the variable does not already have a value. If the variable already has a value, defvar does not override the existing value. Second, defvar has a documentation string.

You can see the current value of a variable, any variable, by using the describe-variable function, which is usually invoked by typing C-h v. If you type C-h v and then kill-ring (followed by ^{⟨}\underline{\overline{RET}}^{⟩}) when prompted, you will see what is in your current kill ring—this may be quite a lot! Conversely, if you have been doing nothing this Emacs session except read this document, you may have nothing in it. At the end of the ‘*Help*’ buffer, you will see the documentation for kill-ring:

     Documentation:
     List of killed text sequences.

The kill ring is defined by a defvar in the following way:

     (defvar kill-ring nil
       "List of killed text sequences.")

In this variable definition, the variable is given an initial value of nil, which makes sense, since if you have saved nothing, you want nothing back if you give a yank command. The documentation string is written just like the documentation string of a defun. As with the documentation string of the defun, the first line of the documentation should be a complete sentence, since some commands, like apropos, print only the first line of documentation. Succeeding lines should not be indented; otherwise they look odd when you use C-h v (describe-variable).

Most variables are internal to Emacs, but some are intended as options that you can readily set with the edit-options command. (These settings last only for the duration of an editing session; to set a value permanently, write a ‘.emacs’ file. See Chapter 16, “Your ‘.emacs’ File”, page 149.)

A readily settable variable is distinguished from others in Emacs by an asterisk, ‘*’, in the first column of its documentation string.

For example:

     (defvar line-number-mode nil
       "*Non-nil means display line number in mode line.")

This means that you can use the edit-options command to change the value of line-number-

mode.

Of course, you can also change the value of line-number-mode by evaluating it within a setq expression, like this:

     (setq line-number-mode t)

See Section 1.9.2, “Using setq”, page 13.

8.5 copy-region-as-kill

The copy-region-as-kill function copies a region of text from a buffer and saves it in a variable called the kill-ring.

If you call copy-region-as-kill immediately after a kill-region command, Emacs appends the newly copied text to the previously copied text. This means that if you yank back the text, you get it all, from both this and the previous operation. On the other hand, if some other command precedes the copy-region-as-kill, the function copies the text into a separate entry in the kill ring.

Here is the complete text of the version 18 copy-region-as-kill, formatted for clarity with several comments added:

     (defun copy-region-as-kill (beg end)
       "Save the region as if killed, but don't kill it."
       (interactive "r")
       (if (eq last-command 'kill-region)

;; then-part: Combine newly copied text

;; with previously copied text.

           (kill-append (buffer-substring beg end) (< end beg))

;; else-part: Add newly copied text as a new element ;; to the kill ring and shorten the kill ring if necessary.

         (setq kill-ring
               (cons (buffer-substring beg end) kill-ring))
         (if (> (length kill-ring) kill-ring-max)
             (setcdr (nthcdr (1- kill-ring-max) kill-ring) nil)))
       (setq this-command 'kill-region)
       (setq kill-ring-yank-pointer kill-ring))

As usual, this function can be divided into its component parts:

     (defun copy-region-as-kill (argument-list)

"documentation..."

       (interactive "r")
       body...)

The arguments are beg and end and the function is interactive with "r", so the two arguments must refer to the beginning and end of the region. If you have been reading though this document from the beginning, understanding these parts of a function is almost becoming routine.

The documentation is somewhat confusing unless you remember that the word ‘kill’ has a meaning different from its usual meaning.

The body of the function starts with an if clause. What this clause does is distinguish between two different situations: whether or not this command is executed immediately after a previous kill-region command. In the first case, the new region is appended to the previously copied text. Otherwise, it is inserted into the beginning of the kill ring as a separate piece of text from the previous piece.

The last two lines of the function are two setq expressions. One of them sets the variable

this-command to kill-region and the other sets the variable kill-ring-yank-pointer to

point to the kill ring.

The body of copy-region-as-kill merits discussion in detail.

8.5.1 The Body of copy-region-as-kill

The copy-region-as-kill function is written so that two or more kills in a row combine their text into a single entry. If you yank back the text from the kill ring, you get it all in one piece. Moreover, kills that kill forward from the current position of the cursor are added to the end of the previously copied text and commands that copy text backwards add it to the beginning of the previously copied text. This way, the words in the text stay in the proper order.

The function makes use of two variables that keep track of the current and previous Emacs command. The two variables are this-command and last-command.

Normally, whenever a function is executed, Emacs sets the value of this-command to the function being executed (which in this case would be copy-region-as-kill). At the same time, Emacs sets the value of last-command to the previous value of this-command. However, the copy-region-as-kill command is different; it sets the value of this-command to killregion, which is the name of the function that calls copy-region-as-kill.

In the first part of the body of the copy-region-as-kill function, an if expression determines whether the value of last-command is kill-region. If so, the then-part of the if expression is evaluated; it uses the kill-append function to concatenate the text copied at this call to the function with the text already in the first element (the car) of the kill ring. On the

other hand, if the value of last-command is not kill-region, then the copy-region-as-kill

function attaches a new element to the kill ring.

The if expression reads as follows; it uses eq, which is a function we have not yet seen:

     (if (eq last-command 'kill-region)

;; then-part

         (kill-append (buffer-substring beg end) (< end beg))

The eq function tests whether its first argument is the same Lisp object as its second argument. The eq function is similar to the equal function in that it is used to test for equality, but differs in that it determines whether two representations are actually the same object inside the computer, but with different names. equal determines whether the structure and contents of two expressions are the same.

The kill-append function

The kill-append function looks like this:

     (defun kill-append (string before-p)
       (setcar kill-ring
               (if before-p
                   (concat string (car kill-ring))
                   (concat (car kill-ring) string))))

We can look at this function in parts. The setcar function uses concat to concatenate the new text to the car of the kill ring. Whether it prepends or appends the text depends on the results of an if expression:

     (if before-p                            ; if-part
         (concat string (car kill-ring))     ; then-part
       (concat (car kill-ring) string))      ; else-part

If the region being killed is before the region that was killed in the last command, then it should be prepended before the material that was saved in the previous kill; and conversely, if the killed text follows what was just killed, it should be appended after the previous text. The if expression depends on the predicate before-p to decide whether the newly saved text should be put before or after the previously saved text.

The symbol before-p is the name of one of the arguments to kill-append. When the kill-append function is evaluated, it is bound to the value returned by evaluating the actual argument. In this case, this is the expression (< end beg). This expression does not directly determine whether the killed text in this command is located before or after the kill text of the last command; what is does is determine whether the value of the variable end is less than the value of the variable beg. If it is, it means that the user is most likely heading towards the beginning of the buffer. Also, the result of evaluating the predicate expression, (< end beg), will be true and the text will be prepended before the previous text. On the other hand, if the value of the variable end is greater than the value of the variable beg, the text will be appended after the previous text.

When the newly saved text will be prepended, then the string with the new text will be concatenated before the old test:

     (concat string (car kill-ring))

But if the text will be appended, it will be concatenated after the old text:

     (concat (car kill-ring) string))

To understand how this works, we first need to review the concat function. The concat function links together or unites two strings of text. The result is a string. For example:

     (concat "abc" "def")
          ⇒ "abcdef"
     (concat "new "
             (car '("first element" "second element")))
          ⇒ "new first element"
     (concat (car
             '("first element" "second element")) " modified")
          ⇒ "first element modified"

We can now make sense of kill-append: it modifies the contents of the kill ring. The kill ring is a list, each element of which is saved text. The setcar function actually changes the first element of this list. It does this by using concat to replace the old first element of the kill ring (the car of the kill ring) with a new first element made by concatenating together the old saved text and the newly saved text. The newly saved text is put in front of the old text or after the old text, depending on whether it is in front of or after the old text in the buffer from which it is cut. The whole concatenation becomes the new first element of the kill ring.

Incidentally, this is what the beginning of my current kill ring looks like:

     ("concatenating together" "saved text" "element" ...

The else-part of copy-region-as-kill

Now, back to the explanation of copy-region-as-kill:

If the last command is not kill-region, then instead of calling kill-append, it calls the else-part of the following code:

(if true-or-false-test

what-is-done-if-test-returns-true ;; else-part

       (setq kill-ring
             (cons (buffer-substring beg end) kill-ring))
       (if (> (length kill-ring) kill-ring-max)
           (setcdr (nthcdr (1- kill-ring-max) kill-ring) nil)))

The else-part of copy-region-as-kill  77

The setq line of the else-part sets the new value of the kill ring to what results from adding the string being killed to the old kill ring.

We can see how this works with a little example:

     (setq example-list '("here is a clause" "another clause"))

After evaluating this expression with C-x C-e, you can evaluate example-list and see what it returns:

     example-list
          ⇒ ("here is a clause" "another clause")

Now, we can add a new element on to this list by evaluating the following expression:

     (setq example-list (cons "a third clause" example-list))

When we evaluate example-list, we find its value is:

     example-list
          ⇒ ("a third clause" "here is a clause" "another clause")

Thus, the third clause was added to the list by cons.

This is exactly similar to what the setq and cons do in the function, except that buffersubstring is used to pull out a copy of a region of text and hand it to the cons. Here is the line again:

     (setq kill-ring (cons (buffer-substring beg end) kill-ring))

The next segment of the else-part of copy-region-as-kill is another if clause. This if clause keeps the kill ring from growing too long. It reads as follows:

     (if (> (length kill-ring) kill-ring-max)
         (setcdr (nthcdr (1- kill-ring-max) kill-ring) nil)))

This code checks whether the length of the kill ring is greater than the maximum permitted length. This is the value of kill-ring-max (which is 30, by default). If the length of the kill ring is too long, then this code sets the last element of the kill ring to nil. It does this by using two functions, nthcdr and setcdr.

We looked at setcdr earlier (see Section 7.5, “setcdr”, page 64). It sets the cdr of a list, just as setcar sets the car of a list. In this case, however, setcdr will not be setting the cdr of the whole kill ring; the nthcdr function is used to cause it to set the cdr of the next to last element of the kill ring—this means that since the cdr of the next to last element is the last element of the kill ring, it will set the last element of the kill ring.

The nthcdr function works by repeatedly taking the cdr of a list—it takes the cdr of the cdr of the cdr … It does this N times and returns the results.

Thus, if we had a four element list that was supposed to be three elements long, we could set the cdr of the next to last element to nil, and thereby shorten the list.

You can see this by evaluating the following three expressions in turn. First set the value of trees to (maple oak pine birch), then set the cdr of its second cdr to nil and then find the value of trees:

     (setq trees '(maple oak pine birch))
          ⇒ (maple oak pine birch)

     (setcdr (nthcdr 2 trees) nil)
          ⇒ nil
     trees
          ⇒ (maple oak pine)

(The value returned by the setcdr expression is nil since that is what the cdr is set to.)

To repeat, in copy-region-as-kill, the nthcdr function takes the cdr a number of times that is one less than the maximum permitted size of the kill ring and sets the cdr of that element (which will be the rest of the elements in the kill ring) to nil. This prevents the kill ring from growing too long.

The next to last line of the copy-region-as-kill function is

     (setq this-command 'kill-region)

This line is not part of either the inner or the outer if expression, so it is evaluated every time copy-region-as-kill is called. Here we find the place where this-command is set to killregion. As we saw earlier, when the next command is given, the variable last-command will be given this value.

Finally, the last line of the copy-region-as-kill function is:

     (setq kill-ring-yank-pointer kill-ring)

The kill-ring-yank-pointer is a global variable that is set to be the kill-ring.

Even though the kill-ring-yank-pointer is called a ‘pointer’, it is a variable just like the kill ring. However, the name has been chosen to help humans understand how the variable is used. The variable is used in functions such as yank and yank-pop (see Chapter 10, “Yanking Text Back”, page 85).

This leads us to the code for bringing back text that has been cut out of the buffer—the yank commands. However, before discussing the yank commands, it is better to learn how lists are implemented in a computer. This will make clear such mysteries as the use of the term ‘pointer’.

8.6 Review

Here is a brief summary of some recently introduced functions.

car

cdr car returns the first element of a list; cdr returns the second and subsequent elements of a list. For example:

               (car '(1 2 3 4 5 6 7))
                    ⇒ 1
               (cdr '(1 2 3 4 5 6 7))
                    ⇒ (2 3 4 5 6 7)

cons cons constructs a list by prepending its first argument to its second argument. For example:

               (cons 1 '(2 3 4)) ⇒ (1 2 3 4)

nthcdr Return the result of taking cdr ‘n’ times on a list. The ‘rest of the rest’ as it were.

For example:

     (nthcdr 3 '(1 2 3 4 5 6 7)) ⇒ (4 5 6 7)

setcar

setcdr setcar changes the first element of a list; setcdr changes the second and subsequent elements of a list. For example:

     (setq triple '(1 2 3))
     (setcar triple '37)
     triple

          ⇒ (37 2 3)

     (setcdr triple '("foo" "bar"))
     triple

          ⇒ (37 "foo" "bar")

progn Evaluate each argument in sequence and then return the value of the last. For example:

     (progn 1 2 3 4) ⇒ 4

save-restriction

Record whatever narrowing is in effect in the current buffer, if any, and restore that narrowing after evaluating the arguments.

search-forward

Search for a string, and if the string is found, move point. Takes four arguments:

1. The string to search for.

2. Optionally, the limit of the search.

3. Optionally, what to do if the search fails, return nil or an error message.

4. Optionally, how many times to repeat the search; if negative, the search goes backwards.

kill-region
delete-region
copy-region-as-kill

kill-region cuts the text between point and mark from the buffer and stores that text in the kill ring, so you can get it back by yanking. delete-region removes the text between point and mark from the buffer and throws it away. You cannot get it back. copy-region-as-kill copies the text between point and mark into the kill ring, from which you can get it by yanking. The function does not cut or remove the text from the buffer.

8.7 Searching Exercises

• Write an interactive function that searches for a string. If the search finds the string, leave point after it and display a message that says “Found!”. (Do not use search-forward for the name of this function; if you do, you will overwrite the existing version of searchforward that comes with Emacs. Use a name such as test-search instead.)

• Write a function that prints the third element of the kill ring in the echo area, if any; if the kill ring does not contain a third element, print an appropriate message.

• As of version 19.29, copy-region-as-kill no longer sets this-command. What are the consequences of this change? What do you suppose motivated it?

9 How Lists are Implemented

In Lisp, atoms are recorded in a straightforward fashion; if the implementation is not straightforward in practice, it is, nonetheless, straightforward in theory. The atom ‘rose’, for example, is recorded as the four contiguous letters ‘r’, ‘o’, ‘s’, ‘e’. A list, on the other hand, is kept differently. The mechanism is equally simple, but it takes a moment to get used to the idea. A list is kept using a series of pairs of pointers. In the series, the first pointer in each pair points to an atom or to another list, and the second pointer in each pair points to the next pair, or to the symbol nil, which marks the end of the list.

A pointer itself is quite simply the electronic address of what is pointed to. Hence, a list is kept as a series of electronic addresses.

For example, the list (rose violet buttercup) has three elements, ‘rose’, ‘violet’, and ‘buttercup’. In the computer, the electronic address of ‘rose’ is recorded in a segment of computer memory along with the address that gives the electronic address of where the atom ‘violet’ is located; and that address (the one that tells where ‘violet’ is located) is kept along with an address that tells where the address for the atom ‘buttercup’ is located.

This sounds more complicated than it is and is easier seen in a diagram:

         __________________
        |___|___|--> |___|___|--> |___|___|--> nil
          |            |            |
          |            |            |
           --> rose     --> violet   --> buttercup

In the diagram, each box represents a word of computer memory that holds a Lisp object, usually in the form of a memory address. The boxes, i.e. the addresses, are in pairs. Each arrow points to what the address is the address of, either an atom or another pair of addresses. The first box is the electronic address of ‘rose’ and the arrow points to ‘rose’; the second box is the address of the next pair of boxes, the first part of which is the address of ‘violet’ and the second part of which is the address of the next pair. The very last box points to the symbol nil, which marks the end of the list.

When a variable is set to a list with a function such as setq, it stores the address of the first box in the variable. Thus, evaluation of the expression

     (setq bouquet '(rose violet buttercup))

creates a situation like this:

     bouquet
          |
          |     __________________
           --> |___|___|--> |___|___|--> |___|___|--> nil
                 |            |            |
                 |            |            |
                  --> rose     --> violet   --> buttercup

In this case, the symbol bouquet holds the address of the first pair of boxes. Indeed, the symbol bouquet consists of a group of address-boxes, one of which is the address of the printed word ‘bouquet’, a second of which is the address of a function definition attached to the symbol, if any, a third of which is the address of the first pair of address-boxes for the list (rose violet

buttercup), and so on.

This same list can be illustrated in a different sort of box notation like this:

bouquet
|
|    --------------       ---------------       ----------------
|   | car   | cdr  |     | car    | cdr  |     | car     | cdr  |
-->| rose  |   o------->| violet |   o------->| butter- |  nil |
|       |      |     |        |      |     | cup     |      |
--------------       ---------------       ----------------

In an earlier section, I suggested that you might imagine a symbol as being a chest of drawers. The function definition is put in one drawer, the value in another, and so on. What is put in the drawer holding the value can be changed without affecting the contents of the drawer holding the function definition, and vice-versa. Actually, what is put in each drawer is the address of the value or function definition. It is as if you found an old chest in the attic, and in one of its drawers you found a map giving you directions to where the buried treasure lies.

(In addition to its name, symbol definition, and variable value, a symbol has a ‘drawer’ for a property list which can be used to record other information. Property lists are not discussed here; see section “Property Lists” in The GNU Emacs Lisp Reference Manual.)

Here is a fanciful representation:

   Chest of Drawers             Contents of Drawers
 ---------------------
|                     |
|    symbol name      |             bouquet
|                     |
 ---------------------
|                     |
|  symbol definition  |             [none]
|                     |
 ---------------------
|                     |
|   variable value    |             (rose violet buttercup)
|                     |
 ---------------------
|                     |
|   property list     |             [not described here]
|                     |
 ---------------------
|/                   \|

If symbol is set to the cdr of a list, the list itself is not changed; the symbol simply has an address further down the list. (In the jargon, car and cdr are ‘non-destructive’.) Thus, evaluation of the following expression

(setq flowers (cdr bouquet))

produces this:

     bouquet        flowers
       |              |
       |     ______|     ____________
        --> |   |   |  --> |   |   |    |   |   |
            |___|___|----> |___|___|--> |___|___|--> nil
              |              |            |
              |              |            |
               --> rose       --> violet   --> buttercup

The value of flowers is (violet buttercup), which is to say, the symbol flowers holds the address of the pair of address-boxes, the first of which holds the address of violet, and the second of which holds the address of buttercup.

A pair of address-boxes is called a cons cell or dotted pair. See section “List Type ” in The GNU Emacs Lisp Reference Manual, and section “Dotted Pair Notation” in The GNU Emacs Lisp Reference Manual, for more information about cons cells and dotted pairs.

The function cons adds a new pair of addresses to the front of a series of addresses like that shown above. For example, evaluating the expression

     (setq bouquet (cons 'lilly bouquet))

produces:

     bouquet                       flowers
       |                             |
       |     ____________|     ____________
        --> |   |   |      |   |   |  --> |   |   |     |   |   |
            |___|___|----> |___|___|----> |___|___|---->|___|___|--> nil
              |              |              |             |
              |              |              |             |
               --> lilly      --> rose       --> violet    --> buttercup

However, this does not change the value of the symbol flowers, as you can see by evaluating the following,

     (eq (cdr (cdr bouquet)) flowers)

which returns t for true.

Until it is reset, flowers still has the value (violet buttercup); that is, it has the address of the cons cell whose first address is of violet. Also, this does not alter any of the pre-existing cons cells; they are all still there.

Thus, in Lisp, to get the cdr of a list, you just get the address of the next cons cell in the series; to get the car of a list, you get the address of the first element of the list; to cons a new element on a list, you add a new cons cell to the front of the list. That is all there is to it! The underlying structure of Lisp is brilliantly simple!

And what does the last address in a series of cons cells refer to? It is the address of the empty list, of nil.

In summary, when a Lisp variable is set to a value, it is provided with the address of the list to which the variable refers.

9.1 Exercise

Set flowers to violet and buttercup. Cons two more flowers on to this list and set this new list to more-flowers. Set the car of flowers to a fish. What does the more-flowers list now contain?

The kill-ring-yank-pointer Variable  85

10 Yanking Text Back

Whenever you cut text out of a buffer with a ‘kill’ command in GNU Emacs, you can bring it back with a ‘yank’ command. The text that is cut out of the buffer is put in the kill ring and the yank commands insert the appropriate contents of the kill ring back into a buffer (not necessarily the original buffer).

A simple C-y (yank) command inserts the first item from the kill ring into the current buffer. If the C-y command is followed immediately by M-y, the first element is replaced by the second element. Successive M-y commands replace the second element with the third, fourth, or fifth element, and so on. When the last element in the kill ring is reached, it is replaced by the first element and the cycle is repeated. (Thus the kill ring is called a ‘ring’ rather than just a ‘list’. However, the actual data structure that holds the text is a list. See Appendix B, “Handling the Kill Ring”, page 171, for the details of how the list is handled as a ring.)

10.1 Kill Ring Overview

The kill ring is a list of textual strings. This is what it looks like:

     ("some text" "a different piece of text" "yet more text")

If this were the contents of my kill ring and I pressed C-y, the string of characters saying ‘some text’ would be inserted in this buffer where my cursor is located.

The yank command is also used for duplicating text by copying it. The copied text is not cut from the buffer, but a copy of it is put on the kill ring and is inserted by yanking it back.

Three functions are used for bringing text back from the kill ring: yank, which is usually bound to C-y; yank-pop, which is usually bound to M-y; and rotate-yank-pointer, which is used by the two other functions.

These functions refer to the kill ring through a variable called the kill-ring-yank-pointer. Indeed, the insertion code for both the yank and yank-pop functions is:

     (insert (car kill-ring-yank-pointer))

To begin to understand how yank and yank-pop work, it is first necessary to look at the

kill-ring-yank-pointer variable and the rotate-yank-pointer function.

10.2 The kill-ring-yank-pointer Variable

kill-ring-yank-pointer is a variable, just as kill-ring is a variable. It points to something by being bound to the value of what it points to, like any other Lisp variable.

Thus, if the value of the kill ring is:

     ("some text" "a different piece of text" "yet more text")

and the kill-ring-yank-pointer points to the second clause, the value of kill-ring-yank-

pointer is:
     ("a different piece of text" "yet more text")

As explained in the previous chapter (see Chapter 9, “List Implementation”, page 81), the computer does not keep two different copies of the text being pointed to by both the kill-ring and the kill-ring-yank-pointer. The words “a different piece of text” and “yet more text” are not duplicated. Instead, the two Lisp variables point to the same pieces of text. Here is a diagram:

     kill-ring     kill-ring-yank-pointer
         |               |
         |      ______|     ____________
          ---> |   |   |  --> |   |   |    |   |   |
               |___|___|----> |___|___|--> |___|___|--> nil
                 |              |            |
                 |              |            |
                 |              |             --> "yet more text"
                 |              |
                 |               --> "a different piece of text
                 |
                  --> "some text"

Both the variable kill-ring and the variable kill-ring-yank-pointer are pointers. But the kill ring itself is usually described as if it were actually what it is composed of. The killring is spoken of as if it were the list rather than that it points to the list. Conversely, the kill-ring-yank-pointer is spoken of as pointing to a list.

These two ways of talking about the same thing sound confusing at first but make sense on reflection. The kill ring is generally thought of as the complete structure of data that holds the information of what has recently been cut out of the Emacs buffers. The kill-ring-yank-pointer on the other hand, serves to indicate—that is, to ‘point to’—that part of the kill ring of which the first element (the car) will be inserted.

The rotate-yank-pointer function changes the element in the kill ring to which the killring-yank-pointer points; when the pointer is set to point to the next element beyond the end of the kill ring, it automatically sets it to point to the first element of the kill ring. This is how the list is transformed into a ring. The rotate-yank-pointer function itself is not difficult, but contains many details. It and the much simpler yank and yank-pop functions are described in an appendix. See Appendix B, “Handling the Kill Ring”, page 171.

10.3 Exercises with yank and nthcdr

• Using C-h v (describe-variable), look at the value of your kill ring. Add several items to your kill ring; look at its value again. Using M-y (yank-pop), move all the way around the kill ring. How many items were in your kill ring? Find the value of kill-ring-max. Was your kill ring full, or could you have kept more blocks of text within it?

• Using nthcdr and car, construct a series of expressions to return the first, second, third, and fourth elements of a list.

11 Loops and Recursion

Emacs Lisp has two primary ways to cause an expression, or a series of expressions, to be evaluated repeatedly: one uses a while loop, and the other uses recursion.

Repetition can be very valuable. For example, to move forward four sentences, you need only write a program that will move forward one sentence and then repeat the process four times. Since a computer does not get bored or tired, such repetitive action does not have the deleterious effects that excessive or the wrong kinds of repetition can have on humans.

11.1 while

The while special form tests whether the value returned by evaluating its first argument is true or false. This is similar to what the Lisp interpreter does with an if; what the interpreter does next, however, is different.

In a while expression, if the value returned by evaluating the first argument is false, the Lisp interpreter skips the rest of the expression (the body of the expression) and does not evaluate it. However, if the value is true, the Lisp interpreter evaluates the body of the expression and then again tests whether the first argument to while is true or false. If the value returned by evaluating the first argument is again true, the Lisp interpreter again evaluates the body of the expression.

The template for a while expression looks like this:

(while true-or-false-test

       body...)

So long as the true-or-false-test of the while expression returns a true value when it is evaluated, the body is repeatedly evaluated. This process is called a loop since the Lisp interpreter repeats the same thing again and again, like an airplane doing a loop. When the result of evaluating the true-or-false-test is false, the Lisp interpreter does not evaluate the rest of the while expression and ‘exits the loop’.

Clearly, if the value returned by evaluating the first argument to while is always true, the body following will be evaluated again and again … and again … forever. Conversely, if the value returned is never true, the expressions in the body will never be evaluated. The craft of writing a while loop consists of choosing a mechanism such that the true-or-false-test returns true just the number of times that you want the subsequent expressions to be evaluated, and then have the test return false.

The value returned by evaluating a while is the value of the true-or-false-test. An interesting consequence of this is that a while loop that evaluates without error will return nil or false regardless of whether it has looped 1 or 100 times or none at all. A while expression that evaluates successfully never returns a true value! What this means is that while is always evaluated for its side effects, which is to say, the consequences of evaluating the expressions within the body of the while loop. This makes sense. It is not the mere act of looping that is desired, but the consequences of what happens when the expressions in the loop are repeatedly evaluated.

11.1.1 A while Loop and a List

A common way to control a while loop is to test whether a list has any elements. If it does, the loop is repeated; but if it does not, the repetition is ended. Since this is an important technique, we will create a short example to illustrate it.

A simple way to test whether a list has elements is to evaluate the list: if it has no elements, it is an empty list and will return the empty list, (), which is a synonym for nil or false. On the other hand, a list with elements will return those elements when it is evaluated. Since Lisp considers as true any value that is not nil, a list that returns elements will test true in a while loop.

For example, you can set the variable empty-list to nil by evaluating the following setq expression:

     (setq empty-list ())

After evaluating the setq expression, you can evaluate the variable empty-list in the usual way, by placing the cursor after the symbol and typing C-x C-e; nil will appear in your echo area:

     empty-list

On the other hand, if you set a variable to be a list with elements, the list will appear when you evaluate the variable, as you can see by evaluating the following two expressions:

     (setq animals '(giraffe gazelle lion tiger))
     animals

Thus, to create a while loop that tests whether there are any items in the list animals, the first part of the loop will be written like this:

     (while animals
            ...

When the while tests its first argument, the variable animals is evaluated. It returns a list. So long as the list has elements, the while considers the results of the test to be true; but when the list is empty, it considers the results of the test to be false.

To prevent the while loop from running forever, some mechanism needs to be provided to empty the list eventually. An oft-used technique is to have one of the subsequent forms in the while expression set the value of the list to be the cdr of the list. Each time the cdr function is evaluated, the list will be made shorter, until eventually only the empty list will be left. At this point, the test of the while loop will return false, and the arguments to the while will no longer be evaluated.

For example, the list of animals bound to the variable animals can be set to be the cdr of the original list with the following expression:

     (setq animals (cdr animals))

If you have evaluated the previous expressions and then evaluate this expression, you will see (gazelle lion tiger) appear in the echo area. If you evaluate the expression again, (lion tiger) will appear in the echo area. If you evaluate it again and yet again, (tiger) appears and then the empty list, shown by nil.

A template for a while loop that uses the cdr function repeatedly to cause the true-or-false-test eventually to test false looks like this:

(while test-whether-list-is-empty body... set-list-to-cdr-of-list)

This test and use of cdr can be put together in a function that goes through a list and prints each element of the list on a line of its own.

11.1.2 An Example: print-elements-of-list

The print-elements-of-list function illustrates a while loop with a list.

The function requires several lines for its output. Since the echo area is only one line, we cannot illustrate how it works in the same way we have been illustrating functions in the past, by evaluating them inside Info. Instead, you need to copy the necessary expressions to your ‘*scratch*’ buffer and evaluate them there. You can copy the expressions by marking the beginning of the region with C-^{⟨}\underline{\overline{SPC}}^{⟩} (set-mark-command), moving the cursor to the end of the region and then copying the region using M-w (copy-region-as-kill). In the ‘*scratch*’ buffer, you can yank the expressions back by typing C-y (yank).

After you have copied the expressions to the ‘*scratch*’ buffer, evaluate each expression in turn. Be sure to evaluate the last expression, (print-elements-of-list animals), by typing C-u C-x C-e, that is, by giving an argument to eval-last-sexp. This will cause the result of the evaluation to be printed in the ‘*scratch*’ buffer instead of being printed in the echo area. (Otherwise you will see something like this in your echo area:

^Jgiraffe^J^Jgazelle^J^Jlion^J^Jtiger^Jnil, in which each ‘^J’ stands for the newline

that in the ‘*scratch*’ buffer puts each word on its own line. You can evaluate these expressions right now in the Info buffer, if you like, to see this effect.)

     (setq animals '(giraffe gazelle lion tiger))
     (defun print-elements-of-list (list)
       "Print each element of LIST on a line of its own."
       (while list
         (print (car list))
         (setq list (cdr list))))
     (print-elements-of-list animals)

When you evaluate the three expressions in sequence in the ‘*scratch*’ buffer, this will be printed in the buffer:

     giraffe
     gazelle
     lion
     tiger
     nil

Each element of the list is printed on a line of its own (that is what the function print does) and then the value returned by the function is printed. Since the last expression in the function is the while loop, and since while loops always return nil, a nil is printed after the last element of the list.

11.1.3 A Loop with an Incrementing Counter

A loop is not useful unless it stops when it ought. Besides controlling a loop with a list, a common way of stopping a loop is to write the first argument as a test that returns false when the correct number of repetitions are complete. This means that the loop must have a counter—an expression that counts how many times the loop repeats itself.

The test can be an expression such as (< count desired-number) which returns t for true if the value of count is less than the desired-number of repetitions and nil for false if the value of count is equal to or is greater than the desired-number. The expression that increments the count can be a simple setq such as (setq count (1+ count)), where 1+ is a built-in function in Emacs Lisp that adds 1 to its argument. (The expression (1+ count) has the same result as (+ count 1), but is easier for a human to read.)

The template for a while loop controlled by an incrementing counter looks like this:

set-count-to-initial-value

     (while (< count desired-number)         ; true-or-false-test

body...

       (setq count (1+ count)))              ; incrementer

Note that you need to set the initial value of count; usually it is set to 1.

Example with incrementing counter

Suppose you are playing on the beach and decide to make a triangle of pebbles, putting one pebble in the first row, two in the second row, three in the third row and so on, like this:

• • • • • • • • • •

(About 2500 years ago, Pythagoras and others developed the beginnings of number theory by considering questions such as this.)

Suppose you want to know how many pebbles you will need to make a triangle with 7 rows?

Clearly, what you need to do is add up the numbers from 1 to 7. There are two ways to do this; start with the smallest number, one, and add up the list in sequence, 1, 2, 3, 4 and so on; or start with the largest number and add the list going down: 7, 6, 5, 4 and so on. Because both mechanisms illustrate common ways of writing while loops, we will create two examples, one counting up and the other counting down. In this first example, we will start with 1 and add 2, 3, 4 and so on.

If you are just adding up a short list of numbers, the easiest way to do it is to add up all the numbers at once. However, if you do not know ahead of time how many numbers your list will have, or if you want to be prepared for a very long list, then you need to design your addition so that what you do is repeat a simple process many times instead of doing a more complex process once.

For example, instead of adding up all the pebbles all at once, what you can do is add the number of pebbles in the first row, 1, to the number in the second row, 2, and then add the total of those two rows to the third row, 3. Then you can add the number in the fourth row, 4, to the total of the first three rows; and so on.

The critical characteristic of the process is that each repetitive action is simple. In this case, at each step we add only two numbers, the number of pebbles in the row and the total already found. This process of adding two numbers is repeated again and again until the last row has been added to the total of all the preceding rows. In a more complex loop the repetitive action might not be so simple, but it will be simpler than doing everything all at once.

The parts of the function definition

The preceding analysis gives us the bones of our function definition: first, we will need a variable that we can call total that will be the total number of pebbles. This will be the value returned by the function.

Second, we know that the function will require an argument: this argument will be the total number of rows in the triangle. It can be called number-of-rows.

Finally, we need a variable to use as a counter. We could call this variable counter, but a better name is row-number. That is because what the counter does is count rows, and a program should be written to be as understandable as possible.

When the Lisp interpreter first starts evaluating the expressions in the function, the value of total should be set to zero, since we have not added anything to it. Then the function should add the number of pebbles in the first row to the total, and then add the number of pebbles in the second to the total, and then add the number of pebbles in the third row to the total, and so on, until there are no more rows left to add.

Both total and row-number are used only inside the function, so they can be declared as local variables with let and given initial values. Clearly, the initial value for total should be 0. The initial value of row-number should be 1, since we start with the first row. This means that the let statement will look like this:

       (let ((total 0)
             (row-number 1))
         body...)

After the internal variables are declared and bound to their initial values, we can begin the while loop. The expression that serves as the test should return a value of t for true so long as the row-number is less than or equal to the number-of-rows. (If the expression tests true only so long as the row number is less than the number of rows in the triangle, the last row will never be added to the total; hence the row number has to be either less than or equal to the number of rows.)

Lisp provides the <= function that returns true if the value of its first argument is less than or equal to the value of its second argument and false otherwise. So the expression that the while will evaluate as its test should look like this:

     (<= row-number number-of-rows)

The total number of pebbles can be found by repeatedly adding the number of pebbles in a row to the total already found. Since the number of pebbles in the row is equal to the row number, the total can be found by adding the row number to the total. (Clearly, in a more complex situation, the number of pebbles in the row might be related to the row number in a more complicated way; if this were the case, the row number would be replaced by the appropriate expression.)

     (setq total (+ total row-number))

What this does is set the new value of total to be equal to the sum of adding the number of pebbles in the row to the previous total.

After setting the value of total, the conditions need to be established for the next repetition of the loop, if there is one. This is done by incrementing the value of the row-number variable, which serves as a counter. After the row-number variable has been incremented, the true-or-false-test at the beginning of the while loop tests whether its value is still less than or equal to the value of the number-of-rows and if it is, adds the new value of the row-number variable to the total of the previous repetition of the loop.

The built-in Emacs Lisp function 1+ adds 1 to a number, so the row-number variable can be incremented with this expression:

     (setq row-number (1+ row-number))

Putting the function definition together

We have created the parts for the function definition; now we need to put them together.

First, the contents of the while expression:

     (while (<= row-number number-of-rows)   ; true-or-false-test
       (setq total (+ total row-number))
       (setq row-number (1+ row-number)))    ; incrementer

Along with the let expression varlist, this very nearly completes the body of the function definition. However, it requires one final element, the need for which is somewhat subtle.

The final touch is to place the variable total on a line by itself after the while expression. Otherwise, the value returned by the whole function is the value of the last expression that is evaluated in the body of the let, and this is the value returned by the while, which is always

nil.

This may not be evident at first sight. It almost looks as if the incrementing expression is the last expression of the whole function. But that expression is part of the body of the while; it is the last element of the list that starts with the symbol while. Moreover, the whole of the while loop is a list within the body of the let.

In outline, the function will look like this:

(defun name-of-function (argument-list)

"documentation..."

(let (varlist)

(while (true-or-false-test) body-of-while... ) ... ) ; Need final expression here.

The result of evaluating the let is what is going to be returned by the defun since the let is not embedded within any containing list, except for the defun as a whole. However, if the while is the last element of the let expression, the function will always return nil. This is not what we want! Instead, what we want is the value of the variable total. This is returned by simply placing the symbol as the last element of the list starting with let. It gets evaluated after the preceding elements of the list are evaluated, which means it gets evaluated after it has been assigned the correct value for the total.

It may be easier to see this by printing the list starting with let all on one line. This format makes it evident that the varlist and while expressions are the second and third elements of the list starting with let, and the total is the last element:

(let (varlist) (while (true-or-false-test) body-of-while... ) total)

Putting everything together, the triangle function definition looks like this:

     (defun triangle (number-of-rows)    ; Version with

; incrementing counter.

       "Add up the number of pebbles in a triangle.
     The first row has one pebble, the second row two pebbles,
     the third row three pebbles, and so on.
     The argument is NUMBER-OF-ROWS."

       (let ((total 0)
             (row-number 1))
         (while (<= row-number number-of-rows)
           (setq total (+ total row-number))
           (setq row-number (1+ row-number)))
         total))

After you have installed triangle by evaluating the function, you can try it out. Here are two examples:

     (triangle 4)
     (triangle 7)

The sum of the first four numbers is 10 and the sum of the first seven numbers is 28.

11.1.4 Loop with a Decrementing Counter

Another common way to write a while loop is to write the test so that it determines whether a counter is greater than zero. So long as the counter is greater than zero, the loop is repeated. But when the counter is equal to or less than zero, the loop is stopped. For this to work, the counter has to start out greater than zero and then be made smaller and smaller by one of the forms that is evaluated repeatedly.

The test will be an expression such as (> counter 0) which returns t for true if the value of counter is greater than zero, and nil for false if the value of counter is equal to or less than zero. The expression that makes the number smaller and smaller can be a simple setq such as (setq counter (1- counter)), where 1- is a built-in function in Emacs Lisp that subtracts 1 from its argument.

The template for a decrementing while loop looks like this:

     (while (> counter 0)                    ; true-or-false-test

body...

       (setq counter (1- counter)))          ; decrementer

Example with decrementing counter

To illustrate a loop with a decrementing counter, we will rewrite the triangle function so the counter decreases to zero.

This is the reverse of the earlier version of the function. In this case, to find out how many pebbles are needed to make a triangle with 3 rows, add the number of pebbles in the third row, 3, to the number in the preceding row, 2, and then add the total of those two rows to the row that precedes them, which is 1.

Likewise, to find the number of pebbles in a triangle with 7 rows, add the number of pebbles in the seventh row, 7, to the number in the preceding row, which is 6, and then add the total of those two rows to the row that precedes them, which is 5, and so on. As in the previous example, each addition only involves adding two numbers, the total of the rows already added up and the number of pebbles in the row that is being added to the total. This process of adding two numbers is repeated again and again until there are no more pebbles to add.

We know how many pebbles to start with: the number of pebbles in the last row is equal to the number of rows. If the triangle has seven rows, the number of pebbles in the last row is 7. Likewise, we know how many pebbles are in the preceding row: it is one less than the number in the row.

The parts of the function definition

We start with three variables: the total number of rows in the triangle; the number of pebbles in a row; and the total number of pebbles, which is what we want to calculate. These variables

can be named number-of-rows, number-of-pebbles-in-row, and total, respectively.

Both total and number-of-pebbles-in-row are used only inside the function and are declared with let. The initial value of total should, of course, be zero. However, the initial value of number-of-pebbles-in-row should be equal to the number of rows in the triangle, since the addition will start with the longest row.

This means that the beginning of the let expression will look like this:

     (let ((total 0)
           (number-of-pebbles-in-row number-of-rows))
       body...)

The total number of pebbles can be found by repeatedly adding the number of pebbles in a row to the total already found, that is, by repeatedly evaluating the following expression:

     (setq total (+ total number-of-pebbles-in-row))
After the number-of-pebbles-in-row is added to the total, the number-of-pebbles-in-row

should be decremented by one, since the next time the loop repeats, the preceding row will be added to the total.

The number of pebbles in a preceding row is one less than the number of pebbles in a row, so the built-in Emacs Lisp function 1- can be used to compute the number of pebbles in the preceding row. This can be done with the following expression:

     (setq number-of-pebbles-in-row
           (1- number-of-pebbles-in-row))

Finally, we know that the while loop should stop making repeated additions when there are no pebbles in a row. So the test for the while loop is simply:

     (while (> number-of-pebbles-in-row 0)

Putting the function definition together

Putting these expressions together, we have a function definition that looks like this:

;;; First subtractive version.

     (defun triangle (number-of-rows)
       "Add up the number of pebbles in a triangle."
       (let ((total 0)
             (number-of-pebbles-in-row number-of-rows))
         (while (> number-of-pebbles-in-row 0)
           (setq total (+ total number-of-pebbles-in-row))
           (setq number-of-pebbles-in-row
                 (1- number-of-pebbles-in-row)))
         total))

As written, this function works.

However, it turns out that one of the local variables, number-of-pebbles-in-row, is unneeded!

When the triangle function is evaluated, the symbol number-of-rows will be bound to a number, giving it an initial value. That number can be changed in the body of the function as if it were a local variable, without any fear that such a change will effect the value of the variable outside of the function. This is a very useful characteristic of Lisp; it means that the variable

number-of-rows can be used anywhere in the function where number-of-pebbles-in-row is

used.

Here is a second version of the function written a bit more cleanly:

     (defun triangle (number)                ; Second version.
       "Return sum of numbers 1 through NUMBER inclusive."
       (let ((total 0))
         (while (> number 0)
           (setq total (+ total number))
           (setq number (1- number)))
         total))

In brief, a properly written while loop will consist of three parts:

1. A test that will return false after the loop has repeated itself the correct number of times.

2. An expression the evaluation of which will return the value desired after being repeatedly evaluated.

3. An expression to change the value passed to the true-or-false-test so that the test returns false after the loop has repeated itself the right number of times.

11.2 Recursion

A recursive function contains code that tells itself to evaluate itself. When the function evaluates itself, it again finds the code that tells itself to evaluate itself, so the function evaluates itself again … and again … A recursive function will keep telling itself to evaluate itself again forever unless it is also provided with a stop condition.

A recursive function typically contains a conditional expression which has three parts:

1. A true-or-false-test that determines whether the function is called again, here called the do-again-test.

2. The name of the function.

3. An expression that causes the conditional expression to test false after the correct number of repetitions, here called the next-step-expression.

Recursive functions can be much simpler than any other kind of function. Indeed, when people first start to use them, they often look so mysteriously simple as to be incomprehensible. Like riding a bicycle, reading a recursive function definition takes a certain knack which is hard at first but then seems simple.

A template for a recursive function looks like this:

(defun name-of-recursive-function (argument-list)

"documentation..."

body...

(if do-again-test

(name-of-recursive-function next-step-expression)))

Each time the recursive function is evaluated, an argument is bound to the value of the nextstep-expression; and that value is used in the do-again-test. The next-step-expression is designed so that the do-again-test returns false when the function should no longer be repeated.

The do-again-test is sometimes called the stop condition, since it stops the repetitions when it tests false.

11.2.1 Recursion with a List

The example of a while loop that printed the elements of a list of numbers can be written recursively. Here is the code, including an expression to set the value of the variable animals to a list.

This example must be copied to the ‘*scratch*’ buffer and each expression must be evaluated

there. Use C-u C-x C-e to evaluate the (print-elements-recursively animals) expression

so that the results are printed in the buffer; otherwise the Lisp interpreter will try to squeeze the results into the one line of the echo area.

Also, place your cursor immediately after the last closing parenthesis of the print-elements-recursively function, before the comment. Otherwise, the Lisp interpreter will try to evaluate the comment.

     (setq animals '(giraffe gazelle lion tiger))
     (defun print-elements-recursively (list)
       "Print each element of LIST on a line of its own.
     Uses recursion."
       (print (car list))                  ; body

(if list ; do-again-test

           (print-elements-recursively     ; recursive call

(cdr list)))) ; next-step-expression

     (print-elements-recursively animals)

The print-elements-recursively function first prints the first element of the list, the car of the list. Then, if the list is not empty, the function invokes itself, but gives itself as its argument, not the whole list, but the second and subsequent elements of the list, the cdr of the list.

When this evaluation occurs, the function prints the first element of the list it receives as its argument (which is the second element of the original list). Then, the if expression is evaluated and when true, the function calls itself with the cdr of the list it is invoked with, which (the second time around) is the cdr of the cdr of the original list.

Each time the function invokes itself, it invokes itself on a shorter version of the original list. Eventually, the function invokes itself on an empty list. The print function prints the empty list as nil. Next, the conditional expression tests the value of list. Since the value of list is nil, the if expression tests false so the then-part is not evaluated. The function as a whole then returns nil. Consequently, you see nil twice when you evaluate the function.

   When you evaluate (print-elements-recursively animals) in the ‘*scratch*’ buffer, you

see this result:

     giraffe
     gazelle
     lion
     tiger
     nil
     nil

(The first nil is the value of the empty list that is printed; the second nil is the value returned by the whole function.)

11.2.2 Recursion in Place of a Counter

The triangle function described in a previous section can also be written recursively. It looks like this:

     (defun triangle-recursively (number)
       "Return the sum of the numbers 1 through NUMBER inclusive.
     Uses recursion."
       (if (= number 1)                    ; do-again-test

1 ; then-part

         (+ number                         ; else-part
            (triangle-recursively          ; recursive call

(1- number))))) ; next-step-expression

     (triangle-recursively 7)

You can install this function by evaluating it and then try it by evaluating (trianglerecursively 7). (Remember to put your cursor immediately after the last parenthesis of the function definition, before the comment.)

To understand how this function works, let’s consider what happens in the various cases when the function is passed 1, 2, 3, or 4 as the value of its argument.

First, what happens if the value of the argument is 1?

The function has an if expression after the documentation string. It tests whether the value of number is equal to 1; if so, Emacs evaluates the then-part of the if expression, which returns the number 1 as the value of the function. (A triangle with one row has one pebble in it.)

Suppose, however, that the value of the argument is 2. In this case, Emacs evaluates the else-part of the if expression.

The else-part consists of an addition, the recursive call to triangle-recursively and a decrementing action; and it looks like this:

     (+ number (triangle-recursively (1- number)))

When Emacs evaluates this expression, the innermost expression is evaluated first; then the other parts in sequence. Here are the steps in detail:

Step 1 Evaluate the innermost expression.

The innermost expression is (1- number) so Emacs decrements the value of number from 2 to 1.

Step 2 Evaluate the triangle-recursively function.

It does not matter that this function is contained within itself. Emacs passes the result Step 1 as the argument used by this instance of the triangle-recursively function In this case, Emacs evaluates triangle-recursively with an argument of 1. This means that this evaluation of triangle-recursively returns 1.

Step 3 Evaluate the value of number.

The variable number is the second element of the list that starts with +; its value is 2.

Step 4 Evaluate the + expression.

The + expression receives two arguments, the first from the evaluation of number (Step 3) and the second from the evaluation of triangle-recursively (Step 2). The result of the addition is the sum of 2 plus 1, and the number 3 is returned, which is correct. A triangle with two rows has three pebbles in it.

An argument of 3

Suppose that triangle-recursively is called with an argument of 3.

Step 1 Evaluate the do-again-test.

The if expression is evaluated first. This is the do-again test and returns false, so the else-part of the if expression is evaluated. (Note that in this example, the do-again-test causes the function to call itself when it tests false, not when it tests true.)

Step 2 Evaluate the innermost expression of the else-part.

The innermost expression of the else-part is evaluated, which decrements 3 to 2. This is the next-step-expression.

Step 3 Evaluate the triangle-recursively function.

The number 2 is passed to the triangle-recursively function. We know what happens when Emacs evaluates triangle-recursively with an argument of 2. After going through the sequence of actions described earlier, it returns a value of 3. So that is what will happen here.

Step 4 Evaluate the addition.

3 will be passed as an argument to the addition and will be added to the number with which the function was called, which is 3.

The value returned by the function as a whole will be 6.

Now that we know what will happen when triangle-recursively is called with an argument of 3, it is evident what will happen if it is called with an argument of 4:

In the recursive call, the evaluation of

(triangle-recursively (1- 4))

will return the value of evaluating

(triangle-recursively 3)

which is 6 and this value will be added to 4 by the addition in the third line.

The value returned by the function as a whole will be 10.

Each time triangle-recursively is evaluated, it evaluates a version of itself with a smaller argument, until the argument is small enough so that it does not evaluate itself.

11.2.3 Recursion Example Using cond

The version of triangle-recursively described earlier is written with the if special form. It can also be written using another special form called cond. The name of the special form cond is an abbreviation of the word ‘conditional’.

Although the cond special form is not used as often in the Emacs Lisp sources as if, it is used often enough to justify explaining it.

The template for a cond expression looks like this:

     (cond body...)

where the body is a series of lists.

Written out more fully, the template looks like this:

     (cond

((first-true-or-false-test first-consequent)

(second-true-or-false-test second-consequent) (third-true-or-false-test third-consequent)

       ...)

When the Lisp interpreter evaluates the cond expression, it evaluates the first element (the car or true-or-false-test) of the first expression in a series of expressions within the body of the

cond.

If the true-or-false-test returns nil the rest of that expression, the consequent, is skipped and the true-or-false-test of the next expression is evaluated. When an expression is found whose true-or-false-test returns a value that is not nil, the consequent of that expression is evaluated. The consequent can be one or more expressions. If the consequent consists of more than one expression, the expressions are evaluated in sequence and the value of the last one is returned. If the expression does not have a consequent, the value of the true-or-false-test is returned.

If none of the true-or-false-tests test true, the cond expression returns nil.

Written using cond, the triangle function looks like this:

     (defun triangle-using-cond (number)
       (cond ((<= number 0) 0)
             ((= number 1) 1)
             ((> number 1)
              (+ number (triangle-using-cond (1- number))))))

In this example, the cond returns 0 if the number is less than or equal to 0, it returns 1 if the

number is 1 and it evaluates (+ number (triangle-using-cond (1- number))) if the number

is greater than 1.

11.3 Looping Exercise

• Write a function similar to triangle in which each row has a value which is the square of the row number. Use a while loop.

• Write a function similar to triangle that multiplies instead of adds the values.

• Rewrite these two functions recursively. Rewrite these functions using cond.

• Write a function for Texinfo mode that creates an index entry at the beginning of a paragraph for every ‘@dfn’ within the paragraph. (In a Texinfo file, ‘@dfn’ marks a definition.

100 Chapter 11: Loops and Recursion For more information, see “Indicating Definitions, Commands, etc.” in Texinfo, The GNU Documentation Format.)

12 Regular Expression Searches

Regular expression searches are used extensively in GNU Emacs. The two functions,

forward-sentence and forward-paragraph illustrate these searches well.

Regular expression searches are described in section “Regular Expression Search” in The GNU Emacs Manual, as well as in section “Regular Expressions” in The GNU Emacs Lisp Reference Manual. In writing this chapter, I am presuming that you have at least a mild acquaintance with them. The major point to remember is that regular expressions permit you to search for patterns as well as for literal strings of characters. For example, the code in forward-sentence searches for the pattern of possible characters that could mark the end of a sentence, and moves point to that spot.

Before looking at the code for the forward-sentence function, it is worth considering what the pattern that marks the end of a sentence must be. The pattern is discussed in the next section; following that is a description of the regular expression search function, re-search-forward. The forward-sentence function is described in the section following. Finally, the forward-paragraph function is described in the last section of this chapter. forward-paragraph is a complex function that introduces several new features.

12.1 The Regular Expression for sentence-end

The symbol sentence-end is bound to the pattern that marks the end of a sentence. What should this regular expression be?

Clearly, a sentence may be ended by a period, a question mark, or an exclamation mark. Indeed, only clauses that end with one of those three characters should be considered the end of a sentence. This means that the pattern should include the character set:

     [.?!]

However, we do not want forward-sentence merely to jump to a period, a question mark, or an exclamation mark, because such a character might be used in the middle of a sentence. A period, for example, is used after abbreviations. So other information is needed.

According to convention, you type two spaces after every sentence, but only one space after a period, a question mark, or an exclamation mark in the body of a sentence. So a period, a question mark, or an exclamation mark followed by two spaces is a good indicator of an end of sentence. However, in a file, the two spaces may instead be a tab or the end of a line. This means that the regular expression should include these three items as alternatives. This group of alternatives will look like this:

     \\($\\| \\|  \\)
            ^^^
           TAB  SPC

Here, ‘$’ indicates the end of the line, and I have pointed out where the tab and two spaces are inserted in the expression. Both are inserted by putting the actual characters into the expression.

Two backslashes, ‘\\’, are required before the parentheses and vertical bars: the first backslash to quote the following backslash in Emacs; and the second to indicate that the following character, the parenthesis or the vertical bar, is special.

Also, a sentence may be followed by one or more carriage returns, like this:

     [ ]*

Like tabs and spaces, a carriage return is inserted into a regular expression by inserting it literally. The asterisk indicates that the ^{⟨}\underline{\overline{RET}}^{⟩} is repeated zero or more times.

But a sentence end does not consist only of a period, a question mark or an exclamation mark followed by appropriate space: a closing quotation mark or a closing brace of some kind may precede the space. Indeed more than one such mark or brace may precede the space. These require a expression that looks like this:

     []\"')}]*

In this expression, the first ‘]’ is the first character in the expression; the second character is ‘"’, which is preceded by a ‘\’ to tell Emacs the ‘"’ is not special. The last three characters are ‘'’, ‘)’, and ‘}’.

All this suggests what the regular expression pattern for matching the end of a sentence should be; and, indeed, if we evaluate sentence-end we find that it returns the following value:

     sentence-end
          ⇒ "[.?!][]\"')}]*\\($\\|     \\|  \\)[
     ]*"

12.2 The re-search-forward Function

The re-search-forward function is very like the search-forward function. (See Section 8.1.3, “The search-forward Function”, page 67.)

re-search-forward searches for a regular expression. If the search is successful, it leaves point immediately after the last character in the target. If the search is backwards, it leaves point just before the first character in the target. You may tell re-search-forward to return t for true. (Moving point is therefore a ‘side effect’.)

Like search-forward, the re-search-forward function takes four arguments:

1. The first argument is the regular expression that the function searches for. The regular expression will be a string between quotations marks.

2. The optional second argument limits how far the function will search; it is a bound, specified as a position in the buffer.

3. The optional third argument specifies how the function responds to failure: nil as the third argument causes the function to signal an error (and print a message) when the search fails; any other value causes it to return nil if the search fails and t if the search succeeds.

4. The optional fourth argument is the repeat count. A negative repeat count causes research-forward to search backwards.

The template for re-search-forward looks like this:

     (re-search-forward "regular-expression"

limit-of-search

what-to-do-if-search-fails repeat-count)

The second, third, and fourth arguments are optional. However, if you want to pass a value to either or both of the last two arguments, you must also pass a value to all the preceding arguments. Otherwise, the Lisp interpreter will mistake which argument you are passing the value to.

In the forward-sentence function, the regular expression will be the value of the variable

sentence-end, namely:

forward-sentence  103

     "[.?!][]\"')}]*\\($\\|  \\|  \\)[ ]*"

The limit of the search will be the end of the paragraph (since a sentence cannot go beyond a paragraph). If the search fails, the function will return nil; and the repeat count will be provided by the argument to the forward-sentence function.

12.3 forward-sentence

The command to move the cursor forward a sentence is a straightforward illustration of how to use regular expression searches in Emacs Lisp. Indeed, the function looks longer and more complicated than it is; this is because the function is designed to go backwards as well as forwards; and, optionally, over more than one sentence. The function is usually bound to the key command M-e.

Here is the code for forward-sentence:

     (defun forward-sentence (&optional arg)
       "Move forward to next sentence-end.  With argument, repeat.
     With negative argument, move backward repeatedly to sentence-beginning.
     Sentence ends are identified by the value of sentence-end
     treated as a regular expression.  Also, every paragraph boundary
     terminates sentences as well."
       (interactive "p")
       (or arg (setq arg 1))
       (while (< arg 0)
         (let ((par-beg
                (save-excursion (start-of-paragraph-text) (point))))
           (if (re-search-backward
                (concat sentence-end "[^\t\n]") par-beg t)
               (goto-char (1- (match-end 0)))
             (goto-char par-beg)))
         (setq arg (1+ arg)))
       (while (> arg 0)
         (let ((par-end
                (save-excursion (end-of-paragraph-text) (point))))
           (if (re-search-forward sentence-end par-end t)
               (skip-chars-backward " \t\n")
             (goto-char par-end)))
         (setq arg (1- arg))))

The function looks long at first sight and it is best to look at its skeleton first, and then its muscle. The way to see the skeleton is to look at the expressions that start in the left-most columns:

     (defun forward-sentence (&optional arg)

"documentation..."

       (interactive "p")
       (or arg (setq arg 1))
       (while (< arg 0)

body-of-while-loop

       (while (> arg 0)

body-of-while-loop

This looks much simpler! The function definition consists of documentation, an interactive expression, an or expression, and while loops.

Let’s look at each of these parts in turn.

We note that the documentation is thorough and understandable.

The function has an interactive "p" declaration. This means that the processed prefix argument, if any, is passed to the function as its argument. (This will be a number.) If the function is not passed an argument (it is optional) then the argument arg will be bound to 1. When forward-sentence is called non-interactively without an argument, arg is bound to nil.

The or expression handles the prefix argument. What it does is either leave the value of arg as it is, but only if arg is bound to a value; or it sets the value of arg to 1, in the case when arg is bound to nil.

The while loops

Two while loops follow the or expression. The first while has a true-or-false-test that tests true if the prefix argument for forward-sentence is a negative number. This is for going backwards. The body of this loop is similar to the body of the second while clause, but it is not exactly the same. We will skip this while loop and concentrate on the second while loop.

The second while loop is for moving point forward. Its skeleton looks like this:

(while (> arg 0) ; true-or-false-test

(let varlist

(if (true-or-false-test) then-part

else-part

       (setq arg (1- arg))))     ; while loop decrementer

The while loop is of the decrementing kind. (See Section 11.1.4, “A Loop with a Decrementing Counter”, page 93.) It has a true-or-false-test that tests true so long as the counter (in this case, the variable arg) is greater than zero; and it has a decrementer that subtracts 1 from the value of the counter every time the loop repeats.

If no prefix argument is given to forward-sentence, which is the most common way the command is used, this while loop will run once, since the value of arg will be 1.

The body of the while loop consists of a let expression, which creates and binds a local variable, and has, as its body, an if expression.

The body of the while loop looks like this:

     (let ((par-end
            (save-excursion (end-of-paragraph-text) (point))))
       (if (re-search-forward sentence-end par-end t)
           (skip-chars-backward " \t\n")
         (goto-char par-end)))

The let expression creates and binds the local variable par-end. As we shall see, this local variable is designed to provide a bound or limit to the regular expression search. If the search fails to find a proper sentence ending in the paragraph, it will stop on reaching the end of the paragraph.

But first, let us examine how par-end is bound to the value of the end of the paragraph. What happens is that the let sets the value of par-end to the value returned when the Lisp interpreter evaluates the expression

     (save-excursion (end-of-paragraph-text) (point))

In this expression, (end-of-paragraph-text) moves point to the end of the paragraph, (point) returns the value of point, and then save-excursion restores point to its original position. Thus, the let binds par-end to the value returned by the save-excursion expression, which is the position of the end of the paragraph. (The (end-of-paragraph-text) function uses forward-paragraph, which we will discuss shortly.)

Emacs next evaluates the body of the let, which is an if expression that looks like this:

     (if (re-search-forward sentence-end par-end t) ; if-part
         (skip-chars-backward " \t\n")              ; then-part
       (goto-char par-end)))                        ; else-part

The if tests whether its first argument is true and if so, evaluates its then-part; otherwise, the Emacs Lisp interpreter evaluates the else-part. The true-or-false-test of the if expression is the regular expression search.

It may seem odd to have what looks like the ‘real work’ of the forward-sentence function buried here, but this is a common way this kind of operation is carried out in Lisp.

The regular expression search

The re-search-forward function searches for the end of the sentence, that is, for the pattern defined by the sentence-end regular expression. If the pattern is found—if the end of the sentence is found—then the re-search-forward function does two things:

1. The re-search-forward function carries out a side effect, which is to move point to the end of the occurrence found.

2. The re-search-forward function returns a value of true. This is the value received by the if, and means that the search was successful.

The side effect, the movement of point, is completed before the if function is handed the value returned by the successful conclusion of the search.

When the if function receives the value of true from a successful call to re-search-forward, the if evaluates the then-part, which is the expression (skip-chars-backward " \t\n"). This expression moves backwards over any blank spaces, tabs or carriage returns until a printed character is found and then leaves point after the character. Since point has already been moved to the end of the pattern that marks the end of the sentence, this action leaves point right after the closing printed character of the sentence, which is usually a period.

On the other hand, if the re-search-forward function fails to find a pattern marking the end of the sentence, the function returns false. The false then causes the if to evaluate its third argument, which is (goto-char par-end): it moves point to the end of the paragraph.

Regular expression searches are exceptionally useful and the pattern illustrated by research-forward, in which the search is the test of an if expression, is handy. You will see or write code incorporating this pattern often.

12.4 forward-paragraph: a Goldmine of Functions

The forward-paragraph function moves point forward to the end of the paragraph. It is usually bound to M-} and makes use of a number of functions that are important in themselves,

including let*, match-beginning, and looking-at.

The function definition for forward-paragraph is considerably longer than the function definition for forward-sentence because it works with a paragraph, each line of which may begin with a fill prefix.

A fill prefix consists of a string of characters that are repeated at the beginning of each line. For example, in Lisp code, it is a convention to start each line of a paragraph-long comment with ‘;;; ’. In Text mode, four blank spaces make up another common fill prefix, creating an indented paragraph. (See section “Fill Prefix” in The GNU Emacs Manual, for more information about fill prefixes.)

The existence of a fill prefix means that in addition to being able to find the end of a paragraph whose lines begin on the left-most column, the forward-paragraph function must be able to find the end of a paragraph when all or many of the lines in the buffer begin with the fill prefix.

Moreover, it is sometimes practical to ignore a fill prefix that exists, especially when blank lines separate paragraphs. This is an added complication.

Rather than print all of the forward-paragraph function, we will only print parts of it. Read without preparation, the function can be daunting!

In outline, the function looks like this:

     (defun forward-paragraph (&optional arg)

"documentation..."

       (interactive "p")
       (or arg (setq arg 1))
       (let*

varlist (while (< arg 0) ; backward-moving-code

           ...
           (setq arg (1+ arg)))

(while (> arg 0) ; forward-moving-code

           ...
           (setq arg (1- arg)))))

The first parts of the function are routine: the function’s argument list consists of one optional argument. Documentation follows.

The lower case ‘p’ in the interactive declaration means that the processed prefix argument, if any, is passed to the function. This will be a number, and is the repeat count of how many paragraphs point will move. The or expression in the next line handles the common case when no argument is passed to the function, which occurs if the function is called from other code rather than interactively. This case was described earlier. (See Section 12.3, “forward-sentence”, page 103.) Now we reach the end of the familiar part of this function.

The let* expression

The next line of the forward-paragraph function begins a let* expression. This is a different kind of expression than we have seen so far. The symbol is let* not let.

The let* special form is like let except that Emacs sets each variable in sequence, one after another, and variables in the latter part of the varlist can make use of the values to which Emacs set variables in the earlier part of the varlist.

In the let* expression in this function, Emacs binds two variables: fill-prefix-regexp and paragraph-separate. The value to which paragraph-separate is bound depends on the

value of fill-prefix-regexp.

Let’s look at each in turn. The symbol fill-prefix-regexp is set to the value returned by evaluating the following list:

     (and fill-prefix
          (not (equal fill-prefix ""))
          (not paragraph-ignore-fill-prefix)
          (regexp-quote fill-prefix))

This is an expression whose first element is the function and.

The and function evaluates each of its arguments until one of the arguments returns a value of nil, in which case the and expression returns nil; however, if none of the arguments returns a value of nil, the value resulting from evaluating the last argument is returned. (Since such a value is not nil, it is considered true in Lisp.) In other words, an and expression returns a true value only if all its arguments are true.

In this case, the variable fill-prefix-regexp is bound to a non-nil value only if the following four expressions produce a true (i.e., a non-nil) value when they are evaluated; otherwise,

fill-prefix-regexp is bound to nil.
fill-prefix

When this variable is evaluated, the value of the fill prefix, if any, is returned. If there is no fill prefix, this variable returns nil.

(not (equal fill-prefix "")

This expression checks whether an existing fill prefix is an empty string, that is, a string with no characters in it. An empty string is not a useful fill prefix.

(not paragraph-ignore-fill-prefix)

This expression returns nil if the variable paragraph-ignore-fill-prefix has been turned on by being set to a true value such as t.

(regexp-quote fill-prefix)

This is the last argument to the and function. If all the arguments to the and are true, the value resulting from evaluating this expression will be returned by the and expression and bound to the variable fill-prefix-regexp,

The result of evaluating this and expression successfully is that fill-prefix-regexp will be bound to the value of fill-prefix as modified by the regexp-quote function. What regexpquote does is read a string and return a regular expression that will exactly match the string and match nothing else. This means that fill-prefix-regexp will be set to a value that will exactly match the fill prefix if the fill prefix exists. Otherwise, the variable will be set to nil.

The second local variable in the let* expression is paragraph-separate. It is bound to the value returned by evaluating the expression:

     (if fill-prefix-regexp
         (concat paragraph-separate
                 "\\|^" fill-prefix-regexp "[ \t]*$")
       paragraph-separate)))

This expression shows why let* rather than let was used. The true-or-false-test for the if depends on whether the variable fill-prefix-regexp evaluates to nil or some other value.

If fill-prefix-regexp does not have a value, Emacs evaluates the else-part of the if expression and binds paragraph-separate to its local value. (paragraph-separate is a regular expression that matches what separates paragraphs.)

But if fill-prefix-regexp does have a value, Emacs evaluates the then-part of the if expression and binds paragraph-separate to a regular expression that includes the fillprefix-regexp as part of the pattern.

Specifically, paragraph-separate is set to the original value of the paragraph separate regular expression concatenated with an alternative expression that consists of the fill-prefix-regexp followed by a blank line. The ‘^’ indicates that the fill-prefix-regexp must begin a line, and the optional whitespace to the end of the line is defined by "[ \t]*$".) The ‘\\|’ defines this portion of the regexp as an alternative to paragraph-separate.

Now we get into the body of the let*. The first part of the body of the let* deals with the case when the function is given a negative argument and is therefore moving backwards. We will skip this section.

The forward motion while loop

The second part of the body of the let* deals with forward motion. It is a while loop that repeats itself so long as the value of arg is greater than zero. In the most common use of the function, the value of the argument is 1, so the body of the while loop is evaluated exactly once, and the cursor moves forward one paragraph.

This part handles three situations: when point is between paragraphs, when point is within a paragraph and there is a fill prefix, and when point is within a paragraph and there is no fill prefix.

The while loop looks like this:

     (while (> arg 0)
       (beginning-of-line)

;; between paragraphs

       (while (prog1 (and (not (eobp))
                          (looking-at paragraph-separate))
                (forward-line 1)))

;; within paragraphs, with a fill prefix

       (if fill-prefix-regexp

;; There is a fill prefix; it overrides paragraph-start.

           (while (and (not (eobp))
                       (not (looking-at paragraph-separate))
                       (looking-at fill-prefix-regexp))
             (forward-line 1))

;; within paragraphs, no fill prefix

         (if (re-search-forward paragraph-start nil t)
             (goto-char (match-beginning 0))
           (goto-char (point-max))))
       (setq arg (1- arg)))

We can see immediately that this is a decrementing counter while loop, using the expression (setq (1- arg)) as the decrementer. The body of the loop consists of three expressions:

;; between paragraphs

     (beginning-of-line)
     (while

body-of-while)

;; within paragraphs, with fill prefix (if true-or-false-test then-part

;; within paragraphs, no fill prefix else-part

When the Emacs Lisp interpreter evaluates the body of the while loop, the first thing it does is evaluate the (beginning-of-line) expression and move point to the beginning of the line. Then there is an inner while loop. This while loop is designed to move the cursor out of the blank space between paragraphs, if it should happen to be there. Finally there is an if expression that actually moves point to the end of the paragraph.

Between paragraphs

First, let us look at the inner while loop. This loop handles the case when point is between paragraphs; it uses three functions that are new to us: prog1, eobp and looking-at.

• prog1 is similar to the progn function, except that prog1 evaluates its arguments in sequence and then returns the value of its first argument as the value of the whole expression. (progn returns the value of its last argument as the value of the expression.) The second and subsequent arguments to prog1 are evaluated only for their side effects.

• eobp is an abbreviation of ‘End Of Buffer P’ and is a function that returns true if point is at the end of the buffer.

• looking-at is a function that returns true if the text following point matches the regular expression passed looking-at as its argument.

The while loop we are studying looks like this:

     (while (prog1 (and (not (eobp))
                        (looking-at paragraph-separate))
                   (forward-line 1)))

This is a while loop with no body! The true-or-false-test of the loop is the expression:

     (prog1 (and (not (eobp))
                 (looking-at paragraph-separate))
            (forward-line 1)))

The first argument to the prog1 is the and expression. It has within in it a test of whether point is at the end of the buffer and also a test of whether the pattern following point matches the regular expression for separating paragraphs.

If the cursor is not at the end of the buffer and if the characters following the cursor mark the separation between two paragraphs, then the and expression is true. After evaluating the and expression, the Lisp interpreter evaluates the second argument to prog1, which is forwardline. This moves point forward one line. The value returned by the prog1 however, is the value of its first argument, so the while loop continues so long as point is not at the end of the buffer and is between paragraphs. When, finally, point is moved to a paragraph, the and expression tests false. Note however, that the forward-line command is carried out anyhow. This means that when point is moved from between paragraphs to a paragraph, it is left at the beginning of the second line of the paragraph.

Within paragraphs

The next expression in the outer while loop is an if expression. The Lisp interpreter evaluates the then-part of the if when the fill-prefix-regexp variable has a value other than nil, and it evaluates the else-part when the value of if fill-prefix-regexp is nil, that is, when there is no fill prefix.

No fill prefix

It is simplest to look at the code for the case when there is no fill prefix first. This code consists of yet another inner if expression, and reads as follows:

     (if (re-search-forward paragraph-start nil t)
         (goto-char (match-beginning 0))
       (goto-char (point-max)))

This expression actually does the work that most people think of as the primary purpose of the forward-paragraph command: it causes a regular expression search to occur that searches forward to the start of the next paragraph and if it is found, moves point there; but if the start of another paragraph if not found, it moves point to the end of the accessible region of the buffer.

The only unfamiliar part of this is the use of match-beginning. This is another function that is new to us. The match-beginning function returns a number specifying the location of the start of the text that was matched by the last regular expression search.

The match-beginning function is used here because of a characteristic of a forward search: a successful forward search, regardless of whether it is a plain search or a regular expression search, will move point to the end of the text that is found. In this case, a successful search will move point to the end of the pattern for paragraph-start, which will be the beginning of the next paragraph rather than the end of the current one.

However, we want to put point at the end of the current paragraph, not at the beginning of the next one. The two positions may be different, because there may be several blank lines between paragraphs.

When given an argument of 0, match-beginning returns the position that is the start of the text that the most recent regular expression search matched. In this case, the most recent regular expression search is the one looking for paragraph-start, so match-beginning returns the beginning position of the pattern, rather than the end of the pattern. The beginning position is the end of the paragraph.

(Incidentally, when passed a positive number as an argument, the match-beginning function will place point at that parenthesized expression in the last regular expression. It is a useful function.)

With a fill prefix

The inner if expression just discussed is the else-part of an enclosing if expression which tests whether there is a fill prefix. If there is a fill prefix, the then-part of this if is evaluated. It looks like this:

     (while (and (not (eobp))
                 (not (looking-at paragraph-separate))
                 (looking-at fill-prefix-regexp))
       (forward-line 1))

What this expression does is move point forward line by line so long as three conditions are true:

1. Point is not at the end of the buffer.

2. The text following point does not separate paragraphs.

3. The pattern following point is the fill prefix regular expression.

The last condition may be puzzling, until you remember that point was moved to the beginning of the line early in the forward-paragraph function. This means that if the text has a fill prefix, the looking-at function will see it.

Summary

In summary, when moving forward, the forward-paragraph function does the following:

• Move point to the beginning of the line.

• Skip over lines between paragraphs.

• Check whether there is a fill prefix, and if there is:

— Go forward line by line so long as the line is not a paragraph separating line.

• But if there is no fill prefix,

— Search for the next paragraph start pattern.

— Go to the beginning of the paragraph start pattern, which will be the end of the previous paragraph.

— Or else go to the end of the accessible portion of the buffer.

For review, here is the code we have just been discussing, formatted for clarity:

(interactive "p")
(or arg (setq arg 1))
(let* (
(fill-prefix-regexp
(and fill-prefix (not (equal fill-prefix ""))
(not paragraph-ignore-fill-prefix)
(regexp-quote fill-prefix)))
(paragraph-separate
(if fill-prefix-regexp
(concat paragraph-separate
"\\|^"
fill-prefix-regexp
"[ \t]*$")
paragraph-separate)))

backward-moving-code (omitted) ...

(while (> arg 0) ; forward-moving-code

(beginning-of-line)
(while (prog1 (and (not (eobp))
   (looking-at paragraph-separate))
(forward-line 1)))
(if fill-prefix-regexp
(while (and (not (eobp))  ; then-part
(not (looking-at paragraph-separate))
(looking-at fill-prefix-regexp))
(forward-line 1))

; else-part: the inner-if

           (if (re-search-forward paragraph-start nil t)
               (goto-char (match-beginning 0))
             (goto-char (point-max))))
         (setq arg (1- arg)))))        ; decrementer

The full definition for the forward-paragraph function not only includes this code for going forwards, but also code for going backwards.

If you are reading this inside of GNU Emacs and you want to see the whole function, you can type M-. (find-tag) and the name of the function when prompted for it. If the find-tag function first asks you for the name of a ‘TAGS’ table, give it the name of the ‘TAGS’ file in your ‘emacs/src’ directory, which will have a pathname such as

‘/usr/local/lib/emacs/19.23/src/TAGS’. (The exact path to the ‘emacs/src’ directory

depends on how your copy of Emacs was installed. If you don’t know the path, you can sometimes find out by typing C-h i to enter Info and then typing C-x C-f to see the path to the ‘emacs/info’ directory. The path to the ‘TAGS’ file is often the corresponding ‘emacs/src’ path; sometimes, however, Info files are stored elsewhere.)

You can also create your own ‘TAGS’ file for directories that lack one.

12.5 Create Your Own ‘TAGS’ File

You can create your own ‘TAGS’ file to help you jump to sources. For example, if you have a large number of files in your ‘~/emacs’ directory, as I do—I have 137 ‘.el’ files in it, of which I load 17— you will find it easier to jump to specific functions if you create a ‘TAGS’ file for that directory than if you search for the function name with grep or some other tool.

You can create a ‘TAGS’ file by calling the etags program that comes as a part of the Emacs distribution. Usually, etags is compiled and installed when Emacs is built. (etags is not an Emacs Lisp function or a part of Emacs; it is a C program.)

To create a ‘TAGS’ file, first switch to the directory in which you want to create the file. In Emacs you can do this with the M-x cd command, or by visiting a file in the directory, or by listing the directory with C-x d (dired). Then type

     M-! etags *.el

to create a ‘TAGS’ file. The etags program takes all the usual shell ‘wildcards’. For example, if you have two directories for which you want a single ‘TAGS file’, type the command like this, where ‘../elisp/’ is the second directory:

     M-! etags  *.el ../elisp/*.el

Type

     M-! etags --help

to see a list of the options accepted by etags.

The etags program handles Emacs Lisp, Common Lisp, Scheme, C, Fortran, Pascal, LaTeX, and most assemblers. The program has no switches for specifying the language; it recognizes the language in an input file according to its file name and contents.

Also, ‘etags’ is very helpful when you are writing code yourself and want to refer back to functions you have already written. Just run etags again at intervals as you write new functions, so they become part of the ‘TAGS’ file.

12.6 Review

Here is a brief summary of some recently introduced functions.

while Repeatedly evaluate the body of the expression so long as the first element of the body tests true. Then return nil. (The expression is evaluated only for its side effects.) For example:

     (let ((foo 2))
       (while (> foo 0)
         (insert (format "foo is %d.\n" foo))
         (setq foo (1- foo))))
          ⇒  foo is 2.
                  foo is 1.
                  nil

(The insert function inserts its arguments at point; the format function returns a string formatted from its arguments the way message formats its arguments; \n produces a new line.)

re-search-forward

Search for a pattern, and if the pattern is found, move point to rest just after it. Takes four arguments, like search-forward:

1. A regular expression that specifies the pattern to search for.

2. Optionally, the limit of the search.

3. Optionally, what to do if the search fails, return nil or an error message.

4. Optionally, how many times to repeat the search; if negative, the search goes backwards.

let* Bind some variables locally to particular values, and then evaluate the remaining arguments, returning the value of the last one. While binding the local variables, use the local values of variables bound earlier, if any. For example:

     (let* ((foo 7)
           (bar (* 3 foo)))
       (message "`bar' is %d." bar))
          ⇒ `bar' is 21.
match-beginning

Return the position of the start of the text found by the last regular expression search.

looking-at

Return t for true if the text after point matches the argument, which should be a regular expression.

eobp Return t for true if point is at the end of the accessible part of a buffer. The end of the accessible part is the end of the buffer if the buffer is not narrowed; it is the end of the narrowed part if the buffer is narrowed.

prog1 Evaluate each argument in sequence and then return the value of the first.

For example:

           (prog1 1 2 3 4) ⇒ 1

12.7 Exercises with re-search-forward

• Write a function to search for a regular expression that matches two or more blank lines in sequence.

• Write a function to search for duplicated words, such as ‘the the’. See section “Syntax of Regular Expressions” in The GNU Emacs Manual, for information on how to write a regexp (a regular expression) to match a string that is composed of two identical halves. You can devise several regexps; some are better than others. The function I use is described in an appendix, along with several regexps. See Appendix A, “the-the Duplicated Words Function”, page 169.

The count-words-region Function  115

13 Counting: Repetition and Regexps

Repetition and regular expression searches are powerful tools that you often use when you write code in Emacs Lisp. This chapter illustrates the use of regular expression searches through the construction of word count commands using while loops and recursion.

The standard Emacs distribution contains a function for counting the number of lines within a region. However, there is no corresponding function for counting words.

Certain types of writing ask you to count words. Thus, if you write an essay, you may be limited to 800 words; if you write a novel, you may discipline yourself to write 1000 words a day. It seems odd to me that Emacs lacks a word count command. Perhaps people use Emacs mostly for code or types of documentation that do not require word counts; or perhaps they restrict themselves to the operating system word count command, wc. Alternatively, people may follow the publishers’ convention and compute a word count by dividing the number of characters in a document by five. In any event, here are commands to count words.

13.1 The count-words-region Function

A word count command could count words in a line, paragraph, region, or buffer. What should the command cover? You could design the command to count the number of words in a complete buffer. However, the Emacs tradition encourages flexibility—you may want to count words in just a section, rather than all of a buffer. So it makes more sense to design the command to count the number of words in a region. Once you have a count-words-region command, you can, if you wish, count words in a whole buffer by marking it with C-x h (mark-whole-buffer).

Clearly, counting words is a repetitive act: starting from the beginning of the region, you count the first word, then the second word, then the third word, and so on, until you reach the end of the region. This means that word counting is ideally suited to recursion or to a while loop.

First, we will implement the word count command with a while loop, then with recursion. The command will, of course, be interactive.

The template for an interactive function definition is, as always:

(defun name-of-function (argument-list) "documentation..." (interactive-expression...)

       body...)

What we need to do is fill in the slots.

The name of the function should be self-explanatory and similar to the existing countlines-region name. This makes the name easier to remember. count-words-region is a good choice.

The function counts words within a region. This means that the argument list must contain symbols that are bound to the two positions, the beginning and end of the region. These two positions can be called ‘beginning’ and ‘end’ respectively. The first line of the documentation should be a single sentence, since that is all that is printed as documentation by a command such as apropos. The interactive expression will be of the form ‘(interactive "r")’, since that will cause Emacs to pass the beginning and end of the region to the function’s argument list. All this is routine.

The body of the function needs to be written so as to do three tasks: first to set up conditions under which the while loop can count words, second to run the while loop, and, third, to send a message to the user.

When a user calls count-words-region, point may be at the beginning or the end of the region. However, the counting process must start at the beginning of the region. This means we will want to put point there if it is not already there. Executing (goto-char beginning) ensures this. Of course, we will want to return point to its expected position when the function finishes its work. For this reason, the body must be enclosed in a save-excursion expression.

The central part of the body of the function consists of a while loop in which one expression jumps point forward word by word, and another expression counts those jumps. The true-or-false-test of the while loop should test true so long as point should jump forward, and false when point is at the end of the region.

We could use (forward-word 1) as the expression for moving point forward word by word, but it is easier to see what Emacs identifies as a ‘word’ if we use a regular expression search.

A regular expression search that finds the pattern for which it is searching leaves point after the last character matched. This means that a succession of successful word searches will move point forward word by word.

As a practical matter, we want the regular expression search to jump over whitespace and punctuation between words as well as over the words themselves. A regexp that refuses to jump over interword whitespace would never jump more than one word! This means that the regexp should include the whitespace and punctuation that follows a word, if any, as well as the word itself. (A word may end a buffer and not have any following whitespace or punctuation, so that part of the regexp must be optional.)

Thus, what we want for the regexp is a pattern defining one or more word constituent characters followed, optionally, by one or more characters that are not word constituents. The regular expression for this is:

     \w+\W*

The buffer’s syntax table determines which characters are and are not word constituents. (See Section 14.2, “What Constitutes a Word or Symbol?”, page 126, for more about syntax. Also, see section “The Syntax Table” in The GNU Emacs Manual, and, section “Syntax Tables” in The GNU Emacs Lisp Reference Manual.)

The search expression looks like this:

     (re-search-forward "\\w+\\W*")

(Note that paired backslashes precede the ‘w’ and ‘W’. A single backslash has special meaning to the Emacs Lisp interpreter. It indicates that the following character is interpreted differently than usual. For example, the two characters, ‘\n’, stand for ‘newline’, rather than for a backslash followed by ‘n’. Two backslashes in a row stand for an ordinary, ‘unspecial’ backslash.)

We need a counter to count how many words there are; this variable must first be set to 0 and then incremented each time Emacs goes around the while loop. The incrementing expression is simply:

     (setq count (1+ count))

Finally, we want to tell the user how many words there are in the region. The message function is intended for presenting this kind of information to the user. The message has to be phrased so that it reads properly regardless of how many words there are in the region: we don’t want to say that “there are 1 words in the region”. The conflict between singular and plural is ungrammmatical. We can solve this problem by using a conditional expression that evaluates different messages depending on the number of words in the region. There are three possibilities: no words in the region, one word in the region, and more than one word. This means that the cond special form is appropriate.

All this leads to the following function definition:

;;; First version; has bugs!

(defun count-words-region (beginning end)
  "Print number of words in the region.
Words are defined as at least one word-constituent
character followed by at least one character that
is not a word-constituent.  The buffer's syntax
table determines which characters these are."
  (interactive "r")
  (message "Counting words in region ... ")

;;; 1. Set up appropriate conditions.

  (save-excursion
    (goto-char beginning)
    (let ((count 0))

;;; 2. Run the while loop.

      (while (< (point) end)
        (re-search-forward "\\w+\\W*")
        (setq count (1+ count)))

;;; 3. Send a message to the user.

      (cond ((zerop count)
             (message
              "The region does NOT have any words."))
            ((= 1 count)
             (message
              "The region has 1 word."))
            (t
             (message
              "The region has %d words." count))))))

As written, the function works, but not in all circumstances.

13.1.1 The Whitespace Bug in count-words-region

The count-words-region command described in the preceding section has two bugs, or rather, one bug with two manifestations. First, if you mark a region containing only whitespace in the middle of some text, the count-words-region command tells you that the region contains one word! Second, if you mark a region containing only whitespace at the end of the buffer or the accessible portion of a narrowed buffer, the command displays an error message that looks like this:

Search failed: "\\w+\\W*"

If you are reading this in Info in GNU Emacs, you can test for these bugs yourself.

First, evaluate the function in the usual manner to install it.

If you wish, you can also install this keybinding by evaluating it, too:

(global-set-key "\C-c=" 'count-words-region)

To conduct the first test, set mark and point to the beginning and end of the following line and then type C-c = (or M-x count-words-region if you have not bound C-c =):

         one   two  three

Emacs will tell you, correctly, that the region has three words.

Repeat the test, but place mark at the beginning of the line and place point just before the word ‘one’. Again type the command C-c = (or M-x count-words-region). Emacs should tell you that the region has no words, since it is composed only of the whitespace at the beginning of the line. But instead Emacs tells you that the region has one word!

For the third test, copy the sample line to the end of the ‘*scratch*’ buffer and then type several spaces at the end of the line. Place mark right after the word ‘three’ and point at the end of line. (The end of the line will be the end of the buffer.) Type C-c = (or M-x countwords-region) as you did before. Again, Emacs should tell you that the region has no words, since it is composed only of the whitespace at the end of the line. Instead, Emacs displays an error message saying ‘Search failed’.

The two bugs stem from the same problem.

Consider the first manifestation of the bug, in which the command tells you that the whitespace at the beginning of the line contains one word. What happens is this: The M-x countwords-region command moves point to the beginning of the region. The while tests whether the value of point is smaller than the value of end, which it is. Consequently, the regular expression search looks for and finds the first word. It leaves point after the word. count is set to one. The while loop repeats; but this time the value of point is larger than the value of end, the loop is exited; and the function displays a message saying the number of words in the region is one. In brief, the regular expression search looks for and finds the word even though it is outside the marked region.

In the second manifestation of the bug, the region is whitespace at the end of the buffer. Emacs says ‘Search failed’. What happens is that the true-or-false-test in the while loop tests true, so the search expression is executed. But since there are no more words in the buffer, the search fails.

In both manifestations of the bug, the search extends or attempts to extend outside of the region.

The solution is to limit the search to the region—this is a fairly simple action, but as you may have come to expect, it is not quite as simple as you might think.

As we have seen, the re-search-forward function takes a search pattern as its first argument. But in addition to this first, mandatory argument, it accepts three optional arguments. The optional second argument bounds the search. The optional third argument, if t, causes the function to return nil rather than signal an error if the search fails. The optional fourth argument is a repeat count. (In Emacs, you can get a function’s documentation by typing C-h f, the name of the function, and then ^{⟨}\underline{\overline{RET}}^{⟩}.)

In the count-words-region definition, the value of the end of the region is held by the variable end which is passed as an argument to the function. Thus, we can add end as an argument to the regular expression search expression:

     (re-search-forward "\\w+\\W*" end)

However, if you make only this change to the count-words-region definition and then test the new version of the definition on a stretch of whitespace, you will receive an error message

saying ‘Search failed’.

What happens is this: the search is limited to the region, and fails as you expect because there are no word-constituent characters in the region. Since it fails, we receive an error message. But we do not want to receive an error message in this case; we want to receive the message that "The region does NOT have any words."

The solution to this problem is to provide re-search-forward with a third argument of t, which causes the function to return nil rather than signal an error if the search fails.

However, if you make this change and try it, you will see the message “Counting words in region ... ” and … you will keep on seeing that message …, until you type C-g (keyboard-

quit).

Here is what happens: the search is limited to the region, as before, and it fails because there are no word-constituent characters in the region, as expected. Consequently, the research-forward expression returns nil. It does nothing else. In particular, it does not move point, which it does as a side effect if it finds the search target. After the re-search-forward expression returns nil, the next expression in the while loop is evaluated. This expression increments the count. Then the loop repeats. The true-or-false-test tests true because the value of point is still less than the value of end, since the re-search-forward expression did not move point. … and the cycle repeats …

The count-words-region definition requires yet another modification, to cause the trueor-false-test of the while loop to test false if the search fails. Put another way, there are two conditions that must be satisfied in the true-or-false-test before the word count variable is incremented: point must still be within the region and the search expression must have found a word to count.

Since both the first condition and the second condition must be true together, the two expressions, the region test and the search expression, can be joined with an and function and embedded in the while loop as the true-or-false-test, like this:

     (and (< (point) end) (re-search-forward "\\w+\\W*" end t))

(For information about and, see Section 12.4, “forward-paragraph: a Goldmine of Functions”, page 105.)

The re-search-forward expression returns t if the search succeeds and as a side effect moves point. Consequently, as words are found, point is moved through the region. When the search expression fails to find another word, or when point reaches the end of the region, the true-or-false-test tests false, the while loop exists, and the count-words-region function displays one or other of its messages.

After incorporating these final changes, the count-words-region works without bugs (or at least, without bugs that I have found!). Here is what it looks like:

;;; Final version: while

     (defun count-words-region (beginning end)
       "Print number of words in the region."
       (interactive "r")
       (message "Counting words in region ... ")

;;; 1. Set up appropriate conditions.

       (save-excursion
         (let ((count 0))
           (goto-char beginning)

;;; 2. Run the while loop.

           (while (and (< (point) end)
                       (re-search-forward "\\w+\\W*" end t))
             (setq count (1+ count)))

;;; 3. Send a message to the user.

           (cond ((zerop count)
                  (message
                   "The region does NOT have any words."))
                 ((= 1 count)
                  (message
                   "The region has 1 word."))
                 (t
                  (message
                   "The region has %d words." count))))))

13.2 Count Words Recursively

You can write the function for counting words recursively as well as with a while loop. Let’s see how this is done.

First, we need to recognize that the count-words-region function has three jobs: it sets up the appropriate conditions for counting to occur; it counts the words in the region; and it sends a message to the user telling how many words there are.

If we write a single recursive function to do everything, we will receive a message for every recursive call. If the region contains 13 words, we will receive thirteen messages, one right after the other. We don’t want this! Instead, we must write two functions to do the job, one of which (the recursive function) will be used inside of the other. One function will set up the conditions and display the message; the other will return the word count.

Let us start with the function that causes the message to be displayed. We can continue to

call this count-words-region.

This is the function that the user will call. It will be interactive. Indeed, it will be similar to our previous versions of this function, except that it will call recursive-count-words to determine how many words are in the region.

We can readily construct a template for this function, based on our previous versions:

;; Recursive version; uses regular expression search

     (defun count-words-region (beginning end)

"documentation..."

(interactive-expression...)

;;; 1. Set up appropriate conditions.

(explanatory message) (set-up functions...

;;; 2. Count the words. recursive call

;;; 3. Send a message to the user.

message providing word count))

The definition looks straightforward, except that somehow, the count returned by the recursive call must be passed to the message displaying the word count. A little thought suggests that this can be done by making use of a let expression: we can bind a variable in the varlist of a let expression to the number of words in the region, as returned by the recursive call; and then the cond expression, using binding, can display the value to the user.

Often, one thinks of the binding within a let expression as somehow secondary to the ‘primary’ work of a function. But in this case, what you might consider the ‘primary’ job of the function, counting words, is done within the let expression.

Using let, the function definition looks like this:

     (defun count-words-region (beginning end)
       "Print number of words in the region."
       (interactive "r")

;;; 1. Set up appropriate conditions.

       (message "Counting words in region ... ")
       (save-excursion
         (goto-char beginning)

;;; 2. Count the words.

         (let ((count (recursive-count-words end)))

;;; 3. Send a message to the user.

           (cond ((zerop count)
                  (message
                   "The region does NOT have any words."))
                 ((= 1 count)
                  (message
                   "The region has 1 word."))
                 (t
                  (message
                   "The region has %d words." count))))))

Next, we need to write the recursive counting function.

A recursive function has at least three parts: the ‘do-again-test’, the ‘next-step-expression’, and the recursive call.

The do-again-test determines whether the function will or will not be called again. Since we are counting words in a region and can use a function that moves point forward for every word, the do-again-test can check whether point is still within the region. The do-again-test should find the value of point and determine whether point is before, at, or after the value of the end of the region. We can use the point function to locate point. Clearly, we must pass the value of the end of the region to the recursive counting function as an argument.

In addition, the do-again-test should also test whether the search finds a word. If it does not, the function should not call itself again.

The next-step-expression changes a value so that when the recursive function is supposed to stop calling itself, it stops. More precisely, the next-step-expression changes a value so that at the right time, the do-again-test stops the recursive function from calling itself again. In this case, the next-step-expression can be the expression that moves point forward word by word.

The third part of a recursive function is the recursive call.

Somewhere, also, we also need a part that does the ‘work’ of the function, a part that does the counting. A vital part!

But already, we have an outline of the recursive counting function:

     (defun recursive-count-words (region-end)

"documentation..." do-again-test next-step-expression recursive call)

Now we need to fill in the slots. Let’s start with the simplest cases first: if point is at or beyond the end of the region, there cannot be any words in the region, so the function should return zero. Likewise, if the search fails, there are no words to count, so the function should return zero.

On the other hand, if point is within the region and the search succeeds, the function should call itself again.

Thus, the do-again-test should look like this:

     (and (< (point) region-end)
          (re-search-forward "\\w+\\W*" region-end t))

Note that the search expression is part of the do-again-test—the function returns t if its search succeeds and nil if it fails. (See Section 13.1.1, “The Whitespace Bug in count-words-region”, page 117, for an explanation of how re-search-forward works.)

The do-again-test is the true-or-false test of an if clause. Clearly, if the do-again-test succeeds, the then-part of the if clause should call the function again; but if it fails, the else-part should return zero since either point is outside the region or the search failed because there were no words to find.

But before considering the recursive call, we need to consider the next-step-expression. What is it? Interestingly, it is the search part of the do-again-test.

In addition to returning t or nil for the do-again-test, re-search-forward moves point forward as a side effect of a successful search. This is the action that changes the value of point so that the recursive function stops calling itself when point completes its movement through the region. Consequently, the re-search-forward expression is the next-step-expression.

In outline, then, the body of the recursive-count-words function looks like this:

(if do-again-test-and-next-step-combined ;; then recursive-call-returning-count

;; else return-zero)

How to incorporate the mechanism that counts?

If you are not used to writing recursive functions, a question like this can be troublesome. But it can and should be approached systematically.

We know that the counting mechanism should be associated in some way with the recursive call. Indeed, since the next-step-expression moves point forward by one word, and since a recursive call is made for each word, the counting mechanism must be an expression that adds one to the value returned by a call to recursive-count-words.

Consider several cases:

• If there are two words in the region, the function should return a value resulting from adding one to the value returned when it counts the first word, plus the number returned when it counts the remaining words in the region, which in this case is one.

• If there is one word in the region, the function should return a value resulting from adding one to the value returned when it counts that word, plus the number returned when it counts the remaining words in the region, which in this case is zero.

• If there are no words in the region, the function should return zero.

From the sketch we can see that the else-part of the if returns zero for the case of no words. This means that the then-part of the if must return a value resulting from adding one to the value returned from a count of the remaining words.

The expression will look like this, where 1+ is a function that adds one to its argument.

     (1+ (recursive-count-words region-end))

The whole recursive-count-words function will then look like this:

     (defun recursive-count-words (region-end)

"documentation..."

;;; 1. do-again-test

       (if (and (< (point) region-end)
                (re-search-forward "\\w+\\W*" region-end t))

;;; 2. then-part: the recursive call

           (1+ (recursive-count-words region-end))

;;; 3. else-part

         0))

Let’s examine how this works:

If there are no words in the region, the else part of the if expression is evaluated and consequently the function returns zero.

If there is one word in the region, the value of point is less than the value of region-end and the search succeeds. In this case, the true-or-false-test of the if expression tests true, and the then-part of the if expression is evaluated. The counting expression is evaluated. This expression returns a value (which will be the value returned by the whole function) that is the sum of one added to the value returned by a recursive call.

Meanwhile, the next-step-expression has caused point to jump over the first (and in this case only) word in the region. This means that when (recursive-count-words region-end) is evaluated a second time, as a result of the recursive call, the value of point will be equal to or greater than the value of region end. So this time, recursive-count-words will return zero. The zero will be added to one, and the original evaluation of recursive-count-words will return one plus zero, which is one, which is the correct amount.

Clearly, if there are two words in the region, the first call to recursive-count-words returns one added to the value returned by calling recursive-count-words on a region containing the remaining word—that is, it adds one to one, producing two, which is the correct amount.

Similarly, if there are three words in the region, the first call to recursive-count-words returns one added to the value returned by calling recursive-count-words on a region containing the remaining two words—and so on and so on.

With full documentation the two functions look like this: The recursive function:

(defun recursive-count-words (region-end)
  "Number of words between point and REGION-END."

;;; 1. do-again-test

  (if (and (< (point) region-end)
           (re-search-forward "\\w+\\W*" region-end t))

;;; 2. then-part: the recursive call

      (1+ (recursive-count-words region-end))

;;; 3. else-part

    0))

The wrapper:

;;; Recursive version

(defun count-words-region (beginning end)
  "Print number of words in the region.
Words are defined as at least one word-constituent
character followed by at least one character that is
not a word-constituent.  The buffer's syntax table
determines which characters these are."
  (interactive "r")
  (message "Counting words in region ... ")
  (save-excursion
    (goto-char beginning)
    (let ((count (recursive-count-words end)))
      (cond ((zerop count)
             (message
              "The region does NOT have any words."))
            ((= 1 count)
             (message "The region has 1 word."))
            (t
             (message
              "The region has %d words." count))))))

13.3 Exercise: Counting Punctuation

Using a while loop, write a function to count the number of punctuation marks in a region— period, comma, semicolon, colon, exclamation mark, question mark. Do the same using recursion.

14 Counting Words in a defun

Our next project is to count the number of words in a function definition. Clearly, this can be done using some variant of count-word-region. See Chapter 13, “Counting Words: Repetition and Regexps”, page 115. If we are just going to count the words in one definition, it is easy enough to mark the definition with the C-M-h (mark-defun) command, and then call

count-word-region.

However, I am more ambitious: I want to count the words and symbols in every definition in the Emacs sources and then print a graph that shows how many functions there are of each length: how many contain 40 to 49 words or symbols, how many contain 50 to 59 words or symbols, and so on. I have often been curious how long a typical function is, and this will tell.

Described in one phrase, the histogram project is daunting; but divided into numerous small steps, each of which we can take one at a time, the project becomes less fearsome. Let us consider what the steps must be:

• First, write a function to count the words in one definition. This includes the problem of handling symbols as well as words.

• Second, write a function to list the numbers of words in each function in a file. This function

    can use the count-words-in-defun function.

• Third, write a function to list the numbers of words in each function in each of several files. This entails automatically finding the various files, switching to them, and counting the words in the definitions within them.

• Fourth, write a function to convert the list of numbers that we created in step three to a form that will be suitable for printing as a graph.

• Fifth, write a function to print the results as a graph.

This is quite a project! But if we take each step slowly, it will not be difficult.

14.1 What to Count?

When we first start thinking about how to count the words in a function definition, the first question is (or ought to be) what are we going to count? When we speak of ‘words’ with respect to a Lisp function definition, we are actually speaking, in large part, of ‘symbols’. For example, the following multiply-by-seven function contains the five symbols defun, multiply-by-seven, number, *, and 7. In addition, in the documentation string, it contains the four words ‘Multiply’, ‘NUMBER’, ‘by’, and ‘seven’. The symbol ‘number’ is repeated, so the definition contains a total of ten words and symbols.

     (defun multiply-by-seven (number)
       "Multiply NUMBER by seven."
       (* 7 number))

However, if we mark the multiply-by-seven definition with C-M-h (mark-defun), and then call count-words-region on it, we will find that count-words-region claims the definition has eleven words, not ten! Something is wrong!

The problem is twofold: count-words-region does not count the ‘*’ as a word, and it counts the single symbol, multiply-by-seven, as containing three words. The hyphens are treated as if they were interword spaces rather than intraword connectors: ‘multiply-by-seven’ is counted as if it were written ‘multiply by seven’.

The cause of this confusion is the regular expression search within the count-words-region definition that moves point forward word by word. In the canonical version of count-words-region, the regexp is:

     "\\w+\\W*"

This regular expression is a pattern defining one or more word constituent characters possibly followed by one or more characters that are not word constituents. What is meant by ‘word constituent characters’ brings us to the issue of syntax, which is worth a section of its own.

14.2 What Constitutes a Word or Symbol?

Emacs treats different characters as belonging to different syntax categories. For example, the regular expression, ‘\\w+’, is a pattern specifying one or more word constituent characters. Word constituent characters are members of one syntax category. Other syntax categories include the class of punctuation characters, such as the period and the comma, and the class of whitespace characters, such as the blank space and the tab character. (For more information, see section “The Syntax Table” in The GNU Emacs Manual, and, section “Syntax Tables” in The GNU Emacs Lisp Reference Manual.)

Syntax tables specify which characters belong to which categories. Usually, a hyphen is not specified as a ‘word constituent character’. Instead, it is specified as being in the ‘class of characters that are part of symbol names but not words.’ This means that the count-words-region function treats it in the same way it treats an interword white space, which is why

count-words-region counts ‘multiply-by-seven’ as three words.

There are two ways to cause Emacs to count ‘multiply-by-seven’ as one symbol: modify the syntax table or modify the regular expression.

We could redefine a hyphen as a word constituent character by modifying the syntax table that Emacs keeps for each mode. This action would serve our purpose, except that a hyphen is merely the most common character within symbols that is not typically a word constituent character; there are others, too.

Alternatively, we can redefine the regular expression used in the count-words definition so as to include symbols. This procedure has the merit of clarity, but the task is a little tricky.

The first part is simple enough: the pattern must match “at least one character that is a word or symbol constituent”. Thus:

     \\(\\w\\|\\s_\\)+

The ‘\\(’ is the first part of the grouping construct that includes the ‘\\w’ and the ‘\\s_’ as alternatives, separated by the ‘\\|’. The ‘\\w’ matches any word-constituent character and the ‘\\s_’ matches any character that is part of a symbol name but not a word-constituent character. The ‘+’ following the group indicates that the word or symbol constituent characters must be matched at least once.

However, the second part of the regexp is more difficult to design. What we want is to follow the first part with “optionally one or more characters that are not constituents of a word or symbol”. At first, I thought I could define this with the following:

     \\(\\W\\|\\S_\\)*"

The upper case ‘W’ and ‘S’ match characters that are not word or symbol constituents. Unfortunately, this expression matches any character that is either not a word constituent or not a symbol constituent. This matches any character!

I then noticed that every word or symbol in my test region was followed by white space (blank space, tab, or newline). So I tried placing a pattern to match one or more blank spaces after the pattern for one or more word or symbol constituents. This failed, too. Words and symbols are often separated by whitespace, but in actual code parentheses may follow symbols and punctuation may follow words. So finally, I designed a pattern in which the word or symbol

The count-words-in-defun Function  127

constituents are followed optionally by characters that are not white space and then followed optionally by white space.

Here is the full regular expression:

     "\\(\\w\\|\\s_\\)+[^\t\n]*[ \t\n]*"

14.3 The count-words-in-defun Function

We have seen that there are several ways to write a count-word-region function. To write a count-words-in-defun, we need merely adapt one of these versions.

The version that uses a while loop is easy to understand, so I am going to adapt that. Because count-words-in-defun will be part of a more complex program, it need not be interactive and it need not display a message but just return the count. These considerations simplify the definition a little.

On the other hand, count-words-in-defun will be used within a buffer that contains function definitions. Consequently, it is reasonable to ask that the function determine whether it is called when point is within a function definition, and if it is, to return the count for that definition. This adds complexity to the definition, but saves us from needing to pass arguments to the function.

These considerations lead us to prepare the following template:

     (defun count-words-in-defun ()

"documentation..." (set up...

(while loop...) return count)

As usual, our job is to fill in the slots.

First, the set up.

We are presuming that this function will be called within a buffer containing function definitions. Point will either be within a function definition or not. For count-words-in-defun to work, point must move to the beginning of the definition, a counter must start at zero, and the counting loop must stop when point reaches the end of the definition.

The beginning-of-defun function searches backwards for an opening delimiter such as a ‘(’ at the beginning of a line, and moves point to that position, or else to the limit of the search. In practice, this means that beginning-of-defun moves point to the beginning of an enclosing or preceding function definition, or else to the beginning of the buffer. We can use beginning-of-defun to place point where we wish to start.

The while loop requires a counter to keep track of the words or symbols being counted. A let expression can be used to create a local variable for this purpose, and bind it to an initial value of zero.

The end-of-defun function works like beginning-of-defun except that it moves point to the end of the definition. end-of-defun can be used as part of an expression that determines the position of the end of the definition.

The set up for count-words-in-defun takes shape rapidly: first we move point to the beginning of the definition, then we create a local variable to hold the count, and, finally, we record the position of the end of the definition so the while loop will know when to stop looping.

The code looks like this:

     (beginning-of-defun)
     (let ((count 0)
           (end (save-excursion (end-of-defun) (point))))

The code is simple. The only slight complication is likely to concern end: it is bound to the position of the end of the definition by a save-excursion expression that returns the value of point after end-of-defun temporarily moves it to the end of the definition.

The second part of the count-words-in-defun, after the set up, is the while loop.

The loop must contain an expression that jumps point forward word by word and symbol by symbol, and another expression that counts the jumps. The true-or-false-test for the while loop should test true so long as point should jump forward, and false when point is at the end of the definition. We have already redefined the regular expression for this (see Section 14.2, “Syntax”, page 126), so the loop is straightforward:

     (while (and (< (point) end)
                 (re-search-forward
                  "\\(\\w\\|\\s_\\)+[^\t\n]*[ \t\n]*" end t)
       (setq count (1+ count)))

The third part of the function definition returns the count of words and symbols. This part is the last expression within the body of the let expression, and can be, very simply, the local variable count, which when evaluated returns the count.

Put together, the count-words-in-defun definition looks like this:

     (defun count-words-in-defun ()
       "Return the number of words and symbols in a defun."
       (beginning-of-defun)
       (let ((count 0)
             (end (save-excursion (end-of-defun) (point))))
         (while
             (and (< (point) end)
                  (re-search-forward
                   "\\(\\w\\|\\s_\\)+[^\t\n]*[ \t\n]*"
                   end t))
           (setq count (1+ count)))
         count))

How to test this? The function is not interactive, but it is easy to put a wrapper around the function to make it interactive; we can use almost the same code as for the recursive version of

count-words-region:

;;; Interactive version.

     (defun count-words-defun ()
       "Number of words and symbols in a function definition."
       (interactive)
       (message
        "Counting words and symbols in function definition ... ")

       (let ((count (count-words-in-defun)))
         (cond
          ((zerop count)
           (message
            "The definition does NOT have any words or symbols."))
          ((= 1 count)
           (message
            "The definition has 1 word or symbol."))
          (t
           (message
            "The definition has %d words or symbols." count)))))

Let’s re-use C-c = as a convenient keybinding:

     (global-set-key "\C-c=" 'count-words-defun)
   Now we can try out count-words-defun: install both count-words-in-defun and count-

words-defun, and set the keybinding, and then place the cursor within the following definition:

     (defun multiply-by-seven (number)
       "Multiply NUMBER by seven."
       (* 7 number))
          ⇒ 10

Success! The definition has 10 words and symbols.

The next problem is to count the numbers of words and symbols in several definitions within a single file.

14.4 Count Several defuns Within a File

A file such as ‘simple.el’ may have 80 or more function definitions within it. Our long term goal is to collect statistics on many files, but as a first step, our immediate goal is to collect statistics on one file.

The information will be a series of numbers, each number being the length of a function definition. We can store the numbers in a list.

We know that we will want to incorporate the information regarding one file with information about many other files; this means that the function for counting definition lengths within one file need only return the lengths’ list. It need not and should not display any messages.

The word count commands contain one expression to jump point forward word by word and another expression to count the jumps. The definitions’ lengths’ function can be designed to work the same way, with one expression to jump point forward definition by definition and another expression to construct the lengths’ list.

This statement of the problem makes it elementary to write the function definition. Clearly, we will start the count at the beginning of the file, so the first command will be (goto-char (point-min)). Next, we start the while loop; and the true-or-false test of the loop can be a regular expression search for the next function definition—so long as the search succeeds, point is moved forward and then the body of the loop is evaluated. The body needs an expression that constructs the lengths’ list. cons, the list construction command, can be used to create the list. That is almost all there is to it.

Here is what this fragment of code looks like:

     (goto-char (point-min))
     (while (re-search-forward "^(defun" nil t)
       (setq lengths-list
             (cons (count-words-in-defun) lengths-list)))

What we have left out is the mechanism for finding the file that contains the function definitions.

In previous examples, we either used this, the Info file, or we switched back and forth to some other buffer, such as the ‘*scratch*’ buffer.

Finding a file is a new process that we have not yet discussed.

14.5 Find a File

To find a file in Emacs, you use the C-x C-f (find-file) command. This command is almost, but not quite right for the lengths problem.

Let’s look at the source for find-file (you can use the find-tag command to find the source of a function):

     (defun find-file (filename)
       "Edit file FILENAME.
     Switch to a buffer visiting file FILENAME,
     creating one if none already exists."
       (interactive "FFind file: ")
       (switch-to-buffer (find-file-noselect filename)))

The definition possesses short but complete documentation and an interactive specification that prompts you for a file name when you use the command interactively. The body of the

definition contains two functions, find-file-noselect and switch-to-buffer.

According to its documentation as shown by C-h f (the describe-function command), the find-file-noselect function reads the named file into a buffer and returns the buffer. However, the buffer is not selected. Emacs does not switch its attention (or yours if you are using find-file-noselect) to the named buffer. That is what switch-to-buffer does: it switches the buffer to which Emacs attention is directed; and it switches the buffer displayed in the window to the new buffer. We have discussed buffer switching elsewhere. (See Section 2.3, “Switching Buffers”, page 17.)

In this histogram project, we do not need to display each file on the screen as the program determines the length of each definition within it. Instead of employing switch-to-buffer, we can work with set-buffer, which redirects the attention of the computer program to a different buffer but does not redisplay it on the screen. So instead of calling on find-file to do the job, we must write our own expression.

   The task is easy: use find-file-noselect and set-buffer.

14.6 lengths-list-file in Detail

The core of the lengths-list-file function is a while loop containing a function to move point forward ‘defun by defun’ and a function to count the number of words and symbols in each defun. This core must be surrounded by functions that do various other tasks, including finding the file, and ensuring that point starts out at the beginning of the file. The function definition looks like this:

lengths-list-file in Detail  131

     (defun lengths-list-file (filename)
       "Return list of definitions' lengths within FILE.
     The returned list is a list of numbers.
     Each number is the number of words or
     symbols in one function definition."
       (message "Working on `%s' ... " filename)
       (save-excursion
         (let ((buffer (find-file-noselect filename))
               (lengths-list))
           (set-buffer buffer)
           (setq buffer-read-only t)
           (widen)
           (goto-char (point-min))
           (while (re-search-forward "^(defun" nil t)
             (setq lengths-list
                   (cons (count-words-in-defun) lengths-list)))
           (kill-buffer buffer)
           lengths-list)))

The function is passed one argument, the name of the file on which it will work. It has four lines of documentation, but no interactive specification. Since people worry that a computer is broken if they don’t see anything going on, the first line of the body is a message.

The next line contains a save-excursion that returns Emacs attention to the current buffer when the function completes. This is useful in case you embed this function in another function that presumes point is restored to the original buffer.

In the varlist of the let expression, Emacs finds the file and binds the local variable buffer to the buffer containing the file. At the same time, Emacs creates lengths-list as a local variable.

Next, Emacs switches its attention to the buffer.

In the following line, Emacs makes the buffer read-only. Ideally, this line is not necessary. None of the functions for counting words and symbols in a function definition should change the buffer. Besides, the buffer is not going to be saved, even if it were changed. This line is entirely the consequence of great, perhaps excessive, caution. The reason for the caution is that this function and those it calls work on the sources for Emacs and it is very inconvenient if they are inadvertently modified. It goes without saying that I did not realize a need for this line until an experiment went awry and started to modify my Emacs source files …

Next comes a call to widen the buffer if it is narrowed. This function is usually not needed— Emacs creates a fresh buffer if none already exists; but if a buffer visiting the file already exists Emacs returns that one. In this case, the buffer may be narrowed and must be widened. If we wanted to be fully ‘user-friendly’, we would arrange to save the restriction and the location of point, but we won’t.

The (goto-char (point-min)) expression moves point to the beginning of the buffer.

Then comes a while loop in which the ‘work’ of the function is carried out. In the loop, Emacs determines the length of each definition and constructs a lengths’ list containing the information.

Emacs kills the buffer after working through it. This is to save space inside of Emacs. My version of Emacs 19 contains over 300 source files of interest. Another function will apply lengths-list-file to each of them. If Emacs visits all of them and deletes none, my computer may run out of virtual memory.

Finally, the last expression within the let expression is the lengths-list variable; its value is returned as the value of the whole function.

You can try this function by installing it in the usual fashion. Then place your cursor after the following expression and type C-x C-e (eval-last-sexp).

     (lengths-list-file "../lisp/debug.el")

(You may need to change the pathname of the file; the one here works if this Info file and the Emacs sources are in neighboring places, such as /usr/local/emacs/info and /usr/local/emacs/lisp. To change the expression, copy it to the ‘*scratch*’ buffer and edit it. Then evaluate it.)

On my version of Emacs, the lengths’ list for ‘debug.el’ takes seven seconds to produce and looks like this:

     (75 41 80 62 20 45 44 68 45 12 34 235)

Note that the length of the last definition in the file is first in the list.

14.7 Count Words in defuns in Different Files

In the previous section, we created a function that returns a list of the lengths of each definition in a file. Now, we want to define a function to return a master list of the lengths of the definitions in a list of files.

Working on each of a list of files is a repetitious act, so we can use either a while loop or recursion.

The design using a while loop is routine. The argument passed the function is a list of files. As we saw earlier (see Section 11.1.1, “Loop Example”, page 87), you can write a while loop so that the body of the loop is evaluated if such a list contains elements, but to exit the loop if the list is empty. For this design to work, the body of the loop must contain an expression that shortens the list each time the body is evaluated, so that eventually the list is empty. The usual technique is to set the value of the list to the value of the cdr of the list each time the body is evaluated.

The template looks like this:

(while test-whether-list-is-empty body... set-list-to-cdr-of-list)

Also, we remember that a while loop returns nil (the result of evaluating the true-or-false-test), not the result of any evaluation within its body. (The evaluations within the body of the loop are done for their side effects.) However, the expression that sets the lengths’ list is part of the body—and that is the value that we want returned by the function as a whole. To do this, we enclose the while loop within a let expression, and arrange that the last element of the let expression contains the value of the lengths’ list. (See “Loop Example with an Incrementing Counter”, page 90.)

These considerations lead us directly to the function itself:

     ;;; Use while loop.
     (defun lengths-list-many-files (list-of-files)
       "Return list of lengths of defuns in LIST-OF-FILES."

       (let (lengths-list)

;;; true-or-false-test

         (while list-of-files
           (setq lengths-list
                 (append
                  lengths-list

;;; Generate a lengths’ list.

                  (lengths-list-file
                   (expand-file-name (car list-of-files)))))

;;; Make files’ list shorter.

           (setq list-of-files (cdr list-of-files)))

;;; Return final value of lengths’ list.

         lengths-list))

expand-file-name is a built-in function that converts a file name to its absolute, long, path name form. Thus,

     debug.el

becomes

     /usr/local/emacs/lisp/debug.el

The only other new element of this function definition is the as yet unstudied function append, which merits a short section for itself.

14.7.1 The append Function

The append function attaches one list to another. Thus,

     (append '(1 2 3 4) '(5 6 7 8))

produces the list

     (1 2 3 4 5 6 7 8)

This is exactly how we want to attach two lengths’ lists produced by lengths-list-file to each other. The results contrast with cons,

     (cons '(1 2 3 4) '(5 6 7 8))

which constructs a new list in which the first argument to cons becomes the first element of the new list:

     ((1 2 3 4) 5 6 7 8)

14.8 Recursively Count Words in Different Files

Besides a while loop, you can work on each of a list of files with recursion. A recursive version of lengths-list-many-files is short and simple.

The recursive function has the usual parts: the ‘do-again-test’, the ‘next-step-expression’, and the recursive call. The ‘do-again-test’ determines whether the function should call itself again, which it will do if the list-of-files contains any remaining elements; the ‘next-step-expression’ resets the list-of-files to the cdr of itself, so eventually the list will be empty; and the recursive call calls itself on the shorter list. The complete function is shorter than this description!

     (defun recursive-lengths-list-many-files (list-of-files)
       "Return list of lengths of each defun in LIST-OF-FILES."
       (if list-of-files                     ; do-again-test
           (append
            (lengths-list-file
             (expand-file-name (car list-of-files)))
            (recursive-lengths-list-many-files
             (cdr list-of-files)))))

In a sentence, the function returns the lengths’ list for the first of the list-of-files appended to the result of calling itself on the rest of the list-of-files.

Here is a test of recursive-lengths-list-many-files, along with the results of running lengths-list-file on each of the files individually.

   Install recursive-lengths-list-many-files and lengths-list-file, if necessary, and

then evaluate the following expressions. You may need to change the files’ pathnames; those here work when this Info file and the Emacs sources are located in their customary places. To change the expressions, copy them to the ‘*scratch*’ buffer, edit them, and then evaluate them.

The results are shown after the ‘⇒’. (These results are for files from Emacs Version 18.57; files from other versions of Emacs may produce different results.)

     (lengths-list-file
      "../lisp/macros.el")
          ⇒ (176 154 86)
     (lengths-list-file
      "../lisp/mailalias.el")
          ⇒ (116 122 265)
     (lengths-list-file
      "../lisp/makesum.el")
          ⇒ (85 179)
     (recursive-lengths-list-many-files
      '("../lisp/macros.el"
        "../lisp/mailalias.el"
        "../lisp/makesum.el"))
            ⇒ (176 154 86 116 122 265 85 179)

The recursive-lengths-list-many-files function produces the output we want.

The next step is to prepare the data in the list for display in a graph.

14.9 Prepare the Data for Display in a Graph

The recursive-lengths-list-many-files function returns a list of numbers. Each number records the length of a function definition. What we need to do now is transform this data into a list of numbers suitable for generating a graph. The new list will tell how many functions definitions contain less than 10 words and symbols, how many contain between 10 and 19 words and symbols, how many contain between 20 and 29 words and symbols, and so on.

In brief, we need to go through the lengths’ list produced by the recursive-lengths-list-many-files function and count the number of defuns within each range of lengths, and produce a list of those numbers.

Based on what we have done before, we can readily foresee that it should not be too hard to write a function that ‘cdrs’ down the lengths’ list, looks at each element, determines which length range it is in, and increments a counter for that range.

However, before beginning to write such a function, we should consider the advantages of sorting the lengths’ list first, so the numbers are ordered from smallest to largest. First, sorting will make it easier to count the numbers in each range, since two adjacent numbers will either be in the same length range or in adjacent ranges. Second, by inspecting a sorted list, we can discover the highest and lowest number, and thereby determine the largest and smallest length range that we will need.

14.9.1 Sorting Lists

Emacs contains a function to sort lists, called (as you might guess) sort. The sort function takes two arguments, the list to be sorted, and a predicate that determines whether the first of two list elements is “less” than the second.

As we saw earlier (see Section 1.8.4, “Using the Wrong Type Object as an Argument”, page 10), a predicate is a function that determines whether some property is true or false. The sort function will reorder a list according to whatever property the predicate uses; this means that sort can be used to sort non-numeric lists by non-numeric criteria—it can, for example, alphabetize a list.

The < function is used when sorting a numeric list. For example,

     (sort '(4 8 21 17 33 7 21 7) '<)

produces this:

     (4 7 7 8 17 21 21 33)

(Note that in this example, both the arguments are quoted so that the symbols are not evaluated before being passed to sort as arguments.)

Sorting the list returned by the recursive-lengths-list-many-files function is straightforward:

     (sort
      (recursive-lengths-list-many-files
       '("../lisp/macros.el"
         "../lisp/mailalias.el"
         "../lisp/makesum.el"))
      '<)

which produces:

     (85 86 116 122 154 176 179 265)

(Note that in this example, the first argument to sort is not quoted, since the expression must be evaluated so as to produce the list that is passed to sort.)

14.9.2 Making a List of Files

The recursive-lengths-list-many-files function requires a list of files as its argument. For our test examples, we constructed such a list by hand; but the Emacs Lisp source directory is too large for us to do for that. Instead, we need to use the directory-files function to construct a list for us.

The directory-files function takes three arguments: the first argument is the name of a directory, a string; a non-nil second argument causes the function to return the files’ absolute pathnames; and the third argument is a selector. If it contains a regular expression (rather than nil), only pathnames that match that regular expression are returned.

Thus, on my system,

     (length
      (directory-files "../lisp" t "\\.el$"))

tells me that my version 19.25 Lisp sources directory contains 307 ‘.el’ files.

An expression to sort the list returned by recursive-lengths-list-many-files looks like this:

     (sort
      (recursive-lengths-list-many-files
       (directory-files "../lisp" t "\\.el$"))
      '<)

Our immediate goal is to generate a list that tells us how many function definitions contain fewer than 10 words and symbols, how many contain between 10 and 19 words and symbols, how many contain between 20 and 29 words and symbols, and so on. With a sorted list of numbers, this is easy: count how many elements of the list are smaller than 10, then, after moving past the numbers just counted, count how many are smaller than 20, then, after moving past the numbers just counted, count how many are smaller than 30, and so on. Each of the numbers, 10, 20, 30, 40, and the like, is one larger than the top of that range. We can call the list of such numbers the top-of-ranges list.

If we wanted to, we could generate this list automatically, but it is simpler to write a list manually. Here it is:

     (defvar top-of-ranges
      '(10  20  30  40  50
        60  70  80  90 100
       110 120 130 140 150
       160 170 180 190 200
       210 220 230 240 250
       260 270 280 290 300)
      "List specifying ranges for `defuns-per-range'.")

To change the ranges, we edit this list.

Next, we need to write the function that creates the list of the number of definitions within each range. Clearly, this function must take the sorted-lengths and the top-of-ranges lists as arguments.

The defuns-per-range function must do two things again and again: it must count the number of definitions within a range specified by the current top-of-range value; and it must shift to the next higher value in the top-of-ranges list after counting the number of definitions in the current range. Since each of these actions is repetitive, we can use while loops for the job. One loop counts the number of definitions in the range defined by the current top-of-range value, and the other loop selects each of the top-of-range values in turn.

Several entries of the sorted-lengths list are counted for each range; this means that the loop for the sorted-lengths list will be inside the loop for the top-of-ranges list, like a small gear inside a big gear.

The inner loop counts the number of definitions within the range. It is a simple counting loop of the type we have seen before. (See Section 11.1.3, “A loop with an incrementing counter”, page 89.) The true-or-false test of the loop tests whether the value from the sorted-lengths list is smaller than the current value of the top of the range. If it is, the function increments the counter and tests the next value from the sorted-lengths list.

The inner loop looks like this:

(while length-element-smaller-than-top-of-range

       (setq number-within-range (1+ number-within-range))
       (setq sorted-lengths (cdr sorted-lengths)))

The outer loop must start with the lowest value of the top-of-ranges list, and then be set to each of the succeeding higher values in turn. This can be done with a loop like this:

     (while top-of-ranges

body-of-loop...

       (setq top-of-ranges (cdr top-of-ranges)))

Put together, the two loops look like this:

     (while top-of-ranges

;; Count the number of elements within the current range. (while length-element-smaller-than-top-of-range

         (setq number-within-range (1+ number-within-range))
         (setq sorted-lengths (cdr sorted-lengths)))

;; Move to next range.

       (setq top-of-ranges (cdr top-of-ranges)))

In addition, in each circuit of the outer loop, Emacs should record the number of definitions within that range (the value of number-within-range) in a list. We can use cons for this purpose. (See Section 7.2, “cons”, page 60.)

The cons function works fine, except that the list it constructs will contain the number of definitions for the highest range at its beginning and the number of definitions for the lowest range at its end. This is because cons attaches new elements of the list to the beginning of the list, and since the two loops are working their way through the lengths’ list from the lower end first, the defuns-per-range-list will end up largest number first. But we will want to print our graph with smallest values first and the larger later. The solution is to reverse the order of the defuns-per-range-list. We can do this using the nreverse function, which reverses the order of a list.

For example,

     (nreverse '(1 2 3 4))

produces:

     (4 3 2 1)

Note that the nreverse function is “destructive”—that is, it changes the list to which it is applied; this contrasts with the car and cdr functions, which are non-destructive. In this case, we do not want the original defuns-per-range-list, so it does not matter that it is destroyed. (The reverse function provides a reversed copy of a list, leaving the original list as is.)

Put all together, the defuns-per-range looks like this:

(defun defuns-per-range (sorted-lengths top-of-ranges)
"SORTED-LENGTHS defuns in each TOP-OF-RANGES range."
(let ((top-of-range (car top-of-ranges))
  (number-within-range 0)
  defuns-per-range-list)

;; Outer loop.

(while top-of-ranges

;; Inner loop.

(while (and

;; Need number for numeric test.

        (car sorted-lengths)
        (< (car sorted-lengths) top-of-range))

;; Count number of definitions within current range.

  (setq number-within-range (1+ number-within-range))
  (setq sorted-lengths (cdr sorted-lengths)))

;; Exit inner loop but remain within outer loop.

(setq defuns-per-range-list
      (cons number-within-range defuns-per-range-list))
(setq number-within-range 0)      ; Reset count to zero.

;; Move to next range.

(setq top-of-ranges (cdr top-of-ranges))

;; Specify next top of range value.

(setq top-of-range (car top-of-ranges)))

;; Exit outer loop and count the number of defuns larger than ;; the largest top-of-range value.

(setq defuns-per-range-list
    (cons
     (length sorted-lengths)
     defuns-per-range-list))

;; Return a list of the number of definitions within each range, ;; smallest to largest.

(nreverse defuns-per-range-list)))

The function is straightforward except for one subtle feature. The true-or-false test of the inner loop looks like this:

(and (car sorted-lengths)
(< (car sorted-lengths) top-of-range))

instead of like this:

(< (car sorted-lengths) top-of-range)

The purpose of the test is to determine whether the first item in the sorted-lengths list is less than the value of the top of the range.

The simple version of the test works fine unless the sorted-lengths list has a nil value. In that case, the (car sorted-lengths) expression function returns nil. The < function cannot compare a number to nil, which is an empty list, so Emacs signals an error and stops the function from attempting to continue to execute.

The sorted-lengths list always becomes nil when the counter reaches the end of the list. This means that any attempt to use the defuns-per-range function with the simple version of the test will fail.

We solve the problem by using the (car sorted-lengths) expression in conjunction with the and expression. The (car sorted-lengths) expression returns a non-nil value so long as the list has at least one number within it, but returns nil if the list is empty. The and expression first evaluates the (car sorted-lengths) expression, and if it is nil, returns false without evaluating the < expression. But if the (car sorted-lengths) expression returns a non-nil value, the and expression evaluates the < expression, and returns that value as the value of the and expression.

This way, we avoid an error. See Section 12.4, “forward-paragraph: a Goldmine of Functions”, page 105, for more information about and.

Here is a short test of the defuns-per-range function. First, evaluate the expression that binds (a shortened) top-of-ranges list to the list of values, then evaluate the expression for binding the sorted-lengths list, and then evaluate the defuns-per-range function.

;; (Shorter list than we will use later.)

     (setq top-of-ranges
      '(110 120 130 140 150
        160 170 180 190 200))
     (setq sorted-lengths
           '(85 86 110 116 122 129 154 176 179 200 265 300 300))
     (defuns-per-range sorted-lengths top-of-ranges)

The list returned looks like this:

     (2 2 2 0 0 1 0 2 0 0 4)

Indeed, there are two elements of the sorted-lengths list smaller than 110, two elements between 110 and 119, two elements between 120 and 129, and so on. There are four elements with a value of 200 or larger.

140 Chapter 14: Counting Words in a defun

15 Readying a Graph

Our goal is to construct a graph showing the numbers of function definitions of various lengths in the Emacs lisp sources.

As a practical matter, if you were creating a graph, you would probably use a program such as gnuplot to do the job. (gnuplot is nicely integrated into GNU Emacs.) In this case, however, we create one from scratch, and in the process we will reaquaint ourselves with some of what we learned before and learn more.

In this chapter, we will first write a simple graph printing function. This first definition will be a prototype, a rapidly written function that enables us to reconnoiter this unknown graph-making territory. We will discover dragons, or find that they are myth. After scouting the terrain, we will feel more confident and enhance the function to label the axes automatically.

Since Emacs is designed to be flexible and work with all kinds of terminals, including character-only terminals, the graph will need to be made from one of the ‘typewriter’ symbols. An asterisk will do; as we enhance the graph-printing function, we can make the choice of symbol a user option.

We can call this function graph-body-print; it will take a numbers-list as its only argument. At this stage, we will not label the graph, but only print its body.

The graph-body-print function inserts a vertical column of asterisks for each element in the numbers-list. The height of each line is determined by the value of that element of the

numbers-list.

Inserting columns is a repetitive act; that means that this function can be written either with a while loop or recursively.

Our first challenge is to discover how to print a column of asterisks. Usually, in Emacs, we print characters onto a screen horizontally, line by line, by typing. We have two routes we can follow: write our own column-insertion function or discover whether one exists in Emacs.

To see whether there is one in Emacs, we can use the M-x apropos command. This command is like the C-h a (command-apropos) command, except that the latter finds only those functions that are commands. The M-x apropos command lists all symbols that match a regular expression, including functions that are not interactive.

What we want to look for is some command that prints or inserts columns. Very likely, the name of the function will contain either the word ‘print’ or the word ‘insert’ or the word

‘column’. Therefore, we can simply type M-x apropos RET print\|insert\|column RET and

look at the result. On my system, this command takes quite some time, and then produces a list of 79 functions and variables. Scanning down the list, the only function that looks as if it might do the job is insert-rectangle. Indeed, this is the function we want; its documentation says:

     insert-rectangle:
     Insert text of RECTANGLE with upper left corner at point.
     RECTANGLE's first line is inserted at point,
     its second line is inserted at a point vertically under point, etc.
     RECTANGLE should be a list of strings.

We can run a quick test, to make sure it does what we expect of it.

Here is the result of placing the cursor after the insert-rectangle expression and typing C-u C-x C-e (eval-last-sexp). The function inserts the strings ‘"first"’, ‘"second"’, and ‘"third"’ at and below point. Also the function returns nil.

     (insert-rectangle '("first" "second" "third"))first
                                                   second
                                                   third
     nil

Of course, we won’t be inserting the text of the insert-rectangle expression itself into the buffer in which we are making the graph, but will call the function from our program. We shall, however, have to make sure that point is in the buffer at the place where the insert-rectangle function will insert its column of strings.

If you are reading this in Info, you can see how this works by switching to another buffer, such as the ‘*scratch*’ buffer, placing point somewhere in the buffer, typing M-:, typing the insert-rectangle expression into the minibuffer at the prompt, and then typing ^{⟨}\underline{\overline{RET}}^{⟩}. This causes Emacs to evaluate the expression in the minibuffer, but to use as the value of point the position of point in the ‘*scratch*’ buffer. (M-: is the keybinding for eval-expression.)

We find when we do this that point ends up at the end of the last inserted line—that is to say, this function moves point as a side-effect. If we were to repeat the command, with point at this position, the next insertion would be below and to the right of the previous insertion. We don’t want this! If we are going to make a bar graph, the columns need to be beside each other.

So we discover that each cycle of the column-inserting while loop must reposition point to the place we want it, and that place will be at the top, not the bottom, of the column. Moreover, we remember that when we print a graph, we do not expect all the columns to be the same height. This means that the top of each column may be at a different height from the previous one. We cannot simply reposition point to the same line each time, but moved over to the right—or perhaps we can…

We are planning to make the columns of the bar graph out of asterisks. The number of asterisks in the column is the number specified by the current element of the numbers-list. We need to construct a list of asterisks of the right length for each call to insert-rectangle. If this list consists solely of the requisit number of asterisks, then we will have position point the right number of lines above the base for the graph to print correctly. This could be difficult.

Alternatively, if we can figure out some way to pass insert-rectangle a list of the same length each time, then we can place point on the same line each time, but move it over one column to the right for each new column. If we do this, however, some of the entries in the list passed to insert-rectangle must be blanks rather than asterisks. For example, if the maximum height of the graph is 5, but the height of the column is 3, then insert-rectangle requires an argument that looks like this:

     (" " " " "*" "*" "*")

This last proposal is not so difficult, so long as we can determine the column height. There are two ways for us to specify the column height: we can arbitrarily state what it will be, which would work fine for graphs of that height; or we can search through the list of numbers and use the maximum height of the list as the maximum height of the graph. If the latter operation were difficult, then the former procedure would be easiest, but there is a function built into Emacs that determines the maximum of its arguments. We can use that function. The function is called max and it returns the largest of all its arguments, which must be numbers. Thus, for example,

     (max  3 4 6 5 7 3)

returns 7. (A corresponding function called min returns the smallest of all its arguments.)

However, we cannot simply call max on the numbers-list; the max function expects numbers as its argument, not a list of numbers. Thus, the following expression,

     (max  '(3 4 6 5 7 3))

produces the following error message;

     Wrong type of argument:  integer-or-marker-p, (3 4 6 5 7 3)

We need a function that passes a list of arguments to a function. This function is apply. This function ‘applies’ its first argument (a function) to its remaining arguments, the last of which may be a list.

For example,

     (apply 'max 3 4 7 3 '(4 8 5))

returns 8.

(Incidentally, I don’t know how you would learn of this function without a book such as this. It is possible to discover other functions, like search-forward or insert-rectangle, by guessing at a part of their names and then using apropos. Even though its base in metaphor is clear—‘apply’ its first argument to the rest—I doubt a novice would come up with that particular word when using apropos or other aid. Of course, I could be wrong; after all, the function was first named by someone who had to invent it.)

The second and subsequent arguments to apply are optional, so we can use apply to call a function and pass the elements of a list to it, like this, which also returns 8:

     (apply 'max '(4 8 5))
   This latter way is how we will use apply. The recursive-lengths-list-many-files func-

tion returns a numbers’ list to which we can apply max (we could also apply max to the sorted numbers’ list; it does not matter whether the list is sorted or not.)

Hence, the operation for finding the maximum height of the graph is this:

     (setq max-graph-height (apply 'max numbers-list))

Now we can return to the question of how to create a list of strings for a column of the graph. Told the maximum height of the graph and the number of asterisks that should appear in the column, the function should return a list of strings for the insert-rectangle command to insert.

Each column is made up of asterisks or blanks. Since the function is passed the value of the height of the column and the number of asterisks in the column, the number of blanks can be found by subtracting the number of asterisks from the height of the column. Given the number of blanks and the number of asterisks, two while loops can be used to construct the list:

;;; First version.

     (defun column-of-graph (max-graph-height actual-height)
       "Return list of strings that is one column of a graph."
       (let ((insert-list nil)
             (number-of-top-blanks
              (- max-graph-height actual-height)))

;; Fill in asterisks.

         (while (> actual-height 0)
           (setq insert-list (cons "*" insert-list))
           (setq actual-height (1- actual-height)))

;; Fill in blanks.

         (while (> number-of-top-blanks 0)
           (setq insert-list (cons " " insert-list))
           (setq number-of-top-blanks
                 (1- number-of-top-blanks)))

;; Return whole list.

         insert-list))

If you install this function and then evaluate the following expression you will see that it returns the list as desired:

     (column-of-graph 5 3)

returns

     (" " " " "*" "*" "*")

As written, column-of-graph contains a major flaw: the symbols used for the blank and for the marked entries in the column are ‘hard-coded’ as a space and asterisk. This is fine for a prototype, but you, or another user, may wish to use other symbols. For example, in testing the graph function, you many want to use a period in place of the space, to make sure the point is being repositioned properly each time the insert-rectangle function is called; or you might want to substitute a ‘+’ sign or other symbol for the asterisk. You might even want to make a graph-column that is more than one display column wide. The program should be more flexible. The way to do that is to replace the blank and the asterisk with two variables that we can call graph-blank and graph-symbol and define those variables separately.

Also, the documentation is not well written. These considerations lead us to the second version of the function:

     (defvar graph-symbol "*"
       "String used as symbol in graph, usually an asterisk.")
     (defvar graph-blank " "
       "String used as blank in graph, usually a blank space.
     graph-blank must be the same number of columns wide
     as graph-symbol.")

(For an explanation of defvar, see Section 8.4, “Initializing a Variable with defvar”, page 72.)

;;; Second version.

     (defun column-of-graph (max-graph-height actual-height)
       "Return list of MAX-GRAPH-HEIGHT strings;
     ACTUAL-HEIGHT are graph-symbols.
     The graph-symbols are contiguous entries at the end
     of the list.
     The list will be inserted as one column of a graph.
     The strings are either graph-blank or graph-symbol."
       (let ((insert-list nil)
             (number-of-top-blanks
              (- max-graph-height actual-height)))
         ;; Fill in graph-symbols.
         (while (> actual-height 0)
           (setq insert-list (cons graph-symbol insert-list))
           (setq actual-height (1- actual-height)))

The graph-body-print Function  145

         ;; Fill in graph-blanks.
         (while (> number-of-top-blanks 0)
           (setq insert-list (cons graph-blank insert-list))
           (setq number-of-top-blanks
                 (1- number-of-top-blanks)))

;; Return whole list.

         insert-list))

If we wished, we could rewrite column-of-graph a third time to provide optionally for a line graph as well as for a bar graph. This would not be hard to do. One way to think of a line graph is that it is no more than a bar graph in which the part of each bar that is below the top is blank. To construct a column for a line graph, the function first constructs a list of blanks that is one shorter than the value, then it uses cons to attach a graph symbol to the list; then it uses cons again to attach the ‘top blanks’ to the list.

It is easy to see how to write such a function, but since we don’t need it, we will not do it. But the job could be done, and if it were done, it would be done with column-of-graph. Even more important, it is worth noting that few changes would have to be made anywhere else. The enhancement, if we ever wish to make it, is simple.

Now, finally, we come to our first actual graph printing function. This prints the body of a graph, not the labels for the vertical and horizontal axes, so we can call this graph-body-print.

15.1 The graph-body-print Function

After our preparation in the preceding section, the graph-body-print function is straightforward. The function will print column after column of asterisks and blanks, using the elements of a numbers’ list to specify the number of asterisks in each column. This is a repetitive act, which means we can use a decrementing while loop or recursive function for the job. In this section, we will write the definition using a while loop.

The column-of-graph function requires the height of the graph as an argument, so we should determine and record that as a local variable.

This leads us to the following template for the while loop version of this function:

     (defun graph-body-print (numbers-list)

"documentation..."

       (let ((height  ...
              ...))
         (while numbers-list

insert-columns-and-reposition-point

           (setq numbers-list (cdr numbers-list)))))

We need to fill in the slots of the template.

Clearly, we can use the (apply 'max numbers-list) expression to determine the height of the graph.

The while loop will cycle through the numbers-list one element at a time. As it is shortened by the (setq numbers-list (cdr numbers-list)) expression, the car of each instance of the list is the value of the argument for column-of-graph.

At each cycle of the while loop, the insert-rectangle function inserts the list returned by column-of-graph. Since the insert-rectangle function moves point to the lower right of the inserted rectangle, we need to save the location of point at the time the rectangle is inserted, move back to that position after the rectangle is inserted, and then move horizontally to the next place from which insert-rectangle is called.

If the inserted columns are one character wide, as they will be if single blanks and asterisks are used, the repositioning command is simply (forward-char 1); however, the width of a column may be greater than one. This means that the repositioning command should be written

(forward-char symbol-width). The symbol-width itself is the length of a graph-blank and

can be found using the expression (length graph-blank). The best place to bind the symbolwidth variable to the value of the width of graph column is in the varlist of the let expression.

These considerations lead to the following function definition:

     (defun graph-body-print (numbers-list)
       "Print a bar graph of the NUMBERS-LIST.
     The numbers-list consists of the Y-axis values."
       (let ((height (apply 'max numbers-list))
             (symbol-width (length graph-blank))
             from-position)
         (while numbers-list
           (setq from-position (point))
           (insert-rectangle
            (column-of-graph height (car numbers-list)))
           (goto-char from-position)
           (forward-char symbol-width)

;; Draw graph column by column.

           (sit-for 0)
           (setq numbers-list (cdr numbers-list)))

;; Place point for X axis labels.

         (forward-line height)
         (insert "\n")
     ))

The one unexpected expression in this function is the (sit-for 0) expression in the while loop. This expression makes the graph printing operation more interesting to watch than it would be otherwise. The expression causes Emacs to ‘sit’ or do nothing for a zero length of time and then redraw the screen. Placed here, it causes Emacs to redraw the screen column by column. Without it, Emacs would not redraw the screen until the function exits.

We can test graph-body-print with a short list of numbers.

 1. Install graph-symbol, graph-blank, column-of-graph and graph-body-print.

2. Copy the following expression:

         (graph-body-print '(1 2 3 4 6 4 3 5 7 6 5 2 3))

3. Switch to the ‘*scratch*’ buffer and place the cursor where you want the graph to start.

 4. Type M-: (eval-expression).

5. Yank the graph-body-print expression into the minibuffer with C-y (yank).

6. Press ^{⟨}\underline{\overline{RET}}^{⟩} to evaluate the graph-body-print expression.

Emacs will print a graph like this:

The recursive-graph-body-print Function  147

                         *
                     *   **
                     *  ****
                    *** ****
                   ********* *
                  ************
                 *************

15.2 The recursive-graph-body-print Function

The graph-body-print function may also be written recursively. In this case, it is divided into two parts: an outside ‘wrapper’ that uses a let expression to determine the values of several variables that need only be found once, such as the maximum height of the graph, and an inside function that is called recursively to print the graph.

The ‘wrapper’ is uncomplicated:

     (defun recursive-graph-body-print (numbers-list)
       "Print a bar graph of the NUMBERS-LIST.
     The numbers-list consists of the Y-axis values."
       (let ((height (apply 'max numbers-list))
             (symbol-width (length graph-blank))
             from-position)
         (recursive-graph-body-print-internal
          numbers-list
          height
          symbol-width)))

The recursive function is a little more difficult. It has four parts: the ‘do-again-test’, the printing code, the recursive call, and the ‘next-step-expression’. The ‘do-again-test’ is an if expression that determines whether the numbers-list contains any remaining elements; if it does, the function prints one column of the graph using the printing code and calls itself again. The function calls itself again according to the value produced by the ‘next-step-expression’ which causes the call to act on a shorter version of the numbers-list.

     (defun recursive-graph-body-print-internal
       (numbers-list height symbol-width)
       "Print a bar graph.
     Used within recursive-graph-body-print function."
       (if numbers-list
           (progn
             (setq from-position (point))
             (insert-rectangle
              (column-of-graph height (car numbers-list)))
             (goto-char from-position)
             (forward-char symbol-width)

(sit-for 0) ; Draw graph column by column.

             (recursive-graph-body-print-internal
              (cdr numbers-list) height symbol-width))))

After installation, this expression can be tested; here is a sample:

     (recursive-graph-body-print '(3 2 5 6 7 5 3 4 6 4 3 2 1))
   Here is what recursive-graph-body-print produces:
                     *
                    **   *
                   ****  *
                   **** ***
                 * *********
                 ************
                 *************
   Either of these two functions, graph-body-print or recursive-graph-body-print, create

the body of a graph.

15.3 Need for Printed Axes

A graph needs printed axes, so you can orient yourself. For a do-once project, it may be reasonable to draw the axes by hand using Emacs’s Picture mode; but a graph drawing function may be used more than once.

For this reason, I have written enhancements to the basic print-graph-body function that automatically print labels for the horizontal and vertical axes. Since the label printing functions do not contain much new material, I have placed their description in an appendix. See Appendix C, “A Graph with Labelled Axes”, page 181.

15.4 Exercise

Write a line graph version of the graph printing functions.

16 Your ‘.emacs’ File

“You don’t have to like Emacs to like it” – this seemingly paradoxical statement is the secret of GNU Emacs. The plain, ‘out of the box’ Emacs is a generic tool. Most people who use it, customize it to suit themselves.

GNU Emacs is mostly written in Emacs Lisp; this means that by writing expressions in Emacs Lisp you can change or extend Emacs.

There are those who appreciate Emacs’s default configuration. After all, Emacs starts you in C mode when you edit a C file, starts you in Fortran mode when you edit a Fortran file, and starts you in Fundamental mode when you edit an unadorned file. This all makes sense, if you do not know who is going to use Emacs. Who knows what a person hopes to do with an unadorned file? Fundamental mode is the right default for such a file, just as C mode is the right default for editing C code. But when you do know who is going to use Emacs—you, yourself—then it makes sense to customize Emacs.

For example, I seldom want Fundamental mode when I edit an otherwise undistinguished file; I want Text mode. This is why I customize Emacs: so it suits me.

You can customize and extend Emacs by writing or adapting a ‘~/.emacs’ file. This is your personal initialization file; its contents, written in Emacs Lisp, tell Emacs what to do.

This chapter describes a simple ‘.emacs’ file; for more information, see section “The Init File” in The GNU Emacs Manual, and section “The Init File” in The GNU Emacs Lisp Reference Manual.

16.1 Site-wide Initialization Files

In addition to your personal initialization file, Emacs automatically loads various site-wide initialization files, if they exist. These have the same form as your ‘.emacs’ file, but are loaded by everyone.

Two site-wide initialization files, ‘site-load.el’ and ‘site-init.el’, are loaded into Emacs and then ‘dumped’ if a ‘dumped’ version of Emacs is created, as is most common. (Dumped copies of Emacs load more quickly. However, once a file is loaded and dumped, a change to it does not lead to a change in Emacs unless you load it yourself or re-dump Emacs. See section “Building Emacs” in The GNU Emacs Lisp Reference Manual, and the ‘INSTALL’ file.)

Three other site-wide initialization files are loaded automatically each time you start Emacs, if they exist. These are ‘site-start.el’, which is loaded before your ‘.emacs’ file, and ‘default.el’, and the terminal type file, which are both loaded after your ‘.emacs’ file.

Settings and definitions in your ‘.emacs’ file will overwrite conflicting settings and definitions in a ‘site-start.el’ file, if it exists; but the settings and definitions in a ‘default.el’ or terminal type file will overwrite those in your ‘.emacs’ file. (You can prevent interference from a terminal type file by setting term-file-prefix to nil. See Section 16.10, “A Simple Extension”, page 156.)

The ‘INSTALL’ file that comes in the distribution contains descriptions of the ‘site-init.el’

and ‘site-load.el’ files.

The ‘loadup.el’, ‘startup.el’, and ‘loaddefs.el’ files control loading. These files are in the ‘lisp’ directory of the Emacs distribution and are worth perusing.

The ‘loaddefs.el’ file contains a good many suggestions as to what to put into your own ‘.emacs’ file, or into a site-wide initialization file.

16.2 Setting Variables for One Session

My copy of Emacs version 19.23 has 392 options that you can set with the edit-options command. These ‘options’ are no more than variables such as we have seen earlier and defined

using defvar.

Emacs determines whether a variable is intended to be easily settable by looking at the first character in its documentation string; if the first character is an asterisk, ‘*’, the variable is a user-settable option. (See Section 8.4, “Initializing a Variable with defvar”, page 72.)

The edit-options command lists all the variables in Emacs that the people who wrote the Emacs Lisp libraries think ought to be readily settable. It provides an easy-to-use interface for resetting these variables.

On the other hand, options set using edit-options are set only for the duration of your editing session. The new values are not saved between sessions. Each time Emacs starts, it reads the original defvar value in its source code. To carry a changed setting from one session to the next, you need to use a setq expression within a ‘.emacs’ file or other file that you load every time you start a session.

For me, the major use of the edit-options command is to suggest variables I might want to set in my ‘.emacs’ file. I urge you to look through the list.

See section “Editing Variable Values” in The GNU Emacs Manual, for more information.

16.3 Beginning a ‘.emacs’ File

When you start Emacs, it loads your ‘.emacs’ file unless you tell it not to by specifying ‘-q’ on the command line. (The emacs -q command gives you a plain, out-of-the-box Emacs.)

A ‘.emacs’ file contains Lisp expressions. Often, these are no more than expressions to set values; sometimes they are function definitions.

See section “The Init File ‘~/.emacs’” in The GNU Emacs Manual, for a short description of initialization files.

This chapter goes over some of the same ground, but is a walk among extracts from a complete, long-used ‘.emacs’ file—my own.

The first part of the file consists of comments: reminders to myself. By now, of course, I remember these things, but when I started, I did not.

     ;;;; Bob's .emacs file
     ; Robert J. Chassell
     ; 26 September 1985

Look at that date! I started this file a long time ago. I have been adding to it ever since.

     ; Each section in this file is introduced by a
     ; line beginning with four semicolons; and each
     ; entry is introduced by a line beginning with
     ; three semicolons.

This describes the usual conventions for comments in Emacs Lisp. Everything on a line that follows a semicolon is a comment. Two, three, and four semicolons are used as section and subsection markers. (See section “Comments” in The GNU Emacs Lisp Reference Manual, for more about comments.)

;;;; The Help Key
; Control-h is the help key;
; after typing control-h, type a letter to
; indicate the subject about which you want help.
; For an explanation of the help facility,
; type control-h three times in a row.

Just remember: type C-h three times for help.

; To find out about any mode, type control-h m
; while in that mode.  For example, to find out
; about mail mode, enter mail mode and then type
; control-h m.

‘Mode help’, as I call this, is very helpful. Usually, it tells you all you need to know.

Of course, you don’t need to include comments like these in your ‘.emacs’ file. I included them in mine because I kept forgetting about Mode help or the conventions for comments—but I was able to remember to look here to remind myself.

16.4 Text and Auto Fill Mode

Now we come to the part that ‘turns on’ Text mode and Auto Fill mode.

;;; Text mode and Auto Fill mode
; The next two lines put Emacs into Text mode
; and Auto Fill mode, and are for writers who
; want to start writing prose rather than code.
(setq default-major-mode 'text-mode)
(add-hook 'text-mode-hook 'turn-on-auto-fill)

Here is the first part of this ‘.emacs’ file that does something besides remind a forgetful human!

The first of the two lines in parentheses tells Emacs to turn on Text mode when you find a file, unless that file should go into some other mode, such as C mode.

When Emacs reads a file, it looks at the extension to the file name, if any. (The extension is the part that comes after a ‘.’.) If the file ends with a ‘.c’ or ‘.h’ extension then Emacs turns on C mode. Also, Emacs looks at first nonblank line of the file; if the line says ‘-*- C -*-’, Emacs turns on C mode. Emacs possesses a list of extensions and specifications that it uses automatically. In addition, Emacs looks near the last page for a per-buffer, “local variables list”, if any.

See sections “How Major Modes are Chosen” and “Local Variables in Files” in The GNU Emacs Manual, for information.

Now, back to the ‘.emacs’ file.

Here is the line again; how does it work?

(setq default-major-mode 'text-mode)

This line is a short, but complete Emacs Lisp expression.

We are already familiar with setq. It sets the following variable, default-major-mode, to the subsequent value, which is text-mode. The single quote mark before text-mode tells Emacs to deal directly with the text-mode variable, not with whatever it might stand for. See Section 1.9, “Setting the Value of a Variable”, page 12, for a reminder of how setq works. The main point is that there is no difference between the procedure you use to set a value in your ‘.emacs’ file and the procedure you use anywhere else in Emacs.

Here is the second line:

     (add-hook 'text-mode-hook 'turn-on-auto-fill)

In this line, the add-hook command, adds turn-on-auto-fill to the variable called textmode-hook. turn-on-auto-fill is the name of a program, that, you guessed it!, turns on Auto Fill mode.

Every time Emacs turns on Text mode, Emacs runs the commands ‘hooked’ onto Text mode. So every time Emacs turns on Text mode, Emacs also turns on Auto Fill mode.

In brief, the first line causes Emacs to enter Text mode when you edit a file, unless the file name extension, first non-blank line, or local variables tell Emacs otherwise.

Text mode among other actions, sets the syntax table to work conveniently for writers. In Text mode, Emacs considers an apostrophe as part of a word like a letter; but Emacs does not consider a period or a space as part of a word. Thus, M-f moves you over ‘it's’. On the other hand, in C mode, M-f stops just after the ‘t’ of ‘it's’.

The second line causes Emacs to turn on Auto Fill mode when it turns on Text mode. In Auto Fill mode, Emacs automatically breaks a line that is too wide and brings the excessively wide part of the line down to the next line. Emacs breaks lines between words, not within them.

When Auto Fill mode is turned off, lines continue to the right as you type them. Depending on how you set the value of truncate-lines, the words you type either disappear off the right side of the screen, or else are shown, in a rather ugly and unreadable manner, as a continuation line on the screen.

16.5 Mail Aliases

Here is a setq to ‘turn on’ mail aliases, along with more reminders.

     ;;; Mail mode
     ; To enter mail mode, type `C-x m'
     ; To enter RMAIL (for reading mail),
     ; type `M-x rmail'
     (setq mail-aliases t)

This setq command sets the value of the variable mail-aliases to t. Since t means true, the line says, in effect, “Yes, use mail aliases.”

Mail aliases are convenient short names for long email addresses or for lists of email addresses. The file where you keep your ‘aliases’ is ‘~/.mailrc’. You write an alias like this:

     alias geo george@foobar.wiz.edu

When you write a message to George, address it to ‘geo’; the mailer will automatically expand ‘geo’ to the full address.

16.6 Indent Tabs Mode

By default, Emacs inserts tabs in place of multiple spaces when it formats a region. (For example, you might indent many lines of text all at once with the indent-region command.)

Tabs look fine on a terminal or with ordinary printing, but they produce badly indented output when you use TEX or Texinfo since TEX ignores tabs.

The following turns off Indent Tabs mode:

     ;;; Prevent Extraneous Tabs
     (setq-default indent-tabs-mode nil)

Note that this line uses setq-default rather than the setq command that we have seen before. The setq-default command sets values only in buffers that do not have their own local values for the variable.

See sections “Tabs vs. Spaces” and “Local Variables in Files” in The GNU Emacs Manual.

16.7 Some Keybindings

Now for some personal keybindings:

     ;;; Compare windows
     (global-set-key "\C-cw" 'compare-windows)

compare-windows is a nifty command that compares the text in your current window with text in the next window. It makes the comparison by starting at point in each window, moving over text in each window as far as they match. I use this command all the time.

This also shows how to set a key globally, for all modes.

The command is global-set-key. It is followed by the keybinding. In a ‘.emacs’ file, the keybinding is written as shown: \C-c stands for ‘control-c’, which means ‘press the control key and the c key at the same time’. The w means ‘press the w key’. The keybinding is surrounded by double quotation marks. In documentation, you would write this as C-c w. (If you were binding a ^{⟨}\underline{\overline{META}}^{⟩} key, such as M-c, rather than a ^{⟨}\underline{\overline{CTL}}^{⟩} key, you would write \M-c. See section “Rebinding Keys in Your Init File” in The GNU Emacs Manual, for details.)

The command invoked by the keys is compare-windows. Note that compare-windows is preceded by a single quote; otherwise, Emacs would first try to evaluate the symbol to determine its value.

These three things, the double quotation marks, the backslash before the ‘C’, and the single quote mark are necessary parts of keybinding that I tend to forget. Fortunately, I have come to remember that I should look at my existing ‘.emacs’ file, and adapt what is there.

As for the keybinding itself: C-c w. This combines the prefix key, C-c, with a single character, in this case, w. This set of keys, C-c followed by a single character, is strictly reserved for individuals’ own use. If you ever write an extension to Emacs, please avoid taking any of these keys for public use. Create a key like C-c C-w instead. Otherwise, we will run out of ‘own’ keys.

Here is another keybinding, with a comment:

     ;;; Keybinding for `occur'
     ; I use occur a lot, so let's bind it to a key:
     (global-set-key "\C-co" 'occur)

The occur command shows all the lines in the current buffer that contain a match for a regular expression. Matching lines are shown in a buffer called ‘*Occur*’. That buffer serves as a menu to jump to occurrences.

Here is how to unbind a key, so it does not work:

     ;;; Unbind `C-x f'
     (global-unset-key "\C-xf")

There is a reason for this unbinding: I found I inadvertently typed C-x f when I meant to type C-x C-f. Rather than find a file, as I intended, I accidentally set the width for filled text, almost always to a width I did not want. Since I hardly ever reset my default width, I simply unbound the key.

The following rebinds an existing key:

     ;;; Rebind `C-x C-b' for `buffer-menu'
     (global-set-key "\C-x\C-b" 'buffer-menu)

By default, C-x C-b runs the list-buffers command. This command lists your buffers in another window. Since I almost always want to do something in that window, I prefer the buffer-menu command, which not only lists the buffers, but moves point into that window.

16.8 Loading Files

Many people in the GNU Emacs community have written extensions to Emacs. As time goes by, these extensions are often included in new releases. For example, the Calendar and Diary packages are now part of the standard Emacs version 19 distribution; they were not part of the standard Emacs version 18 distribution.

(Calc, which I consider a vital part of Emacs, would be part of the standard distribution except that it is so large it is packaged separately.)

You can use a load command to evaluate a complete file and thereby install all the functions and variables in the file into Emacs. For example:

     (load "~/emacs/kfill")

This evaluates, i.e. loads, the ‘kfill.el’ file (or if it exists, the faster, byte compiled ‘kfill.elc’ file) from the ‘emacs’ sub-directory of your home directory.

(‘kfill.el’ was adapted from Kyle E. Jones’ ‘filladapt.el’ package by Bob Weiner and “provides no muss, no fuss word wrapping and filling of paragraphs with hanging indents, included text from news and mail messages, and Lisp, C++, PostScript or shell comments.” I use it all the time and hope it is incorporated into the standard distribution.)

If you load many extensions, as I do, then instead of specifying the exact location of the extension file, as shown above, you can specify that directory as part of Emacs’s load-path. Then, when Emacs loads a file, it will search that directory as well as its default list of directories. (The default list is specified in ‘paths.h’ when Emacs is built.)

The following command adds your ‘~/emacs’ directory to the existing load path:

     ;;; Emacs Load Path
     (setq load-path (cons "~/emacs" load-path))

Incidentally, load-library is an interactive interface to the load function. The complete function looks like this:

A Simple Extension: line-to-top-of-window  155

     (defun load-library (library)
       "Load the library named LIBRARY.
     This is an interface to the function `load'."
       (interactive "sLoad library: ")
       (load library))

The name of the function, load-library, comes from the use of ‘library’ as a conventional synonym for ‘file’. The source for the load-library command is in the ‘files.el’ library.

Another interactive command that does a slightly different job is load-file. See section “Libraries of Lisp Code for Emacs” in The GNU Emacs Manual, for information on the distinction between load-library and this command.

16.9 Autoloading

Instead of installing a function by loading the file that contains it, or by evaluating the function definition, you can make the function available but not actually install it until it is first called. This is called autoloading.

When you execute an autoloaded function, Emacs automatically evaluates the file that contains the definition, and then calls the function.

Emacs starts quicker with autoloaded functions, since their libraries are not loaded right away; but you need to wait a moment when you first use such a function, while its containing file is evaluated.

Rarely used functions are frequently autoloaded. The ‘loaddefs.el’ library contains hundreds of autoloaded functions, from bookmark-set to wordstar-mode. Of course, you may come to use a ‘rare’ function frequently. In this case, you should load that function’s file with a load expression in your ‘.emacs’ file.

In my ‘.emacs’ file for Emacs version 19.23, I load 17 libraries that contain functions that would otherwise be autoloaded. (Actually, it would have been better to include these files in my ‘dumped’ Emacs when I built it, but I forgot. See section “Building Emacs” in The GNU Emacs Lisp Reference Manual, and the ‘INSTALL’ file for more about dumping.)

You may also want to include autoloaded expressions in your ‘.emacs’ file. autoload is a built-in function that takes up to five arguments, the final three of which are optional. The first argument is the name of the function to be autoloaded; the second is the name of the file to be loaded. The third argument is documentation for the function, and the fourth tells whether the function can be called interactively. The fifth argument tells what type of object—autoload can handle a keymap or macro as well as a function (the default is a function).

Here is a typical example:

     (autoload 'html-helper-mode
       "html-helper-mode" "Edit HTML documents" t)

This expression autoloads the html-helper-mode function from the ‘html-helper-mode.el’ file (or, if it exists, from the byte compiled file ‘html-helper-mode.elc’.) The file must be located in a directory specified by load-path. The documentation says that this is a mode to help you edit documents written in the HyperText Markup Language. You can call this mode interactively by typing M-x html-helper-mode. (You need to duplicate the function’s regular documentation in the autoload expression because the regular function is not yet loaded, so its documentation is not available.)

See section “Autoload” in The GNU Emacs Lisp Reference Manual, for more information.

16.10 A Simple Extension: line-to-top-of-window

Here is a simple extension to Emacs that moves the line point is on to the top of the window. I use this all the time, to make text easier to read.

You can put the following code into a separate file and then load it from your ‘.emacs’ file, or you can include it within your ‘.emacs’ file.

Here is the definition:

     ;;; Line to top of window;
     ;;; replace three keystroke sequence  C-u 0 C-l
     (defun line-to-top-of-window ()
       "Move the line point is on to top of window."
       (interactive)
       (recenter 0))

Now for the keybinding.

Although most of an Emacs version 18 ‘.emacs’ file works with version 19, there are some differences (also, of course, there are new features in Emacs 19).

In version 19 Emacs, you can write a function key like this: ‘[f6]’. In version 18, you must specify the key strokes sent by the keyboard when you press that function key. For example, a Zenith 29 keyboard sends ^{⟨}\underline{\overline{ESC P}}^{⟩} when I press its sixth function key; an Ann Arbor Ambassador keyboard sends ^{⟨}\underline{\overline{ESC O F}}^{⟩}. Write these keystrokes as ‘\eP’ and ‘\eOF’, respectively.

In my version 18 ‘.emacs’ file, I bind line-to-top-of-window to a key that depends on the type of terminal:

     (defun z29-key-bindings ()
       "Function keybindings for Z29 terminal."
       ;; ...
       (global-set-key "\eP" 'line-to-top-of-window))
     (defun aaa-key-bindings ()
       "Function keybindings for Ann Arbor Ambassador"
       ;; ...
       (global-set-key "\eOF" 'line-to-top-of-window))

(You can find out what a function key sends by typing the function key, and then typing C-h l (view-lossage) which displays the last 100 input keystrokes.)

After specifying the key bindings, I evaluate an expression that chooses among keybindings, depending on the type of terminal I am using. However, before doing that, I turn off the predefined, default terminal-specific keybindings, which overwrite bindings in the ‘.emacs’ if they clash.

     ;;; Turn Off Predefined Terminal Keybindings
     ; The following turns off the predefined
     ; terminal-specific keybindings such as the
     ; vt100 keybindings in lisp/term/vt100.el.
     ; If there are no predefined terminal
     ; keybindings, or if you like them,
     ; comment this out.
     (setq term-file-prefix nil)

Here is the selection expression itself:

     (let ((term (getenv "TERM")))
       (cond
        ((equal term "z29") (z29-key-bindings))
        ((equal term "aaa") (aaa-key-bindings))
        (t (message
            "No binding for terminal type %s."
            term))))

In Emacs version 19, function keys (as well as mouse button events and non-ascii characters) are written within square brackets, without quotation marks. I bind line-to-top-of-window to my ^{⟨}\underline{\overline{F6}}^{⟩} function key like this:

     (global-set-key [f6] 'line-to-top-of-window)

Much simpler!

For more information, see section “Rebinding Keys in Your Init File” in The GNU Emacs Manual.

If you run both Emacs 18 and Emacs 19, you can select which code to evaluate with the following conditional:

     (if (string=
          (int-to-string 18)
          (substring (emacs-version) 10 12))
         ;; evaluate version 18 code
         (progn
            ... )
       ;; else evaluate version 19 code
       ...

16.11 Keymaps

Emacs uses keymaps to record which keys call which commands. Specific modes, such as C mode or Text mode, have their own keymaps; the mode-specific keymaps override the global map that is shared by all buffers.

The global-set-key function binds, or rebinds, the global keymap. For example, the following binds the key C-c C-l to the function line-to-top-of-window:

     (global-set-key "\C-c\C-l" 'line-to-top-of-window))

Mode-specific keymaps are bound using the define-key function, which takes a specific keymap as an argument, as well as the key and the command. For example, my ‘.emacs’ file contains the following expression to bind the texinfo-insert-@group command to C-c C-c g:

     (define-key texinfo-mode-map "\C-c\C-cg"
       'texinfo-insert-@group)

The texinfo-insert-@group function itself is a little extension to Texinfo mode that inserts ‘@group’ into a Texinfo file. I use this command all the time and prefer to type the three strokes C-c C-c g rather than the six strokes @ g r o u p. (‘@group’ and its matching ‘@end group’ are commands that keep all enclosed text together on one page; many multi-line examples in this book are surrounded by ‘@group ... @end group’.)

Here is the texinfo-insert-@group function definition:

(defun texinfo-insert-@group ()
  "Insert the string @group in a Texinfo buffer."
  (interactive)
  (beginning-of-line)
  (insert "@group\n"))

(Of course, I could have used Abbrev mode to save typing, rather than write a function to insert a word; but I prefer key strokes consistent with other Texinfo mode key bindings.)

You will see numerous define-key expressions in ‘loaddefs.el’ as well as in the various mode libraries, such as ‘c-mode.el’ and ‘lisp-mode.el’.

See section “Customizing Key Bindings” in The GNU Emacs Manual, and section “Keymaps” in The GNU Emacs Lisp Reference Manual, for more information about keymaps.

16.12 X11 Colors

You can specify colors when you use Emacs version 19 with the MIT X Windowing system. (All the previous examples should work with both Emacs version 18 and Emacs version 19; this works only with Emacs version 19.)

I hate the default colors and specify my own.

Most of my specifications are in various X initialization files. I wrote notes to myself in my ‘.emacs’ file to remind myself what I did:

;; I use TWM for window manager;
;; my ~/.xsession file specifies:
;    xsetroot -solid navyblue -fg white

Actually, the root of the X window is not part of Emacs at all, but I like the reminder anyhow.

;; My ~/.Xresources file specifies:
;     XTerm*Background:    sky blue
;     XTerm*Foreground:    white
;     emacs*geometry:      =80x40+100+0
;     emacs*background:    blue
;     emacs*foreground:    grey97
;     emacs*cursorColor:   white
;     emacs*pointerColor:  white

Here are the expressions in my ‘.emacs’ file that set values:

;;; Set highlighting colors for isearch and drag
(set-face-foreground 'highlight "white" )
(set-face-background 'highlight "slate blue")
(set-face-background 'region    "slate blue")
(set-face-background
 'secondary-selection "turquoise")

;; Set calendar highlighting colors
(setq calendar-load-hook
  '(lambda ()
     (set-face-foreground 'diary-face   "skyblue")
     (set-face-background 'holiday-face "slate blue")
     (set-face-foreground 'holiday-face "white")))

The various shades of blue soothe my eye and prevent me from seeing the screen flicker.

16.13 V19 Miscellaneous

Here are a few miscellaneous settings for version 19 Emacs:

− Automatically resize the minibuffer as needed:

(resize-minibuffer-mode 1)
(setq resize-minibuffer-mode t)

− Turn on highlighting for search strings:

(setq search-highlight t)

− Set every frame to show a menu bar and to come forward when you move the mouse onto

it.

(setq default-frame-alist
      '((menu-bar-lines . 1)
        (auto-lower . t)
        (auto-raise . t)))

− Set the shape and color of the mouse cursor:

; Cursor shapes are defined in
; `/usr/include/X11/cursorfont.h';
; for example, the `target' cursor is number 128;
; the `top_left_arrow' cursor is number 132.
(let ((mpointer (x-get-resource "*mpointer"
                                "*emacs*mpointer")))
  ;; If you have not set your mouse pointer
  ;;     then sent it, otherwise leave as is:
  (if (eq mpointer nil)
      (setq mpointer "132")) ; top_left_arrow
  (setq x-pointer-shape (string-to-int mpointer))
  (set-mouse-color "white"))

16.14 A Modified Mode Line

Finally, a feature I really like: a modified mode line.

Since I sometimes work over a network, I replaced the ‘Emacs: ’ that is normally written on the left hand side of the mode line by the name of the system—otherwise, I forget which machine I am using. In addition, I list the default directory lest I lose track of where I am, and I specify the line point is on, with ‘Line’ spelled out. My ‘.emacs’ file looks like this:

(setq mode-line-system-identification
(substring (system-name) 0
         (string-match "\\..+" (system-name))))

     (setq default-mode-line-format
           (list ""
                 'mode-line-modified
                 "<"
                 'mode-line-system-identification
                 "> "
                 "%14b"
                 " "
                 'default-directory
                 " "
                 "%[("
                 'mode-name
                 'minor-mode-alist
                 "%n"
                 'mode-line-process
                 ")%]--"
                  "Line %l--"
                 '(-3 . "%P")
                 "-%-"))
     ;; Start with new default.
     (setq mode-line-format default-mode-line-format)

I set the default mode line format so as to permit various modes, such as Info, to override it. Many elements in the list are self-explanatory: mode-line-modified is a variable the tells whether the buffer has been modified, mode-name tells the name of the mode, and so on.

The ‘"%14b"’ displays the current buffer name (using the buffer-name function with which we are familiar); the ‘14’ specifies the maximum number of characters that will be displayed. When a name has fewer characters, whitespace is added to fill out to this number. ‘%[’ and ‘%]’ cause a pair of square brackets to appear for each recursive editing level. ‘%n’ says ‘Narrow’ when narrowing is in effect. ‘%P’ tells you the percentage of the buffer that is above the bottom of the window, or ‘Top’, ‘Bottom’, or ‘All’. (A lower case ‘p’ tell you the percentage above the top of the window.) ‘%-’ inserts enough dashes to fill out the line.

In and after Emacs version 19.29, you can use frame-title-format to set the title of an Emacs frame. This variable has the same structure as mode-line-format.

Mode line formats are described in section “Mode Line Format” in The GNU Emacs Lisp Reference Manual.

Remember, “You don’t have to like Emacs to like it” — your own Emacs can have different colors, different commands, and different keys than a default Emacs.

On the other hand, if you want to bring up a plain ‘out of the box’ Emacs, with no customization, type:

     emacs -q

This will start an Emacs that does not load your ‘~/.emacs’ initialization file. A plain, default Emacs. Nothing more.

debug  161

17 Debugging

GNU Emacs has two debuggers, debug and edebug. The first is built into the internals of Emacs and is always with you; the second is an extension to Emacs that has become part of the standard distribution in version 19.

Both debuggers are described extensively in section “Debugging Lisp Programs” in The GNU Emacs Lisp Reference Manual. In this chapter, I will walk through a short example of each.

17.1 debug

Suppose you have written a function definition that is intended to return the sum of the numbers 1 through a given number. (This is the triangle function discussed earlier. See “Example with Decrementing Counter”, page 93, for a discussion.)

However, your function definition has a bug. You have mistyped ‘1=’ for ‘1-’. Here is the broken definition:

     (defun triangle-bugged (number)
       "Return sum of numbers 1 through NUMBER inclusive."
       (let ((total 0))
         (while (> number 0)
           (setq total (+ total number))
           (setq number (1= number)))      ; Error here.
         total))

If you are reading this in Info, you can evaluate this definition in the normal fashion. You will see triangle-bugged appear in the echo area.

Now evaluate the triangle-bugged function with an argument of 4:

     (triangle-bugged 4)

You will produce an error message that says:

     Symbol's function definition is void: 1=

In practice, for a bug as simple as this, this error message will tell you what you need to know to correct the definition. However, suppose you are not quite certain what is going on?

You can turn on debugging by setting the value of debug-on-error to t:

     (setq debug-on-error t)

This causes Emacs to enter the debugger next time it encounters an error.

You can turn off debug-on-error by setting it to nil:

     (setq debug-on-error nil)

Set debug-on-error to t and evaluate the following:

     (triangle-bugged 4)

This time, Emacs will create a buffer called ‘*Backtrace*’ that looks like this:

     ---------- Buffer: *Backtrace* ----------
     Signalling: (void-function 1=)
       (1= number))
       (setq number (1= number)))
       (while (> number 0) (setq total (+ total number))
             (setq number (1= number))))
       (let ((total 0)) (while (> number 0) (setq total ...)
             (setq number ...)) total))
       triangle-bugged(4)
       eval((triangle-bugged 4))
       eval-last-sexp(nil)
     * call-interactively(eval-last-sexp)
     ---------- Buffer: *Backtrace* ----------

(I have reformatted this example slightly; the debugger does not fold long lines.)

You read the ‘*Backtrace*’ buffer from the bottom up; it tells you what Emacs did that led to the error. In this case, what Emacs did was make an interactive call to C-x C-e (evallast-sexp), which led to the evaluation of the triangle-bugged expression. Each line above tells you what the Lisp interpreter evaluated next.

The third line from the top of the buffer is

     (setq number (1= number))

Emacs tried to evaluate this expression; in order to do so, it tried to evaluate the inner expression shown on the second line from the top:

     (1= number)

This is where the error occurred; as the top line says:

     Signalling: (void-function 1=)

You can correct the mistake, re-evaluate the function definition, and then run your test again.

If you are reading this in Info, you can now turn off debug-on-error by setting it to nil:

     (setq debug-on-error nil)

17.2 debug-on-entry

A second way to start debug on a function is to enter the debugger when you call the function. You can do this by calling debug-on-entry.

Type:

     M-x debug-on-entry RET triangle-bugged RET

Now, evaluate the following:

     (triangle-bugged 5)

Emacs will create a ‘*Backtrace*’ buffer and tell you that it is beginning to evaluate the

triangle-bugged function:

debug-on-quit and (debug)  163

---------- Buffer: *Backtrace* ----------
Entering:
* triangle-bugged(5)
  eval((triangle-bugged 5))
  eval-last-sexp(nil)
* call-interactively(eval-last-sexp)
---------- Buffer: *Backtrace* ----------

In the ‘*Backtrace*’ buffer, type d. Emacs will evaluate the first expression in trianglebugged; the buffer will look like this:

---------- Buffer: *Backtrace* ----------
Beginning evaluation of function call form:
* (let ((total 0)) (while (> number 0) (setq total ...)
        (setq number ...)) total))
  triangle-bugged(5)
* eval((triangle-bugged 5))
  eval-last-sexp(nil)
* call-interactively(eval-last-sexp)
---------- Buffer: *Backtrace* ----------

Now, type d again, eight times, slowly. Each time you type d, Emacs will evaluate another expression in the function definition. Eventually, the buffer will look like this:

---------- Buffer: *Backtrace* ----------
Beginning evaluation of function call form:
* (setq number (1= number)))
* (while (> number 0) (setq total (+ total number))
        (setq number (1= number))))
* (let ((total 0)) (while (> number 0)
        (setq total ...) (setq number ...)) total))
  triangle-bugged(5)
* eval((triangle-bugged 5))
  eval-last-sexp(nil)
* call-interactively(eval-last-sexp)
---------- Buffer: *Backtrace* ----------

Finally, after you type d two more times, Emacs will reach the error, and the top two lines of the ‘*Backtrace*’ buffer will look like this:

---------- Buffer: *Backtrace* ----------
Signalling: (void-function 1=)
* (1= number))
...
---------- Buffer: *Backtrace* ----------

By typing d, you were able to step through the function.

You can quit a ‘*Backtrace*’ buffer by typing q; this quits the trace, but does not cancel

debug-on-entry.

To cancel the effect of debug-on-entry, call cancel-debug-on-entry and the name of the function, like this:

M-x cancel-debug-on-entry RET triangle-debugged RET

(If you are reading this in Info, cancel debug-on-entry now.)

17.3 debug-on-quit and (debug)

In addition to setting debug-on-error or calling debug-on-entry, there are two other ways to start debug.

You can start debug whenever you type C-g (keyboard-quit) by setting the variable debugon-quit to t. This is useful for debugging infinite loops.

Or, you can insert a line that says (debug) into your code where you want the debugger to start, like this:

     (defun triangle-bugged (number)
       "Return sum of numbers 1 through NUMBER inclusive."
       (let ((total 0))
         (while (> number 0)
           (setq total (+ total number))

(debug) ; Start debugger.

           (setq number (1= number)))      ; Error here.
         total))

The debug function is described in detail in section “The Lisp Debugger” in The GNU Emacs Lisp Reference Manual.

17.4 The edebug Source Level Debugger

Edebug normally displays the source of the code you are debugging, with an arrow at the left that shows which line you are currently executing.

You can walk through the execution of a function, line by line, or run quickly until reaching a breakpoint where execution stops.

Edebug is described in section “Edebug” in The GNU Emacs Lisp Reference Manual.

Here is a bugged function definition for triangle-recursively. See Section 11.2.2, “Recursion in place of a counter”, page 97, for a review of it. This example is presented without indentation to the left of the defun, as explained below.

(defun triangle-recursively-bugged (number)
  "Return sum of numbers 1 through NUMBER inclusive.
Uses recursion."
  (if (= number 1)
      1
    (+ number
       (triangle-recursively-bugged
        (1= number)))))               ; Error here.

Normally, you would install this definition by positioning your cursor after the function’s closing parenthesis and typing C-x C-e (eval-last-sexp) or else by positioning your cursor within the definition and typing C-M-x (eval-defun). (By default, the eval-defun command works only in Emacs Lisp mode or in Lisp Interactive mode.)

However, to prepare this function definition for Edebug, you must first instrument the code using a different command. In Emacs version 19, you can do this by positioning your cursor within the definition and typing the following:

     M-x edebug-defun RET

This will cause Emacs to load Edebug automatically if it is not already loaded, and properly instrument the function. (After loading Edebug, you can use its standard keybindings, such as

C-u C-M-x (eval-defun with a prefix argument) for edebug-defun.)

In Emacs version 18, you need to load Edebug yourself; you can do this by putting the appropriate load command in your ‘.emacs’ file.

If you are reading this in Info, you can instrument the triangle-recursively-bugged function shown above. edebug-defun fails to locate the bounds of a definition whose defun line is indented; so the example is presented without the usual spaces to the left of the defun.

After instrumenting the function, place your cursor after the following expression and type

C-x C-e (eval-last-sexp):
     (triangle-recursively-bugged 3)

You will be jumped back to the source for triangle-recursively-bugged and the cursor positioned at the beginning of the if line of the function. Also, you will see an arrow at the left hand side of that line that looks like this: ‘=>’. The arrow marks the line where the function is executing.

     =>⋆(if (= number 1)

In the example, the location of point is displayed with a star, ‘⋆’ (in Info, it is displayed as ‘-!-’).

If you now press ^{⟨}\underline{\overline{SPC}}^{⟩}, point will move to the next expression to be executed; the line will look like this:

     =>(if ⋆(= number 1)

As you continue to press ^{⟨}\underline{\overline{SPC}}^{⟩}, point will move from expression to expression. At the same time, whenever an expression returns a value, that value will be displayed in the echo area. For example, after you move point past number, you will see the following:

     Result: 3 = C-c

This means the value of number is 3, which is ascii ^{⟨}\underline{\overline{CTL-C}}^{⟩} (the third letter of the alphabet).

You can continue moving through the code until you reach the line with the error. Before evaluation, that line looks like this:

     =>  ⋆(1= number)))))               ; Error here.

When you press ^{⟨}\underline{\overline{SPC}}^{⟩} once again, you will produce an error message that says:

     Symbol's function definition is void: 1=

This is the bug.

Press ‘q’ to quit Edebug.

To remove instrumentation from a function definition, simply re-evaluate it with a command that does not instrument it. For example, you could place your cursor after the definition’s closing parenthesis and type C-x C-e.

Edebug does a great deal more than walk with you through a function. You can set it so it races through on its own, stopping only at an error or at specified stopping points; you can cause it to display the changing values of various expressions; you can find out how many times a function is called, and more.

Edebug is described in section “Edebug” in The GNU Emacs Lisp Reference Manual.

17.5 Debugging Exercises

• Install the count-words-region function and then cause it to enter the built-in debugger when you call it. Run the command on a region containing two words. You will need to press d a remarkable number of times. On your system, is a ‘hook’ called after the command finishes? (For information on hooks, see section “Command Loop Overview” in The GNU Emacs Lisp Reference Manual.)

• Copy count-words-region into the ‘*scratch*’ buffer, remove white space before the defun line if necessary, instrument the function for Edebug, and walk through its execution. The function does not need to have a bug, although you can introduce one if you wish. If the function lacks a bug, the walk-through completes without problems.

• While running Edebug, type ? to see a list of all the Edebug commands. (The globaledebug-prefix is usually C-x X, i.e. ^{⟨}\underline{\overline{CTL}}^{⟩}-x followed by an upper case X; use this prefix for commands made outside of the Edebug debugging buffer.)

• In the Edebug debugging buffer, use the p (edebug-bounce-point) command to see where in the region the count-words-region is working.

• Move point to some spot further down function and then type the h (edebug-goto-here) command to jump to that location.

• Use the t (edebug-trace-mode) command to cause Edebug to walk through the function on its own; use an upper case T for edebug-Trace-fast-mode.

• Set a breakpoint, then run Edebug in Trace mode until it reaches the stopping point.

18 Conclusion

We have now reached the end of this Introduction. You have now learned enough about programming in Emacs Lisp to set values, to write simple ‘.emacs’ files for yourself and your friends, and write simple customizations and extensions to Emacs.

This is a place to stop. Or, if you wish, you can now go onward, and teach yourself.

You have learned some of the basic nuts and bolts of programming. But only some. There are a great many more brackets and hinges that are easy to use that we have not touched.

A path you can follow right now lies among the sources to GNU Emacs and in The GNU Emacs Lisp Reference Manual.

The Emacs Lisp sources are an adventure. When you read the sources and come across a function or expression that is unfamiliar, you need to figure out or find out what it does.

Go to the Reference Manual. It is a thorough, complete, and fairly easy-to-read description of Emacs Lisp. It is written not only for experts, but for people who know what you know. (The Reference Manual comes with the standard GNU Emacs distribution. Like this introduction, it comes as a Texinfo source file, so you can read it on-line and as a typeset, printed book.)

Go to the other on-line help that is part of GNU Emacs: the on-line documentation for all functions, and find-tags, the program that takes you to sources.

Here is an example of how I explore the sources. Because of its name, ‘simple.el’ is the file I looked at first, a long time ago. As it happens some of the functions in ‘simple.el’ are complicated, or at least look complicated at first sight. The first function, for example, looks complicated. This is the open-line function.

You may want to walk through this function slowly, as we did with the forward-sentence function. (See Section 12.3, “forward-sentence”, page 103.) Or you may want to skip that function and look at another, such as split-line. You don’t need to read all the functions. According to count-words-in-defun, the split-line function contains 27 words and symbols.

Even though it is short, split-line contains four expressions we have not studied: skip-

chars-forward, indent-to, insert, and ‘?\n’.

Consider the insert function. (It is mentioned in passing in the review section in Chapter 12, “Regular Expression Searches”, page 101.) In Emacs, you can find out more about insert by typing C-h f (describe-function) and the name of the function. This gives you the function documentation. You can look at its source using find-tag, which is bound to M-. (this is not so helpful in this case; the function is a primitive written in C rather than Lisp). Finally, you can find out what the Reference Manual has to say by visiting the manual in Info, and typing i (Info-index) and the name of the function, or by looking up insert in the index to a printed copy of the manual.

Similarly, you can find out what is meant by ‘?\n’. You can try using Info-index with ‘?\n’. It turns out that this action won’t help; but don’t give up. If you search the index for ‘\n’ without the ‘?’, you will be taken directly to the relevant section of the manual. (See section “Character Type” in The GNU Emacs Lisp Reference Manual. ‘?\n’ stands for the newline character.)

You may be able to guess what is done by skip-chars-forward and indent-to; or you can look them up, too. (Incidentally, the describe-function function itself is in ‘help.el’; it is one of those long, but decipherable functions. Its definition illustrates how to customize the interactive expression without using the standard character codes; and it shows how to create a temporary buffer.)

Other interesting source files include ‘paragraphs.el’, ‘loaddefs.el’, and ‘loadup.el’. The ‘paragraphs.el’ file includes short, easily understood functions as well as longer ones. The ‘loaddefs.el’ file contains the many standard autoloads and many keymaps. I have never looked at it all; only at parts. ‘loadup.el’ is the file that loads the standard parts of Emacs; it tells you a great deal about how Emacs is built. (See section “Building Emacs” in The GNU Emacs Lisp Reference Manual, for more about building.)

As I said, you have learned some nuts and bolts; however, and very importantly, we have hardly touched major aspects of programming; I have said nothing about how to sort information, except to use the predefined sort function; I have said nothing about how to store information, except to use variables and lists; I have said nothing about how to write programs that write programs. These are topics for another, and different kind of book, a different kind of learning.

What you have done is learn enough for much practical work with GNU Emacs. What you have done is get started. This is the end of a beginning.

Appendix A The the-the Function

Sometimes when you you write text, you duplicate words—as with “you you” near the beginning of this sentence. I find that most frequently, I duplicate “the’; hence, I call the function for detecting duplicated words, the-the.

As a first step, you could use the following regular expression to search for duplicates:

     \\(\\w+[ \t\n]+\\)\\1

This regexp matches one or more word-constituent characters followed by one or more spaces, tabs, or newlines. However, it does not detect duplicated words on different lines, since the ending of the first word, the end of the line, is different from the ending of the second word, a space. (For more information about regular expressions, see Chapter 12, “Regular Expression Searches”, page 101, as well as section “Syntax of Regular Expressions” in The GNU Emacs Manual, and section “Regular Expressions” in The GNU Emacs Lisp Reference Manual.)

You might try searching just for duplicated word-constituent characters but that does not work since the pattern detects doubles such as the two occurrences of ‘th’ in ‘with the’.

Another possible regular expression is for word-constituent characters that are followed by non-word-constituent characters. Here, ‘\\w+’ matches one or more word-constituent characters and ‘\\W*’ matches zero or more non-word-constituent characters.

     \\(\\(\\w+\\)\\W*\\)\\1

Again, not useful.

Here is the pattern that I use. It is not perfect, but good enough. ‘\\b’ matches the empty string, provided it is at the beginning or end of a word; ‘[^@ \n\t]+’ matches one or more occurrences of any characters that are not an @-sign, space, newline, or tab.

     \\b\\([^@ \n\t]+\\)[ \n\t]+\\1\\b

One can write more complicated expressions, but I found that this expression is good enough, so I use it.

Here is the the-the function, as I include it in my ‘.emacs’ file, along with a handy global key binding:

     (defun the-the ()
       "Search forward for for a duplicated word."
       (interactive)
       (message "Searching for for duplicated words ...")
       (push-mark)
       ;; This regexp is not perfect
       ;; but is fairly good over all:
       (if (re-search-forward
            "\\b\\([^@ \n\t]+\\)[ \n\t]+\\1\\b" nil 'move)
           (message "Found duplicated word.")
         (message "End of buffer")))
     ;; Bind `the-the' to  C-c \
     (global-set-key "\C-c\\" 'the-the)

Here is test text:

     one two two three four five
     five six seven

170 Appendix A: The the-the Function

You can substitute the other regular expressions shown above in the function definition and try each of them on this list.

The rotate-yank-pointer Function  171

Appendix B Handling the Kill Ring

The kill ring is a list that is transformed into a ring by the workings of the rotate-yank-

pointer function. The yank and yank-pop commands use the rotate-yank-pointer function.

This appendix describes the rotate-yank-pointer function as well as both the yank and the yank-pop commands.

B.1 The rotate-yank-pointer Function

The rotate-yank-pointer function changes the element in the kill ring to which kill-ring-yank-pointer points. For example, it can change kill-ring-yank-pointer from pointing to the second element to point to the third element.

   Here is the code for rotate-yank-pointer:
     (defun rotate-yank-pointer (arg)
       "Rotate the yanking point in the kill ring."
       (interactive "p")
       (let ((length (length kill-ring)))
         (if (zerop length)

;; then-part

             (error "Kill ring is empty")

;; else-part

           (setq kill-ring-yank-pointer
                 (nthcdr (% (+ arg
                               (- length
                                  (length
                                   kill-ring-yank-pointer)))
                            length)
                         kill-ring)))))

The function looks complex, but as usual, it can be understood by taking it apart piece by piece. First look at it in skeletal form:

     (defun rotate-yank-pointer (arg)
       "Rotate the yanking point in the kill ring."
       (interactive "p")

(let varlist

         body...)

This function takes one argument, called arg. It has a brief documentation string; and it is interactive with a small ‘p’, which means that the argument must be a processed prefix passed to the function as a number.

The body of the function definition is a let expression, which itself has a body as well as a varlist.

The let expression declares a variable that will be only usable within the bounds of this function. This variable is called length and is bound to a value that is equal to the number of items in the kill ring. This is done by using the function called length. (Note that this function has the same name as the variable called length; but one use of the word is to name the function and the other is to name the variable. The two are quite distinct. Similarly, an English speaker will distinguish between the meanings of the word ‘ship’ when he says: "I must ship this package immediately." and "I must get aboard the ship immediately.")

The function length tells the number of items there are in a list, so (length kill-ring) returns the number of items there are in the kill ring.

B.1.1 The Body of rotate-yank-pointer

The body of rotate-yank-pointer is a let expression and the body of the let expression is an if expression.

The purpose of the if expression is to find out whether there is anything in the kill ring. If the kill ring is empty, the error function stops evaluation of the function and prints a message in the echo area. On the other hand, if the kill ring has something in it, the work of the function is done.

Here is the if-part and then-part of the if expression:

     (if (zerop length)                      ; if-part
         (error "Kill ring is empty")        ; then-part
       ...

If there is not anything in the kill ring, its length must be zero and an error message sent to the user: ‘Kill ring is empty’. The if expression uses the function zerop which returns true if the value it is testing is zero. When zerop tests true, the then-part of the if is evaluated. The then-part is a list starting with the function error, which is a function that is similar to the message function (see Section 1.8.5, “message”, page 11), in that it prints a one-line message in the echo area. However, in addition to printing a message, error also stops evaluation of the function within which it is embedded. In this case, this means that the rest of the function will not be evaluated if the length of the kill ring is zero.

(In my opinion, it is slightly misleading, at least to humans, to use the term ‘error’ as the name of this function. A better term would be ‘cancel’. Strictly speaking, of course, you cannot point to, much less rotate a pointer to a list that has no length, so from the point of view of the computer, the word ‘error’ is correct. But a human expects to attempt this sort of thing, if only to find out whether the kill ring is full or empty. This is an act of exploration.

(From the human point of view, the act of exploration and discovery is not necessarily an error, and therefore should not be labeled as one, even in the bowels of a computer. As it is, the code in Emacs implies that a human who is acting virtuously, by exploring his or her environment, is making an error. This is bad. Even though the computer takes the same steps as it does when there is an ‘error’, a term such as ‘cancel’ would have a clearer connotation.)

The else-part of the if expression

The else-part of the if expression is dedicated to setting the value of kill-ring-yank-pointer when the kill ring has something in it. The code looks like this:

     (setq kill-ring-yank-pointer
           (nthcdr (% (+ arg
                         (- length
                            (length kill-ring-yank-pointer)))
                      length)
                   kill-ring)))))

This needs some examination. Clearly, kill-ring-yank-pointer is being set to be equal to some cdr of the kill ring, using the nthcdr function that is described in an earlier section. (See Section 8.5, “copy-region-as-kill”, page 73.) But exactly how does it do this?

Before looking at the details of the code let’s first consider the purpose of the rotate-yank-pointer function.

   The rotate-yank-pointer function changes what kill-ring-yank-pointer points to. If

kill-ring-yank-pointer starts by pointing to the first element of a list, a call to rotateyank-pointer causes it to point to the second element; and if kill-ring-yank-pointer points to the second element, a call to rotate-yank-pointer causes it to point to the third element. (And if rotate-yank-pointer is given an argument greater than 1, it jumps the pointer that many elements.)

   The rotate-yank-pointer function uses setq to reset what the kill-ring-yank-pointer

points to. If kill-ring-yank-pointer points to the first element of the kill ring, then, in the simplest case, the rotate-yank-pointer function must cause it to point to the second element. Put another way, kill-ring-yank-pointer must be reset to have a value equal to the cdr of the kill ring.

That is, under these circumstances,

     (setq kill-ring-yank-pointer
        ("some text" "a different piece of text" "yet more text"))
     (setq kill-ring
        ("some text" "a different piece of text" "yet more text"))

the code should do this:

     (setq kill-ring-yank-pointer (cdr kill-ring))

As a result, the kill-ring-yank-pointer will look like this:

     kill-ring-yank-pointer
          ⇒ ("a different piece of text" "yet more text"))

The actual setq expression uses the nthcdr function to do the job.

As we have seen before (see Section 7.3, “nthcdr”, page 62), the nthcdr function works by repeatedly taking the cdr of a list—it takes the cdr of the cdr of the cdr …

The two following expressions produce the same result:

     (setq kill-ring-yank-pointer (cdr kill-ring))
     (setq kill-ring-yank-pointer (nthcdr 1 kill-ring))

In the rotate-yank-pointer function, however, the first argument to nthcdr is a rather complex looking expression with lots of arithmetic inside of it:

     (% (+ arg
           (- length
              (length kill-ring-yank-pointer)))
        length)

As usual, we need to look at the most deeply embedded expression first and then work our way towards the light.

The most deeply embedded expression is (length kill-ring-yank-pointer). This finds the length of the current value of the kill-ring-yank-pointer. (Remember that the killring-yank-pointer is the name of a variable whose value is a list.)

The measurement of the length is inside the expression:

     (- length (length kill-ring-yank-pointer))

In this expression, the first length is the variable that was assigned the length of the kill ring in the let statement at the beginning of the function. (One might think this function would be clearer if the variable length were named length-of-kill-ring instead; but if you look at the text of the whole function, you will see that it is so short that naming this variable length is not a bother, unless you are pulling the function apart into very tiny pieces as we are doing here.)

   So the line (- length (length kill-ring-yank-pointer)) tells the difference between the

length of the kill ring and the length of the list whose name is kill-ring-yank-pointer.

To see how all this fits into the rotate-yank-pointer function, let’s begin by analyzing the case where kill-ring-yank-pointer points to the first element of the kill ring, just as kill-ring does, and see what happens when rotate-yank-pointer is called with an argument of 1.

In this case, the variable length and the value of the expression (length kill-ring-yank-pointer will be the same since the variable length is the length of the kill ring and the killring-yank-pointer is pointing to the whole kill ring. Consequently, the value of

     (- length (length kill-ring-yank-pointer))

will be zero. Since the value of arg will be 1, this will mean that the value of the whole expression

     (+ arg (- length (length kill-ring-yank-pointer)))

will be 1.

Consequently, the argument to nthcdr will be found as the result of the expression

     (% 1 length)

The % remainder function

To understand (% 1 length), we need to understand %. According to its documentation (which I just found by typing C-h f ^{⟨}\underline{\overline{%}}^{⟩} ^{⟨}\underline{\overline{RET}}^{⟩}), the % function returns the remainder of its first argument divided by its second argument. For example, the remainder of 5 divided by 2 is 1. (2 goes into 5 twice with a remainder of 1.)

What surprises people who don’t often do arithmetic is that a smaller number can be divided by a larger number and have a remainder. In the example we just used, 5 was divided by 2. We can reverse that and ask, what is the result of dividing 2 by 5? If you can use fractions, the answer is obviously 2/5 or .4; but if, as here, you can only use whole numbers, the result has to be something different. Clearly, 5 can go into 2 zero times, but what of the remainder? To see what the answer is, consider a case that has to be familiar from childhood:

• 5 divided by 5 is 1 with a remainder of 0;

• 6 divided by 5 is 1 with a remainder of 1;

• 7 divided by 5 is 1 with a remainder of 2.

• Similarly, 10 divided by 5 is 2 with a remainder of 0;

• 11 divided by 5 is 2 with a remainder of 1;

• 12 divided by 5 is 1 with a remainder of 2.

By considering the cases as parallel, we can see that

• zero divided by 5 must be zero with a remainder of zero;

• 1 divided by 5 must be zero with a remainder of 1;

• 2 divided by 5 must be zero with a remainder of 2;

and so on.

So, in this code, if the value of length is 5, then the result of evaluating

     (% 1 5)

is 1. (I just checked this by placing the cursor after the expression and typing C-x C-e. Indeed, 1 is printed in the echo area.)

Using % in rotate-yank-pointer

When the kill-ring-yank-pointer points to the beginning of the kill ring, and the argument passed to rotate-yank-pointer is 1, the % expression returns 1:

     (- length (length kill-ring-yank-pointer))
          ⇒ 0

therefore,

     (+ arg (- length (length kill-ring-yank-pointer)))
          ⇒ 1

and consequently:

     (% (+ arg (- length (length kill-ring-yank-pointer)))
        length)
          ⇒ 1

regardless of the value of length.

As a result of this, the setq kill-ring-yank-pointer expression simplifies to:

     (setq kill-ring-yank-pointer (nthcdr 1 kill-ring))

What it does is now easy to understand. Instead of pointing as it did to the first element of the kill ring, the kill-ring-yank-pointer is set to point to the second element.

Clearly, if the argument passed to rotate-yank-pointer is two, then the kill-ring-yank-pointer is set to (nthcdr 2 kill-ring); and so on for different values of the argument.

Similarly, if the kill-ring-yank-pointer starts out pointing to the second element of the kill ring, it length is shorter than the length of the kill ring by 1, so the computation of the remainder is based on the expression (% (+ arg 1) length). This means that the kill-ring-yank-pointer is moved from the second element of the kill ring to the third element if the argument passed to rotate-yank-pointer is 1.

Pointing to the last element

The final question is, what happens if the kill-ring-yank-pointer is set to the last element of the kill ring? Will a call to rotate-yank-pointer mean that nothing more can be taken from the kill ring? The answer is no. What happens is different and useful. The kill-ring-yank-pointer is set to point to the beginning of the kill ring instead.

Let’s see how this works by looking at the code, assuming the length of the kill ring is 5

and the argument passed to rotate-yank-pointer is 1. When the kill-ring-yank-pointer

points to the last element of the kill ring, its length is 1. The code looks like this:

     (% (+ arg (- length (length kill-ring-yank-pointer))) length)

When the variables are replaced by their numeric values, the expression looks like this:

     (% (+ 1 (- 5 1)) 5)

This expression can be evaluated by looking at the most embedded inner expression first and working outwards: The value of (- 5 1) is 4; the sum of (+ 1 4) is 5; and the remainder of dividing 5 by 5 is zero. So what rotate-yank-pointer will do is

     (setq kill-ring-yank-pointer (nthcdr 0 kill-ring))

which will set the kill-ring-yank-pointer to point to the beginning of the kill ring.

So what happens with successive calls to rotate-yank-pointer is that it moves the killring-yank-pointer from element to element in the kill ring until it reaches the end; then it jumps back to the beginning. And this is why the kill ring is called a ring, since by jumping back to the beginning, it is as if the list has no end! (And what is a ring, but an entity with no end?)

B.2 yank

After learning about rotate-yank-pointer, the code for the yank function is almost easy. It has only one tricky part, which is the computation of the argument to be passed to rotate-

yank-pointer.

The code looks like this:

     (defun yank (&optional arg)
       "Reinsert the last stretch of killed text.
     More precisely, reinsert the stretch of killed text most
     recently killed OR yanked.
     With just C-U as argument, same but put point in front
     (and mark at end).  With argument n, reinsert the nth
     most recently killed stretch of killed text.
     See also the command \\[yank-pop]."
       (interactive "*P")
       (rotate-yank-pointer (if (listp arg) 0
                              (if (eq arg '-) -1
                                (1- arg))))
       (push-mark (point))
       (insert (car kill-ring-yank-pointer))
       (if (consp arg)
           (exchange-point-and-mark)))

Glancing over this code, we can understand the last few lines readily enough. The mark is pushed, that is, remembered; then the first element (the car) of what the kill-ring-yank-pointer points to is inserted; and then, if the argument passed the function is a cons, point and mark are exchanged so the point is put in the front of the inserted text rather than at the end. This option is explained in the documentation. The function itself is interactive with "*P". This means it will not work on a read-only buffer, and that the unprocessed prefix argument is passed to the function.

Passing the argument

The hard part of yank is understanding the computation that determines the value of the argument passed to rotate-yank-pointer. Fortunately, it is not so difficult as it looks at first sight.

What happens is that the result of evaluating one or both of the if expressions will be a number and that number will be the argument passed to rotate-yank-pointer.

Laid out with comments, the code looks like this:

     (if (listp arg)                         ; if-part

0 ; then-part (if (eq arg '-) ; else-part, inner if

-1 ; inner if’s then-part (1- arg)))) ; inner if’s else-part

This code consists of two if expression, one the else-part of the other.

The first or outer if expression tests whether the argument passed to yank is a list. Oddly enough, this will be true if yank is called without an argument—because then it will be passed the value of nil for the optional argument and an evaluation of (listp nil) returns true! So, if no argument is passed to yank, the argument passed to rotate-yank-pointer inside of yank is zero. This means the pointer is not moved and the first element to which kill-ring-yank-pointer points is inserted, as we expect. Similarly, if the argument for yank is C-u, this will be read as a list, so again, a zero will be passed to rotate-yank-pointer. (C-u produces an unprocessed prefix argument of (4), which is a list of one element.) At the same time, later in the function, this argument will be read as a cons so point will be put in the front and mark at the end of the insertion. (The P argument to interactive is designed to provide these values for the case when an optional argument is not provided or when it is C-u.)

The then-part of the outer if expression handles the case then there is no argument or when it is C-u. The else-part handles the other situations. The else-part is itself another if expression.

The inner if expression tests whether the argument is a minus sign. (This is done by pressing the ^{⟨}\underline{\overline{META}}^{⟩} and - keys at the same time, or the ^{⟨}\underline{\overline{ESC}}^{⟩} key and then the - key). In this case, the rotate-yank-pointer function is passed -1 as an argument. This moves the kill-ring-yank-pointer backwards, which is what is desired.

If the true-or-false-test of the inner if expression is false (that is, if the argument is not a minus sign), the else-part of the expression is evaluated. This is the expression (1- arg). Because of the two if expressions, it will only occur when the argument is a positive number or when it is a negative number (not just a minus sign on its own). What (1- arg) does is decrement the number and return it. (The 1- function subtracts one from its argument.) This means that if the argument to rotate-yank-pointer is 1, it is reduced to zero, which means the first element to which kill-ring-yank-pointer points is yanked back, as you would expect.

Passing a negative argument

Finally, the question arises, what happens if either the remainder function, %, or the nthcdr function is passed a negative argument, as they quite well may?

The answers can be found by a quick test. When (% -1 5) is evaluated, a negative number is returned; and if nthcdr is called with a negative number, it returns the same value as if it were called with a first argument of zero. This can be seen be evaluating the following code.

Here the ‘⇒’ points to the result of evaluating the code preceding it. This was done by positioning the cursor after the code and typing C-x C-e (eval-last-sexp) in the usual fashion. You can do this if you are reading this in Info inside of GNU Emacs.

     (% -1 5) ⇒ -1
     (setq animals '(cats dogs elephants))
          ⇒ (cats dogs elephants)
     (nthcdr 1 animals)
          ⇒ (dogs elephants)
     (nthcdr 0 animals)
          ⇒ (cats dogs elephants)
     (nthcdr -1 animals)
          ⇒ (cats dogs elephants)

So, if a minus sign or a negative number is passed to yank, the kill-ring-yank-point is rotated backwards until it reaches the beginning of the list. Then it stays there. Unlike the other case, when it jumps from the end of the list to the beginning of the list, making a ring, it stops. This makes sense. You often want to get back to the most recently clipped out piece of text, but you don’t usually want to insert text from as many as thirty kill commands ago. So you need to work through the ring to get to the end, but won’t cycle around it inadvertently if you are trying to come back to the beginning.

Incidentally, any number passed to yank with a minus sign preceding it will be treated as −1. This is evidently a simplification for writing the program. You don’t need to jump back towards the beginning of the kill ring more than one place at a time and doing this is easier than writing a function to determine the magnitude of the number that follows the minus sign.

B.3 yank-pop

After understanding yank, the yank-pop function is easy. Leaving out the documentation to save space, it looks like this:

     (defun yank-pop (arg)
       (interactive "*p")
       (if (not (eq last-command 'yank))
           (error "Previous command was not a yank"))
       (setq this-command 'yank)
       (let ((before (< (point) (mark))))
         (delete-region (point) (mark))
         (rotate-yank-pointer arg)
         (set-mark (point))
         (insert (car kill-ring-yank-pointer))
         (if before (exchange-point-and-mark))))

The function is interactive with a small ‘p’ so the prefix argument is processed and passed to the function. The command can only be used after a previous yank; otherwise an error message is sent. This check uses the variable last-command which is discussed elsewhere. (See Section 8.5, “copy-region-as-kill”, page 73.)

The let clause sets the variable before to true or false depending whether point is before or after mark and then the region between point and mark is deleted. This is the region that was just inserted by the previous yank and it is this text that will be replaced. Next the kill-

yank-pop  179

ring-yank-pointer is rotated so that the previously inserted text is not reinserted yet again. Mark is set at the beginning of the place the new text will be inserted and then the first element to which kill-ring-yank-pointer points is inserted. This leaves point after the new text. If in the previous yank, point was left before the inserted text, point and mark are now exchanged so point is again left in front of the newly inserted text. That is all there is to it!

180 Appendix B: Handling the Kill Ring

Appendix C A Graph with Labelled Axes

Printed axes help you understand a graph. They convey scale. In an earlier chapter (see Chapter 15, “Readying a Graph”, page 141), we wrote the code to print the body of a graph. Here we write the code for print and labelling vertical and horizontal axes, along with the body itself.

Since insertions fill a buffer to the right and below point, the new graph printing function should first print the Y or vertical axis, then the body of the graph, and finally the X or horizontal axis. This sequence lays out for us the contents of the function:

1. Set up code.

2. Print Y axis.

3. Print body of graph.

4. Print X axis.

Here is an example of how a finished graph should look:

         10 -
                       *
                       *  *
                       *  **
                       *  ***
          5 -      *   *******
                 * *** *******
                 *************
               ***************
          1 - ****************
              |   |    |    |
              1   5   10   15

In this graph, both the vertical and the horizontal axes are labeled with numbers. However, in some graphs, the horizontal axis is time and would be better labeled with months, like this:

          5 -      *
                 * ** *
                 *******
               ********** **
          1 - **************
              |    ^|
              Jan  June   Jan

Indeed, with a little thought, we can easily come up with a variety of vertical and horizontal labelling schemes. Our task could become complicated. But complications breed confusion. Rather than permit this, it is better choose a simple labelling scheme for our first effort, and to modify or replace it later.

These considerations suggest the following outline for the print-graph function:

     (defun print-graph (numbers-list)

"documentation..."

       (let ((height  ...
             ...))
         (print-Y-axis height ... )
         (graph-body-print numbers-list)
         (print-X-axis ... )))

We can work on each part of the print-graph function definition in turn.

C.1 The print-graph Varlist

In writing the print-graph function, the first task is to write the varlist in the let expression. (We will leave aside for the moment any thoughts about making the function interactive or about the contents of its documentation string.)

The varlist should set several values. Clearly, the top of the label for the vertical axis must be at least the height of the graph, which means that we must obtain this information here. Note that the print-graph-body function also requires this information. There is no reason to calculate the height of the graph in two different places, so we should change for print-graph-body from the way we defined it earlier to take advantage of the calculation.

Similarly, both the function for printing the X axis labels and the print-graph-body function need to learn the value of the width of each symbol. We can perform the calculation here and change the definition for print-graph-body from the way we defined it in the previous chapter.

The length of the label for the horizontal axis must be at least as long as the graph. However, this information is used only in the function that prints the horizontal axis, so it does not need to be calculated here.

These thoughts lead us directly to the following form for the varlist in the let for print-

graph:
     (let ((height (apply 'max numbers-list)) ; First version.
           (symbol-width (length graph-blank)))

As we shall see, this expression is not quite right.

C.2 The print-Y-axis Function

The job of the print-Y-axis function is to print a label for the vertical axis that looks like this:

         10 -

          5 -

          1 -

The function should be passed the height of the graph, and then should construct and insert the appropriate numbers and marks.

It is easy enough to see in the figure what the Y axis label should look like; but to say in words, and then to write a function definition to do the job is another matter. It is not quite true to say that we want a number and a tick every five lines: there are only three lines between the ‘1’ and the ‘5’ (lines 2, 3, and 4), but four lines between the ‘5’ and the ‘10’ (lines 6, 7, 8, and 9). It is better to say that we want a number and a tick mark on the base line (number 1) and then that we want a number and a tick on the fifth line from the bottom and on every line that is a multiple of five.

The next issue is what height the label should be. Suppose the maximum height of tallest column of the graph is seven. Should the highest label on the Y axis be ‘5 -’, and should the graph stick up above the label? Or should the highest label be ‘7 -’, and mark the peak of the graph? Or should the highest label be 10 -, which is a multiple of five, and be higher than the topmost value of the graph?

The latter form is preferred. Most graphs are drawn within rectangles whose sides are an integral number of steps long—5, 10, 15, and so on for a step distance of five. But as soon as we decide to use a step height for the vertical axis, we discover that the simple expression in the varlist for computing the height is wrong. The expression is (apply 'max numbers-list). This returns the precise height, not the maximum height plus whatever is necessary to round up to the nearest multiple of five. A more complex expression is required.

As usual in cases like this, a complex problem becomes simpler if it is divided into several smaller problems.

First, consider the case when the highest value of the graph is an integral multiple of five— when it is 5, 10, 15 ,or some higher multiple of five. In this case, we can use this value as the Y axis height.

A fairly simply way to determine whether a number is a multiple of five is to divide it by five and see if the division results in a remainder. If there is no remainder, the number is a multiple of five. Thus, seven divided by five has a remainder of two, and seven is not an integral multiple of five. Put in slightly different language, more reminiscent of the classroom, five goes into seven once, with a remainder of two. However, five goes into ten twice, with no remainder: ten is an integral multiple of five.

C.2.1 Side Trip: Compute a Remainder

In Lisp, the function for computing a remainder is %. The function returns the remainder of its first argument divided by its second argument. As it happens, % is a function in Emacs Lisp that you cannot discover using apropos: you find nothing if you type M-x apropos ^{⟨}\underline{\overline{RET}}^{⟩} remainder ^{⟨}\underline{\overline{RET}}^{⟩}. The only way to learn of the existence of % is to read about it in a book such as this or in the Emacs Lisp sources. The % function is used in the code for rotate-yank-pointer, which is described in an appendix. (See Section B.1.1, “The Body of rotate-yank-pointer”, page 172.)

You can try the % function by evaluating the following two expressions:

     (% 7 5) (% 10 5)

The first expression returns 2 and the second expression returns 0.

To test whether the returned value is zero or some other number, we can use the zerop function. This function returns t if its argument, which must be a number, is zero.

     (zerop (% 7 5))
          ⇒ nil
     (zerop (% 10 5))
          ⇒ t

Thus, the following expression will return t if the height of the graph is evenly divisible by five:

     (zerop (% height 5))

(The value of height, of course, can be found from (apply 'max numbers-list).)

On the other hand, if the value of height is not a multiple of five, we want to reset the value to the next higher multiple of five. This is straightforward arithmetic using functions with which we are already familiar. First, we divide the value of height by five to determine how many times five goes into the number. Thus, five goes into twelve twice. If we add one to this quotient and multiply by five, we will obtain the value of the next multiple of five that is larger than the height. Five goes into twelve twice. Add one to two, and multiply by five; the result is fifteen, with is the next multiple of five that is higher than twelve. The Lisp expression for this is:

     (* (1+ (/ height 5)) 5)

For example, if you evaluate the following, the result is 15:

     (* (1+ (/ 12 5)) 5)

All through this discussion, we have been using ‘five’ as the value for spacing labels on the Y axis; but we may want to use some other value. For generality, we should replace ‘five’ with a variable to which we can assign a value. The best name I can think of for this variable is Y-axis-label-spacing. Using this term, and an if expression, we produce the following:

     (if (zerop (% height Y-axis-label-spacing))
         height

;; else

       (* (1+ (/ height Y-axis-label-spacing))
          Y-axis-label-spacing))

This expression returns the value of height itself if the height is an even multiple of the value of the Y-axis-label-spacing or else it computes and returns a value of height that is equal to the next higher multiple of the value of the Y-axis-label-spacing.

We can now include this expression in the let expression of the print-graph function (after first setting the value of Y-axis-label-spacing):

     (defvar Y-axis-label-spacing 5
       "Number of lines from one Y axis label to next.")
     ...
     (let* ((height (apply 'max numbers-list))
            (height-of-top-line
             (if (zerop (% height Y-axis-label-spacing))
                 height

;; else

               (* (1+ (/ height Y-axis-label-spacing))
                  Y-axis-label-spacing)))
            (symbol-width (length graph-blank))))
     ...

(Note use of the let* function: the initial value of height is computed once by the (apply 'max numbers-list) expression and then the resulting value of height is used to compute its final value. See “The let* expression”, page 106, for more about let*.)

C.2.2 Construct a Y Axis Element

When we print the vertical axis, we want to insert strings such as ‘5 -’ and ‘10 - ’ every five lines. Moreover, we want the numbers and dashes to line up, so shorter numbers must be padded with leading spaces. If some of the strings use two digit numbers, the strings with single digit numbers must include a leading blank space before the number.

To figure out the length of the number, the length function is used. But the length function works only with a string, not with a number. So the number has to be converted from being a number to being a string. This is done with the int-to-string function. For example,

     (length (int-to-string 35))
          ⇒ 2
     (length (int-to-string 100))
          ⇒ 3

In addition, in each label, each number is followed by a string such as ‘ - ’, which we will call the Y-axis-tic marker. This variable is defined with defvar:

     (defvar Y-axis-tic " - "
        "String that follows number in a Y axis label.")

The length of the Y label is the sum of the length of the Y axis tick mark and the length of the number of the top of the graph.

     (length (concat (int-to-string height) Y-axis-tic)))

This value will be calculated by the print-graph function in its varlist as full-Y-label-width and passed on. (Note that we did not think to include this in the varlist when we first proposed it.)

To make a complete vertical axis label, a tick mark is concatenated with a number; and the two together may be preceded by one or more spaces depending on how long the number is. The label consists of three parts: the (optional) leading spaces, the number, and the tic mark. The function is passed the value of the number for the specific row, and the value of the width of the top line, which is calculated (just once) by print-graph.

     (defun Y-axis-element (number full-Y-label-width)
       "Construct a NUMBERed label element.
     A numbered element looks like this `  5 - ',
     and is padded as needed so all line up with
     the element for the largest number."
       (let* ((leading-spaces
              (- full-Y-label-width
                 (length
                  (concat (int-to-string number)
                          Y-axis-tic)))))
         (concat
          (make-string leading-spaces ? )
          (int-to-string number)
          Y-axis-tic)))

The Y-axis-element function concatenates together the leading spaces, if any; the number, as a string; and the tick mark.

To figure out how many leading spaces the label will need, the function subtracts the actual length of the label—the length of the number plus the length of the tic mark—from the desired label width.

Blank spaces are inserted using the make-string function. This function takes two arguments: the first tells it how long the string will be and the second is a symbol for the character to insert, in a special format. In this case, the format is a question mark followed by a blank space, like this, ‘? ’. See section “Character Type” in The GNU Emacs Lisp Reference Manual, for a description of the syntax for characters.

The int-to-string function is used in the concatenation expression, to convert the number to a string that is concatenated with the leading spaces and the tic mark.

C.2.3 Create a Y Axis Column

The preceding functions provide all the tools needed to construct a function that generates a list of numbered and blank strings to insert as the label for the vertical axis:

     (defun Y-axis-column (height width-of-label)
       "Construct list of Y axis labels and blank strings.
     For HEIGHT of line above base and WIDTH-OF-LABEL."
       (let (Y-axis)
         (while (> height 1)
           (if (zerop (% height Y-axis-label-spacing))

;; Insert label.

               (setq Y-axis
                     (cons
                      (Y-axis-element height width-of-label)
                      Y-axis))

;; Else, insert blanks.

             (setq Y-axis
                   (cons
                    (make-string width-of-label ? )
                    Y-axis)))
           (setq height (1- height)))

;; Insert base line.

         (setq Y-axis
               (cons (Y-axis-element 1 width-of-label) Y-axis))
         (nreverse Y-axis)))

In this function, we start with the value of height and repetitively subtract one from its value. After each subtraction, we test to see whether the value is an integral multiple of the Y-axis-label-spacing. If it is, we construct a numbered label using the Y-axis-element function; if not, we construct a blank label using the make-string function. The base line consists of the number one followed by a tic mark.

C.2.4 Final Version of print-Y-axis

The list constructed by the Y-axis-column function is passed to the print-Y-axis function, which inserts the list as a column.

     (defun print-Y-axis
       (height full-Y-label-width &optional vertical-step)
       "Insert Y axis using HEIGHT and FULL-Y-LABEL-WIDTH.
     Height must be the  maximum height of the graph.
     Full width is the width of the highest label element.
     Optionally, print according to VERTICAL-STEP."
     ;; Value of height and full-Y-label-width
     ;; are passed by `print-graph'.

       (let ((start (point)))
         (insert-rectangle
          (Y-axis-column height full-Y-label-width vertical-step))

;; Place point ready for inserting graph.

         (goto-char start)

;; Move point forward by value of full-Y-label-width

         (forward-char full-Y-label-width)))

The print-Y-axis uses the insert-rectangle function to insert the Y axis labels created by the Y-axis-column function. In addition, it places point at the correct position for printing the body of the graph.

   You can test print-Y-axis:

1. Install

         Y-axis-label-spacing
         Y-axis-tic
         Y-axis-element
         Y-axis-columnprint-Y-axis

2. Copy the following expression:

         (print-Y-axis 12 5)

3. Switch to the ‘*scratch*’ buffer and place the cursor where you want the axis labels to start.

 4. Type M-: (eval-expression).

5. Yank the graph-body-print expression into the minibuffer with C-y (yank).

6. Press ^{⟨}\underline{\overline{RET}}^{⟩} to evaluate the expression.

Emacs will print labels vertically, the top one being ‘10 - ’. (The print-graph function will pass the value of height-of-top-line, which in this case would be 15.)

C.3 The print-X-axis Function

X axis labels are much like Y axis labels, except that the tics are on a line above the numbers. Labels should look like this:

         |   |    |    | 1   5   10   15

The first tic is under the first column of the graph and is preceded by several blank spaces. These spaces provide room in rows above for the Y axis labels. The second, third, fourth, and subsequent tics are all spaced equally, according to the value of X-axis-label-spacing.

The second row of the X axis consists of numbers, preceded by several blank spaces and also separated according to the value of the variable X-axis-label-spacing.

The value of the variable X-axis-label-spacing should itself be measured in units of symbol-width, since you may want to change the width of the symbols that you are using to print the body of the graph without changing the ways the graph is labeled.

The print-X-axis function is constructed in more or less the same fashion as the print-Y-axis function except that it has two lines: the line of tic marks and the numbers. We will write a separate function to print each line and then combine them within the print-X-axis function.

This is a three step process:

1. Write a function to print the X axis tic marks, print-X-axis-tic-line.

2. Write a function to print the X numbers, print-X-axis-numbered-line.

3. Write a function to print both lines, the print-X-axis function, using print-X-axis-tic-

    line and print-X-axis-numbered-line.

C.3.1 X Axis Tic Marks

The first function should print the X axis tic marks. We must specify the tic marks themselves and their spacing:

     (defvar X-axis-label-spacing
       (if (boundp 'graph-blank)
           (* 5 (length graph-blank)) 5)
       "Number of units from one X axis label to next.")

(Note that the value of graph-blank is set by another defvar. The boundp predicate checks whether it has already been set; boundp returns nil if it has not. If graph-blank were unbound and we did not use this conditional construction, we would receive an error message saying

‘Symbol's value as variable is void’.)
     (defvar X-axis-tic-symbol "|"
       "String to insert to point to a column in X axis.")

The goal is to make a line that looks like this:

            |   |    |    |

The first tic is indented so that it is under the first column, which is indented to provide space for the Y axis labels.

A tic element consists of the blank spaces that stretch from one tic to the next plus a tic symbol. The number of blanks is determined by the width of the tic symbol and the X-axis-

label-spacing.

The code looks like this:

     ;;; X-axis-tic-element
     ...
     (concat
      (make-string

;; Make a string of blanks.

       (-  (* symbol-width X-axis-label-spacing)
           (length X-axis-tic-symbol))
       ? )

;; Concatenate blanks with tic symbol.

      X-axis-tic-symbol)
     ...

Next, we determine how many blanks are needed to indent the first tic mark to the first column of the graph. This uses the value of full-Y-label-width passed it by the print-graph function.

The code to make X-axis-leading-spaces looks like this:

;; X-axis-leading-spaces
...
(make-string full-Y-label-width ? )
...

We also need to determine the length of the horizontal axis, which is the length of the numbers list, and the number of tics in the horizontal axis:

;; X-length
...
(length numbers-list)
;; tic-width
...
(* symbol-width X-axis-label-spacing)
;; number-of-X-tics
(if (zerop (% (X-length tic-width)))
    (/ (X-length tic-width))
  (1+ (/ (X-length tic-width))))

All this leads us directly to the function for printing the X axis tic line:

(defun print-X-axis-tic-line
  (number-of-X-tics X-axis-leading-spaces X-axis-tic-element)
  "Print tics for X axis."
    (insert X-axis-leading-spaces)
    (insert X-axis-tic-symbol)  ; Under first column.

;; Insert second tic in the right spot.

    (insert (concat
             (make-string
              (-  (* symbol-width X-axis-label-spacing)

;; Insert white space up to second tic symbol.

                  (* 2 (length X-axis-tic-symbol)))
              ? )
             X-axis-tic-symbol))

;; Insert remaining tics.

    (while (> number-of-X-tics 1)
      (insert X-axis-tic-element)
      (setq number-of-X-tics (1- number-of-X-tics))))

The line of numbers is equally straightforward:

First, we create a numbered element with blank spaces before each number:

(defun X-axis-element (number)
  "Construct a numbered X axis element."
  (let ((leading-spaces
         (-  (* symbol-width X-axis-label-spacing)
             (length (int-to-string number)))))
    (concat (make-string leading-spaces ? )
            (int-to-string number))))

Next, we create the function to print the numbered line, starting with the number “1” under the first column:

     (defun print-X-axis-numbered-line
       (number-of-X-tics X-axis-leading-spaces)
       "Print line of X-axis numbers"
       (let ((number X-axis-label-spacing))
         (insert X-axis-leading-spaces)
         (insert "1")
         (insert (concat
                  (make-string

;; Insert white space up to next number.

                   (-  (* symbol-width X-axis-label-spacing) 2)
                   ? )
                  (int-to-string number)))

;; Insert remaining numbers.

         (setq number (+ number X-axis-label-spacing))
         (while (> number-of-X-tics 1)
           (insert (X-axis-element number))
           (setq number (+ number X-axis-label-spacing))
           (setq number-of-X-tics (1- number-of-X-tics)))))
   Finally, we need to write the print-X-axis that uses print-X-axis-tic-line and print-
X-axis-numbered-line.

The function must determine the local values of the variables used by both print-X-axis-tic-line and print-X-axis-numbered-line, and then it must call them. Also, it must print the carriage return that separates the two lines.

The function consists of a varlist that specifies five local variables, and calls to each of the two line printing functions:

     (defun print-X-axis (numbers-list)
       "Print X axis labels to length of NUMBERS-LIST."
       (let* ((leading-spaces
               (make-string full-Y-label-width ? ))
            ;; symbol-width is provided by graph-body-print
            (tic-width (* symbol-width X-axis-label-spacing))
            (X-length (length numbers-list))
            (X-tic
             (concat
              (make-string

;; Make a string of blanks.

               (-  (* symbol-width X-axis-label-spacing)
                   (length X-axis-tic-symbol))
               ? )

;; Concatenate blanks with tic symbol.

              X-axis-tic-symbol))
            (tic-number
             (if (zerop (% X-length tic-width))
                 (/ X-length tic-width)
               (1+ (/ X-length tic-width)))))

      (print-X-axis-tic-line tic-number leading-spaces X-tic)
      (insert "\n")
      (print-X-axis-numbered-line tic-number leading-spaces)))
You can test print-X-axis:
1. Install X-axis-tic-symbol, X-axis-label-spacing, print-X-axis-tic-line, as well as
 X-axis-element, print-X-axis-numbered-line, and print-X-axis.

2. Copy the following expression:

      (progn
       (let ((full-Y-label-width 5)
             (symbol-width 1))
         (print-X-axis
          '(1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16))))

3. Switch to the ‘*scratch*’ buffer and place the cursor where you want the axis labels to start.

4. Type M-: (eval-expression).

5. Yank the test expression into the minibuffer with C-y (yank).

6. Press ^{⟨}\underline{\overline{RET}}^{⟩} to evaluate the expression.

Emacs will print the horizontal axis like this:

       |   |    |    |    | 1   5   10   15   20

C.4 Printing the Whole Graph

Now we are nearly ready to print the whole graph.

The function to print the graph with the proper labels follows the outline we created earlier (see Appendix C, “A Graph with Labelled Axes”, page 181), but with additions.

Here is the outline:

  (defun print-graph (numbers-list)

"documentation..."

    (let ((height  ...
          ...))
      (print-Y-axis height ... )
      (graph-body-print numbers-list)
      (print-X-axis ... )))

The final version is different from what we planned in two ways: first, it contains additional values calculated once in the varlist; second, it carries an option to specify the labels’ increment per row. This latter feature turns out to be essential; otherwise a graph may have more rows than fit on a display or on a sheet of paper.

This new feature requires a change to the Y-axis-column function, to add vertical-step to it. The function looks like this:

;;; Final version.

(defun Y-axis-column
  (height width-of-label &optional vertical-step)
  "Construct list of labels for Y axis.
HEIGHT is maximum height of graph.
WIDTH-OF-LABEL is maximum width of label.
VERTICAL-STEP, an option, is a positive integer
that specifies how much a Y axis label increments
for each line.  For example, a step of 5 means
that each line is five units of the graph."
  (let (Y-axis
        (number-per-line (or vertical-step 1)))
    (while (> height 1)
      (if (zerop (% height Y-axis-label-spacing))

;; Insert label.

          (setq Y-axis
                (cons
                 (Y-axis-element
                  (* height number-per-line)
                  width-of-label)
                 Y-axis))

;; Else, insert blanks.

        (setq Y-axis
              (cons
               (make-string width-of-label ? )
               Y-axis)))
      (setq height (1- height)))

;; Insert base line.

    (setq Y-axis (cons (Y-axis-element
                        (or vertical-step 1)
                        width-of-label)
                       Y-axis))
    (nreverse Y-axis)))

The values for the maximum height of graph and the width of a symbol are computed by print-graph in its let expression; so graph-body-print must be changed to accept them.

;;; Final version.

(defun graph-body-print (numbers-list height symbol-width)
  "Print a bar graph of the NUMBERS-LIST.
The numbers-list consists of the Y-axis values.
HEIGHT is maximum height of graph.
SYMBOL-WIDTH is number of each column."
  (let (from-position)
    (while numbers-list
      (setq from-position (point))
      (insert-rectangle
       (column-of-graph height (car numbers-list)))
      (goto-char from-position)
      (forward-char symbol-width)

Testing print-graph  193

;; Draw graph column by column.

      (sit-for 0)
      (setq numbers-list (cdr numbers-list)))

;; Place point for X axis labels.

    (forward-line height)
    (insert "\n")))

Finally, the code for the print-graph function:

;;; Final version.

(defun print-graph
  (numbers-list &optional vertical-step)
  "Print labelled bar graph of the NUMBERS-LIST.
The numbers-list consists of the Y-axis values.
Optionally, VERTICAL-STEP, a positive integer,
specifies how much a Y axis label increments for
each line.  For example, a step of 5 means that
each row is five units."
  (let* ((symbol-width (length graph-blank))

;; height is both the largest number ;; and the number with the most digits.

         (height (apply 'max numbers-list))
         (height-of-top-line
          (if (zerop (% height Y-axis-label-spacing))
              height

;; else

            (* (1+ (/ height Y-axis-label-spacing))
               Y-axis-label-spacing)))
         (vertical-step (or vertical-step 1))
         (full-Y-label-width
          (length
           (concat
            (int-to-string
             (* height-of-top-line vertical-step))
            Y-axis-tic))))
    (print-Y-axis
     height-of-top-line full-Y-label-width vertical-step)
    (graph-body-print
     numbers-list height-of-top-line symbol-width)
    (print-X-axis numbers-list)))

C.4.1 Testing print-graph

We can test the print-graph function with a short list of numbers

1. Install the final versions of Y-axis-column, graph-body-print, and print-graph (in ad-

dition to the rest of the code.)

2. Copy the following expression:

    (print-graph '(3 2 5 6 7 5 3 4 6 4 3 2 1))

3. Switch to the ‘*scratch*’ buffer and place the cursor where you want the axis labels to start.

 4. Type M-: (eval-expression).

5. Yank the test expression into the minibuffer with C-y (yank).

6. Press ^{⟨}\underline{\overline{RET}}^{⟩} to evaluate the expression.

Emacs will print a graph that looks like this:

     10 -

              *
             **   *
      5 -   ****  *
            **** ***
          * *********
          ************
      1 - *************
          |   |    |    |
          1   5   10   15

On the other hand, if you pass print-graph a vertical-step value of 2, by evaluating this expression:

     (print-graph '(3 2 5 6 7 5 3 4 6 4 3 2 1) 2)

The graph looks like this:

     20 -

              *
             **   *
     10 -   ****  *
            **** ***
          * *********
          ************
      2 - *************
          |   |    |    |
          1   5   10   15

(A question: is the ‘2’ on the bottom of the vertical axis a bug or a feature? If you think it is a bug, and should be a ‘1’ instead, (or even a ‘0’), you can modify the sources.)

C.4.2 Graphing Numbers of Words and Symbols

Now for the graph for which all this code was written: a graph that shows how many function definitions contain fewer than 10 words and symbols, how many contain between 10 and 19 words and symbols, how many contain between 20 and 29 words and symbols, and so on.

This is a multi-step process. First make sure you have loaded all the requisit code.

It is a good idea to reset the value of top-of-ranges in case you have sent it to some different value. You can evaluate the following:

     (setq top-of-ranges
      '(10  20  30  40  50
        60  70  80  90 100
       110 120 130 140 150
       160 170 180 190 200
       210 220 230 240 250
       260 270 280 290 300)

Next create a list of the number of words and symbols in each range.

Evaluate the following:

     (setq list-for-graph
            (defuns-per-range
              (sort
               (recursive-lengths-list-many-files
                (directory-files "/usr/local/emacs/lisp"
                                 t ".+el$"))
               '<)
              top-of-ranges))

On my machine, this takes about an hour. It looks though 303 Lisp files in my copy of Emacs version 19.23. After all that computing, the list-for-graph has this value:

     (537 1027 955 785 594 483 349 292 224 199 166 120 116 99 90 80 67 48 52 45 41 33 28 26 25 20 12 28 11 13 220)

This means that my copy of Emacs has 537 function definitions with fewer than 10 words or symbols in them, 1,027 function definitions with 10 to 19 words or symbols in them, 955 function definitions with 20 to 29 words or symbols in them, and so on.

Clearly, just by looking at this list we can see that most function definitions contain ten to thirty words and symbols.

Now for printing. We do not want to print a graph that is 1,030 lines high … Instead, we should print a graph that is fewer than twenty-five lines high. A graph that height can be displayed on almost any monitor, and easily printed on a sheet of paper.

This means that each value in list-for-graph must be reduced to one-fiftieth it present value.

Here is a short function to do just that, using two functions we have not yet seen, mapcar

and lambda.
     (defun one-fiftieth (full-range)
       "Return list, each number one-fiftieth of previous."
      (mapcar '(lambda (arg) (/ arg 50)) full-range))

A lambda Expression

lambda is the symbol for an anonymous function, a function without a name. Every time you use an anonymous function, you need to include its whole body.

Thus,

     (lambda (arg) (/ arg 50))

is a function definition that says ‘return the value resulting from dividing whatever is passed to me as arg by 50’.

Earlier, for example, we had a function multiply-by-seven; it multiplied its argument by 7. This function is similar, except it divides its argument by 50; and, it has no name. The anonymous equivalent of multiply-by-seven is:

     (lambda (number) (* 7 number))

(See Section 3.1, “The defun Special Form”, page 21.)

If we want to multiply 3 by 7, we can write:

     (multiply-by-seven 3)
      \_______________/ ^
              |         |
           function  argument

This expression returns 21.

Similarly, we can write:

     ((lambda (number) (* 7 number)) 3)
      \____________________________/ ^
                    |                |
           anonymous function     argument

If we want to divide 100 by 50, we can write:

     ((lambda (arg) (/ arg 50)) 100)
      \______________________/  \_/
                  |              |
         anonymous function   argument

This expression returns 2. The 100 is passed to the function, which divides that number by 50.

See section “Lambda Expressions” in The GNU Emacs Lisp Reference Manual, for more about lambda. Lisp and lambda expressions derive from the Lambda Calculus.

The mapcar Function

mapcar is a function that calls its first argument with each element of its second argument, in turn. The second argument must be a sequence.

For example,

     (mapcar '1+ '(2 4 6)) ⇒ (3 5 7)

The function 1+ which adds one to its argument, is executed on each element of the list, and a new list is returned.

Contrast this with apply, which applies its first argument to all the remaining. (See Chapter 15, “Readying a Graph”, page 141, for a explanation of apply.)

In the definition of one-fiftieth, the first argument is the anonymous function:

     (lambda (arg) (/ arg 50))

and the second argument is full-range, which will be bound to list-for-graph.

The whole expression looks like this:

     (mapcar '(lambda (arg) (/ arg 50)) full-range))

See section “Mapping Functions” in The GNU Emacs Lisp Reference Manual, for more about

mapcar.

Using the one-fiftieth function, we can generate a list in which each element is one-fiftieth the size of the corresponding element in list-for-graph.

     (setq fiftieth-list-for-graph
           (one-fiftieth list-for-graph))

The resulting list looks like this:

     (10 20 19 15 11 9 6 5 4 3 3 2 2 1 1 1 1 0 1 0 0 0 0 0 0 0 0 0 0 0 4)

This we are almost ready to print! (We also notice the loss of information: many of the higher ranges are 0, meaning that fewer than 50 defuns had that many words or symbols—but not necessarily meaning that none had that many words or symbols.)

Another Bug … Most Insidious

I said ‘almost ready to print’! Of course, there is a bug in the print-graph function …

It has a vertical-step option, but not a horizontal-step option. The top-of-range scale

goes from 10 to 300 by tens. But the print-graph function will print only by ones.

This is a classic example of what some consider the most insidious type of bug, the bug of omission. This is not the kind of bug you can find by studying the code, for it is not in the code; it is an omitted feature. Your best actions are to try your program early and often; and try to arrange, as much as you can, to write code that is easy to understand and easy to change. Try to be aware, whenever you can, that whatever you have written, will be rewritten, if not soon, eventually. A hard maxim to follow.

It is the print-X-axis-numbered-line function that needs the work; and then the print-X-axis and the print-graph functions need to be adapted. Not much needs to be done; there is one nicety: the numbers ought to line up under the tic marks. This takes a little thought.

   Here is the corrected print-X-axis-numbered-line:
     (defun print-X-axis-numbered-line
       (number-of-X-tics X-axis-leading-spaces
        &optional horizontal-step)
       "Print line of X-axis numbers"
       (let ((number X-axis-label-spacing)
             (horizontal-step (or horizontal-step 1)))

(insert X-axis-leading-spaces)

;; Delete extra leading spaces.

(delete-char
 (- (1-
     (length (int-to-string horizontal-step)))))
(insert (concat
         (make-string

;; Insert white space.

          (-  (* symbol-width
                 X-axis-label-spacing)
              (1-
               (length
                (int-to-string horizontal-step)))
              2)
          ? )
         (int-to-string
          (* number horizontal-step))))

;; Insert remaining numbers.

(setq number (+ number X-axis-label-spacing))
(while (> number-of-X-tics 1)
  (insert (X-axis-element
           (* number horizontal-step)))
  (setq number (+ number X-axis-label-spacing))
  (setq number-of-X-tics (1- number-of-X-tics)))))

If you are reading this in Info, you can see the new versions of print-X-axis print-graph and evaluate them. If you are reading this in a printed book, you can see the changed lines here (the full text is too much to print).

(defun print-X-axis (numbers-list horizontal-step)
...
(print-X-axis-numbered-line
 tic-number leading-spaces horizontal-step))
(defun print-graph
(numbers-list
&optional vertical-step horizontal-step)
...
(print-X-axis numbers-list horizontal-step))

C.4.3 The Printed Graph

When made and installed, you can call the print-graph command like this:

(print-graph fiftieth-list-for-graph 50 10)

Here is the graph:

1000 -  *
**
**
**
**
750 -  ***
***
***
***
****
500 - *****
******
******
******
*******
250 - ********
*********                     *
***********                   *
*************                 *
50 - ***************** *           *
|   |    |    |    |    |    |    |
10  50  100  150  200  250  300  350

The largest group of functions contain 10 – 19 words and symbols each.

200 Appendix C: A Graph with Labelled Axes

Index 201

Index

%

% (remainder function) . . . . . . . . . . . . . . . . . . . . . . . . 183

(

(debug) in code. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164

*

* (multiplication) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

* for read-only buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 ‘*Backtrace*’ buffer . . . . . . . . . . . . . . . . . . . . . . . . . . 162 ‘*scratch*’ buffer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

.

‘.emacs’ file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 ‘.emacs’ file, beginning of . . . . . . . . . . . . . . . . . . . . . . 150

/

/ (division) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

>

> (greater than) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

<

<= (less than or equal) . . . . . . . . . . . . . . . . . . . . . . . . . . 91

A

add-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 Anonymous function . . . . . . . . . . . . . . . . . . . . . . . . . . 195 append-to-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 apply . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 apropos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Argument as local variable . . . . . . . . . . . . . . . . . . . . . . 94 ‘argument’ defined. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 ‘argument list’ defined . . . . . . . . . . . . . . . . . . . . . . . . 22 Argument, wrong type of . . . . . . . . . . . . . . . . . . . . . . . 10 Arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Arguments’ data types . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Arguments, variable number of . . . . . . . . . . . . . . . . . . 10 Asterisk for read-only buffer . . . . . . . . . . . . . . . . . . . . 46 Auto Fill mode turned on. . . . . . . . . . . . . . . . . . . . . . 152 autoload . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 Automatic mode selection . . . . . . . . . . . . . . . . . . . . . 151 Axis, print horizontal. . . . . . . . . . . . . . . . . . . . . . . . . . 187 Axis, print vertical . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182

B

Backtrace buffer for debugging . . . . . . . . . . . . . . . . . 162 beginning-of-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . 50 ‘bind’ defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 ‘body’ defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Body of graph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Buffer size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Buffer, history of word. . . . . . . . . . . . . . . . . . . . . . . . . . 16 buffer-file-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 buffer-menu, bound to key. . . . . . . . . . . . . . . . . . . . 154 buffer-name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Bug, most insidious type. . . . . . . . . . . . . . . . . . . . . . . 197 Byte compiling. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

C

C language primitives . . . . . . . . . . . . . . . . . . . . . . . . . . 21 C, a digression into. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 ‘call’ defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 cancel-debug-on-entry . . . . . . . . . . . . . . . . . . . . . . . 163 car, introduced. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 cdr, introduced. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 Changing a function definition . . . . . . . . . . . . . . . . . . 23 Clipping text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 Code installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 ‘command’ defined. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Comments in Lisp code . . . . . . . . . . . . . . . . . . . . . . . . . 23 Common Lisp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . x compare-windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 concat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 cond . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 Conditional ’twixt Emacs 18 and 19 . . . . . . . . . . . . 157 Conditional with if . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 cons, introduced. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 copy-region-as-kill . . . . . . . . . . . . . . . . . . . . . . . . . . 73 copy-to-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 Count words recursively . . . . . . . . . . . . . . . . . . . . . . . 120 count-words-in-defun . . . . . . . . . . . . . . . . . . . . . . . . 128 count-words-region . . . . . . . . . . . . . . . . . . . . . . . . . . 115 Counting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Counting words in a defun . . . . . . . . . . . . . . . . 125, 127 current-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Customizing your ‘.emacs’ file . . . . . . . . . . . . . . . . . 149 Cutting and storing text . . . . . . . . . . . . . . . . . . . . . . . . 65

D

Data types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 debug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 debug-on-entry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 debug-on-error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 debug-on-quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 debugging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161

default-mode-line-format. . . . . . . . . . . . . . . . . . . . 159

‘default.el’ init file . . . . . . . . . . . . . . . . . . . . . . . . . . 149 Definition installation . . . . . . . . . . . . . . . . . . . . . . . . . . 23 Definition writing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Definition, how to change . . . . . . . . . . . . . . . . . . . . . . . 23 defun . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 defvar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 delete-region. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 Deleting text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 describe-function . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 describe-function, introduced. . . . . . . . . . . . . . . . . 37 Digression into C. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 directory-files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 Division . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 Duplicated words function . . . . . . . . . . . . . . . . . . . . . 169

E

edebug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 edit-options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 edit-options, introduced . . . . . . . . . . . . . . . . . . . . . . 73 Else . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 Emacs version, choosing . . . . . . . . . . . . . . . . . . . . . . . 157 ‘empty list’ defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 eobp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 eq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 eq (example of use) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 equal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 Erasing text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 Error for symbol without value . . . . . . . . . . . . . . . . . . . 8 Error message generation . . . . . . . . . . . . . . . . . . . . . . . . 4 etags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 ‘evaluate’ defined. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Evaluating inner lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Evaluation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Evaluation practice. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ‘expression’ defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

F

Falsehood and truth in Lisp. . . . . . . . . . . . . . . . . . . . . 32 Find a File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 Find function documentation. . . . . . . . . . . . . . . . . . . . 37 Find source of function . . . . . . . . . . . . . . . . . . . . . . . . . 37 find-tags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 Flowers in a field . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Focusing attention (narrowing) . . . . . . . . . . . . . . . . . . 55 ‘form’ defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Formatting convention. . . . . . . . . . . . . . . . . . . . . . . . . . 42 Formatting help. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 forward-paragraph . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 forward-sentence . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 ‘function’ defined. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ‘function definition’ defined . . . . . . . . . . . . . . . . . . 21 Function definition installation . . . . . . . . . . . . . . . . . . 23 Function definition writing . . . . . . . . . . . . . . . . . . . . . . 21 Function definition, how to change . . . . . . . . . . . . . . 23 Functions, primitive . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

G

Generate an error message . . . . . . . . . . . . . . . . . . . . . . . 4 Getting a buffer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Global set key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 global-set-key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 global-unset-key . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 Graph prototype . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Graph, printing all . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 graph-body-print . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 graph-body-print Final version. . . . . . . . . . . . . . . . 192

H

Handling the kill ring. . . . . . . . . . . . . . . . . . . . . . . . . . 171 Help typing lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Horizontal axis printing. . . . . . . . . . . . . . . . . . . . . . . . 187

I

if . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 ‘if-part’ defined. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 indent-tabs-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 Indentation for formatting . . . . . . . . . . . . . . . . . . . . . . 42 Initialization file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 Initializing a variable . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 Inner list evaluation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 insert-buffer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 insert-buffer-substring . . . . . . . . . . . . . . . . . . . . . . 40 Insidious type of bug . . . . . . . . . . . . . . . . . . . . . . . . . . 197 Install a Function Definition . . . . . . . . . . . . . . . . . . . . 23 Install code permanently. . . . . . . . . . . . . . . . . . . . . . . . 26 int-to-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 interactive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 ‘interactive function’ defined . . . . . . . . . . . . . . . . . 15 Interactive functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 Interactive options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 interactive, example use of . . . . . . . . . . . . . . . . . . . 46 Interpreter, Lisp, explained . . . . . . . . . . . . . . . . . . . . . . 3 Interpreter, what it does . . . . . . . . . . . . . . . . . . . . . . . . . 5

K

Key setting globally . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 Key unbinding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 Keymaps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 Keyword . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 Kill ring handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 Kill ring overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 kill-append. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 kill-region. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 Killing text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

L

lambda . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 lengths-list-file . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 lengths-list-many-files. . . . . . . . . . . . . . . . . . . . . 132 let . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 let expression sample . . . . . . . . . . . . . . . . . . . . . . . . . . 28 let expression, parts of . . . . . . . . . . . . . . . . . . . . . . . . . 27 let variables uninitialized . . . . . . . . . . . . . . . . . . . . . . 28 Library, as term for ‘file’ . . . . . . . . . . . . . . . . . . . . . . . . 37 line-to-top-of-window . . . . . . . . . . . . . . . . . . . . . . . 156 Lisp Atoms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Lisp history. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . x Lisp interpreter, explained . . . . . . . . . . . . . . . . . . . . . . . 3 Lisp interpreter, what it does . . . . . . . . . . . . . . . . . . . . 5 Lisp Lists. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 list-buffers, rebound . . . . . . . . . . . . . . . . . . . . . . . 154 list-options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 Lists in a computer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 load-library. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 load-path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 Loading files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 ‘local variable’ defined . . . . . . . . . . . . . . . . . . . . . . . 27 Local variables list, per-buffer, . . . . . . . . . . . . . . . . . 151 Location of point. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 looking-at . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 Loops . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 Loops and recursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

M

Maclisp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . x Mail aliases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 make-string. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 mapcar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 mark-whole-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 match-beginning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 message . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 Mode line format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 Mode selection, automatic . . . . . . . . . . . . . . . . . . . . . 151 Motion by sentence and paragraph . . . . . . . . . . . . . 101

N

Narrowing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 ‘narrowing’ defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 nil . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 nil, history of word . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 nreverse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 nthcdr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62, 73

O

occur . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 optional . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 Optional arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 Options for interactive . . . . . . . . . . . . . . . . . . . . . . . 25 Options introduced. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Options, easily settable . . . . . . . . . . . . . . . . . . . . . . . . 150 or . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 other-buffer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

P

Paragraphs, movement by . . . . . . . . . . . . . . . . . . . . . 101 Parts of let expression . . . . . . . . . . . . . . . . . . . . . . . . . 27 Passing information to functions. . . . . . . . . . . . . . . . . . 8 Pasting text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 Patterns, searching for. . . . . . . . . . . . . . . . . . . . . . . . . 101 Per-buffer, local variables list . . . . . . . . . . . . . . . . . . 151 Permanent code installation. . . . . . . . . . . . . . . . . . . . . 26 point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 ‘point’ defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Point location. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Point, mark, buffer preservation . . . . . . . . . . . . . . . . . 33 Practicing evaluation . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Preserving point, mark, and buffer . . . . . . . . . . . . . . 33 Primitive functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Primitives written in C . . . . . . . . . . . . . . . . . . . . . . . . . 21 Print horizontal axis . . . . . . . . . . . . . . . . . . . . . . . . . . 187 Print vertical axis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 print-elements-of-list . . . . . . . . . . . . . . . . . . . . . . . 89

print-elements-recursively . . . . . . . . . . . . . . . . . . 96

print-graph Final version. . . . . . . . . . . . . . . . . . . . . 193 print-graph varlist . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 print-X-axis. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190

print-X-axis-numbered-line . . . . . . . . . . . . . . . . . 190

print-X-axis-tic-line . . . . . . . . . . . . . . . . . . . . . . . 189 print-Y-axis. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 Printing the whole graph . . . . . . . . . . . . . . . . . . . . . . 191 prog1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 progn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Program, running one . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Prototype graph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141

R

re-search-forward . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 Read-only buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 Readying a graph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Rebinding keys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 Recursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 Recursion and loops . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 recursive-count-words . . . . . . . . . . . . . . . . . . . . . . . 124

recursive-graph-body-print . . . . . . . . . . . . . . . . . 147
recursive-lengths-list-many-files . . . . . . . . . . 134

Recursively counting words . . . . . . . . . . . . . . . . . . . . 120 regexp-quote. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 Region, what it is . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 Regular expression searches . . . . . . . . . . . . . . . . . . . . 101 Regular expressions for word counting . . . . . . . . . . 115 Remainder function, % . . . . . . . . . . . . . . . . . . . . . . . . . 183 Repetition (loops) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 Repetition for word counting. . . . . . . . . . . . . . . . . . . 115 Retrieving text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 reverse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 Ring, making a list like a . . . . . . . . . . . . . . . . . . . . . . 171 rotate-yank-pointer. . . . . . . . . . . . . . . . . . . . . . 85, 171 Run a program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

S

Sample let expression . . . . . . . . . . . . . . . . . . . . . . . . . . 28 save-excursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 save-restriction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 search-forward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 Searches, illustrating . . . . . . . . . . . . . . . . . . . . . . . . . . 101 sentence-end. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 Sentences, movement by . . . . . . . . . . . . . . . . . . . . . . . 101 set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 set-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 setcar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 setcdr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 setq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Setting a key globally . . . . . . . . . . . . . . . . . . . . . . . . . 153 Setting value of variable . . . . . . . . . . . . . . . . . . . . . . . . 12 ‘side effect’ defined. . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Simple extension in ‘.emacs’ file. . . . . . . . . . . . . . . . 156

simplified-beginning-of-buffer . . . . . . . . . . . . . . 37

‘site-init.el’ init file . . . . . . . . . . . . . . . . . . . . . . . . 149 ‘site-load.el’ init file . . . . . . . . . . . . . . . . . . . . . . . . 149 Size of buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 sort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 Source level debugger. . . . . . . . . . . . . . . . . . . . . . . . . . 164 Special form . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Special form of defun. . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Storing and cutting text . . . . . . . . . . . . . . . . . . . . . . . . 65 ‘string’ defined. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 switch-to-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Switching to a buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Symbol names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Symbol without value error . . . . . . . . . . . . . . . . . . . . . . 8 Symbolic expressions, introduced . . . . . . . . . . . . . . . . . 1 Syntax categories and tables . . . . . . . . . . . . . . . . . . . 126

T

Tabs, preventing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 ‘TAGS’ file, create own . . . . . . . . . . . . . . . . . . . . . . . . . 112 TAGS table, specifying . . . . . . . . . . . . . . . . . . . . . . . . . 37 Text between double quotation marks. . . . . . . . . . . . . 2 Text Mode turned on. . . . . . . . . . . . . . . . . . . . . . . . . . 151 Text retrieval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 the-the . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 ‘then-part’ defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 top-of-ranges . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 triangle-bugged . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 triangle-recursively . . . . . . . . . . . . . . . . . . . . . . . . . 97 Truth and falsehood in Lisp. . . . . . . . . . . . . . . . . . . . . 32 Types of data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

U

Unbinding key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 Uninitialized let variables . . . . . . . . . . . . . . . . . . . . . . 28

V

Variable initialization. . . . . . . . . . . . . . . . . . . . . . . . . . . 72 Variable number of arguments. . . . . . . . . . . . . . . . . . . 10 Variable, setting value . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ‘varlist’ defined. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 Version of Emacs, choosing . . . . . . . . . . . . . . . . . . . . 157 Vertical axis printing . . . . . . . . . . . . . . . . . . . . . . . . . . 182

W

what-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 while . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 Whitespace in lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Whole graph printing. . . . . . . . . . . . . . . . . . . . . . . . . . 191 Widening. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Widening, example of . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Word counting in a defun . . . . . . . . . . . . . . . . . . . . . 125 Words and symbols in defun . . . . . . . . . . . . . . . . . . . 125 Words, counted recursively. . . . . . . . . . . . . . . . . . . . . 120 Words, duplicated. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 Writing a function definition . . . . . . . . . . . . . . . . . . . . 21 Wrong type of argument . . . . . . . . . . . . . . . . . . . . . . . . 10

X

X axis printing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 X-axis-element . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189

Y

Y axis printing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 Y-axis-column . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 Y-axis-column Final version. . . . . . . . . . . . . . . . . . . 191 Y-axis-label-spacing . . . . . . . . . . . . . . . . . . . . . . . . 184 Y-axis-tic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 yank . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85, 176 yank-pop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178

Z

zap-to-char. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 zerop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172

About the Author

Robert J. Chassell has worked with GNU Emacs since 1985. He writes and edits, teaches Emacs and Emacs Lisp, and is a director and the Secretary/Treasurer of the Free Software Foundation, Inc. He has an abiding interest in social and economic history and flies his own airplane.

Short Contents

Preface. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ix 1 List Processing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Practicing Evaluation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3 How To Write Function Definitions . . . . . . . . . . . . . . . . . . . . . 21 4 A Few Buffer–Related Functions . . . . . . . . . . . . . . . . . . . . . . . 37 5 A Few More Complex Functions . . . . . . . . . . . . . . . . . . . . . . . 45 6 Narrowing and Widening . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 7 car, cdr, cons: Fundamental Functions . . . . . . . . . . . . . . . . 59 8 Cutting and Storing Text. . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 9 How Lists are Implemented . . . . . . . . . . . . . . . . . . . . . . . . . . 81 10 Yanking Text Back . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 11 Loops and Recursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 12 Regular Expression Searches . . . . . . . . . . . . . . . . . . . . . . . . . 101 13 Counting: Repetition and Regexps . . . . . . . . . . . . . . . . . . . . 115 14 Counting Words in a defun. . . . . . . . . . . . . . . . . . . . . . . . . 125 15 Readying a Graph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 16 Your ‘.emacs’ File. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 17 Debugging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 18 Conclusion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 Appendix A The the-the Function . . . . . . . . . . . . . . . . . . . . . 169 Appendix B Handling the Kill Ring . . . . . . . . . . . . . . . . . . . . . . 171 Appendix C A Graph with Labelled Axes . . . . . . . . . . . . . . . . . . 181 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 ii Programming in Emacs Lisp

Table of Contents

Preface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ix

For Whom This is Written . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ix Lisp History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . x A Note for Novices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . x Thank You. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . xi

1 List Processing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Lisp Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.1.1 Lisp Atoms. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.1.2 Whitespace in Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.1.3 GNU Emacs Helps You Type Lists . . . . . . . . . . . . . . . . . . . 3

1.2 Run a Program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3 Generate an Error Message. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.4 Symbol Names and Function Definitions. . . . . . . . . . . . . . . . . . . . . . . 4 1.5 The Lisp Interpreter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

1.5.1 Byte Compiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.6 Evaluation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

1.6.1 Evaluating Inner Lists. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.7 Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

1.7.1 Error Message for a Symbol Without a Value . . . . . . . . . 8 1.8 Arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

1.8.1 Arguments’ Data Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 1.8.2 An Argument as the Value of a Variable or List . . . . . . . 9 1.8.3 Variable Number of Arguments . . . . . . . . . . . . . . . . . . . . . 10 1.8.4 Using the Wrong Type Object as an Argument. . . . . . . 10 1.8.5 The message Function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

1.9 Setting the Value of a Variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 1.9.1 Using set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 1.9.2 Using setq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 1.9.3 Counting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

1.10 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 1.11 Exercises . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

2 Practicing Evaluation . . . . . . . . . . . . . . . . . . . . . . . . 15

2.1 Buffer Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.2 Getting Buffers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 2.3 Switching Buffers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.4 Buffer Size and the Location of Point . . . . . . . . . . . . . . . . . . . . . . . . 18 2.5 Exercise . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

3 How To Write Function Definitions . . . . . . . . . . . 21

3.1 The defun Special Form . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 3.2 Install a Function Definition. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

3.2.1 Change a Function Definition . . . . . . . . . . . . . . . . . . . . . . . 23 3.3 Make a Function Interactive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

3.3.1 An Interactive multiply-by-seven. . . . . . . . . . . . . . . . . . 25 3.4 Different Options for interactive. . . . . . . . . . . . . . . . . . . . . . . . . . . 25 3.5 Install Code Permanently . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 3.6 let . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

3.6.1 The Parts of a let Expression . . . . . . . . . . . . . . . . . . . . . . 27 3.6.2 Sample let Expression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 3.6.3 Uninitialized Variables in a let Statement . . . . . . . . . . . 28

3.7 The if Special Form . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

3.7.1 The type-of-animal Function in Detail . . . . . . . . . . . . . 30 3.8 If–then–else Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 3.9 Truth and Falsehood in Lisp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 3.10 save-excursion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

3.10.1 Template for a save-excursion Expression. . . . . . . . . 33 3.11 Review . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 3.12 Exercises . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

4 A Few Buffer–Related Functions. . . . . . . . . . . . . . 37

4.1 Finding More Information. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.2 A Simplified beginning-of-buffer Definition . . . . . . . . . . . . . . . . 37 4.3 The Definition of mark-whole-buffer . . . . . . . . . . . . . . . . . . . . . . . 39

4.3.1 Body of mark-whole-buffer . . . . . . . . . . . . . . . . . . . . . . . 39 4.4 The Definition of append-to-buffer . . . . . . . . . . . . . . . . . . . . . . . . 40

4.4.1 The append-to-buffer Interactive Expression . . . . . . . 41 4.4.2 The Body of append-to-buffer . . . . . . . . . . . . . . . . . . . . 41

      4.4.3 save-excursion in append-to-buffer . . . . . . . . . . . . . . 42

4.5 Review . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 4.6 Exercises . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

5 A Few More Complex Functions . . . . . . . . . . . . . . 45

5.1 The Definition of copy-to-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 5.2 The Definition of insert-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

5.2.1 The Interactive Expression in insert-buffer . . . . . . . . 46 A Read-only Buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 ‘b’ in an Interactive Expression . . . . . . . . . . . . . . . . . . . . . 46

5.2.2 The Body of the insert-buffer Function . . . . . . . . . . . 47 5.2.3 insert-buffer With an if Instead of an or . . . . . . . . . 47 5.2.4 The or in the Body . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 5.2.5 The let Expression in insert-buffer . . . . . . . . . . . . . . 49

5.3 Complete Definition of beginning-of-buffer . . . . . . . . . . . . . . . . 50 5.3.1 Optional Arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 5.3.2 beginning-of-buffer with an Argument . . . . . . . . . . . 51

What happens in a large buffer . . . . . . . . . . . . . . . . . . . . . 51 What happens in a small buffer . . . . . . . . . . . . . . . . . . . . 52

5.3.3 The Complete beginning-of-buffer . . . . . . . . . . . . . . . 53 5.4 Review . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 5.5 &optional Argument Exercise. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

6 Narrowing and Widening . . . . . . . . . . . . . . . . . . . . 55

6.1 The save-restriction Special Form . . . . . . . . . . . . . . . . . . . . . . . . 55 6.2 what-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 6.3 Exercise with Narrowing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

7 car, cdr, cons: Fundamental Functions . . . . . . . . 59

7.1 car and cdr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 7.2 cons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60

7.2.1 Find the Length of a List: length . . . . . . . . . . . . . . . . . . 61 7.3 nthcdr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 7.4 setcar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 7.5 setcdr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 7.6 Exercise . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

8 Cutting and Storing Text . . . . . . . . . . . . . . . . . . . . 65

8.1 zap-to-char. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 8.1.1 The interactive Expression . . . . . . . . . . . . . . . . . . . . . . . 66 8.1.2 The Body of zap-to-char. . . . . . . . . . . . . . . . . . . . . . . . . . 66 8.1.3 The search-forward Function. . . . . . . . . . . . . . . . . . . . . . 67 8.1.4 The progn Function. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 8.1.5 Summing up zap-to-char. . . . . . . . . . . . . . . . . . . . . . . . . . 68 8.1.6 The Version 18 Implementation . . . . . . . . . . . . . . . . . . . . . 68

The body of the progn expression . . . . . . . . . . . . . . . . . . 70

8.2 kill-region. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 8.3 delete-region: A Digression into C. . . . . . . . . . . . . . . . . . . . . . . . . 71 8.4 Initializing a Variable with defvar. . . . . . . . . . . . . . . . . . . . . . . . . . . 72 8.5 copy-region-as-kill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

8.5.1 The Body of copy-region-as-kill . . . . . . . . . . . . . . . . . 74 The kill-append function . . . . . . . . . . . . . . . . . . . . . . . . . 75 The else-part of copy-region-as-kill . . . . . . . . . . . . . 76

8.6 Review . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 8.7 Searching Exercises . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

9 How Lists are Implemented . . . . . . . . . . . . . . . . . . 81

9.1 Exercise . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

10 Yanking Text Back . . . . . . . . . . . . . . . . . . . . . . . . . 85

10.1 Kill Ring Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 10.2 The kill-ring-yank-pointer Variable. . . . . . . . . . . . . . . . . . . . . 85 10.3 Exercises with yank and nthcdr . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86

11 Loops and Recursion . . . . . . . . . . . . . . . . . . . . . . . 87

11.1 while . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 11.1.1 A while Loop and a List . . . . . . . . . . . . . . . . . . . . . . . . . . 87 11.1.2 An Example: print-elements-of-list . . . . . . . . . . . . 89 11.1.3 A Loop with an Incrementing Counter. . . . . . . . . . . . . . 89

Example with incrementing counter. . . . . . . . . . . . . . . . . 90 The parts of the function definition . . . . . . . . . . . . . . . . . 91 Putting the function definition together . . . . . . . . . . . . . 92

11.1.4 Loop with a Decrementing Counter . . . . . . . . . . . . . . . . 93 Example with decrementing counter . . . . . . . . . . . . . . . . 93 The parts of the function definition . . . . . . . . . . . . . . . . . 94 Putting the function definition together . . . . . . . . . . . . . 94

11.2 Recursion. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 11.2.1 Recursion with a List . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 11.2.2 Recursion in Place of a Counter. . . . . . . . . . . . . . . . . . . . 97

An argument of 3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 11.2.3 Recursion Example Using cond . . . . . . . . . . . . . . . . . . . . 99

11.3 Looping Exercise . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

12 Regular Expression Searches . . . . . . . . . . . . . . . 101

12.1 The Regular Expression for sentence-end . . . . . . . . . . . . . . . . . 101 12.2 The re-search-forward Function. . . . . . . . . . . . . . . . . . . . . . . . . 102 12.3 forward-sentence . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

The while loops . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 The regular expression search . . . . . . . . . . . . . . . . . . . . . 105

12.4 forward-paragraph: a Goldmine of Functions. . . . . . . . . . . . . . 105 The let* expression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 The forward motion while loop . . . . . . . . . . . . . . . . . . . 108 Between paragraphs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 Within paragraphs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 No fill prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 With a fill prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111

12.5 Create Your Own ‘TAGS’ File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 12.6 Review . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 12.7 Exercises with re-search-forward . . . . . . . . . . . . . . . . . . . . . . . . 114

13 Counting: Repetition and Regexps . . . . . . . . . 115

13.1 The count-words-region Function. . . . . . . . . . . . . . . . . . . . . . . . 115

13.1.1 The Whitespace Bug in count-words-region. . . . . . 117 13.2 Count Words Recursively . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 13.3 Exercise: Counting Punctuation . . . . . . . . . . . . . . . . . . . . . . . . . . . 124

14 Counting Words in a defun . . . . . . . . . . . . . . . . . 125

14.1 What to Count?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 14.2 What Constitutes a Word or Symbol?. . . . . . . . . . . . . . . . . . . . . . 126 14.3 The count-words-in-defun Function . . . . . . . . . . . . . . . . . . . . . 127 14.4 Count Several defuns Within a File . . . . . . . . . . . . . . . . . . . . . . . 129 14.5 Find a File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 14.6 lengths-list-file in Detail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 14.7 Count Words in defuns in Different Files . . . . . . . . . . . . . . . . . . 132

14.7.1 The append Function . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 14.8 Recursively Count Words in Different Files . . . . . . . . . . . . . . . . . 133 14.9 Prepare the Data for Display in a Graph . . . . . . . . . . . . . . . . . . . 134

14.9.1 Sorting Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 14.9.2 Making a List of Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

15 Readying a Graph . . . . . . . . . . . . . . . . . . . . . . . . . 141

15.1 The graph-body-print Function . . . . . . . . . . . . . . . . . . . . . . . . . . 145 15.2 The recursive-graph-body-print Function. . . . . . . . . . . . . . . 147 15.3 Need for Printed Axes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 15.4 Exercise . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148

16 Your ‘.emacs’ File . . . . . . . . . . . . . . . . . . . . . . . . . 149

16.1 Site-wide Initialization Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 16.2 Setting Variables for One Session . . . . . . . . . . . . . . . . . . . . . . . . . . 150 16.3 Beginning a ‘.emacs’ File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 16.4 Text and Auto Fill Mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 16.5 Mail Aliases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 16.6 Indent Tabs Mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 16.7 Some Keybindings. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 16.8 Loading Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 16.9 Autoloading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 16.10 A Simple Extension: line-to-top-of-window . . . . . . . . . . . . 156 16.11 Keymaps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 16.12 X11 Colors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 16.13 V19 Miscellaneous. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 16.14 A Modified Mode Line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159

17 Debugging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161

17.1 debug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 17.2 debug-on-entry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 17.3 debug-on-quit and (debug) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 17.4 The edebug Source Level Debugger . . . . . . . . . . . . . . . . . . . . . . . . 164 17.5 Debugging Exercises . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166

18 Conclusion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167

Appendix A The the-the Function. . . . . . . . . . . . . 169

Appendix B Handling the Kill Ring . . . . . . . . . . . 171

B.1 The rotate-yank-pointer Function . . . . . . . . . . . . . . . . . . . . . . . 171

B.1.1 The Body of rotate-yank-pointer . . . . . . . . . . . . . . . 172 The else-part of the if expression . . . . . . . . . . . . . . . . . 172 The % remainder function . . . . . . . . . . . . . . . . . . . . . . . . . 174 Using % in rotate-yank-pointer . . . . . . . . . . . . . . . . . 175 Pointing to the last element . . . . . . . . . . . . . . . . . . . . . . . 175

B.2 yank . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 Passing the argument. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 Passing a negative argument . . . . . . . . . . . . . . . . . . . . . . 177

B.3 yank-pop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178

Appendix C A Graph with Labelled Axes. . . . . . 181

C.1 The print-graph Varlist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 C.2 The print-Y-axis Function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182

C.2.1 Side Trip: Compute a Remainder . . . . . . . . . . . . . . . . . . 183 C.2.2 Construct a Y Axis Element . . . . . . . . . . . . . . . . . . . . . . 184 C.2.3 Create a Y Axis Column . . . . . . . . . . . . . . . . . . . . . . . . . . 186 C.2.4 Final Version of print-Y-axis . . . . . . . . . . . . . . . . . . . . 186

C.3 The print-X-axis Function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187

C.3.1 X Axis Tic Marks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 C.4 Printing the Whole Graph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191

C.4.1 Testing print-graph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 C.4.2 Graphing Numbers of Words and Symbols . . . . . . . . . 194

A lambda Expression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 The mapcar Function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 Another Bug … Most Insidious. . . . . . . . . . . . . . . . . . . 197

C.4.3 The Printed Graph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201

viii Programming in Emacs Lisp